### v2

Instrumented Difference-in-Differences with heterogeneous treatment effectsthanks: I am grateful to Daiji Kawaguchi, Ryo Okui, and Shosei Sakaguchi for their continued guidance and support. All errors are my own.
Sho Miyaji Graduate School of Economics, The University of Tokyo, 7-3-1 Hongo, Bunkyoku, Tokyo 113-0033, Japan; Email: shomiyaji-apple@g.ecc.u-tokyo.ac.jp.
(May 24, 2024)
Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques. This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV). In a simple setting with two periods and two groups, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome. In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET). In contrast to Fuzzy DID design considered in de Chaisemartin and D’Haultfœuille (2018), our DID-IV design does not ex-ante require strong restrictions on the treatment adoption behavior across units, and our target parameter, the LATET, is policy-relevant if the instrument is based on the policy change of interest to the researcher. We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units, which we call staggered DID-IV designs. We propose an estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. We illustrate our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom. In this application, the two-way fixed effects instrumental variable regression, which is the conventional approach to implement staggered DID-IV designs, yields a negative estimate, whereas our estimation method indicates the substantial gain from schooling.
Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
JEL Classification: C21, C23, C26
1   Introduction
Instrumental variable (IV) strategies are widely used for causal inference across many fields in economics. Despite its popularity, empirical researchers commonly face two challenges in employing the IV method in practice. The first challenge is the internal validity of the IV estimate; it is generally difficult to find an instrument that is randomized across units in practice, especially in observational studies. The second challenge is the external validity of the IV estimate; under heterogeneous treatment effects, a linear IV estimand captures the average treatment effect, for those who are affected by the instrument, and this local average treatment effect (LATE) may not be policy-relevant if the instrument does not represent the policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
To enhance at least the external validity, many studies have exploited variation arising from the different timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques. For example, Duflo (2001) estimates returns to schooling in Indonesia, leveraging variation arising from the new school construction program across regions and cohorts as an instrument for education attainment. As another example, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, using variation occurring from the different timing of implementation of school reform across municipalities and cohorts as an instrument for parents’ education attainment. The underlying identification strategies behind these studies, however, are inconsistent with IV designs: these policy changes are not randomized across units in reality, thereby potentially threatening the internal validity of the IV estimates. Rather, the underlying identification strategies are similar to those in difference-in-differences (DID) designs, under which we identify the effect of the policy change (treatment) on the outcome, exploiting the variation of that policy across units and over time.
In this paper, we combine the IV techniques with DID designs and provide a new econometric framework, which we call an instrumented difference-in-differences (DID-IV). In this design, we can achieve both the internal and external validity of the IV estimate if the parallel trends assumptions in the treatment and the outcome are plausible in a given application, and the instrument is based on the policy shift of interest to the researcher. We formalize the target parameter and identifying assumptions in this design, and develop the credible estimation method that is robust to treatment effect heterogeneity in multiple time periods settings.
First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument over two periods (unexposed group), whereas some units become exposed at the second period (exposed group). In this setting, our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects among the units who belong to an exposed group and are induced to treatment by instrument in the second period. Our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups. This framework comes from a simple and motivating observation: in the two-group/two-period ($2\times 2$) settings, a popular estimand scales the DID estimand of the outcome by the DID estimand of the treatment between the two groups, which we call the Wald-DID estimand. Indeed, we show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
In the canonical DID-IV designs, the interpretation of the parallel trends assumption in the outcome is less clear than that of the parallel trends assumption in DID designs because it does not depend on the untreated outcome, but depends on the outcome under no exposure to the instrument, which we call the unexposed outcome. Motivated by this observation, we next clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs. To do so, we first uncover the selection mechanism behind the canonical DID-IV designs. More precisely, we show that in DID-IV settings, time also plays the role of instrument, that is, time also affects one’s treatment choices over time in the absence of policy shock. Following to the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers. For instance, the time compliers do not adopt the treatment in the first period, but become adopt it in the second period without instrument. We then show that the average time trend in the outcome without instrument equals a weighted average of the indirect effect of time on the outcome through treatment in each time group. We call this an expected time gain and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
de Chaisemartin and D’Haultfœuille (2018) (hereafter, ”dCDH” in this section) also consider the same setting in this article, but formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. The main difference between this article and dCDH is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects among the units who belong to an exposed group and start to receive the treatment in the second period. Accordingly, the identifying assumptions in dCDH are also different from those in this article. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET if the treatment effect is stable over time, and propose the alternative estimands that are free from the strong restriction on treatment effect heterogeneity.
Motivated by the disagreement between this article and dCDH, we next compare DID-IV designs to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs. First, we show that dCDH’s identifying assumptions ex-ante impose strong restrictions on the treatment adoption behavior across units, and these restrictions are asymmetric between exposed and unexposed groups. Next, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one parameter measures the treatment effects among the sub-population of the compliers and the other parameter measures the treatment effects among the time compliers. This decomposition result has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET. Second, the SLATET is generally not policy-relevant even when the researchers exploit the policy change of their interest as an instrument for the treatment.
We also revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and demonstrate that while their argument is correct, it is misguided. We first show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome in DID designs. We then argue that this assumption is unsuitable for DID-IV settings for three reasons. First, this assumption generally fails to capture the average time trends of the outcome even in an unexposed group. This is because in DID-IV settings, units are allowed to receive the treatment in the absence of the policy shock (instrument). Second, in most applications of the DID-IV method, it is generally infeasible to impute the untreated potential outcomes (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires that the data contain the units with zero education attainment, which is unrealistic in practice. Finally, in DID-IV settings, we can not directly test this assumption using the pre-exposed period data in general. Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the plausibility of this assumption.
Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units. In this setting, units start exposed to the instrument at a particular point in time and remain exposed to that instrument afterward. Indeed, in most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation. The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g. Oreopoulos (2006), Lundborg et al. (2014), and Meghir et al. (2018)), the phase-in introduction of head starts across states (e.g. Johnson and Jackson (2019)), or the gradual adoption of broadband internet programs across municipalities (e.g. Akerman et al. (2015), Bhuller et al. (2013)). We call the underlying identification strategies staggered DID-IV designs, and establish the target parameter and identifying assumptions. Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument. We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ from the initial exposure to the instrument. Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings and state the interpretation of each assumption.
We extend our DID-IV designs in several directions. We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment. We also consider extensions to repeated cross sections, and triple DID-IV designs. Lastly, we consider the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs in the presence of heterogeneous treatment effects. In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable regressions. In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time. Our proposed method avoids this issue and is robust to treatment effect heterogeneity. Our estimation procedure consists of two steps: we first subset the data for each CLATT and next run the two stage least squares regression in each data set. We call this a stacked two stage least squares (STS) regression and ensure its validity. Following to Callaway and Sant’Anna (2021), we also propose the weighting scheme to construct summary causal parameters from each CLATT, and discuss the procedure of pretrends tests in staggered DID-IV designs.
We illustrate our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the United Kingdom, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment. In this application, we first assess the plausibility of DID-IV identification assumptions implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. We find that the TWFEIV estimate is strictly negative and not significantly different from zero. Actually, the TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting: we treat the already exposed units as controls during the period after both regions were affected by the policy change. Finally, we use our estimation method to reassess returns to schooling in the U.K. We find that our STS estimates are all positive in each relative period after the school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate. Specifically, our weighting estimate indicates the roughly $20\%$ gain from schooling in the U.K.
Overall, this paper provides a new research design by combining the IV techniques with DID designs. Our DID-IV design can be applied to various empirical settings. First, it can cope with the general adoption process of the treatment when we are interested in the effect of the treatment on the outcome in DID designs. For instance, consider the case that some units already receive the treatment in period 00 before the policy shock in period $1$, which distributes the treatment. In this setting, the canonical DID design is infeasible in practice unless we ex ante discard these units from the data. On the other hand, our DID-IV design can be directly applied to this setting and allows us to identify the LATET. In other words, we can view that DID-IV is DID with the noncompliance of the assigned treatment.
Second, our DID-IV design can yield valid counterfactual group to estimate the treatment effect when we have no control group under DID designs. For instance, consider the case that treatment is education attainment and we are interested in returns to schooling as in Duflo (2001). In this setting, the canonical DID design generally fails to construct the control group, as the data would lack the units with zero education attainment in reality. On the other hand, our DID-IV design can yield exposed and unexposed groups if one can exploit the policy shock across units as an instrument for the treatment. Then, by comparing the evolution of the treatment and the outcome between the two groups, one can identify the LATET under this design.
Finally, our DID-IV design can overcome the endogeneity issue in empirical work. In practice, it is common that the treatment adoption is endogenous over time, leading to a failure in identifying the treatment effect, for instance, via simple linear regressions or two-way fixed effects regressions. Our DID-IV design can cope with the common identification challenges by leveraging variation arising from the policy shock across units as an instrument for the treatment. Importantly, in this design, we do not require the random assignment of the policy shock across units; rather, we require only the plausibility of the parallel trends assumptions in the treatment and the outcome, which would enhance both the internal and external validity of the IV estimate in a given application.
The rest of the paper is organized as follows. The next subsection discusses the related literature. Section 2 establishes DID-IV designs in a simple setting with two periods and two groups. Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs. Section 4 contains extensions. Section 5 presents our estimation method. Section 6 presents our empirical application. Section 7 concludes. All proofs are given in the Appendix.
1.1   Related literature
Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018)). We contribute to this literature in three important ways.
The first contribution of this paper is to investigate the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH” in this section). In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption. Hudson et al. (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings. Build on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs; the settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
In this paper, we first formalize $2\times 2$ DID-IV designs and complement de Chaisemartin (2010) and Hudson et al. (2017). Specifically, we introduce the path of the instrument in potential outcomes and uncover the ”hidden assumption” in $2\times 2$ DID-IV designs; this requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument. We also investigate the selection mechanism behind $2\times 2$ DID-IV designs, and clarify the interpretation of the parallel trends assumption in the outcome.
Given our identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs. The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicate the relative advantages of DID-IV over Fuzzy DID designs.
The second contribution of this paper is to extend $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units. In practice, in order to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for the treatment in more than two periods (e.g., Black et al. (2005), Lundborg et al. (2014), and Johnson and Jackson (2019)). No previous studies, however, extend $2\times 2$ DID-IV designs to such settings. Our staggered DID-IV designs ensure the theoretical validity of their identification strategies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects. When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g. Black et al. (2005), Lundborg et al. (2014)). In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time. Our estimation method developed in this paper would serve as an alternative to the TWFEIV estimators and make DID-IV designs more credible in a given application.
Our paper is also related to a recent and growing DID literature. Several recent papers have pointed out the issue of implementing DID designs via two-way fixed effects regressions or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al. (2021); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)). All of these papers have documented that regression coefficients by the conventional approaches may fail to properly summarize treatment effects under DID designs. Some of these papers have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity. None of these papers, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in the timing of policy adoption across units as an instrument for the treatment. In this paper, we formalize the underlying identification strategy as DID-IV designs, and provide the reliable estimation method. Built on the work in this paper, Miyaji (2024) studies the properties of the TWFEIV estimand under staggered DID-IV designs, and point out the issue of implementing this design via TWFEIV regressions.
2   DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in a two-period/two-group setting. We first establish the target parameter and identifying assumptions. We then uncover the selection mechanism behind the canonical DID-IV design, and clarify the interpretation of the parallel assumption in the outcome. We next describe the relationship between DID and DID-IV, and show that DID-IV is DID with the noncompliance of the assigned treatment. At the end of this section, we also describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
2.1   Set up
We introduce the notation we use throughout section 2. We consider a panel data setting with two periods and $N$ units. For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support. For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receives the treatment in period $t$. Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$. Throughout section 2, we assume that $\{Y_{i,0},Y_{i,1},D_{i,0},D_{i,1},Z_{i,0},Z_{i,1}\}_{i=1}^{N}$ are independent and identically distributed (i.i.d).
We introduce the path of the treatment and the instrument. Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ be the instrument path. We assume that no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$. Hereafter, we refer to this as a sharp assignment of the instrument. We denote $E_{i}\in\{0,1\}$ to be the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group). In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment: we assume that the treatment path can take four values with non-zero probability, that is, we have $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or a policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment. For instance, Duflo (2001) estimates returns to schooling in Indonesia, using a new school construction program across regions as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. Hereafter, we refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
We make a no carryover assumption on the potential outcomes $Y_{i,t}(d,z)$.
Assumption 1 (No carryover assumption).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),Y_{i,0}(d,z)=Y_{i,0}(d_{0},z),Y_{i,1}(d,z)=Y_{i,1}(d_{1},z),$
where $d=(d_{0},d_{1})$ is the generic element of the treatment path $D_{i}$.
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several recent papers impose this assumption with settings of a non-staggered treatment; see, e.g., de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021). [^1] When treatment is binary and there exist only two periods, one can explicitly introduce the treatment path in potential outcomes because it takes only four values. We therefore consider this case in Appendix B as an extension. One can also weaken Assumption 1 in other cases such as non-binary, ordered treatments or multiple time periods, but this would requires the cumbersome notation and complicates the definition of our target parameter, thus is beyond scope of this paper.
Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$, which describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$. Specifically, the first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group. Following to the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ to be the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ to be the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ to be the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ to be the defiers.
Henceforth, we keep Assumption 1. In the next section, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
2.2   Target parameter in two time periods
Our target parameter in $2\times 2$ DID-IV designs is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
Definition.
The local average treatment effect on the treated (LATET) in period $t=1$ is
$LATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$. In the DID-IV literature, de Chaisemartin (2010) consider the same target parameter, and Hudson et al. (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$. The LATET has been also proposed in heterogeneous effects IV models with binary instrument; see, e.g., Słoczyński et al. (2022) and Słoczyński (2020).
We define the LATET as our target parameter for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)). Duflo (2001), for instance, are interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”. To answer this question, the author exploits a new school construction program across regions as an instrument for education attainment. In this context, the LATET would be a policy-relevant parameter in the presence of heterogeneous treatment effects. Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT). In section 2.5, we show that both causal parameters measure the treatment effects among the units who are affected by instrument (policy shock) and belong to an exposed group $(E_{i}=1)$.
Remark 1.
de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting we are considering here222de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings. We therefore rewrite their target parameter by the notation we use in this article..
Definition.
The switcher local average treatment effect on the treated (SLATET) is
$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
where the switcher (SW) is the units who become treated in period $1$.
This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW). As we will show in Appendix E, this parameter is different from the LATET and can be decomposed into a weighted average of two different causal parameters. In section 2.6, we provide a detailed discussion of the connections between this article and de Chaisemartin and D’Haultfœuille (2018).
2.3   Identification assumptions in two time periods
This section establishes DID-IV designs in a simple setting with two groups and two periods. In this setting, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
$w_{DID}$ $=\frac{E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]}{E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]}.$
Following to the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand. This estimand is obtained from the IV regression of the outcome on the treatment where the group and post-time dummies are included instruments and the interaction of the two is excluded instrument.
In practice, the Wald-DID estimand $w_{DID}$ is commonly interpreted as measuring the local average treatment effect (LATE) in the presence of heterogeneous treatment effects, but the underlying identifying assumptions are not stated formally. Duflo (2001), for instance, notes that ”if returns to education are not constant, the 2SLS estimates are a weighted average of the returns to education for people who are affected by the instruments (Angrist and Imbens (1995))”.
We consider the following identifying assumptions for the Wald-DID estimand to capture the LATET. Henceforth, we assume that the denominator of the Wald-DID estimand is different from zero without loss of generality.
Assumption 2 (Exclusion restriction for potential outcomes).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{0,1\},Y_{i,t}(d,z)=Y_{i,t}(d).$
This assumption requires that the instrument path does not directly affect potential outcomes other than through treatment. This assumption is common in the IV literature; see e.g., Imbens and Angrist (1994) and Abadie (2003). In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) impose the similar assumption without introducing the instrument path.
Given Assumption 1 and Assumption 2, we can write the observed outcomes $Y_{i,t}$ as
$Y_{i,t}=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0).$
Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path were $z$:
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes. Note that we can not observe different potential outcomes for the same unit at the same time. For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument over two time periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
Assumption 3 (Monotonicity Assumption in period $t=1$).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 8.53581pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way. This assumption implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) make the same assumption. Hereafter, we consider the type of monotonicity assumption which rules out the existence of the defiers $DF^{Z}$.
Assumption 4 (No anticipation in the first stage).
[…]
Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group. This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex-ante not known for all the units in an exposed group. This assumption is ”hidden assumption” in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al. (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways. Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume the no anticipation for all units $i$. Roth et al. (2023) take the intermediate approach: they assume the no anticipation for the treated units. Our no anticipation assumption on potential treatment choices is in line with that of Roth et al. (2023).
Next, we impose the parallel trends assumptions in the treatment and the outcome. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) assume the similar assumptions.
Assumption 5 (Parallel Trends Assumption in the treatment).
$E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=0]=E[D_{i,1}((0,0))-D_{i,0}((0,0))|E_{i}=1].$
Assumption 5 is a parallel trends assumption in the treatment. This assumption requires that the expectation of the treatment between exposed and unexposed groups would have followed the same path if the assignment of the instrument had not occurred. For instance, in Duflo (2001), this assumption requires that the evolution of mean education attainment would have been the same between exposed and unexposed groups if the policy shock had not occurred during two periods.
Assumption 6 (Parallel Trends Assumption in the outcome).
$E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]=E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
Assumption 6 is a parallel trends assumption in the outcome. This assumption requires that the evolution of the unexposed outcome is the same on average between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of the policy shock.
When empirical researchers exploit variation arising from policy shock as an instrument for the treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6. For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on ”the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the local average treatment effects on the treated (LATET) in period $1$.
Theorem 1.
If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$, that is,
$w_{DID}=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
holds.
Proof.
See Appendix A. ∎
We now provide the intuition behind $2\times 2$ DID-IV designs. As we already noted, the Wald-DID estimand scales the DID estimand of the outcome by the DID estimand of the treatment. Motivated by this observation, in $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1). First, in the denominator, we use the DID identification strategy in order to identify the direct effect of the instrument on the treatment: we recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment). Next, in the numerator, we use the DID identification strategy in order to identify the indirect effect of the instrument on the outcome through treatment: we recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome). Finally, by scaling the indirect effect obtained from the numerator by the direct effect obtained from the denominator, we can capture the LATET in period $t=1$.
Figure 1: DID-IV identification strategy in two periods and two groups. Notes: This figure plots the evolution of the expectation of the treatment (First stage DID) and the outcome (Reduced form DID) for exposed and unexposed groups respectively. The dotted line represents the counterfactual trends of the expectation of the treatment and the outcome in an exposed group. The effects of the instrument on the treatment and the outcome are $0.2$ and $2$ respectively, and the local average treatment effect on the treated is $10$.
2.4   Interpreting the parallel trends assumption in the outcome
So far, we have formalized the target parameter and identifying assumptions in $2\times 2$ DID-IV designs. Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand in order to identify the effects of the instrument on the treatment and the outcome. The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
In this section, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs. To do so, we first uncover the selection mechanism behind the canonical DID-IV designs. Specifically, we show that in $2\times 2$ DID-IV designs, time also plays the role of instrument, and investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. We then show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment. We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path were $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
We recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods. In this setting, one can see that time also plays the role of instrument, that is, time also affects one’s treatment choices during the two periods.
Similar to the group variable $G_{i}^{Z}$ introduced in section 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ to be the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ to be the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ to be the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ to be the time defiers. The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods. In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument). On the other hand, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument, that is, their treatment choices are affected by time in the absence of the policy shock.
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, the types painted in gray color in both tables.
In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but can not observe $D_{1}((0,0))$ for each unit (see Table 1). Here, we use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but can not observe $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
Table 1: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=1$ $D_{1}((0,0))=0$
$D_{0}((0,0))=1,D_{1}((0,1))=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}((0,0))=1,D_{1}((0,1))=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 2: Unexposed group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$ $D_{1}((0,1))=0$
$D_{0}((0,0))=1,D_{1}((0,0))=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}((0,0))=1,D_{1}((0,0))=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Given these tables, we now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively:
$\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
$\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
We note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$. At the same time, we also note that $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation. For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$. Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of the policy shock for time group $g^{t}$ in group $E_{i}=e$:
$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$
$\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$
$\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument for the time compliers in an unexposed group, respectively.
Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$): we can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit. On the other hand, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservables in an exposed group ($e=1$) because we can not specify the time group of unit $i$ in that group. For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we can not distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes respectively:
$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},$
$\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (in other words, the stationary conditions hold), these quantities are equal to zero. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms:
$\Delta_{CM^{T},e}$ $=E[Y_{i,1}(0)-Y_{i,0}(0)|CM^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(1)-Y_{i,1}(0)|CM^{T},E_{i}=e]}_{\text{Selection gain among $CM^{T}$}},$
$\Delta_{DF^{T},e}$ $=E[Y_{i,1}(1)-Y_{i,0}(1)|DF^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(0)-Y_{i,1}(1)|DF^{T},E_{i}=e]}_{\text{Selection gain among $DF^{T}$}}.$
The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the expectation of the time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively. The second term represents the selection gain in each time group in group $E_{i}=e$, respectively. The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
The following theorem establishes the relationship between $\Delta_{e}$ and $\Delta_{g^{t},e}$.
Theorem 2.
Suppose Assumptions 1-4 hold, then
$\Delta_{e}=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e},$
where the weight $w_{g^{t},e}$ is non-negative and sum to one, that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
Proof.
See Appendix A. ∎
Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ is equal to a weighted average of the indirect effect of time on the outcome through treatment for each time group $g^{t}$ in group $e$. The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$. In this paper, we call $\Delta_{e}$ an expected time gain in group $e$ hereafter.
In light of the view discussed so far, we can conclude that the parallel trends assumption in the outcome (Assumption 6), i.e., $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups, that is, the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that of DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period. If there exist pre-periods before the exposure to the instrument, we can indirectly check the validity of Assumption 6 as well as Assumption 5, comparing the evolution of the outcome between exposed and unexposed groups. The formal discussion for the pretrend test is provided in section 5.3.
2.5   Relationship between DID and DID-IV
In this section, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with the noncompliance of the assigned treatment.
To see this, we first introduce the notation in DID designs according to section 2.1. Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program, which distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$. In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group). Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group. Hereafter, in this section, we refer to treatment and control groups as exposed and unexposed groups, respectively.
We note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units. First, we assume that all the units are not allowed to receive the treatment in period 00:
$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (1)
Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and should not receive the treatment in period $1$ if that unit belongs to an unexposed group:
$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (2)
These restrictions imply that we have Tables 3-4 in DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions. First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs. Assumption 3 follows from (2) and Assumption 4 follows from (1). Assumption 6 follows from combining (1) with (2).
Second, the ATT, the target parameter in DID designs, is equal to the LATET in the canonical DID set up. This follows from Tables 3 - 4:
Table 3: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
Table 4: Treatment group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
$=LATET.$
We note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$). We therefore define the LATET as our target parameter in DID-IV designs in section 2.2.
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Table 3 and Table 4, we can rewrite the parallel trends assumption in the untreated outcome as follows:
$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
$\iff$ $\Delta_{0}=\Delta_{1},$
where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively in DID designs:
$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
The above discussions clarify that DID-IV is a natural generalization of DID, and the difference between the two designs arises because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$. In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
2.6   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) (Henceforth, ”dCDH” in this section) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting in this article. However, they formalize $2\times 2$ DID-IV designs differently, and call them Fuzzy DID. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID. The detailed discussion is contained in Appendix E, and this section provides the main points of that discussion.
The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see section 2.2 in this article). Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs. Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
In Appendix E, we first show that dCDH’s identifying assumptions ex ante impose strong and asymmetric restrictions on the treatment adoption behavior across units between exposed and unexposed groups. Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Table 5 and Table 6333Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6.. These tables are in line with Table 1 and Table 2 in this article, and clarify the difference between dCDH’s restrictions and ours: we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by monotonicity assumption, whereas dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group. This difference is ciritical for empirical application: in contrast to the monotonicity assumption, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application.
Table 5: Exposed group ($z=1$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 6: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. Specifically, we show that the SLATET is a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $CM^{Z}\land AT^{T}$, which is the sub-population of the compliers $CM^{Z}$, in an exposed group. This decomposition result has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET. Second, this parameter may not be policy-relevant even when the researchers use the policy change of their interest as an instrument for the treatment because it generally contains the treatment effects among the time compliers $CM^{T}$ who are affected by time but not affected by the instrument.
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided. In this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups respectively for the Wald-DID estimand to identify the SLATET.
First, we show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome in DID designs. Recall that in DID-IV designs, we do not assume the parallel trends assumption in the untreated outcome, but assume the parallel trends assumption in the unexposed outcome. Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption generally fails to capture the average time trend of the outcome in an unexposed group. We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET. In the proof of this argument, we also demonstrate that the strong restrictions imposed by dCDH on treatment adoption behavior is also induced by imposing the parallel trends assumption in the untreated outcome.
Next, we argue that the parallel trends assumption in the untreated outcome is unsuitable for DID-IV settings for three reasons. First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group. Second, in most applications of the DID-IV method, we can not impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al. (2005)). For instance, in Black et al. (2005), this assumption requires that the data contain the parents with zero education attainment, which is unrealistic in practice. Finally, in DID-IV settings, we can not generally test this assumption using pre-exposed period data, as some units can adopt the treatment before period 00. Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the validity of this assumption. Surprisingly, we also show that their placebo test coincides with our pretrends test, which we will introduce in section 5.3 for assessing the plausibility of Assumption 5 and Assumption 6, with only a slight modification.
3   DID-IV in multiple time periods
We now extend the $2\times 2$ DID-IV design to multiple periods settings. Specifically, we consider the setting with the staggered adoption of the instrument across units: the exposure to the instrument occurs across units at different times and once units become exposed to the instrument, units remain exposed to that instrument afterward. We call this a staggered DID-IV design, and establish the target parameter and identifying assumptions.
3.1   Set up
We introduce the notation we use throughout section 3 to section 5. We consider a panel data setting with $T$ periods and $N$ units. For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome and $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status. Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$. Throughout section 3 to section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
We make the following assumption about the assignment process of the instrument.
Assumption 7 (Staggered adoption for $Z_{i,t}$).
For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
Assumption 7 requires that once units start exposed to the instrument, units remain exposed to that instrument afterward. In the DID literature, this assumption is made for the adoption process of the treatment and sometimes called ”staggered treatment adoption”, see, e.g., Athey and Imbens (2022), Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
Exploiting variation arising from the staggered adoption of the instrument across units is pervasive in economics. For instance, Black et al. (2005) estimate the causal relationship between parents’ and children’s education by exploiting the different timing of implementation of school reforms across municipalities as an instrument for parental education.
Given Assumption 7, we can uniquely characterize one’s instrument path by the initial exposure date of the instrument, which we denote $E_{i}=\min\{t:Z_{i,t}=1\}$. If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$. Based on the initial exposure period $E_{i}$, we can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start exposed to the instrument in period $E_{i}=e$.
Similar to the two periods setting in section 2.1, we allow the general adoption process for the treatment: treatment can potentially turn on/off repeatedly over time. de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) consider the same setting in the recent DID literature.
Next, we introduce the potential outcomes framework in multiple time periods. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$.
Assumption 7 allows us to rewrite $D_{i,t}(z)$ by the initial exposure date $E_{i}=e$. Let $D_{i,t}^{e}$ denote the potential treatment status in period $t$ if unit $i$ is first exposed to the instrument in period $e$. Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument. Hereafter, we call $D_{i,t}^{\infty}$ the ”never exposed treatment”. Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
$D_{i,t}=D_{i,t}^{\infty}+\sum_{1\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$. Hereafter, we refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage. In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
Next, we introduce the group variable which describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$. Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$. Specifically, the first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$. Following to the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ to be the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ to be the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ to be the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ to be the defiers in period $t$ and the initial exposure date $e$.
Finally, we make a no carryover assumption on potential outcomes $Y_{i,t}(d,z)$.
Assumption 8 (No carryover assumption).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d_{t},z).$
Henceforth, we keep Assumption 7 and 8. In the next section, we define the target parameter in staggered DID-IV designs.
3.2   Target parameter in staggered DID-IV designs
In staggered DID-IV designs, our target parameter is the cohort specific local average treatment effect on the treated (CLATT) defined below.
Definition.
The cohort specific local average treatment effect on the treated (CLATT) at a given relative period $l$ from the initial adoption of the instrument is
$CLATT_{e,e+l}$ $=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
Each CLATT is a natural generalization of the LATET in section 2.2 and suitable for the setting of the staggered instrument adoption. This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$, that is, who are induced to treatment by instrument in period $e+l$. Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$ and the compliers $CM_{e,e+l}$. We note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$. For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
This parameter would be policy-relevant if researchers leverage the policy change of their interest as an instrument for the treatment. Suppose for instance that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment. In this example, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest because the compliers in each relative period after a policy change represents the dynamic effect of that policy on the take up rates of the program, and this parameter can be interpreted as estimating the effect of participation in job training on real earnings among those affected in each relative period.
3.3   Identification assumptions in staggered DID-IV designs
In this section, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
Assumption 9 (Exclusion Restriction in multiple time periods).
$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d)\hskip 8.53581pta.s.$
Assumption 9 extends the exclusion restriction in two time periods (Assumption 2) to multiple period settings. This assumption requires that the instrument path does not directly affect the potential outcome for all time periods and its effects are only through treatment.
Given Assumption 8 and Assumption 9, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$. Following to section 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$:
$Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$ respectively. The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed to the instrument. Hereafter, we refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
Assumption 10 (Monotonicity Assumption in multiple time periods).
$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{S}(E_{i})\hskip 5.69054pt\text{and for all}\hskip 5.69054ptl\geq 0.$
This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$. We recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$. Assumption 10 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$. This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$. Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
Assumption 11 (No anticipation in the first stage).
[…]
Assumption 11 requires that potential treatment choices in any $l$ period before the initial exposure to the instrument is equal to the never exposed treatment. This assumption is a natural generalization of Assumption 4 to multiple period settings and restricts the anticipatory behavior before the initial exposure to the instrument.
Assumption 12 (Parallel Trends Assumption in the treatment in multiple time periods).
[…]
Assumption 12 is a parallel trends assumption in the treatment in multiple periods and multiple cohorts. This assumption requires that the trends of the treatment across cohorts would have followed the same path, on average, if there is no exposure to the instrument. Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs: these studies impose the same type of parallel trends assumption on untreated outcomes with settings of multiple periods and multiple cohorts.
Assumption 13 (Parallel Trends Assumption in the outcome in multiple time periods).
[…]
Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and multiple cohorts. This assumption requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the assignment of the instrument had not occurred. From the discussion in section 2.4, we can interpret that this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
4   Extensions
This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections. It also includes the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs. For more details and the proofs in this section, see Appendix B.
Non-binary, ordered treatment
Up to now, we have considered only the case of a binary treatment. However the same idea can be applied when treatment takes a finite number of ordered values, that is, we have $D_{i,t}\in\{0,1,\dots,J\}$. When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
Definition.
The average causal response on the treated (ACRT) is
$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
where the weight $w_{j}$ is:
$w_{j}=\frac{Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}.$
The average causal response on the treated (ACRT) is a weighted average of the effects of a unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument. The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), but the difference here is that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the average causal response on the treated (ACRT) under the same assumptions in Theorem 1.
Theorem 3.
If Assumptions 1-6 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
$W_{DID}$ $=\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1]$
$=ACRT.$
Proof.
See Appendix B.1. ∎
If we have a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), which is the natural generalization of the ACRT. The identifying assumptions are the same in section 3.3.
Definition.
The cohort specific average causal response on the treated (CACRT) at a given relative period $l$ from the initial adoption of the instrument is
$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
where the weight $w^{e}_{e+l,j}$ is:
$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}.$
Triple DID-IV designs
[…]
Repeated cross sections
In some applications, researchers have no access to panel data and use repeated cross section data [^5] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for the treatment as in Duflo (2001).. In Appendix B.2, we consider repeated cross section data settings and present the identification assumptions in DID-IV designs.
Introducing the treatment path in $2\times 2$ DID-IV designs
In section 2.1, we impose Assumption 1 (No carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment. We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment. In this case, our target parameter slightly changes with the cumbersome notation, formally defined below.
Definition.
The local average treatment effect on the treated with the treatment path (LATET’) is
$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
where $\Delta_{0}$ and $\Delta_{1}$ are:
$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
The weights $w_{0}$ and $w_{1}$ are:
$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
In Appendix B.3, we show that the similar assumptions as in Assumptions 1-6 in section 2.3 are sufficient for the Wald-DID to capture the LATET’.
5   Estimation and Inference
When researchers leverage variation in the timing of policy adoption across units as an instrument for the treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al. (2014), Black et al. (2005), Akerman et al. (2015), and Bhuller et al. (2013)).
Unfortunately, however, recent research has shown that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects. Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative [^6] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.. Miyaji (2024) further shows that the negative weight problem arises if the effect of the instrument on the treatment or the outcome is not stable over time.
In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity. Specifically, we first propose a simple regression-based method for estimating each CLATT. Following to Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT. At the end of this section, we also discuss the pretrends tests for checking the plausibility of parallel trends assumption in the treatment and the outcome.
5.1   Stacked two stage least squares regression
In this section, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Recall that if we have a binary treatment, our target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
$CLATT_{e,e+l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
Our estimation method consists of two steps.
Step 1.
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$).
Researchers should carefully specify the choice set $U$ in practice. If there exists a never exposed cohort $E_{i}=\infty$, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If there exists no never exposed cohort, we can set $U=\max\{E_{i}\}$, i.e., the last exposed cohort. In this case, we can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$. We note that the units in an already exposed cohort $e=1$ are not included in any data sets because we should include the pre-exposed period $e-1$ for estimating each $CLATT_{e,e+l}$.
Step 2.
In each data set, we run the following IV regression and obtain the IV estimator $\hat{\beta}^{e,l}_{IV}$:
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}=\pi^{e,l}_{0}+\pi^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\pi^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\pi^{e,l}(\mathbf{1}\{E_{i}=e\}\cdot\mathbf{1}\{T_{i}=e+l\})+\eta^{e,l}_{i,t},$
where the group indicator $\mathbf{1}\{E_{i}=e\}$ and the post-period indicator $\mathbf{1}\{T_{i}=e+l\}$ are the included instruments and the interaction of the two is the excluded instrument.
Formally, our proposed estimator $\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}$ takes the following form:
$\widehat{CLATT}_{e,e+l}\equiv\hat{\beta}^{e,l}_{IV}=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}},$
where $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are:
$\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
$\hat{\pi}^{e,l}$ $=\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. We note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure above, we call this a stacked two-stage least squares (STS) estimator.
Our STS estimator is related to the DID estimators in staggered DID designs recently proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). Specifically, the DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when we have no covariates and use never treated units as a control group. Their proposed methods avoid the issue of two-way fixed effects estimators in staggered DID designs, and our STS estimator avoids the issue of two-way fixed effects instrumental variable estimators in staggered DID-IV designs.
Theorem 4 below guarantees the validity of our STS estimator.
Theorem 4.
Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal:
[…]
Proof.
See Appendix C. ∎
From Theorem 4, we can also construct the standard error of the STS estimator, using the sample analogue of the asymptotic variance $V(\psi_{i,e,l})$.
Remark 2.
Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data. In Appendix C, we present the influence function of our SLS estimator in repeated cross section settings.
In triple DID-IV designs, one can replace the IV regression in step $2$ with the following one ($g_{i}=A,B$):
$Y_{i,t}$ $=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\beta^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\beta^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\beta^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\beta^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\beta^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
$+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}$ $=\pi^{e,l}_{0}+\pi^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\pi^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\pi^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\pi^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\pi^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\pi^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
$+\pi^{e,l}_{7}\mathbf{1}_{E_{i}=e,T_{i}=e+l,g_{i}=A}+\eta^{e,l}_{i,t},$
where $\mathbf{1}_{A}$ is indicator function and takes one if $A$ is true. In Appendix C, we also present the influence function of our triple DID-IV estimator $\beta_{IV}^{e,l}$.
5.2   Weighting scheme
The previous section provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs. In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$. However, in many cases, researchers may also want to estimate summary causal parameters, which capture the overall effect of a treatment on an outcome. Actually, in staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$. In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
In this section, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
Specifically, we consider the following weighting scheme as in Callaway and Sant’Anna (2021):
$\theta^{IV}=\sum_{e}\sum_{t=1}^{T}w(e,t)\cdot CLATT_{e,t},$
where $w(e,t)$ are some reasonable weighting functions assigned to each $CLATT_{e,t}$.
To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we first define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$. We call this the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$), formally defined below.
Definition.
The cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$) at a given relative period $l$ from the initial adoption of the instrument is
$CAET_{e,e+l}^{1}=E[D_{i,e+l}-D_{i,e+l}^{\infty}|E_{i}=e].$
If treatment is binary, each $CETT_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
$CAET^{1}_{e,e+l}$ $=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]$
$=Pr(CM_{e,e+l}|E_{i}=e).$
In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity. We can employ their framework straightforwardly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
For instance, to aggregate dynamic treatment effects in cohort $e$ over time in staggered DID designs, Callaway and Sant’Anna (2021) consider the following summary measure:
$\theta_{sel}(\tilde{e})=\frac{1}{T-\tilde{e}+1}\sum_{t=\tilde{e}}^{T}CATT(\tilde{e},t),$
where the $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)). Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start exposed to that treatment.
Table 7: Weights for a variety of summary causal parameters
Target Parameter $w(e,t)$
$\theta_{es(l)}^{IV}$ $\mathbf{1}\{e+l\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{es(l,l^{\prime})}^{bal,IV}$ $\mathbf{1}\{e+l^{\prime}\leq T\}\mathbf{1}\{t=e+l\}P(E=e|E+l^{\prime}\leq T)\displaystyle\frac{CAET_{e,e+l}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,e+l}^{1}}$
$\theta_{sel(\tilde{e})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{e=\tilde{e}\}\displaystyle\frac{CAET_{\tilde{e},t}^{1}}{\sum_{t=\tilde{e}}^{T}CAET_{\tilde{e},t}^{1}}$
$\theta_{c(\tilde{t})}^{IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t=\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{c(\tilde{t})}^{cumm,IV}$ $\mathbf{1}\{t\geq e\}\mathbf{1}\{t\leq\tilde{t}\}P(E=e|E\leq t)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{e\in\mathcal{S}(E)}CAET_{e,t}^{1}}$
$\theta_{W}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)/\sum_{e\in\mathcal{S}(E)}\sum_{t=1}^{T}\mathbf{1}\{t\geq e\}P(E=e|E\leq T)$
$\theta_{sel}^{o,IV}$ $\mathbf{1}\{t\geq e\}P(E=e|E\leq T)\displaystyle\frac{CAET_{e,t}^{1}}{\sum_{t=e}^{T}CAET_{e,t}^{1}}$
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter corresponds with that of Callaway and Sant’Anna (2021). We superscript IV to associate each parameter with staggered DID-IV designs.
In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t},$
where the weights are non-negative and sum to one by construction. This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger at the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021). In Table 7, we summarize the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs. We note that our proposed weighting functions are the same if we have non-binary, ordered treatment where our target parameter is the $CACRT_{\tilde{e},t}$.
Estimation and Inference
We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows:
$\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the stacked IV regression in section 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\zeta_{i,e,t}^{w}+o_{p}(1),$
where $\zeta_{i,e,t}^{w}$ represents the influence function, and satisfies $E[\zeta_{i,e,t}^{w}]=0$ and $E[\zeta_{i,e,t}^{w}{\zeta_{i,e,t}^{w}}^{\top}]<\infty$.
Corollary 1 below represents the asymptotic distribution of the plug-in estimator $\hat{\theta}_{IV}$ and ensures its validity.
Corollary 1.
If the assumptions of Theorem 4 hold,
[…]
where the influence function $l_{i}^{\theta_{IV}}$ takes the following form:
$l_{i}^{\theta_{IV}}=\sum_{e}\sum_{t=1}^{T}\left(w(e,t)\cdot\psi_{i,e,t}+\zeta_{i,e,t}^{w}\cdot{CLATT}_{e,t}\right).$
5.3   Pretrend test
In this section, we describe how one can conduct pretrend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
$E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
$\iff$ $E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=0]=E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=1].$ (3)
Similarly, one can assess the plausibility of Assumption 6 between period $t=-1$ and $t=0$ by testing the following null hypothesis
$E[Y_{i,0}-Y_{i,-1}|E_{i}=0]=E[Y_{i,0}-Y_{i,-1}|E_{i}=1]$
$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]=E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (4)
We can generalize these tests to multiple pre-exposed periods settings by using the pretrends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
Remark 3.
When researchers implement the staggered DID-IV design via TWFEIV regressions, they often use the dynamic specifications in the first stage and the reduced form:
$Y_{i,t}=\mu_{i.}+\delta_{t.}+\sum_{l}\alpha_{l}\mathbf{1}\{t-E_{i}=l\}+\epsilon_{i,t},$ (5)
$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\sum_{l}\beta_{l}\mathbf{1}\{t-E_{i}=l\}+\eta_{i,t},$ (6)
where $E_{i}$ represents the initial exposure date. They plot the coefficients $\alpha_{l}$ in equation (5) and $\beta_{l}$ in equation (6) in each relative period $l$, and test whether the coefficients before the initial exposure to the instrument are significantly different from zero or not (e.g., Akerman et al. (2015) and Bhuller et al. (2013)).
However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous. Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and the parallel trends assumption in the untreated outcome are plausible. Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
6   Application
Oreopoulos (2006) estimates returns to schooling using a major education reform in the United Kingdom that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, while the reform was not implemented until 1957 in Northern Ireland. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
In the first part of their analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors [^7] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data” to estimate some local average treatment effects.
Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the two-way fixed effects instrumental variable regression, using the reform as an excluded instrument for education attainment:
$Y_{i,t}=\mu_{n.}+\delta_{.t}+\beta_{IV}D_{i,t}+\epsilon_{i,t},$ (7)
$D_{i,t}=\gamma_{n.}+\zeta_{.t}+\pi Z_{i,t}+\eta_{i,t}.$ (8)
Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change. The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$ respectively. Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect. The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy shock.
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized in reality. Oreopoulos (2006) notes that the reform was implemented with political support, taking into account the cost and benefit of the school reform. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the random variation of the policy change. Actually, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, and potentially vary across cohorts. The TWFEIV estimate, therefore, can potentially fail to aggregate the causal effects and yield a misleading conclusion.
In this section, we first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006). We then estimate the TWFEIV regression in the author’s setting. Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
6.1   Assessing the identifying assumptions in staggered DID-IV designs
We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT) as education attainment is a non-binary, ordered treatment. We can identify each CACRT and its summary measures if the underlying DID-IV identifying assumptions seem plausible.
Exclusion restriction.
It would be plausible, given that the policy reform did not affect log annual earnings other than by increasing education attainment. This assumption may be violated for instance if the reform affected both the quality and quantity of education.
Monotonicity assumption.
It would be automatically satisfied in the author’s setting: the policy reform (instrument) increased the minimum schooling-leaving age from $14$ to $15$, which ensures that there are no defiers during the periods after the policy shock.
No anticipation in the first stage.
It would be plausible that there is no anticipation before the policy reform, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England. This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome. To do so, we use the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively. The results are shown in Figure 2.
Figure 2: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006). Notes: The results for the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy. The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$. The blue lines represent the estimates with pointwise $95\%$ confidence intervals for pre-exposed periods in both panels. These should be equal to zero under the null hypothesis that the parallel trends assumptions in the treatment and the outcome hold. The red lines represent the estimates with pointwise $95\%$ confidence intervals for post-exposed periods in both panels.
Parallel trends assumption in the treatment.
It requires that the expectation of education attainment would have followed the same path between England and North Ireland across cohorts in the absence of the school reform. Panel (a) in Figure 2 plots the results of the interacted two-way fixed effects regression in the first stage along with a $95\%$ pointwise confidence interval. The pre-exposed estimates are not significantly different from zero and indicate the validity of Assumption 12.
Parallel trends assumption in the outcome.
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts if the reform had not occurred. Panel (b) in Figure 2 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 13: though there exists a downward pretrends, the estimates before the exposure to the policy change are not significantly different from zero.
Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all the estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
We note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage. In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
6.2   Illustrating our estimation method
We start by estimating two-way fixed effects instrumental variable regression in the author’s setting. In our estimation, we slightly modify the author’s specification to clearly illustrate the pitfalls of TWFEIV regression. Specifically, Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
The result is shown in Table 8. The TWFEIV estimate is $-0.009$ and it is not significantly different from 00. This implies that on the whole, the returns to schooling in the U.K. may be nearly zero. However, this may be a misleading conclusion: we cannot interpret that the TWFEIV estimand captures the properly weighted average of each CACRT if the effects of the policy reform on education attainment and log annual earnings are not stable across cohorts (Miyaji (2024)).
The TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting: when we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$ because both regions are already exposed to the policy shock during these periods. In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, whereas all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
Figure 3: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006). Notes: The results for returns to schooling under the staggered DID-IV identification strategy. The unexposed group $U$ is North Ireland and the reference period is period $t=1946$. The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
We now employ our proposed method to estimate each CACRT. We first create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units who were aged at $14$ between $1957$ and $1965$ in order to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (C.2) in Appendix C.
Table 8: Returns to schooling in Oreopoulos (2006)
Estimate Standard Error 95% CI
TSLS with fixed effects -0.009 0.039 [-0.085, 0.068]
$\theta_{sel}^{IV}(e)$ 0.240 0.098 [0.047, 0.433]
Notes: Sample size $82790$ observations.
Figure 3 plots the point estimates and the corresponding $95\%$ confidence intervals in each relative period after the school reform. The estimates for each $CACRT_{e,t}$ range from $14\%$ to $38\%$ with wide confidence intervals, and $4$ out of $10$ estimates are statistically significant.
Finally, we estimate the summary causal measure constructed from each $CACRT_{e,t}$. Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$):
$\theta_{sel}^{IV}(e)=\sum_{t=1947}^{1956}\frac{CAET^{1}_{e,t}}{\sum_{t=1947}^{1956}CAET^{1}_{e,t}}CACRT_{e,t}.$
Here, each weight assigned to each $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
The result is shown in Table 8. The estimate is $0.24$ and it is significantly different from zero. The estimated returns to schooling are substantial: this is likely because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family, and have potentially much gain from the school reform.
The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative. Overall, our results indicate that the economic returns of education are substantial in the U.K. and the estimation method matters in staggered DID-IV designs in practice.
7   Conclusion
Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome. In this design, we commonly face two challenges: the internal and external validity of the IV estimate. To overcome at least the external validity, many studies leverage variation occurring from the different timing of policy adoption across units as an instrument for the treatment, and use instrumental variable techniques. The instrument, however, generally fails to satisfy the independence assumption in practice, and the underlying identification strategy is parallel those in DID designs.
In this paper, we provide a new econometric framework, which we call an instrumented difference-in-differences (DID-IV). We start by considering a simple setting with two periods and two groups. In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that under $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). We also clarify the interpretation of the parallel trends assumption in the outcome. Specifically, we show that in DID-IV designs, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups. Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
Next, we consider the DID-IV design in the case where there are more than two time periods and units become exposed to the instrument at different times. We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions. Specifically, our target parameter in staggered DID-IV designs is the cohort specific local average treatment effects on the treated (CLATT). The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
We also provide the estimation method in staggered DID-IV designs, which does not require strong restrictions on treatment effect heterogeneity. In practice, most DID-IV applications implement the staggered DID-IV design by running TWFEIV regressions. In companion paper (Miyaji (2024)), however, we show that the conventional approach is inadequate for summarizing the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time. Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects. We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pretrends test to assess the validity of the parallel trends assumption in the treatment and the outcome.
Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the United Kingdom, exploiting the timing variation of the introduction of school reforms between British and North Ireland. First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible. Next, we estimate TWFEIV regressions and our stacked two stage least squares regressions in each dataset. We find that the TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive. Finally, we estimate the summary causal parameter by exploiting our STS estimates. The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the U.K.
Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment. To avoid the issue of using TWFEIV estimators in this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity. Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
Appendix A Proofs in section 2
Proof of Theorem 1.
Proof.
First we consider the denominator of the Wald-DID estimand. For the denominator, we have
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$ (9)
+$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1)$
$=$ $Pr(CM^{Z}|E=1).$ (10)
where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 (no anticipation in the first stage) and the forth equality follows from Assumption 3 (Monotonicity assumption) and Assumption 5 (Parallel Trends Assumption in the treatment).
Next, to show that the Wald-DID estimand equals to the LATET, it suffices to show that
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$ $=E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1)$
$=LATET\cdot Pr(CM^{Z}|E=1).$
For the numerator, we have
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[(D_{1}((0,1))-D_{1}((0,0)))\cdot(Y_{1}(1)-Y_{1}(0))|E=1]$
$=$ $E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1).$
$=$ $LATET\cdot Pr(CM^{Z}|E=1).$ (11)
The first equality follows from Assumption 1 (No carryover assumption) and Assumption 2 (Exclusion restriction). The second equality follows from the simple manipulation. The third equality follows from Assumption 4. The forth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
Combining the result (11) with (10), we obtain
$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=1]}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1)}{Pr(CM^{Z}|E=1)}$
$=LATET.$
Completing the proof. ∎
Proof of Theorem 2.
Recall that we define $\Delta_{e}$ as follows:
$\Delta_{e}\equiv E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=e].$
Then, we obtain:
$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
Here the first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$. Clearly, the weight $w_{g^{t},e}$ sum to one by construction.
Completing the proof.
Appendix B Extensions
B.1   Non-binary, ordered treatment
Proof of Theorem 3.
Proof.
Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$. We note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction. Here, we can rewrite the observed outcome $Y_{t}$ as follows:
$Y_{t}$ $=Z_{1}\cdot Y_{t}(D_{t}((0,1)))+(1-Z_{1})\cdot Y_{t}(D_{t}((0,0)))$
=$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
where we use Assumption 1 and 2.
For the numerator in the Wald DID estimand, we have:
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1].$
Here the first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
In terms of the $\lambda_{Z,j}^{t}$, we write the previous expression as:
$E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
$=E\left[\sum_{j=0}^{J}Y_{1}(j)\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]+Y_{1}(0)\cdot(\lambda_{1,0}^{1}-\lambda_{0,0}^{1})|E=1\right]$
$=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right].$
where the second equality follows from the fact: $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
We note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
$E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1,E=1]Pr(\lambda_{1,j}^{1}-\lambda_{0,j}^{1}=1|E=1)$
$=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|D_{1}((0,1))\geq j>D_{1}((0,0)),E=1]$
×$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (12)
For the denominator in the Wald-DID estimand, by the similar argument, we first obtain:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
Here the first equality follows from the simple algebra and Assumption 4. The second equality follows from Assumption 5.
Then, we have:
$E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
$=E\left[\sum_{j=0}^{J}j\cdot[\lambda_{1,j}^{1}-\lambda_{1,j+1}^{1}-\lambda_{0,j}^{1}+\lambda_{0,j+1}^{1}]|E=1\right]$
$=E\left[\sum_{j=1}^{J}[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
$=\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (13)
Combining the result (13) with (12), we obtain the desired result. ∎
B.2   Repeated cross sections
In this section, we assume repeated cross section settings and present our identification results.
B.2.1 Two time periods
Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$. Let $Z_{i}$ denote the instrument path for unit $i$: $Z_{i}=(0,0)$ if unit $i$ is not exposed to the instrument and $Z_{i}=(0,0)$ if unit $i$ is exposed to the instrument. Let $E_{i}\in\{0,1\}$ denote the group indicator for unit $i$: $E_{i}=1$ if $Z_{i}=(0,1)$ and $E_{i}=0$ if $Z_{i}=(0,0)$. Let $T_{i}\in\{0,1\}$ denote the binary indicator for time. For all $z\in\mathcal{S}(Z)$, let $Y_{i}(0,z),Y_{i}(1,z)$, and $D_{i}(z)$ denote potential outcomes and potential treatment choices for unit $i$. Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In two periods and two groups settings, our target parameter is the local average treatment effect in period $1$:
$LATET$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,CM^{Z}].$
We make the following identification assumptions for the Wald-DID estimand to capture the LATET. These assumptions are suitable for repeated cross section settings.
Assumption 14 (Exclusion restriction).
$\forall z\in\mathcal{S}(Z),Y_{i}(0,z)=Y_{i}(0),Y_{i}(1,z)=Y_{i}(1).$
Assumption 15 (Monotonicity).
$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1.$
Assumption 16 (No anticipation in the first stage).
[…]
Assumption 17 (Parallel Trends assumption in the treatment).
$E[D_{i}((0,0))|E_{i}=1,T_{i}=1]-E[D_{i}((0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[D_{i}((0,0))|E_{i}=0,T_{i}=1]-E[D_{i}((0,0))|E_{i}=0,T_{i}=0].$
Assumption 18 (Parallel Trends assumption in the outcome).
$E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=1,T_{i}=0]$
$=$ $E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=0].$
Theorem 5 below shows that the Wald-DID estimand is equal to the local average treatment effect on the treated (LATET) in period $1$ under Assumptions 14-18. Here, we assume the binary treatment, but non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus we omit it for brevity). In the proof, we omit the index $i$ to ease the notation.
Theorem 5.
Suppose Assumptions 14-18 hold. If we assume a binary treatment, the Wald-DID estimand corresponds to the local average treatment effect on the treated (LATET) in period $1$:
$w_{DID}$ $\equiv\frac{E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}}{E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}}$
$=E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]$
$=LATET.$
Proof.
First we consider the denominator of the Wald-DID estimand. For the denominator, we have
$E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}$
$=$ $E[D((0,1))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}$
$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,0))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1,T=1)$
$=$ $Pr(CM^{Z}|E=1,T=1).$ (14)
The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the forth equality follows from Assumption 15 and Assumption 17.
Next, for the numerator, we have
$E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}$
$=$ $E[Y(D((0,1)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,0)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
$=$ $E[(D((0,1))-D((0,0)))\cdot(Y(1)-Y(0))|E=1,T=1]$
$=$ $E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1,T=1).$
$=$ $LATET\cdot Pr(CM^{Z}|E=1,T=1).$ (15)
The second equality follows from the simple manipulation. The third equality follows from Assumption 16. The forth equality follows from Assumption 18 and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
Combining the result (14) with (15), we obtain
$w_{DID}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1,T=1)}{Pr(CM^{Z}|E=1,T=1)}$
$=LATET.$
Completing the proof. ∎
B.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We assume Assumption 7 in section 3, that is, we assume the staggered assignment of the instrument across units. In addition, we assume Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In multiple time periods and multiple cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
$CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
$=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
When repeated cross section data are available and we have multiple time periods, our staggered DID-IV designs consist of Assumptions 7, 10, 11 (in section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
Assumption 19 (Parallel Trends Assumption in the treatment in multiple time periods).
[…]
Assumption 20 (Parallel Trends Assumption in the outcome in multiple time periods).
[…]
B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
In this section, we discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. We first consider the two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption). We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting. Finally, we show that a Wald-DID estimand identifies the LATET’ (see section 4) under the suitable assumptions similar to the ones in section 2.3. Henceforth, we consider the same setting as in section 2.1 except that we do not adopt Assumption 1.
First, we make the following no anticipation assumptions on the outcome.
Assumption 21 (No anticipation on the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
[…]
This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((0,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land NT^{T}$. We note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al. (2023) (see Table 3 and Table 4 in section 2.5) [^8] Here, it is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
Assumption 22 (No anticipation on the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
[…]
This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((1,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land DF^{T}$. Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
We can rationalize Assumption 21 and 22 by noting that both the two Assumptions and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units who belong to the exposed group and the compliers $CM^{Z}$.
We now make the parallel trends assumption in the outcome that is suitable for the set up we are considering here. We first define the similar notation as in section 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
$\bar{\Delta}_{AT^{T},e}\equiv E[Y_{i,1}((1,1))-Y_{i,0}((1,1))|AT^{T},E_{i}=e],$
$\bar{\Delta}_{NT^{T},e}\equiv E[Y_{i,1}((0,0))-Y_{i,0}((0,0))|NT^{T},E_{i}=e],$
$\bar{\Delta}_{CM^{T},e}\equiv E[Y_{i,1}((0,1))-Y_{i,0}((0,1))|CM^{T},E_{i}=e],$
$\bar{\Delta}_{DF^{T},e}\equiv E[Y_{i,1}((1,0))-Y_{i,0}((1,0))|DF^{T},E_{i}=e].$
Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups respectively:
$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},$
$\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
where the definition of the weight $w_{g^{t},e}$ is same as in section 2.4.
As we previously noted in section 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups. The following assumption indeed make a such restriction in the current set up.
Assumption 23 (Parallel trend assumption in the outcome).
The expected time gain between exposed and unexposed groups is same, that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
The theorem below shows that a Wald-DID estimand is equal to the target parameter $LATET^{\prime}$ under Assumption 2-5 and Assumption 21-23.
Theorem 6.
Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$, that is, we have
$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]}=LATET^{\prime}.$
Proof.
First, we consider the denominator of the Wald-DID estimand. From the similar argument in the proof of Theorem 1, one can show the following result:
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$ $=Pr(CM^{Z}|E=1)$
$=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (16)
Here the second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
$E[Y_{1}-Y_{0}|E=1]$ $=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{0}((1,1))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
$+\bar{\Delta}_{NT^{T},1}w_{NT^{T},1}+\bar{\Delta}_{DF^{T},1}w_{DF^{T},1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
$=\bar{\Delta}_{1}$
$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$ (17)
The first equality follows from Table 1 and Assumption 2. The second equality follows from Assumption 21, Assumption 22 and simple manipulation. The third equality follows from Table 1. The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
We next decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
$E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}((1,1))-Y_{0}((1,1))|E=0,AT^{T}]Pr(AT^{T}|E=0)$
$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=0,CM^{T}]Pr(CM^{T}|E=0)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=0,NT^{T}]Pr(NT^{T}|E=0)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=0,DF^{T}]Pr(DF^{T}|E=0)$
$=\bar{\Delta}_{0}.$ (18)
Here the first equality follows from Table 2 and the second equality follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
From the result (18) with (17) and Assumption 23, we obtain
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1).$ (19)
Combining the result (19) with (16), we obtain the desired result. ∎
Appendix C Proof in section 4
C.1   Proof of Theorem 4
First, we prove the consistency of our SLS estimator. Recall that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
$\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
$\hat{\pi}^{e,l}$ $=\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty, that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty, i.e., $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator
$\widehat{CLATT}_{e,e+l}$ $=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}}$
$=\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}.$
is also well defined.
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. We note that by the Law of Large Numbers and Slutsky’s theorem, we have
[…]
$=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e],$
where the second equality follows from the Law of Iterated Expectations.
By the similar arguments, we have
[…]
From the Slutsky’s theorem, we have
[…]
Lastly, it remains to show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
For the numerator, we have
$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (21)
The third equality follows from Assumption 8, 9 and 11. The forth equality follows from Assumption 13 and the fifth equality follows from Assumption 10.
For the denominator, by the similar calculations,
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=Pr(CM_{e,e+l}|E_{i}=e),$ (22)
where the final equality follows from Assumption 10 and Assumption 12.
Combining the result (21) with (22), we obtain the desirable result:
$\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}$
$=\frac{E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e)}{Pr(CM_{e,e+l}|E_{i}=e)}$
$=CLATT_{e,e+l}.$
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function.
We define $\theta_{e,l}$ to be
$\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
We use the following fact repeatedly in the derivation. This fact is also found in de Chaisemartin and D’Haultfœuille (2018).
Fact 1.
If
$\sqrt{n}(\hat{A}-A)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}+o_{p}(1),\hskip 8.53581pt\sqrt{n}(\hat{B}-B)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}b_{i}+o_{p}(1),$
we have
$\sqrt{n}\left(\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{a_{i}-(A/B)b_{i}}{B}+o_{p}(1).$
First we derive the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{1}+\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}+\alpha_{e,l}^{2}))$. We note that
$\sqrt{n}(E_{N}[\mathbf{1}\{E_{i}=e\}]-E[\mathbf{1}\{E_{i}=e\}])=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\mathbf{1}\{E_{i}=e\}-E[\mathbf{1}\{E_{i}=e\}]).$
and
$\sqrt{n}(E_{N}[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}]-E[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}])$
$=\frac{1}{\sqrt{n}}\sum_{i}^{n}((Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}-E[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]).$
From the observations and using Fact 1, we obtain
$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\alpha_{e,l}^{1})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}+o_{p}(1).$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\zeta_{i,e,l}+o_{p}(1).$
By symmetry, we also obtain the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$:
$\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}+o_{p}(1).$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\phi_{i,e,l}+o_{p}(1).$
This implies that
$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}-\alpha_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\zeta_{i,e,l}-\phi_{i,e,l})+o_{p}(1)$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}a_{i,e,l}+o_{p}(1).$ (23)
Next we derive the influence function of $\sqrt{n}(\hat{\pi}_{e,l}^{1}+\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}+\pi_{e,l}^{2}))$. From the symmetry, by replacing $Y$ with $D$, we obtain the following:
$\sqrt{n}(\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}-\pi_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\eta_{i,e,l}-\xi_{i,e,l})+o_{p}(1)$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}b_{i,e,l}+o_{p}(1),$ (24)
where we define
$\eta_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]},$
$\xi_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}.$
Using (23), (24) and Fact 1, by some calculations, we obtain
$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\sqrt{n}(\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}})$
$=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1),$
where the influence function $\psi_{i,e,l}$ is:
$\psi_{i,e,l}$
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\left(\zeta_{i,e,l}-\phi_{i,e,l}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}(\eta_{i,e,l}-\xi_{i,e,l})\right)$
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (25)
where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
$V(\psi_{i,e,l})$
$=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
C.2   Repeated cross sections
We also present the influence function of our STS estimator when repeated cross section data are available. [^9] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups.. Let $\theta_{e,l}^{r}$ define
$\theta_{e,l}^{r}\equiv\frac{\alpha_{e,e+l}-\alpha_{e,e-1}-[\beta_{U,e+l}-\beta_{U,e-1}]}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]},$
where
$\alpha_{e,t}=E[Y_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\beta_{U,t}=E[Y_{i}|E_{i}\in U,T_{i}=t],$
$\pi_{e,t}=E[D_{i}|E_{i}=e,T_{i}=t],\hskip 5.69054pt\gamma_{U,t}=E[D_{i}|E_{i}\in U,T_{i}=t]\hskip 5.69054pt(t=e+l,e-1).$
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. By the similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
$\psi_{i,e,l}^{r}$ $=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}-[\gamma_{e,l}^{1}-\gamma_{e,l}^{2}]}$
$\times$ $\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg{)},$ (26)
where $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
C.3   Triple DID-IV designs
We also present the influence function of our triple DID-IV estimator when panel or repeated cross section data are available. Let $\theta_{e,l,g}^{p}$ and $\theta_{e,l,g}^{r}$ define
$\theta_{e,l,g}^{p}\equiv\frac{\alpha_{e,A}-\alpha_{e,B}-[\beta_{U,A}-\beta_{U,B}]}{\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]},$
$\theta_{e,l,g}^{r}\equiv\frac{[(\alpha_{e,e+l,A}-\alpha_{e,e-1,A})-(\alpha_{e,e+l,B}-\alpha_{e,e-l,B})]-[(\beta_{U,e+l,A}-\beta_{U,e-1,A})-(\beta_{U,e+l,B}-\beta_{U,e-1,B})]}{[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]},$
where
$\alpha_{e,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}],$
$\pi_{e,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B),$
and
$\alpha_{e,t,g^{\prime}}=E[Y_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,t,g^{\prime}}=E[Y_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}],$
$\pi_{e,t,g^{\prime}}=E[D_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,t,g^{\prime}}=E[D_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B,\hskip 5.69054ptt=e+l,e-1).$
In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function:
$\psi_{i,e,l,g}^{p}=\frac{1}{DID_{e,l,g}^{p}}(\zeta_{i,e,A}-\zeta_{i,e,B}-\zeta_{i,U,A}+\zeta_{i,U,B}),$
where $DID_{e,l,g}^{p}=\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]$, $\delta_{i,e,l,g}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l,g}^{p}\cdot(D_{i,e+l}-D_{i,e-1})$, and
$\zeta_{i,e,g^{\prime}}=\frac{\mathbf{1}_{e,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}=e,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,g^{\prime}}]},\hskip 5.69054pt\zeta_{i,U,g^{\prime}}=\frac{\mathbf{1}_{U,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}\in U,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,g^{\prime}}]}.$
In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function:
$\psi_{i,e,l,g}^{r}$
$=\frac{1}{DID_{e,l,g}^{r}}([\zeta_{i,e,e+l,A}-\zeta_{i,e,e-1,A}-\zeta_{i,e,e+l,B}+\zeta_{i,e,e-1,B}]-[\zeta_{i,U,e+l,A}-\zeta_{i,U,e-1,A}-\zeta_{i,U,e+l,B}+\zeta_{i,U,e-1,B}]),$
where $DID_{e,l,g}^{r}=[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]$, $\delta_{i,e,l,g}^{r}=Y_{i}-\theta_{e,l,g}^{r}\cdot D_{i}$, and
$\zeta_{i,e,t,g^{\prime}}=\frac{\mathbf{1}_{e,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,t,g^{\prime}}]},$
$\zeta_{i,U,t,g^{\prime}}=\frac{\mathbf{1}_{U,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,t,g^{\prime}}]}.$
Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
In this appendix we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Hereafter, let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
$R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}.$
First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^10] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024). In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
$\hat{\beta}_{IV}$ $=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]}$
$=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]\cdot\widehat{WDID}_{e,t}}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
where we define:
$\widehat{WDID}_{e,t}\equiv\frac{\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^11] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
By the similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
[…]
where $CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e]$. Similarly, we have
[…]
Combining the result (30) with (34) and by the Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (35)
where $\Delta_{e,t}$ is:
$\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
$w_{e,t}^{1}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (36)
$w_{e,t}^{2}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (37)
where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and population residuals for England in cohort $t$ respectively.
In Oreopoulos (2006), we can only identify $CLATT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in North Ireland in $1957$. This implies that each $\Delta_{e,t}$ in equation (35) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
Figure 4 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, using equations (30) - (34).
Figure 4 shows that all the CLATTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted. This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
Figure 4: Decomposition result for the TWFEIV estimand in Oreopoulos (2006). Notes: The figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$. The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
Appendix E Comparing DID-IV to Fuzzy DID
In this appendix we compare DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH”). We assume that the reader has briefly read dCDH. We first explain how dCDH’s identifying assumptions are related to ours in this article. We then point out the issues inherent in Fuzzy DID designs. First, we show that dCDH’s assumptions ex ante impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. Next, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. At the end of this appendix, we also revisit the ”issue” pointed out by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
E.1   Comparing dCDH’s identifying assumptions to ours
In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we will refer to Assumptions 14-18 in this article. For now, we ignore the difference in the definition of target parameter between dCDH and this article.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^12] In repeated cross section settings, the formal definition is provided in Assumption 29.. We note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^13] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation).. In the following, we therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^14] Note that in Appendix B.2 in this article, we construct the group variable $E$ from the instrument path $Z$ in order to uncover Assumption 16 (No anticipation in the first stage). For the detail explanations for each assumption, see dCDH.
Assumption 24 (Fuzzy design).
$E[D|Z=1,T=1]>E[D|Z=1,T=0],$ (38)
$E[D|Z=1,T=1]-E[D|Z=1,T=0]>E[D|Z=0,T=1]-E[D|Z=0,T=0].$ (39)
Assumption 25 (Stable percentage of treated units in an unexposed group).
$1\displaystyle 0<E[D|Z=0,T=1]=E[D|Z=0,T=0]<$ 1 .
Assumption 26 (Treatment participation equation).
[…]
Assumption 27 (Parallel Trends assumption in the untreated outcome).
$E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]=E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0].$
Assumption 28 (Stable treatment effect over time).
$E[Y(1)-Y(0)|Z,T=1,D_{0}(Z)=1]=E[Y(1)-Y(0)|Z,T=0,D_{0}(Z)=1].$
In Assumption 28, we modify the dCDH’s notation in order to ease the interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$, which dCDH define as $D(t)$.
We now state the relationships between dCDH’s assumptions and ours. We first explain the similarities. First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 in this article imply this condition [^15] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article. The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002). Here, we point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26, that is, they assume that the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
Finally, dCDH implicitly impose Assumption 14 (Exclusion Restriction) and Assumption 16 (No anticipation in the first stage) in this article when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 26, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
We next describe the differences between dCDH’s assumptions and ours. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 27), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 25), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 26, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction in order for this estimand to identify the LATET. Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
Remark E.1.1.
When Assumption 25 is violated, dCDH consider the following assumption in order for the Wald-DID estimand to capture the SLATET.
Assumption 29 (Homogeneous treatment effect between exposed and unexposed groups).
$SLATET=SLATET^{\prime},$
where we define161616dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)],$
$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)].$
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the ”homogeneous” treatment effect assumption. In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the ”homogeneous” treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 24-28 as dCDH’s main identifying assumptions.
E.2   Strong and asymmetric restrictions in Fuzzy DID designs
In this section, we show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. To do so, we first prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables as in Table 1 and Table 2 in section 2.4 in this article, which present all the types in exposed and unexposed groups respectively in Fuzzy DID designs, and confirm that dCDH’s assumptions ex ante exclude many types in exposed and unexposed groups and its restrictions is heavily assigned on the latter group. At the end of this section, we provide empirical implications of the distinction between dCDH’s restrictions and ours in this article.
First, we show that Assumption 26 (Treatment Participation equation) implies the following assumption.
Assumption 30 (Monotone change in treatment in exposed and unexposed groups).
For all $z\in\{0,1\}$,
[…]
holds.
Here, recall that we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under $T=t$ in group $Z$.
Lemma 1.
Assumption 26 implies Assumption 30.
Proof.
Fix $z\in\{0,1\}$. Then we have $v_{z1}\leq v_{z0}$ or $v_{z1}\geq v_{z0}$.
Therefore, conditioning $Z=z$, we have
[…]
because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$.
Completing the proof. ∎
Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. We note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
As we have seen in section 2.4 in this article, time also plays the role of instrument in DID-IV settings. From this observation, we can view that in an unexposed group ($z=0$), Assumption 30 corresponds to a monotonicity assumption with respect to time, formally defined below.
Assumption 31 (Monotonicity Assumption w.r.t. time in an unexposed group).
[…]
Assumption 31 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$. In terms of the notation introduced in section 2.4 in this article, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability. Hereafter, we consider the type of Assumption 31 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
Next, we show that dCDH’s assumptions imply that there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
Lemma 2.
Suppose Assumption 25, Assumption 26 and Assumption 31 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
Proof.
Without loss of generality, suppose that Assumption 31 excludes the time defiers $DF^{T}$ in an unexposed group ($z=0$).
Then, Assumption 25 implies:
$E[D|T=1,Z=0]=E[D|T=0,Z=0]$
$\iff$ $Pr(D_{1}(0)>D_{0}(0)|Z=0)=0$
$\iff$ $Pr(CM^{T}|Z=0)=0$ (40)
The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (40) implies that there are no time comliers $CM^{T}$ in an unexposed group ($z=0$). ∎
Finally, we show that the condition (38) in Assumption 24 (Fuzzy design) and Assumption 26 imply that in an exposed group ($z=1$), there are no units switching from treatment to non treatment between period 00 and period $1$. In other words, in Assumption 30, we should have:
$Pr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
Lemma 3.
The condition (38) in Assumption 24 and Assumption 26 imply $Pr(D_{0}(1)\leq D_{1}(1)|Z=1)$=1.
Proof.
From the condition (38) in Assumption 24, we have
$E[D|Z=1,T=1]>E[D|Z=1,T=0]$
$\iff$ $E[D_{1}(1)-D_{0}(1)|Z=1]>0$
$\iff$ $Pr(D_{1}(1)>D_{0}(1)|Z=1)-Pr(D_{0}(1)>D_{1}(1)|Z=1)>0,$ (41)
where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Recall that Assumption 26 implies Assumption 30 from Lemma 1. Suppose that $Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1$ holds. Then, we have $Pr(D_{1}(1)>D_{0}(1)|Z=1)=0$. This contradicts with the condition (41). ∎
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 9 and Table 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^17] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: there exist five and three types in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
The distinction between dCDH’s restrictions and ours is apparent when we compare Table 9 and Table 10 to Table 1 and Table 2 in this article [^18] We note that Tables 1-2 in this article is constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
Table 9: Exposed group ($z=1$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 10: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
These differences between dCDH’s restrictions and ours seem critical in many empirical settings: unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, thus the practitioners may want to avoid imposing these restrictions in empirical work. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications. However, these assumptions give rise to two additional issues. First, when we impose the condition (38) in Assumption 24, we should test the plausibility of this condition from the data in practice because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$. Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group in this article) in our fuzzy setting”, they do not explain how we should proceed when this condition is not met in practice.
Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time, and this complicates our estimation procedure in practice. Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility is not clear, especially when we have multiple periods and multiple groups.
E.3   Decomposing the SLATET
In the previous section, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups. In this section, under these restrictions, we show the decomposition result for dCDH’s target parameter, the SLATET, and state the concerning implications.
The following theorem presents the decomposition result for the SLATET under Fuzzy DID designs, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
Theorem 7.
Suppose Assumptions 24 - 28 hold. Then, we can decompose the SLATET into two different causal parameters:
$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=1,T=1,CM^{T}]Pr(CM^{T}|Z=1,SW)$
$+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW),$ (42)
where the switcher (SW) in an exposed group ($Z=1$) is the units who become treated in time $T=1$.
Proof.
The proof directly follows from Table 9 in an exposed group. ∎
Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other parameter measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
This theorem has two concerning implications. First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters. Although the SLATET indeed captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for the treatment. Suppose that the instrument represents the important policy lever for the researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group. In this situation, the LATET is a policy-relevant parameter by construction. On the other hand, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group. As the fraction of the time compliers $CM^{T}$ increases in that group, the SLATET puts more weight on the treatment effects among the time compliers, and this would contaminate the interpretation of the magnitude of this parameter even in such policy-relevant context.
Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 29. When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 in order for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
We show that this assumption can not be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups. To do so, we first note that if we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2. Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
Table 11: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Then, from Table 11, we can interpret the SLATET’ (see Assumption 29) as the treatment effects among the time compliers in an unexposed group in period $1$:
$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)]$
$=E[Y(1)-Y(0)|Z=0,T=1,CM^{T}].$ (43)
By combining equations (42) and (43), we can conclude that Assumption 29 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups. Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group. This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than one might expect.
E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that while their argument is correct, it is misguided. During this discussion, we also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2. Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) in order to identify the SLATET, and propose alternative estimands that do not rely on this assumption. We first show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article. We next argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
First, we clarify the interpretation of Assumption 28 by introducing the additional notation. This assumption is actually being conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population. For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$. Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
Then, from Table 9 and Table 10, we can decompose Assumption 28 in the following:
$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
$\iff\delta_{0,1,AT^{T}}$ $=\delta_{0,0,AT^{T}},$ (44)
$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]=E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
$\iff$ $\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$=$ $\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1},$ (45)
where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (44) shows that in an unexposed group ($z=0$), Assumption 28 requires the stable treatment effect in the time always-takers $AT^{T}$. Equation (45) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
Conditions (44) and (45) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcomes. In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods (see Table 9 and Table 10). Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (44) and (45).
To see this formally, we use our notation to briefly sketch the proof of dCDH on the Wald-DID estimand identifying SLATET under fuzzy DID designs. First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand. We can decompose this quantity into two terms:
$E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$
$=$ $(\delta_{0,1,AT^{T}}-\delta_{0,0,AT^{T}})\cdot w_{AT^{T},0}+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ in an unexposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (44).
Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows:
$E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$
$=$ $\Big{\{}\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
$-\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}-\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}\Big{\}}$
$+$ $\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1),$
where the type SW (switcher) includes the units who switch from non treatment to treatment during the two periods. The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term. The first term disappears if we assume condition (45). Here, we note that we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 25 and Assumption 26, the Wald-DID estimand captures the SLATET under Assumption 27 (parallel trends assumption in the untreated outcome) and conditions (44) and (45). From the discussions so far, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
Here, we add few remarks to clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units in order for the Wald-DID estimand to identify the SLATET. It turns out that this also stems from their reliance on the parallel trends assumption in the untreated outcome.
Remark E.4.1.
dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10). This identification result seems not intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
Remark E.4.2.
dCDH exclude the time compliers $CM^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for that type, we additionally capture the treatment effects among the time compliers $CM^{T}$ from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which implies that the SLATET in an exposed group is offset by the SLATET’ in an unexposed group.
So far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome. Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group; in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and thus we can not observe the untreated outcomes for these types in an unexposed group.
Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general. For instance, Black et al. (2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment. In the author’s setting, the parallel trends assumption in the untreated outcome requires that the data contain the parents with zero education attainment, which is unrealistic in practice.
Finally, this assumption is generally not testable using the pre-exposed period data in a given application. This seems apparent because some units can already adopt the treatment before period 00, and we can not compare the pretrends of the untreated outcome between exposed and unexposed groups.
Although dCDH propose the placebo test in their appendix, we show that it is incomplete for checking the plausibility of the parallel trends assumption in the untreated outcome. dCDH’s test consists of two steps. First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$:
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=0,$ (46)
$E[D|Z=0,T=0]-E[D|Z=0,T=-1]=0.$ (47)
Second, we test the following null hypothesis:
$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]=E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (48)
Although dCDH ”assume” the conditions (46) and (47), dCDH also note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27. We therefore describe the dCDH’s test in two steps.
dCDH’s test has some drawbacks. First, the conditions (46) and (47), which we should check in the first step, seem strong, but dCDH do not explain how we should proceed when these conditions are violated in practice. Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods; it does not compare the evolution of the mean untreated outcome between exposed and unexposed group before period $T=0$. Indeed, dCDH describe the equality (48) as a necessary condition for Assumption 27 and Assumption 28 to be satisfied under the conditions (46) and (47). Finally, it is unclear how we should conduct this test when we have multiple pre-exposed period data. Although dCDH assume ”for instance that data is available for the period $T=-1$”, they do not provide any guidance on how to perform this test in multiple pre-exposed period settings.
Surprisingly, dCDH’s test corresponds to our pretrends test in section 5.3 in this article if we replace the conditions (46)- (47) in the first step with the weaker one:
$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (49)
Under repeated cross section settings, one can figure out that the conditions (48) - (49) coincide with the conditions (3) -(4) in section 5.3 in this article.

### v4

@@
 
-Instrumented Difference-in-Differences with heterogeneous treatment effectsthanks: I am grateful to Daiji Kawaguchi, Ryo Okui, and Shosei Sakaguchi for their continued guidance and support.
+Instrumented Difference-in-Differences with Heterogeneous Treatment Effectsthanks: This paper won the Kanematsu Prize from the Research Institution for Economics and Business Administration, Kobe University, in 2023.
+It also won the Best Young Scholar Award at SETA (Symposium on Econometrics and Applications) Conference, Taipei, in 2024.
+I am grateful to my advisors, Daiji Kawaguchi and Ryo Okui, for their continued guidance and support.
+I am also thankful to Alberto Abadie, Shakeeb Khan, Masamune Iwasawa, Shoya Ishimaru, Shosei Sakaguchi, Takahide Yanagi, and the three anonymous referees of the Kanematsu Prize for their insightful comments.
+I gratefully acknowledge the support provided by the JSPS KAKENHI Grant JP 24KJ0817.
 All errors are my own.
 Sho Miyaji Graduate School of Economics, The University of Tokyo, 7-3-1 Hongo, Bunkyoku, Tokyo 113-0033, Japan; Email: shomiyaji-apple@g.ecc.u-tokyo.ac.jp.
-(May 24, 2024)
-Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques.
+(July 23, 2024)
+Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment.
 This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
-In a simple setting with two periods and two groups, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome.
 In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET).
-In contrast to Fuzzy DID design considered in de Chaisemartin and D’Haultfœuille (2018), our DID-IV design does not ex-ante require strong restrictions on the treatment adoption behavior across units, and our target parameter, the LATET, is policy-relevant if the instrument is based on the policy change of interest to the researcher.
-We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units, which we call staggered DID-IV designs.
-We propose an estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity.
+In contrast to Fuzzy DID design considered in de Chaisemartin and D’Haultfœuille (2018), our DID-IV design does not ex ante require strong restrictions on treatment adoption behavior across units.
+Additionally, our target parameter, the LATET, is policy-relevant if the instrument is based on the policy change of interest to the researcher.
+We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design, and propose an estimation method that is robust to treatment effect heterogeneity.
 We illustrate our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom.
-In this application, the two-way fixed effects instrumental variable regression, which is the conventional approach to implement staggered DID-IV designs, yields a negative estimate, whereas our estimation method indicates the substantial gain from schooling.
+In this application, the two-way fixed effects instrumental variable regression, which is the conventional approach to implement a staggered DID-IV design, yields a negative estimate, whereas our estimation method indicates a substantial gain from schooling.
 Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
@@
 Despite its popularity, empirical researchers commonly face two challenges in employing the IV method in practice.
-The first challenge is the internal validity of the IV estimate; it is generally difficult to find an instrument that is randomized across units in practice, especially in observational studies.
-The second challenge is the external validity of the IV estimate; under heterogeneous treatment effects, a linear IV estimand captures the average treatment effect, for those who are affected by the instrument, and this local average treatment effect (LATE) may not be policy-relevant if the instrument does not represent the policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
-To enhance at least the external validity, many studies have exploited variation arising from the different timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques.
+The first challenge lies in the internal validity of the IV estimate; finding an instrument that is effectively randomized across units, especially in observational studies, proves to be difficult.
+The second challenge lies in the external validity of the IV estimate; under heterogeneous treatment effects, a linear IV estimand captures the average treatment effect, for those who are affected by the instrument (Imbens and Angrist (1994)).
+However, this local average treatment effect (LATE) may not be policy-relevant if the instrument does not represent the policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
+To enhance at least the external validity of the IV estimate, many studies have exploited variation in the timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques.
 For example, Duflo (2001) estimates returns to schooling in Indonesia, leveraging variation arising from the new school construction program across regions and cohorts as an instrument for education attainment.
@@
 (2005) estimates the causal link between parents’ and children’s education attainment, using variation occurring from the different timing of implementation of school reform across municipalities and cohorts as an instrument for parents’ education attainment.
-The underlying identification strategies behind these studies, however, are inconsistent with IV designs: these policy changes are not randomized across units in reality, thereby potentially threatening the internal validity of the IV estimates.
-Rather, the underlying identification strategies are similar to those in difference-in-differences (DID) designs, under which we identify the effect of the policy change (treatment) on the outcome, exploiting the variation of that policy across units and over time.
-In this paper, we combine the IV techniques with DID designs and provide a new econometric framework, which we call an instrumented difference-in-differences (DID-IV).
-In this design, we can achieve both the internal and external validity of the IV estimate if the parallel trends assumptions in the treatment and the outcome are plausible in a given application, and the instrument is based on the policy shift of interest to the researcher.
-We formalize the target parameter and identifying assumptions in this design, and develop the credible estimation method that is robust to treatment effect heterogeneity in multiple time periods settings.
-First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument over two periods (unexposed group), whereas some units become exposed at the second period (exposed group).
-In this setting, our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects among the units who belong to an exposed group and are induced to treatment by instrument in the second period.
+Notably, however, the underlying identification strategy behind these studies is inconsistent with IV designs: the instrument based on these policy changes is not randomized across units in reality, potentially threatening the internal validity of the IV estimate.
+Rather, the underlying identification strategy is similar to difference-in-differences (DID) designs, under where the effect of the policy shock (treatment) on the outcome is identified by exploiting the variation of that policy change across units and over time.
+In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
+We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity in multiple time periods settings.
+Importantly, our DID-IV design does not require the randomization assumption of the instrument.
+In this design, we can achieve the internal and external validity of the IV estimate if the parallel trends assumptions in the treatment and the outcome are plausible in a given application, and the instrument is based on the policy shift of interest to the researcher.
+First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument over two periods (unexposed group), whereas some units become exposed in the second period (exposed group).
+In this setting, our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period.
 Our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups.
-This framework comes from a simple and motivating observation: in the two-group/two-period ($2\times 2$) settings, a popular estimand scales the DID estimand of the outcome by the DID estimand of the treatment between the two groups, which we call the Wald-DID estimand.
-Indeed, we show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
-In the canonical DID-IV designs, the interpretation of the parallel trends assumption in the outcome is less clear than that of the parallel trends assumption in DID designs because it does not depend on the untreated outcome, but depends on the outcome under no exposure to the instrument, which we call the unexposed outcome.
-Motivated by this observation, we next clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs.
-To do so, we first uncover the selection mechanism behind the canonical DID-IV designs.
-More precisely, we show that in DID-IV settings, time also plays the role of instrument, that is, time also affects one’s treatment choices over time in the absence of policy shock.
-Following to the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers.
+This framework comes from a simple and motivating observation: in the two-group/two-period ($2\times 2$) settings, a popular estimand, which we call the Wald-DID estimand, scales the DID estimand of the outcome by the DID estimand of the treatment between the two groups.
+We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
+In DID-IV designs, the interpretation of the parallel trends assumption in the outcome is less clear than that of the parallel trends assumption in DID designs because it does not depend on the untreated outcome.
+Rather, it depends on the outcome under no exposure to the instrument, which we call an unexposed outcome.
+Motivated by this observation, we then clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
+Specifically, we first uncover the selection mechanism behind DID-IV designs, showing that, in DID-IV settings, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of the policy shock.
+Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers.
 For instance, the time compliers do not adopt the treatment in the first period, but become adopt it in the second period without instrument.
-We then show that the average time trend in the outcome without instrument equals a weighted average of the indirect effect of time on the outcome through treatment in each time group.
-We call this an expected time gain and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
-de Chaisemartin and D’Haultfœuille (2018) (hereafter, ”dCDH” in this section) also consider the same setting in this article, but formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-The main difference between this article and dCDH is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects among the units who belong to an exposed group and start to receive the treatment in the second period.
-Accordingly, the identifying assumptions in dCDH are also different from those in this article.
-Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET if the treatment effect is stable over time, and propose the alternative estimands that are free from the strong restriction on treatment effect heterogeneity.
-Motivated by the disagreement between this article and dCDH, we next compare DID-IV designs to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
-First, we show that dCDH’s identifying assumptions ex-ante impose strong restrictions on the treatment adoption behavior across units, and these restrictions are asymmetric between exposed and unexposed groups.
-Next, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one parameter measures the treatment effects among the sub-population of the compliers and the other parameter measures the treatment effects among the time compliers.
-This decomposition result has two concerning implications.
-First, the interpretation of the SLATET is less clear than that of the LATET.
-Second, the SLATET is generally not policy-relevant even when the researchers exploit the policy change of their interest as an instrument for the treatment.
-We also revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and demonstrate that while their argument is correct, it is misguided.
-We first show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome in DID designs.
+Subsequently, we show that the average time trend in the unexposed outcome equals a weighted average of the indirect effect of time on the outcome through treatment in each time group.
+We call this an expected time gain, and show that the parallel trends assumption in the outcome in DID-IV designs requires the same expected time gain between exposed and unexposed groups.
+de Chaisemartin and D’Haultfœuille (2018) (hereafter, ”dCDH”) also consider the same setting as in this article, but formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
+The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period.
+Accordingly, the identifying assumptions in dCDH are also different from those used here.
+Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET under Fuzzy DID designs if the treatment effect is stable over time, and propose the alternative estimands that are free from the strong restriction on treatment effect heterogeneity.
+Next, motivated by the disagreement between this article and dCDH, we compare DID-IV designs to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
+First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units.
+These restrictions turn out to be asymmetric between exposed and unexposed groups, and difficult to defend from the institutional knowledge of the problem in a given application.
+Then, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers, who are affected by time but not affected by instrument.
+This decomposition result for the SLATET has two concerning implications.
+First, the interpretation of the SLATET is less clear than that of the LATET in DID-IV designs.
+Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
+We also revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and demonstrate that their argument is correct but misguided.
+First, we show that their argument stems from their reliance on the parallel trends assumption in DID designs.
 We then argue that this assumption is unsuitable for DID-IV settings for three reasons.
@@
 Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
-In this setting, units start exposed to the instrument at a particular point in time and remain exposed to that instrument afterward.
+In this setting, units start being exposed to the instrument at a particular point in time and remain exposed to that instrument afterward.
 Indeed, in most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation.
@@
 (2014), and Meghir et al.
-(2018)), the phase-in introduction of head starts across states (e.g.
-Johnson and Jackson (2019)), or the gradual adoption of broadband internet programs across municipalities (e.g.
+(2018)), phased-in introduction of head starts across states (e.g.
+Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g.
 Akerman et al.
@@
 (2013)).
-We call the underlying identification strategies staggered DID-IV designs, and establish the target parameter and identifying assumptions.
-Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument.
-We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ from the initial exposure to the instrument.
+We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
+Specifically, in this design, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument.
+We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument.
 Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings and state the interpretation of each assumption.
@@
 Lastly, we consider the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
-Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs in the presence of heterogeneous treatment effects.
+Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects.
 In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable regressions.
-In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
+In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs.
+Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
 Our proposed method avoids this issue and is robust to treatment effect heterogeneity.
-Our estimation procedure consists of two steps: we first subset the data for each CLATT and next run the two stage least squares regression in each data set.
+The estimation procedure consists of two steps: we subset the data that contain only two cohorts and two periods and then, in each data set, we run the TWFEIV regression.
 We call this a stacked two stage least squares (STS) regression and ensure its validity.
-Following to Callaway and Sant’Anna (2021), we also propose the weighting scheme to construct summary causal parameters from each CLATT, and discuss the procedure of pretrends tests in staggered DID-IV designs.
-We illustrate our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the United Kingdom, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment.
+Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects under staggered DID-IV designs, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$.
+We also discuss the procedure of pre-trend tests to check the validity of the parallel trends assumptions in the treatment and the outcome in DID-IV designs.
+We illustrate our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment.
 In this application, we first assess the plausibility of DID-IV identification assumptions implicitly imposed by Oreopoulos (2006).
@@
 We find that the TWFEIV estimate is strictly negative and not significantly different from zero.
-Actually, the TWFEIV regression performs the ”bad comparisons” (c.f.
-Goodman-Bacon (2021)) in the author’s setting: we treat the already exposed units as controls during the period after both regions were affected by the policy change.
-Finally, we use our estimation method to reassess returns to schooling in the U.K.
-We find that our STS estimates are all positive in each relative period after the school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate.
-Specifically, our weighting estimate indicates the roughly $20\%$ gain from schooling in the U.K.
-Overall, this paper provides a new research design by combining the IV techniques with DID designs.
+The TWFEIV regression performs the ”bad comparisons” (c.f.
+Goodman-Bacon (2021)) in the author’s setting: we treat the already exposed units as controls during the periods after both regions were affected by the policy change.
+Finally, we use our estimation method to reassess returns to schooling in the UK.
+We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate.
+Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
+Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs.
 Our DID-IV design can be applied to various empirical settings.
 First, it can cope with the general adoption process of the treatment when we are interested in the effect of the treatment on the outcome in DID designs.
-For instance, consider the case that some units already receive the treatment in period 00 before the policy shock in period $1$, which distributes the treatment.
-In this setting, the canonical DID design is infeasible in practice unless we ex ante discard these units from the data.
-On the other hand, our DID-IV design can be directly applied to this setting and allows us to identify the LATET.
+For instance, consider that some units already receive the treatment in period 00 before the policy shock, which distributes the treatment in period $1$.
+In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data.
+By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET.
 In other words, we can view that DID-IV is DID with the noncompliance of the assigned treatment.
-Second, our DID-IV design can yield valid counterfactual group to estimate the treatment effect when we have no control group under DID designs.
-For instance, consider the case that treatment is education attainment and we are interested in returns to schooling as in Duflo (2001).
-In this setting, the canonical DID design generally fails to construct the control group, as the data would lack the units with zero education attainment in reality.
-On the other hand, our DID-IV design can yield exposed and unexposed groups if one can exploit the policy shock across units as an instrument for the treatment.
-Then, by comparing the evolution of the treatment and the outcome between the two groups, one can identify the LATET under this design.
+Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs.
+For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001).
+In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality.
+However, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument).
+Once these groups are created under this design, one can identify the LATET by comparing the evolution of the treatment and the outcome between the two groups.
 Finally, our DID-IV design can overcome the endogeneity issue in empirical work.
-In practice, it is common that the treatment adoption is endogenous over time, leading to a failure in identifying the treatment effect, for instance, via simple linear regressions or two-way fixed effects regressions.
-Our DID-IV design can cope with the common identification challenges by leveraging variation arising from the policy shock across units as an instrument for the treatment.
-Importantly, in this design, we do not require the random assignment of the policy shock across units; rather, we require only the plausibility of the parallel trends assumptions in the treatment and the outcome, which would enhance both the internal and external validity of the IV estimate in a given application.
+In practice, due to the endogenous adoption of the treatment, researchers often fail to identify the treatment effect via simple linear regressions or two-way fixed effects regressions.
+Using our DID-IV design allows addressing the common identification challenge by leveraging variation in the timing of policy adoption across units as an instrument for treatment.
+Importantly, in this design, the identifying variation comes not from the random assignment of the policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome.
+This identification strategy would be less demanding than IV designs in some applications, and enhances the internal and external validity of the IV estimate in practice.
 The rest of the paper is organized as follows.
-The next subsection discusses the related literature.
+The following subsection discusses the related literature.
 Section 2 establishes DID-IV designs in a simple setting with two periods and two groups.
@@
 1.1   Related literature
-Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al.
-(2017); de Chaisemartin and D’Haultfœuille (2018)).
-We contribute to this literature in three important ways.
-The first contribution of this paper is to investigate the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH” in this section).
-In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption.
+Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018)) and contributes to this literature in three important ways.
+First, this paper investigates the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018).
+In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption [^1] In biometrics, Ye et al.
+(2023) also consider the ”instrumented difference-in-differences” design.
+However, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to identify this parameter..
 Hudson et al.
 (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings.
-Build on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs; the settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
-In this paper, we first formalize $2\times 2$ DID-IV designs and complement de Chaisemartin (2010) and Hudson et al.
+Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs.
+The settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
+In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al.
 (2017).
-Specifically, we introduce the path of the instrument in potential outcomes and uncover the ”hidden assumption” in $2\times 2$ DID-IV designs; this requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument.
+Specifically, we introduce the path of the instrument in $2\times 2$ DID-IV designs and uncover the ”hidden assumption” in the previous literature.
+This assumption requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument.
 We also investigate the selection mechanism behind $2\times 2$ DID-IV designs, and clarify the interpretation of the parallel trends assumption in the outcome.
-Given our identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
-The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicate the relative advantages of DID-IV over Fuzzy DID designs.
-The second contribution of this paper is to extend $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units.
-In practice, in order to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for the treatment in more than two periods (e.g., Black et al.
+Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
+Specifically, we show that the identifying assumptions in Fuzzy DID designs ex ante requires strong restrictions on treatment adoption behavior across units, and the target parameter in Fuzzy DID designs can be decomposed into a weighted average of two different causal parameters.
+The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicates the relative advantages of DID-IV over Fuzzy DID.
+Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units.
+In practice, to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g., Black et al.
 (2005), Lundborg et al.
 (2014), and Johnson and Jackson (2019)).
-No previous studies, however, extend $2\times 2$ DID-IV designs to such settings.
-Our staggered DID-IV designs ensure the theoretical validity of their identification strategies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
+No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings.
+Our staggered DID-IV designs ensure the theoretical validity of the underlying identification strategy behind these studies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
 Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects.
@@
 (2014)).
-In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time.
-Our estimation method developed in this paper would serve as an alternative to the TWFEIV estimators and make DID-IV designs more credible in a given application.
-Our paper is also related to a recent and growing DID literature.
-Several recent papers have pointed out the issue of implementing DID designs via two-way fixed effects regressions or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al.
+In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time.
+The proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
+Our paper is also related to a growing DID literature.
+Several recent studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al.
 (2021); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)).
-All of these papers have documented that regression coefficients by the conventional approaches may fail to properly summarize treatment effects under DID designs.
-Some of these papers have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity.
-None of these papers, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in the timing of policy adoption across units as an instrument for the treatment.
-In this paper, we formalize the underlying identification strategy as DID-IV designs, and provide the reliable estimation method.
-Built on the work in this paper, Miyaji (2024) studies the properties of the TWFEIV estimand under staggered DID-IV designs, and point out the issue of implementing this design via TWFEIV regressions.
+All of them have documented that regression coefficients by conventional approaches may fail to properly summarize treatment effects under DID designs.
+Some of them have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity.
+None of them, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in policy adoption timing across units as an instrument for treatment.
+This paper formalizes the rationale behind using the TWFEIV regression as an instrumented difference-in-differences design, and provide a reliable estimation method in multiple periods settings.
+Miyaji (2024) builds on this work to study the properties of the TWFEIV estimand under staggered DID-IV design and points out the issue of implementing this design via TWFEIV regressions.
 2   DID-IV in two time periods
-In this section, we formalize an instrumented difference-in-differences (DID-IV) in a two-period/two-group setting.
-We first establish the target parameter and identifying assumptions.
-We then uncover the selection mechanism behind the canonical DID-IV design, and clarify the interpretation of the parallel assumption in the outcome.
-We next describe the relationship between DID and DID-IV, and show that DID-IV is DID with the noncompliance of the assigned treatment.
+In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings.
+We first establish the target parameter and the identifying assumptions in this design.
+We then uncover the selection mechanism behind this design, and clarify the interpretation of the parallel assumption in the outcome.
+We next describe the relationship between DID and DID-IV, indicating that DID-IV is DID with the noncompliance of the assigned treatment.
 At the end of this section, we also describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
@@
 For any random variable $R$, we denote $\mathcal{S}(R)$ to be its support.
-For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receives the treatment in period $t$.
+For each $i\in\{1,\dots,N\}$ and $t\in\{0,1\}$, let $Y_{i,t}$ denote the outcome, and $D_{i,t}\in\{0,1\}$ denote the treatment status: $D_{i,t}=1$ if unit $i$ receives the treatment in period $t$ and $D_{i,t}=0$ if unit $i$ does not receive the treatment.
 Let $Z_{i,t}\in\{0,1\}$ denote the instrument status: $Z_{i,t}=1$ if unit $i$ is exposed to the instrument in period $t$ and $Z_{i,t}=0$ if unit $i$ is not exposed to the instrument in period $t$.
@@
 We introduce the path of the treatment and the instrument.
-Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ be the instrument path.
-We assume that no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$.
-Hereafter, we refer to this as a sharp assignment of the instrument.
-We denote $E_{i}\in\{0,1\}$ to be the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group).
-In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment: we assume that the treatment path can take four values with non-zero probability, that is, we have $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
-In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or a policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment.
+Let $D_{i}=(D_{i,0},D_{i,1})$ be the treatment path and $Z_{i}=(Z_{i,0},Z_{i,1})$ the instrument path.
+We assume no one is exposed to the instrument in period $t=0$: $Z_{i,0}=0$ for all $i$; we refer to this as a sharp assignment of the instrument.
+We denote $E_{i}\in\{0,1\}$ as the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group).
+In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
+In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment.
 For instance, Duflo (2001) estimates returns to schooling in Indonesia, using a new school construction program across regions as an instrument for education attainment.
@@
 Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$.
-Hereafter, we refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
+We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
 We make a no carryover assumption on the potential outcomes $Y_{i,t}(d,z)$.
@@
 Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$.
-In the DID literature, several recent papers impose this assumption with settings of a non-staggered treatment; see, e.g., de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021).
-[^1] When treatment is binary and there exist only two periods, one can explicitly introduce the treatment path in potential outcomes because it takes only four values.
+In the DID literature, several studies have used this assumption in non-staggered treatment settings (e.g.
+de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)).
+[^2] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes because it takes only four values.
 We therefore consider this case in Appendix B as an extension.
-One can also weaken Assumption 1 in other cases such as non-binary, ordered treatments or multiple time periods, but this would requires the cumbersome notation and complicates the definition of our target parameter, thus is beyond scope of this paper.
+One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatments, or multiple time periods, but it would require cumbersome notations and complicate the definition of the target parameter, going beyond scope of this paper.
 Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$, which describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$.
-Specifically, the first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group.
-Following to the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ to be the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ to be the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ to be the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ to be the defiers.
-Henceforth, we keep Assumption 1. In the next section, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
+The first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group.
+Following the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ as the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ as the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ as the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ as the defiers.
+Henceforth, we keep Assumption 1. In the next subsection, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
 2.2   Target parameter in two time periods
-Our target parameter in $2\times 2$ DID-IV designs is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
+In $2\times 2$ DID-IV designs, our target parameter is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
 Definition.
@@
 (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$.
-The LATET has been also proposed in heterogeneous effects IV models with binary instrument; see, e.g., Słoczyński et al.
-(2022) and Słoczyński (2020).
-We define the LATET as our target parameter for two reasons.
+The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g.
+Słoczyński (2020); Słoczyński et al.
+(2022)).
+We define LATET as the target parameter for two reasons.
 First, this parameter would be particularly of interest if the instrument reflects a policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
-Duflo (2001), for instance, are interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”.
+Duflo (2001), for instance, is interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”.
 To answer this question, the author exploits a new school construction program across regions as an instrument for education attainment.
@@
 Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT).
-In section 2.5, we show that both causal parameters measure the treatment effects among the units who are affected by instrument (policy shock) and belong to an exposed group $(E_{i}=1)$.
+Subsection 2.5 shows that both causal parameters measure the treatment effects among the units affected by instrument (policy shock) and belonging to an exposed group $(E_{i}=1)$.
 Remark 1.
-de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting we are considering here222de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings.
-We therefore rewrite their target parameter by the notation we use in this article..
+de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting as here333de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings.
+Therefore, we rewrite their target parameter by the notation used in this article..
 Definition.
@@
 This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW).
-As we will show in Appendix E, this parameter is different from the LATET and can be decomposed into a weighted average of two different causal parameters.
+As shown in Appendix E, this parameter turns out to be different from the LATET and can be decomposed into a weighted average of two different causal parameters.
 In section 2.6, we provide a detailed discussion of the connections between this article and de Chaisemartin and D’Haultfœuille (2018).
 2.3   Identification assumptions in two time periods
-This section establishes DID-IV designs in a simple setting with two groups and two periods.
+This subsection establishes DID-IV designs in a simple setting with two groups and two periods.
 In this setting, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
 $w_{DID}$ $=\frac{E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]}{E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]}.$
-Following to the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand.
+Following the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand.
 This estimand is obtained from the IV regression of the outcome on the treatment where the group and post-time dummies are included instruments and the interaction of the two is excluded instrument.
@@
 In the DID-IV literature, de Chaisemartin (2010) and Hudson et al.
-(2017) impose the similar assumption without introducing the instrument path.
+(2017) impose a similar assumption without introducing the instrument path.
 Given Assumption 1 and Assumption 2, we can write the observed outcomes $Y_{i,t}$ as
@@
 Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes.
-For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path were $z$:
+For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path is $z$:
 $Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
@@
 Note that we can not observe different potential outcomes for the same unit at the same time.
-For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument over two time periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
+For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument during the two periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
 Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
@@
 This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way.
-This assumption implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability.
+It implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability.
 In the DID-IV literature, de Chaisemartin (2010) and Hudson et al.
 (2017) make the same assumption.
-Hereafter, we consider the type of monotonicity assumption which rules out the existence of the defiers $DF^{Z}$.
+Hereafter, we consider the type of monotonicity assumption that rules out the existence of the defiers $DF^{Z}$.
 Assumption 4 (No anticipation in the first stage).
@@
 This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex-ante not known for all the units in an exposed group.
-This assumption is ”hidden assumption” in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al.
+This is ”hidden assumption” in the DID-IV literature: de Chaisemartin (2010) and Hudson et al.
 (2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
 In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways.
-Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume the no anticipation for all units $i$.
+Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume it for all units $i$.
 Roth et al.
-(2023) take the intermediate approach: they assume the no anticipation for the treated units.
-Our no anticipation assumption on potential treatment choices is in line with that of Roth et al.
+(2023) take the intermediate approach: they adopt the no anticipation assumption for the treated units.
+In this paper, the no anticipation assumption on potential treatment choices is in line with that of Roth et al.
 (2023).
-Next, we impose the parallel trends assumptions in the treatment and the outcome. In the DID-IV literature, de Chaisemartin (2010) and Hudson et al. (2017) assume the similar assumptions.
+Next, we impose the parallel trends assumptions in the treatment and the outcome. de Chaisemartin (2010) and Hudson et al. (2017) make the similar assumptions.
 Assumption 5 (Parallel Trends Assumption in the treatment).
@@
 $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]=E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
-Assumption 6 is a parallel trends assumption in the outcome.
-This assumption requires that the evolution of the unexposed outcome is the same on average between exposed and unexposed groups.
-For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of the policy shock.
+Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.
+For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
 When empirical researchers exploit variation arising from policy shock as an instrument for the treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6.
 For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on ”the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
-The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the local average treatment effects on the treated (LATET) in period $1$.
+The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the LATET in period $1$.
 Theorem 1.
-If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$, that is,
+If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
 $w_{DID}=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
@@
 As we already noted, the Wald-DID estimand scales the DID estimand of the outcome by the DID estimand of the treatment.
-Motivated by this observation, in $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1).
-First, in the denominator, we use the DID identification strategy in order to identify the direct effect of the instrument on the treatment: we recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment).
-Next, in the numerator, we use the DID identification strategy in order to identify the indirect effect of the instrument on the outcome through treatment: we recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome).
+Motivated by this observation, under $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1).
+First, in the denominator of the Wald-DID estimand, we use the DID identification strategy to identify the direct effect of the instrument on the treatment.
+We recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment).
+Next, in the numerator of the Wald-DID estimand, we use the DID identification strategy to identify the indirect effect of the instrument on the outcome through treatment.
+We recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome).
 Finally, by scaling the indirect effect obtained from the numerator by the direct effect obtained from the denominator, we can capture the LATET in period $t=1$.
 Figure 1: DID-IV identification strategy in two periods and two groups.
-Notes: This figure plots the evolution of the expectation of the treatment (First stage DID) and the outcome (Reduced form DID) for exposed and unexposed groups respectively.
+Notes: This figure plots the evolution of the expectation of the treatment (First-stage DID) and the outcome (Reduced-form DID) for exposed and unexposed groups, respectively.
 The dotted line represents the counterfactual trends of the expectation of the treatment and the outcome in an exposed group.
@@
 2.4   Interpreting the parallel trends assumption in the outcome
-So far, we have formalized the target parameter and identifying assumptions in $2\times 2$ DID-IV designs.
-Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand in order to identify the effects of the instrument on the treatment and the outcome.
+Thus far, we have formalized the target parameter and the identifying assumptions in $2\times 2$ DID-IV designs.
+Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand to identify the effects of the instrument on the treatment and the outcome.
 The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
-In this section, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs.
+In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs.
 To do so, we first uncover the selection mechanism behind the canonical DID-IV designs.
-Specifically, we show that in $2\times 2$ DID-IV designs, time also plays the role of instrument, and investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-We then show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment.
+Specifically, we investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups.
+Then, we show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment.
 We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
 First, we introduce the additional notation.
-Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path were $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
-We recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods.
-In this setting, one can see that time also plays the role of instrument, that is, time also affects one’s treatment choices during the two periods.
-Similar to the group variable $G_{i}^{Z}$ introduced in section 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ to be the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ to be the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ to be the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ to be the time defiers.
+Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path is $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
+Recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods.
+In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods.
+Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
 The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods.
 In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument).
-On the other hand, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument, that is, their treatment choices are affected by time in the absence of the policy shock.
-By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
+However, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument; that is, their treatment choices are affected by time in the absence of a policy shock.
+By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
 Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs.
-Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, the types painted in gray color in both tables.
-In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but can not observe $D_{1}((0,0))$ for each unit (see Table 1).
-Here, we use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$.
-In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but can not observe $D_{1}((0,1))$ for each unit (see Table 2).
+Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
+In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1).
+We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$.
+In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2).
 If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time.
@@
 $D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-Given these tables, we now consider the interpretation of the parallel trends assumption in the outcome.
-Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively:
+We now consider the interpretation of the parallel trends assumption in the outcome.
+Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in the exposed and unexposed groups, respectively.
 $\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
 $\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
-We note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$.
-At the same time, we also note that $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
+Note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$.
+At the same time, $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
 To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation.
 For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$.
-Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of the policy shock for time group $g^{t}$ in group $E_{i}=e$:
+Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
 $\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$
@@
 $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
-For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument for the time compliers in an unexposed group, respectively.
-Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$): we can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument.
+For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument, respectively, for the time compliers in an unexposed group.
+Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$).
+We can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument.
 For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit.
-On the other hand, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservables in an exposed group ($e=1$) because we can not specify the time group of unit $i$ in that group.
+However, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservable in an exposed group ($e=1$) because we cannot specify the time group of unit $i$ in that group.
 For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we cannot distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
 We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$.
-First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes respectively:
+First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
 $\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},$
@@
 This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods.
-If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (in other words, the stationary conditions hold), these quantities are equal to zero.
-Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms:
+If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (the stationary conditions hold), these quantities are equal to zero.
+Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
 $\Delta_{CM^{T},e}$ $=E[Y_{i,1}(0)-Y_{i,0}(0)|CM^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(1)-Y_{i,1}(0)|CM^{T},E_{i}=e]}_{\text{Selection gain among $CM^{T}$}},$
@@
 $\Delta_{e}=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e},$
-where the weight $w_{g^{t},e}$ is non-negative and sum to one, that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
+where the weight $w_{g^{t},e}$ is non-negative and sums to one; that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
 Proof.
@@
 In this paper, we call $\Delta_{e}$ an expected time gain in group $e$ hereafter.
-In light of the view discussed so far, we can conclude that the parallel trends assumption in the outcome (Assumption 6), i.e., $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups, that is, the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
+Based on the above discussions, we conclude that the parallel trends assumption in the outcome (Assumption 6), that is, $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups; the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
 Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group.
-The underlying idea to capture the former effect is parallel to that of DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
-If there exist pre-periods before the exposure to the instrument, we can indirectly check the validity of Assumption 6 as well as Assumption 5, comparing the evolution of the outcome between exposed and unexposed groups.
-The formal discussion for the pretrend test is provided in section 5.3.
+The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
+In practice, if we have pre-exposed period data, we can indirectly check the validity of Assumption 6 by comparing the evolution of the outcome between exposed and unexposed groups.
+The formal discussion for the pre-trend test is provided in Subsection 5.3.
 2.5   Relationship between DID and DID-IV
-In this section, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with the noncompliance of the assigned treatment.
-To see this, we first introduce the notation in DID designs according to section 2.1.
-Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program, which distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$.
+In this subsection, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with noncompliance of the assigned treatment.
+Following Subsection 2.1, we introduce the notation in DID designs.
+Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$.
 In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group).
 Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group.
-Hereafter, in this section, we refer to treatment and control groups as exposed and unexposed groups, respectively.
-We note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units.
-First, we assume that all the units are not allowed to receive the treatment in period 00:
+We refer to treatment and control groups as exposed and unexposed groups, respectively.
+Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units.
+First, we assume that all units are not allowed to receive the treatment in period 00.
 $D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (1)
-Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and should not receive the treatment in period $1$ if that unit belongs to an unexposed group:
+Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
 $1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (2)
-These restrictions imply that we have Tables 3-4 in DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
+These restrictions imply that we have Tables 3-4 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
 Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions.
@@
 Assumption 6 follows from combining (1) with (2).
-Second, the ATT, the target parameter in DID designs, is equal to the LATET in the canonical DID set up. This follows from Tables 3 - 4:
+Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 3 - 4:
 Table 3: Exposed group (assigned to $z=(0,1)$)
@@
 $=LATET.$
-We note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$).
-We therefore define the LATET as our target parameter in DID-IV designs in section 2.2.
+Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$).
+Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
 Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups.
-From Table 3 and Table 4, we can rewrite the parallel trends assumption in the untreated outcome as follows:
+From Tables 3- 4, we can rewrite the parallel trends assumption in the untreated outcome as:
 $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
@@
 $\iff$ $\Delta_{0}=\Delta_{1},$
-where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively in DID designs:
+where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
 $w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
-The above discussions clarify that DID-IV is a natural generalization of DID, and the difference between the two designs arises because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$.
+Overall, this subsection clarifies that DID-IV is a natural generalization of DID, and the differences between the two designs arise because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$.
 In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
 2.6   Comparing DID-IV to Fuzzy DID
-de Chaisemartin and D’Haultfœuille (2018) (Henceforth, ”dCDH” in this section) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting in this article.
-However, they formalize $2\times 2$ DID-IV designs differently, and call them Fuzzy DID.
+de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture some causal effects under the same setting as in this article.
+However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
 In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID.
-The detailed discussion is contained in Appendix E, and this section provides the main points of that discussion.
-The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see section 2.2 in this article).
+The detailed discussion is given in Appendix E, and this section provides the overview of that discussion.
+The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see Subsection 2.2).
 Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs.
-Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
-In Appendix E, we first show that dCDH’s identifying assumptions ex ante impose strong and asymmetric restrictions on the treatment adoption behavior across units between exposed and unexposed groups.
-Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Table 5 and Table 6333Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6..
-These tables are in line with Table 1 and Table 2 in this article, and clarify the difference between dCDH’s restrictions and ours: we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by monotonicity assumption, whereas dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
-This difference is ciritical for empirical application: in contrast to the monotonicity assumption, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application.
+Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
+In Appendix E, we first show that the identifying assumptions in Fuzzy DID designs ex-ante impose the strong restrictions on the treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
+Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 5-6444Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6., which are in line with Tables 1-2.
+From these tables, one can see that dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
+Note that in our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption.
+The difference between dCDH’s restrictions and ours is critical for empirical application; in contrast to the monotonicity assumption, which excludes the defiers in both groups, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application.
+This is because these restrictions ex ante exclude the units who change the treatment adoption in the absence of a policy shock (instrument).
 Table 5: Exposed group ($z=1$)
@@
 Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
-Specifically, we show that the SLATET is a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $CM^{Z}\land AT^{T}$, which is the sub-population of the compliers $CM^{Z}$, in an exposed group.
-This decomposition result has two concerning implications.
+The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers.
+The other causal parameter measures the treatment effects among the time compliers $CM^{T}$.
+The time compliers switch from non-treatment to treatment in period $1$ in the absence of the policy shock (instrument); that is, they are affected by time but not affected by the instrument.
+This decomposition result for the SLATET has two concerning implications.
 First, the interpretation of the SLATET is less clear than that of the LATET.
-Second, this parameter may not be policy-relevant even when the researchers use the policy change of their interest as an instrument for the treatment because it generally contains the treatment effects among the time compliers $CM^{T}$ who are affected by time but not affected by the instrument.
+Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
 Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
-In this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups respectively for the Wald-DID estimand to identify the SLATET.
-First, we show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome in DID designs.
-Recall that in DID-IV designs, we do not assume the parallel trends assumption in the untreated outcome, but assume the parallel trends assumption in the unexposed outcome.
-Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption generally fails to capture the average time trend of the outcome in an unexposed group.
+During this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups for the Wald-DID estimand to identify the SLATET.
+First, we show that dCDH’s argument regarding the use of the Wald-DID estimand is based on their reliance on the parallel trends assumption in DID designs.
+Recall that in our DID-IV design, we did not make the parallel trends assumption in the untreated outcome, but made the parallel trends assumption in the unexposed outcome.
+Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group.
 We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
-In the proof of this argument, we also demonstrate that the strong restrictions imposed by dCDH on treatment adoption behavior is also induced by imposing the parallel trends assumption in the untreated outcome.
-Next, we argue that the parallel trends assumption in the untreated outcome is unsuitable for DID-IV settings for three reasons.
-First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group.
+In the proof, we also demonstrate that imposing the parallel trends assumption in DID designs yields the strong restrictions on the treatment adoption behavior across units under Fuzzy DID designs.
+Next, after confirming that dCDH’s argument is based on the parallel trends assumption in DID designs, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
+First, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group.
 Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al.
@@
 For instance, in Black et al.
-(2005), this assumption requires that the data contain the parents with zero education attainment, which is unrealistic in practice.
-Finally, in DID-IV settings, we can not generally test this assumption using pre-exposed period data, as some units can adopt the treatment before period 00.
-Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the validity of this assumption.
-Surprisingly, we also show that their placebo test coincides with our pretrends test, which we will introduce in section 5.3 for assessing the plausibility of Assumption 5 and Assumption 6, with only a slight modification.
+(2005), this assumption requires the data to contain parents with zero education attainment, which is unrealistic in practice.
+Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.
+We show that dCDH’s placebo test is incomplete for assessing the validity of this assumption.
+Surprisingly, we also show that their placebo test coincides with our pretrends test with a slight modification, introduced in section 5.3, to assess the plausibility of Assumption 5 and Assumption 6.
 3   DID-IV in multiple time periods
-We now extend the $2\times 2$ DID-IV design to multiple periods settings.
-Specifically, we consider the setting with the staggered adoption of the instrument across units: the exposure to the instrument occurs across units at different times and once units become exposed to the instrument, units remain exposed to that instrument afterward.
-We call this a staggered DID-IV design, and establish the target parameter and identifying assumptions.
+We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
+In reality, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g.
+Black et al.
+(2005), Bhuller et al.
+(2013), Lundborg et al.
+(2014), and Meghir et al.
+(2018)).
+We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
 3.1   Set up
@@
 We consider a panel data setting with $T$ periods and $N$ units.
-For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome and $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status.
-Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$.
+For each $i\in\{1,\dots N\}$ and $t\in\{1,\dots,T\}$, let $Y_{i,t}$ denote the outcome, $D_{i,t}\in\{0,1\}$ denote the treatment status, and $Z_{i,t}\in\{0,1\}$ denote the instrument status.
+Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively.
 Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
@@
 For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
-Assumption 7 requires that once units start exposed to the instrument, units remain exposed to that instrument afterward.
-In the DID literature, this assumption is made for the adoption process of the treatment and sometimes called ”staggered treatment adoption”, see, e.g., Athey and Imbens (2022), Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
-Exploiting variation arising from the staggered adoption of the instrument across units is pervasive in economics.
-For instance, Black et al.
-(2005) estimate the causal relationship between parents’ and children’s education by exploiting the different timing of implementation of school reforms across municipalities as an instrument for parental education.
-Given Assumption 7, we can uniquely characterize one’s instrument path by the initial exposure date of the instrument, which we denote $E_{i}=\min\{t:Z_{i,t}=1\}$.
+Assumption 7 requires that once units start getting exposed to the instrument, units remain exposed to that instrument.
+In the DID literature, several studies make a similar assumption for the treatment and call it the ”staggered treatment adoption” (e.g.
+Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
+Given Assumption 7, one can uniquely characterize one’s instrument path by the initial exposure date of the instrument, denoted as $E_{i}=\min\{t:Z_{i,t}=1\}$.
 If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$.
-Based on the initial exposure period $E_{i}$, we can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start exposed to the instrument in period $E_{i}=e$.
-Similar to the two periods setting in section 2.1, we allow the general adoption process for the treatment: treatment can potentially turn on/off repeatedly over time.
-de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) consider the same setting in the recent DID literature.
+Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start being exposed to the instrument in period $E_{i}=e$.
+Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time.
+de Chaisemartin and D’Haultfœuille (2020) and Imai and Kim (2021) also consider the same setting in the recent DID literature.
 Next, we introduce the potential outcomes framework in multiple time periods.
@@
 Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument.
-Hereafter, we call $D_{i,t}^{\infty}$ the ”never exposed treatment”.
+We call $D_{i,t}^{\infty}$ the ”never exposed treatment”.
 Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
@@
 We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$.
-Hereafter, we refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage.
+We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage.
 In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
-Next, we introduce the group variable which describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$.
+Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$.
 Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$.
-Specifically, the first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$.
-Following to the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ to be the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ to be the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ to be the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ to be the defiers in period $t$ and the initial exposure date $e$.
+The first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$.
+Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
 Finally, we make a no carryover assumption on potential outcomes $Y_{i,t}(d,z)$.
@@
 $=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
-Each CLATT is a natural generalization of the LATET in section 2.2 and suitable for the setting of the staggered instrument adoption.
-This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$, that is, who are induced to treatment by instrument in period $e+l$.
-Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$ and the compliers $CM_{e,e+l}$.
-We note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$.
+Each CLATT is a natural generalization of the LATET in Subsection 2.2 and suitable for the setting of the staggered instrument adoption.
+This parameter measures the treatment effects at a given relative period $l$ from the initial exposure date $E_{i}=e$, for those who belong to cohort $e$, and are the compliers $CM_{e,e+l}$ who are induced to treatment by instrument in period $e+l$.
+Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$.
+Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$.
 For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
-This parameter would be policy-relevant if researchers leverage the policy change of their interest as an instrument for the treatment.
-Suppose for instance that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment.
-In this example, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest because the compliers in each relative period after a policy change represents the dynamic effect of that policy on the take up rates of the program, and this parameter can be interpreted as estimating the effect of participation in job training on real earnings among those affected in each relative period.
+This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher.
+For example, suppose that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment.
+In this setting, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest; the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
 3.3   Identification assumptions in staggered DID-IV designs
-In this section, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
+In this subsection, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
 Assumption 9 (Exclusion restriction in multiple time periods).
@@
 Given Assumption 8 and Assumption 9, we can write the potential outcome $Y_{i,t}(d,z)$ as $Y_{i,t}(d_{t})=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0)$.
-Following to section 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$:
+Following Subsection 2.3, we introduce the potential outcomes in period $t$ if unit $i$ is assigned to the instrument path $z\in\mathcal{S}(Z)$.
 $Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
-Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$ respectively.
-The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed to the instrument.
-Hereafter, we refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
+Since the initial exposure date $E_{i}$ completely characterizes the instrument path, we can write the potential outcomes for cohort $e$ and cohort $\infty$ as $Y_{i,t}(D_{i,t}^{e})$ and $Y_{i,t}(D_{i,t}^{\infty})$, respectively.
+The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed.
+We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
 Assumption 10 (Monotonicity assumption in multiple time periods).
@@
 This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$.
-We recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$.
+Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$.
 Assumption 10 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$.
@@
 […]
-Assumption 12 is a parallel trends assumption in the treatment in multiple periods and multiple cohorts.
-This assumption requires that the trends of the treatment across cohorts would have followed the same path, on average, if there is no exposure to the instrument.
-Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs: these studies impose the same type of parallel trends assumption on untreated outcomes with settings of multiple periods and multiple cohorts.
+Assumption 12 is a parallel trends assumption in the treatment in multiple periods and cohorts.
+It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument.
+Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated outcomes in settings of multiple periods and cohorts.
 Assumption 13 (Parallel trends assumption in the outcome in multiple time periods).
 […]
-Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and multiple cohorts.
-This assumption requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the assignment of the instrument had not occurred.
-From the discussion in section 2.4, we can interpret that this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
+Assumption 13 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts.
+It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned.
+From the discussion in Subsection 2.4, this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
 4   Extensions
 This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections.
-It also includes the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
-For more details and the proofs in this section, see Appendix B.
+It includes the case in which the treatment path was introduced in potential outcomes in $2\times 2$ DID-IV designs.
+For more details and proofs, see Appendix B.
 Non-binary, ordered treatment
 Up to now, we have considered only the case of a binary treatment.
-However the same idea can be applied when treatment takes a finite number of ordered values, that is, we have $D_{i,t}\in\{0,1,\dots,J\}$.
+However the same idea can be applied when treatment takes a finite number of ordered values; that is, we have $D_{i,t}\in\{0,1,\dots,J\}$.
 When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
@@
 $w_{j}=\frac{Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}.$
-The average causal response on the treated (ACRT) is a weighted average of the effects of a unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument.
-The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), but the difference here is that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
-Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the average causal response on the treated (ACRT) under the same assumptions in Theorem 1.
+The ACRT is a weighted average of the effects of an increase of one unit in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument.
+The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
+Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under the same assumptions in Theorem 1.
 Theorem 3.
@@
 See Appendix B.1. ∎
-If we have a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), which is the natural generalization of the ACRT.
-The identifying assumptions are the same in section 3.3.
+With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT.
+The identifying assumptions are the same as in Subsection 3.3.
 Definition.
@@
 Introducing the treatment path in $2\times 2$ DID-IV designs
-In section 2.1, we impose Assumption 1 (No carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment.
+In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment.
 We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment.
-In this case, our target parameter slightly changes with the cumbersome notation, formally defined below.
+In this case, our target parameter changes slightly with some complex notation.
 Definition.
@@
 $w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
-In Appendix B.3, we show that the similar assumptions as in Assumptions 1-6 in section 2.3 are sufficient for the Wald-DID to capture the LATET’.
+Appendix B.3 shows that the similar assumptions to Assumptions 1-6 in Subsection 2.3 are sufficient for the Wald-DID to capture the LATET’.
 5   Estimation and inference
-When researchers leverage variation in the timing of policy adoption across units as an instrument for the treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al.
+When researchers leverage variation in policy adoption timing across units as an instrument for treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al.
 (2014), Black et al.
@@
 (2013)).
-Unfortunately, however, recent research has shown that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects.
-Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative [^6] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally..
-Miyaji (2024) further shows that the negative weight problem arises if the effect of the instrument on the treatment or the outcome is not stable over time.
+The companion paper (Miyaji (2024)), however, shows that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects.
+Specifically, it shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
+[^7] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.
 In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity.
-Specifically, we first propose a simple regression-based method for estimating each CLATT.
-Following to Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT.
-At the end of this section, we also discuss the pretrends tests for checking the plausibility of parallel trends assumption in the treatment and the outcome.
+First, we propose a simple regression-based method for estimating each CLATT.
+Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT.
+At the end of this section, we also discuss the pre-trend tests for checking the plausibility of parallel trends assumption in the treatment and the outcome in DID-IV designs.
 5.1   Stacked two stage least squares regression
-In this section, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs.
-Recall that if we have a binary treatment, our target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
+We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs.
+If we have a binary treatment, the target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
 $CLATT_{e,e+l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
@@
 Researchers should carefully specify the choice set $U$ in practice.
-If there exists a never exposed cohort $E_{i}=\infty$, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$).
-If there exists no never exposed cohort, we can set $U=\max\{E_{i}\}$, i.e., the last exposed cohort.
-In this case, we can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
-We note that the units in an already exposed cohort $e=1$ are not included in any data sets because we should include the pre-exposed period $e-1$ for estimating each $CLATT_{e,e+l}$.
+If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$).
+If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
+Note that the units in an already exposed cohort $e=1$ are not included in any data sets because the pre-exposed period $e-1$ should be included for estimating each $CLATT_{e,e+l}$.
 Step 2.
@@
 Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation.
-We note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period.
-From the estimation procedure above, we call this a stacked two-stage least squares (STS) estimator.
-Our STS estimator is related to the DID estimators in staggered DID designs recently proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
-Specifically, the DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when we have no covariates and use never treated units as a control group.
-Their proposed methods avoid the issue of two-way fixed effects estimators in staggered DID designs, and our STS estimator avoids the issue of two-way fixed effects instrumental variable estimators in staggered DID-IV designs.
+Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period.
+From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator.
+Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
+The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group.
+Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, and our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
 Theorem 4 below guarantees the validity of our STS estimator.
 Theorem 4.
-Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal:
+Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
 […]
@@
 Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data.
-In Appendix C, we present the influence function of our SLS estimator in repeated cross section settings.
+Appendix C presents the influence function of the SLS estimator in repeated cross section settings.
 In triple DID-IV designs, one can replace the IV regression in step $2$ with the following one ($g_{i}=A,B$):
@@
 $+\pi^{e,l}_{7}\mathbf{1}_{E_{i}=e,T_{i}=e+l,g_{i}=A}+\eta^{e,l}_{i,t},$
-where $\mathbf{1}_{A}$ is indicator function and takes one if $A$ is true. In Appendix C, we also present the influence function of our triple DID-IV estimator $\beta_{IV}^{e,l}$.
+where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true. Appendix C presents the influence function of the triple DID-IV estimator $\beta_{IV}^{e,l}$.
 5.2   Weighting scheme
-The previous section provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs.
+The previous subsection provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs.
 In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$.
-However, in many cases, researchers may also want to estimate summary causal parameters, which capture the overall effect of a treatment on an outcome.
-Actually, in staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$.
+However, researchers may also want to estimate summary causal parameters that capture the overall effect of a treatment on an outcome.
+In staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$.
 In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
@@
 where $w(e,t)$ are some reasonable weighting functions assigned to each $CLATT_{e,t}$.
-To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we first define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$.
-We call this the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$), formally defined below.
+To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$, called the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$).
 Definition.
@@
 In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity.
-We can employ their framework straightforwardly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
+We can employ their framework directly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
 For instance, to aggregate dynamic treatment effects in cohort $e$ over time in staggered DID designs, Callaway and Sant’Anna (2021) consider the following summary measure:
 $\theta_{sel}(\tilde{e})=\frac{1}{T-\tilde{e}+1}\sum_{t=\tilde{e}}^{T}CATT(\tilde{e},t),$
-where the $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
-Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start exposed to that treatment.
+where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
+Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start being exposed to it.
 Table 7: Weights for a variety of summary causal parameters
@@
 Each target parameter corresponds with that of Callaway and Sant’Anna (2021).
-We superscript IV to associate each parameter with staggered DID-IV designs.
+IV is added to associate each parameter with staggered DID-IV designs.
 In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
@@
 This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$.
-This weighting scheme would be reasonable in that it is designed to be larger at the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
+This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
 By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021).
@@
 Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs.
-We note that our proposed weighting functions are the same if we have non-binary, ordered treatment where our target parameter is the $CACRT_{\tilde{e},t}$.
+Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
 Estimation and inference
-We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows:
+We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows.
 $\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
@@
 $l_{i}^{\theta_{IV}}=\sum_{e}\sum_{t=1}^{T}\left(w(e,t)\cdot\psi_{i,e,t}+\zeta_{i,e,t}^{w}\cdot{CLATT}_{e,t}\right).$
-5.3   Pretrend test
-In this section, we describe how one can conduct pretrend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
-Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
+5.3   Pre-trend test
+We now describe how one can conduct pre-trend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
+Suppose that data are available for period $t=-1$; one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
 $E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
@@
 $\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]=E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (4)
-We can generalize these tests to multiple pre-exposed periods settings by using the pretrends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al.
+These tests can be generalized to multiple pre-exposed periods settings by using the pre-trend tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al.
 (2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
@@
 However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous.
-Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and the parallel trends assumption in the untreated outcome are plausible.
+Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and parallel trends assumption in the untreated outcome are plausible.
 Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
 6   Application
-Oreopoulos (2006) estimates returns to schooling using a major education reform in the United Kingdom that increased the years of compulsory schooling from 14 to 15.
-Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, while the reform was not implemented until 1957 in Northern Ireland.
+Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15.
+Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland.
 The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K.
 General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
-In the first part of their analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately.
+In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately.
 Due to the imprecision of their standard errors [^8] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K.
-data” to estimate some local average treatment effects.
-Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the two-way fixed effects instrumental variable regression, using the reform as an excluded instrument for education attainment:
+data”.
+Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
 $Y_{i,t}=\mu_{n.}+\delta_{.t}+\beta_{IV}D_{i,t}+\epsilon_{i,t},$ (7)
@@
 Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change.
-The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$ respectively.
+The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$, respectively.
 Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect.
-The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy shock.
-The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized in reality.
-Oreopoulos (2006) notes that the reform was implemented with political support, taking into account the cost and benefit of the school reform.
-This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the random variation of the policy change.
-Actually, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
-In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, and potentially vary across cohorts.
-The TWFEIV estimate, therefore, can potentially fail to aggregate the causal effects and yield a misleading conclusion.
+The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy change.
+The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit.
+This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the IV design.
+Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
+In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, potentially varying across cohorts.
+That the author was conscious of the heterogeneous treatment effects is clear from the title of ”Estimating Average and Local Average Treatment Effects of Education when Compulsory Schooling Laws Really Matter”.
+The TWFEIV estimator, therefore, can potentially fail to aggregate the treatment effects, yielding a misleading conclusion under staggered DID-IV designs.
 In this section, we first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006).
@@
 6.1   Assessing the identifying assumptions in staggered DID-IV designs
-We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006).
-In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT) as education attainment is a non-binary, ordered treatment.
-We can identify each CACRT and its summary measures if the underlying DID-IV identifying assumptions seem plausible.
+We discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006).
+In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment.
+We can identify each CACRT and its summary measure if the underlying DID-IV identifying assumptions seem plausible.
 Exclusion restriction.
@@
 Monotonicity assumption.
-It would be automatically satisfied in the author’s setting: the policy reform (instrument) increased the minimum schooling-leaving age from $14$ to $15$, which ensures that there are no defiers during the periods after the policy shock.
+It would be automatically satisfied in the author’s setting: the policy change (instrument) increased the minimum schooling-leaving age from $14$ to $15$, ensuring that there are no defiers during the periods after the policy shock.
 No anticipation in the first stage.
-It would be plausible that there is no anticipation before the policy reform, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England.
+It would be plausible that there is no anticipation, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England.
 This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
-Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome.
-To do so, we use the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively.
+Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome, using the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively.
 The results are shown in Figure 2.
 Figure 2: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006).
-Notes: The results for the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy.
+Notes: This figure presents the results of the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy.
 The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$.
@@
 Parallel trends assumption in the outcome.
-It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts if the reform had not occurred.
+It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform.
 Panel (b) in Figure 2 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval.
-The pre-exposed estimates seems consistent with Assumption 13: though there exists a downward pretrends, the estimates before the exposure to the policy change are not significantly different from zero.
-Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post exposed periods.
-In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all the estimates are statistically significant.
+The pre-exposed estimates seems consistent with Assumption 13: though an upward pre-trend exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
+Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post-exposed periods.
+In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant.
 In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
-We note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
+Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
 In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
@@
 We start by estimating two-way fixed effects instrumental variable regression in the author’s setting.
-In our estimation, we slightly modify the author’s specification to clearly illustrate the pitfalls of TWFEIV regression.
-Specifically, Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
+To clearly illustrate the pitfalls of TWFEIV regression, in our estimation, we slightly modify the author’s specification.
+Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
 The result is shown in Table 8.
-The TWFEIV estimate is $-0.009$ and it is not significantly different from 00.
-This implies that on the whole, the returns to schooling in the U.K.
-may be nearly zero.
-However, this may be a misleading conclusion: we cannot interpret that the TWFEIV estimand captures the properly weighted average of each CACRT if the effects of the policy reform on education attainment and log annual earnings are not stable across cohorts (Miyaji (2024)).
+The TWFEIV estimate is $-0.009$, not significantly different from 00, which indicates that, on the whole, the returns to schooling in the UK is nearly zero.
+However, this may be a misleading conclusion.
+We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
 The TWFEIV regression performs the ”bad comparisons” (c.f.
-Goodman-Bacon (2021)) in the author’s setting: when we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$ because both regions are already exposed to the policy shock during these periods.
+Goodman-Bacon (2021)) in the author’s setting.
+When we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods.
 In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006).
@@
 Figure 3: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006).
-Notes: The results for returns to schooling under the staggered DID-IV identification strategy.
+Notes: This figure shows the results for returns to schooling under the staggered DID-IV identification strategy.
 The unexposed group $U$ is North Ireland and the reference period is period $t=1946$.
 The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
-We now employ our proposed method to estimate each CACRT.
-We first create data sets by cohort $t$ ($t=1947,\dots,1956$).
+We now employ the proposed method to estimate each CACRT.
+First, we create data sets by cohort $t$ ($t=1947,\dots,1956$).
 Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland.
-We define North Ireland as an unexposed group $U$ and discard the units who were aged at $14$ between $1957$ and $1965$ in order to avoid the bad comparisons.
+We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons.
 We then run TSLS regression in each data set.
@@
 The estimates for each $CACRT_{e,t}$ range from $14\%$ to $38\%$ with wide confidence intervals, and $4$ out of $10$ estimates are statistically significant.
-Finally, we estimate the summary causal measure constructed from each $CACRT_{e,t}$.
-Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$):
+Finally, we estimate the summary causal measure by aggregating each $CACRT_{e,t}$.
+Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$).
 $\theta_{sel}^{IV}(e)=\sum_{t=1947}^{1956}\frac{CAET^{1}_{e,t}}{\sum_{t=1947}^{1956}CAET^{1}_{e,t}}CACRT_{e,t}.$
-Here, each weight assigned to each $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
+Here, each weight assigned to $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
 The result is shown in Table 8.
 The estimate is $0.24$ and it is significantly different from zero.
-The estimated returns to schooling are substantial: this is likely because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family, and have potentially much gain from the school reform.
+The estimated returns to schooling are substantial, perhaps because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family that potentially have much to gain from school reform.
 The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative.
-Overall, our results indicate that the economic returns of education are substantial in the U.K.
-and the estimation method matters in staggered DID-IV designs in practice.
+Overall, our results indicate that the economic returns of education are substantial in the UK, and the estimation method matters in staggered DID-IV designs in practice.
 7   Conclusion
 Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome.
-In this design, we commonly face two challenges: the internal and external validity of the IV estimate.
-To overcome at least the external validity, many studies leverage variation occurring from the different timing of policy adoption across units as an instrument for the treatment, and use instrumental variable techniques.
-The instrument, however, generally fails to satisfy the independence assumption in practice, and the underlying identification strategy is parallel those in DID designs.
-In this paper, we provide a new econometric framework, which we call an instrumented difference-in-differences (DID-IV).
-We start by considering a simple setting with two periods and two groups.
+In this design, however, we commonly face two challenges: the internal and external validity of the IV estimate.
+To overcome at least external validity, many studies have leveraged variation occurring from the different timing of policy adoption across units as an instrument for treatment.
+The instrument based on these policy changes, however, generally fails to satisfy the randomization assumption in IV designs in practice, casting doubt on the internal validity of the IV estimate.
+Rather, the underlying identification strategy is parallel to DID designs, under which we identify the effect of the policy shock on the outcome, leveraging the variation of the exposure to that policy shock across units and over time.
+In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV) by combining the IV techniques with DID designs.
+We start with a simple two periods and two groups setting.
 In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups.
-We show that under $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET).
+We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
 We also clarify the interpretation of the parallel trends assumption in the outcome.
-Specifically, we show that in DID-IV designs, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
+Specifically, we show that in DID-IV settings, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
 Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
-Next, we consider the DID-IV design in the case where there are more than two time periods and units become exposed to the instrument at different times.
+Next, we consider the DID-IV design in more than two time periods with units being exposed to the instrument at different times.
 We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions.
-Specifically, our target parameter in staggered DID-IV designs is the cohort specific local average treatment effects on the treated (CLATT).
+Our target parameter in this design is the cohort specific local average treatment effects on the treated (CLATT).
 The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
@@
 Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects.
-We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pretrends test to assess the validity of the parallel trends assumption in the treatment and the outcome.
-Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the United Kingdom, exploiting the timing variation of the introduction of school reforms between British and North Ireland.
+We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trend tests to assess the validity of the parallel trends assumption in the treatment and the outcome.
+Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland.
 First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible.
-Next, we estimate TWFEIV regressions and our stacked two stage least squares regressions in each dataset.
-We find that the TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive.
+Next, we estimate a TWFEIV regression and our stacked two stage least squares regressions in each dataset.
+The TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive.
 Finally, we estimate the summary causal parameter by exploiting our STS estimates.
-The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the U.K.
+The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the UK.
 Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment.
-To avoid the issue of using TWFEIV estimators in this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity.
+To avoid the issue of using the TWFEIV estimator under this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity.
 Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
@@
 The third equality follows from Assumption 4 (No anticipation in the first stage) and the forth equality follows from Assumption 3 (Monotonicity assumption) and Assumption 5 (Parallel trends assumption in the treatment).
-Next, to show that the Wald-DID estimand equals to the LATET, it suffices to show that
+Next, to show that the Wald-DID estimand equals the LATET, it suffices to show that
 $E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$ $=E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1)$
@@
 The third equality follows from Assumption 4.
-The forth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and the simple calculation.
+The fourth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and simple calculation.
 The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
@@
 $=LATET.$
-Completing the proof. ∎
+This completes the proof. ∎
 Proof of Theorem 2.
@@
 $=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
-Here the first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$.
-Clearly, the weight $w_{g^{t},e}$ sum to one by construction.
-Completing the proof.
+The first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$.
+Clearly, the weight $w_{g^{t},e}$ sums to one by construction.
+This completes the proof.
 Appendix B Extensions
@@
 where we use Assumption 1 and 2.
-For the numerator in the Wald DID estimand, we have:
+For the numerator in the Wald DID estimand, we have
 $E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
@@
 $=E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1].$
-Here the first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
+The first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
 In terms of the $\lambda_{Z,j}^{t}$, we write the previous expression as:
@@
 $=E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right].$
-where the second equality follows from the fact: $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
-We note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
+where the second equality follows from $\lambda_{1,0}^{1}-\lambda_{0,0}^{1}=1$.
+Note that $\lambda_{1,j}^{1}\geq\lambda_{0,j}^{1}$ from Assumption 3 and that $\lambda_{1,j}^{1}$ and $\lambda_{0,j}^{1}$ can take only two values, zero or one. Therefore, we obtain:
 $E\left[\sum_{j=1}^{J}(Y_{1}(j)-Y_{1}(j-1))\cdot[\lambda_{1,j}^{1}-\lambda_{0,j}^{1}]|E=1\right]$
@@
 ×$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (12)
-For the denominator in the Wald-DID estimand, by the similar argument, we first obtain:
+For the denominator in the Wald-DID estimand, by similar argument, we obtain:
 $E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
@@
 $=E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
-Here the first equality follows from the simple algebra and Assumption 4. The second equality follows from Assumption 5.
+The first equality follows from simple algebra and Assumption 4. The second equality follows from Assumption 5.
 Then, we have:
@@
 B.2   Repeated cross sections
-In this section, we assume repeated cross section settings and present our identification results.
+Here, we present our identification results under repeated cross section settings.
 B.2.1 Two time periods
@@
 We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
-In two periods and two groups settings, our target parameter is the local average treatment effect in period $1$:
+In a two-period and two-group setting, our target parameter is the local average treatment effect on the treated in period $1$:
 $LATET$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
@@
 $=$ $E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=0].$
-Theorem 5 below shows that the Wald-DID estimand is equal to the local average treatment effect on the treated (LATET) in period $1$ under Assumptions 14-18.
-Here, we assume the binary treatment, but non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus we omit it for brevity).
+Theorem 5 below shows that the Wald-DID estimand is equal to the LATET in period $1$ under Assumptions 14-18.
+Here, we assume the binary treatment, but the non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus, we omit it for brevity).
 In the proof, we omit the index $i$ to ease the notation.
@@
 $=$ $Pr(CM^{Z}|E=1,T=1).$ (14)
-The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the forth equality follows from Assumption 15 and Assumption 17.
+The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the fourth equality follows from Assumption 15 and Assumption 17.
 Next, for the numerator, we have
@@
 The third equality follows from Assumption 16.
-The forth equality follows from Assumption 18 and the simple calculation.
+The fourth equality follows from Assumption 18 and the simple calculation.
 The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
@@
 $=LATET.$
-Completing the proof. ∎
 B.2.2 Multiple time periods
 Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$.
-We assume Assumption 7 in section 3, that is, we assume the staggered assignment of the instrument across units.
-In addition, we assume Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$.
+We adopt Assumption 7 in Section 3; that is, we assume the staggered assignment of the instrument across units.
+In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$.
 Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to.
@@
 We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
-In multiple time periods and multiple cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
+In multiple time period and multiple cohort settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
 $CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
 $=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
-When repeated cross section data are available and we have multiple time periods, our staggered DID-IV designs consist of Assumptions 7, 10, 11 (in section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
+When repeated cross section data are available and we have multiple time periods, the staggered DID-IV designs consist of Assumptions 7, 10, 11 (in Section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
 Assumption 19 (Parallel Trends Assumption in the treatment in multiple time periods).
@@
 B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
-In this section, we discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$.
-We first consider the two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption).
+We discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$.
+We consider two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption).
 We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting.
-Finally, we show that a Wald-DID estimand identifies the LATET’ (see section 4) under the suitable assumptions similar to the ones in section 2.3.
-Henceforth, we consider the same setting as in section 2.1 except that we do not adopt Assumption 1.
+Finally, we show that a Wald-DID estimand identifies the LATET’ (see Section 4) under suitable assumptions similar to the ones in Subsection 2.3.
+We consider the same setting as in Subsection 2.1 except that we do not adopt Assumption 1.
 First, we make the following no anticipation assumptions on the outcome.
-Assumption 21 (No anticipation on the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
+Assumption 21 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
 […]
-This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((0,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land NT^{T}$.
-We note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al.
-(2023) (see Table 3 and Table 4 in section 2.5) [^8] Here, it is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
-Assumption 22 (No anticipation on the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
+This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
+Note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al.
+(2023) (see Table 3 and Table 4 in Subsection 2.5) [^9] It is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
+Assumption 22 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
 […]
-This assumption requires that the potential outcome at period 00 is same as the baseline outcome $Y_{i,0}((1,0),z)$ given the instrument path $z$, for those who are assigned to the instrument $z=(0,1)$ and belongs to the type $CM^{Z}\land DF^{T}$.
+This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
 Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
-We can rationalize Assumption 21 and 22 by noting that both the two Assumptions and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units who belong to the exposed group and the compliers $CM^{Z}$.
-We now make the parallel trends assumption in the outcome that is suitable for the set up we are considering here.
-We first define the similar notation as in section 2.4.
+We can rationalize Assumption 21 and 22 by noting that both of them and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units belonging to the exposed group and the compliers $CM^{Z}$.
+We now make the parallel trends assumption in the outcome suitable for the set up considered here.
+We first define the similar notation as in Subsection 2.4.
 Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
@@
 $\bar{\Delta}_{DF^{T},e}\equiv E[Y_{i,1}((1,0))-Y_{i,0}((1,0))|DF^{T},E_{i}=e].$
-Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups respectively:
+Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
 $\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},$
 $\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
-where the definition of the weight $w_{g^{t},e}$ is same as in section 2.4.
-As we previously noted in section 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups.
-The following assumption indeed make a such restriction in the current set up.
+where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
+As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups.
+The following assumption makes such a restriction in the current set up.
 Assumption 23 (Parallel trend assumption in the outcome).
-The expected time gain between exposed and unexposed groups is same, that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
+The expected time gain between exposed and unexposed groups is the same; that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
 The theorem below shows that a Wald-DID estimand is equal to the target parameter $LATET^{\prime}$ under Assumption 2-5 and Assumption 21-23.
 Theorem 6.
-Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$, that is, we have
+Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$; that is, we have
 $\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]}=LATET^{\prime}.$
 Proof.
-First, we consider the denominator of the Wald-DID estimand. From the similar argument in the proof of Theorem 1, one can show the following result:
+First, we consider the denominator of the Wald-DID estimand. From the argument in the proof of Theorem 1, one can show the following result:
 $E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$ $=Pr(CM^{Z}|E=1)$
 $=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (16)
-Here the second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
+The second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
 Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
@@
 The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
-We next decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
+Next, we decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
 $E[Y_{1}-Y_{0}|E=0]$
@@
 $=\bar{\Delta}_{0}.$ (18)
-Here the first equality follows from Table 2 and the second equality follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
+The first equality follows from Table 2 and the second follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
 From the result (18) with (17) and Assumption 23, we obtain
@@
 Combining the result (19) with (16), we obtain the desired result. ∎
-Appendix C Proof in section 4
+Appendix C Proofs in Section 4
 C.1   Proof of theorem 4
-First, we prove the consistency of our SLS estimator. Recall that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
+First, we prove the consistency of our SLS estimator. $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
 $\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
@@
 We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined.
-More precisely, we assume that cohort $e$ is non-empty, that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty, i.e., $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$.
+More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$.
 This assumption guarantees that each Wald-DID estimator
@@
 is also well defined.
-First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. We note that by the Law of Large Numbers and Slutsky’s theorem, we have
+First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and Slutsky’s theorem, we have
 […]
@@
 where the second equality follows from the Law of Iterated Expectations.
-By the similar arguments, we have
+By similar argument, we have
 […]
@@
 […]
-From the Slutsky’s theorem, we have
+From Slutsky’s theorem, we have
 […]
-Lastly, it remains to show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
+Lastly, we show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
 For the numerator, we have
@@
 $=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (21)
-The third equality follows from Assumption 8, 9 and 11. The forth equality follows from Assumption 13 and the fifth equality follows from Assumption 10.
-For the denominator, by the similar calculations,
+The third equality follows from Assumption 8, 9 and 11. The fourth equality follows from Assumption 13 and the fifth follows from Assumption 10.
+For the denominator, by similar calculations,
 $E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
@@
 where the final equality follows from Assumption 10 and Assumption 12.
-Combining the result (21) with (22), we obtain the desirable result:
+Combining (21) with (22), we obtain the desirable result:
 $\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}$
@@
 $\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
-We use the following fact repeatedly in the derivation. This fact is also found in de Chaisemartin and D’Haultfœuille (2018).
+The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation.
 Fact 1.
@@
 We also present the influence function of our STS estimator when repeated cross section data are available.
-[^9] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups..
+[^10] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who presented the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups..
 Let $\theta_{e,l}^{r}$ define
@@
 In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$.
-By the similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
+Using similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
 $\psi_{i,e,l}^{r}$ $=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}-[\gamma_{e,l}^{1}-\gamma_{e,l}^{2}]}$
@@
 $\pi_{e,t,g^{\prime}}=E[D_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,t,g^{\prime}}=E[D_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B,\hskip 5.69054ptt=e+l,e-1).$
-In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function:
+In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function.
 $\psi_{i,e,l,g}^{p}=\frac{1}{DID_{e,l,g}^{p}}(\zeta_{i,e,A}-\zeta_{i,e,B}-\zeta_{i,U,A}+\zeta_{i,U,B}),$
@@
 $\zeta_{i,e,g^{\prime}}=\frac{\mathbf{1}_{e,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}=e,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,g^{\prime}}]},\hskip 5.69054pt\zeta_{i,U,g^{\prime}}=\frac{\mathbf{1}_{U,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}\in U,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,g^{\prime}}]}.$
-In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function:
+In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function.
 $\psi_{i,e,l,g}^{r}$
@@
 Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
-In this appendix we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006).
-Hereafter, let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$.
+Here, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006).
+Let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$.
 Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
@@
 and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^12] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
-By the similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
+Using similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
 […]
@@
 […]
-Combining the result (30) with (34) and by the Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
+Combining the result (30) with (34) and by Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
 $\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (35)
@@
 $w_{e,t}^{2}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (37)
-where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and population residuals for England in cohort $t$ respectively.
+where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
 In Oreopoulos (2006), we can only identify $CLATT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in North Ireland in $1957$.
@@
 Figure 4: Decomposition result for the TWFEIV estimand in Oreopoulos (2006).
-Notes: The figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term).
+Notes: This figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term).
 The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$.
@@
 At the end of this appendix, we also revisit the ”issue” pointed out by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
-E.1   Comparing dCDH’s identifying assumptions to ours
+E.1   Comparing the identifying assumptions in dCDH and this article
 In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours.
-Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we will refer to Assumptions 14-18 in this article.
+Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-18.
 For now, we ignore the difference in the definition of target parameter between dCDH and this article.
 dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^13] In repeated cross section settings, the formal definition is provided in Assumption 29..
-We note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^13] This follows from two observations in dCDH.
+In dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^14] This follows from two observations in dCDH.
 First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively.
 Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation)..
-In the following, we therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively.
-[^14] Note that in Appendix B.2 in this article, we construct the group variable $E$ from the instrument path $Z$ in order to uncover Assumption 16 (No anticipation in the first stage).
-For the detail explanations for each assumption, see dCDH.
+We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively.
+[^15] Note that in Appendix B.2, we construct the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage), the hidden assumption in the previous literature.
+For the detailed explanations for each assumption, see dCDH.
 Assumption 24 (Fuzzy design).
@@
 $E[Y(1)-Y(0)|Z,T=1,D_{0}(Z)=1]=E[Y(1)-Y(0)|Z,T=0,D_{0}(Z)=1].$
-In Assumption 28, we modify the dCDH’s notation in order to ease the interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$, which dCDH define as $D(t)$.
+In Assumption 28, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
 We now state the relationships between dCDH’s assumptions and ours.
 We first explain the similarities.
-First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 in this article imply this condition [^15] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
+First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 imply this condition [^16] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
 Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article.
 The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002).
-Here, we point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26, that is, they assume that the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
-Finally, dCDH implicitly impose Assumption 14 (Exclusion Restriction) and Assumption 16 (No anticipation in the first stage) in this article when they formalize potential outcomes and treatment participation equations.
+We point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26; that is, the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
+Finally, dCDH implicitly impose Assumption 14 (Exclusion restriction) and Assumption 16 (No anticipation in the first stage) when they formalize potential outcomes and treatment participation equations.
 Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14.
 As one can see in Assumption 26, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
-We next describe the differences between dCDH’s assumptions and ours.
+The differences between dCDH’s assumptions and ours are as follows.
 First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 27), whereas we assume the parallel trends assumption in the unexposed outcome.
@@
 Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 26, whereas we do not impose such restriction.
-Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction in order for this estimand to identify the LATET.
+Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET.
 Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
 Remark E.1.1.
-When Assumption 25 is violated, dCDH consider the following assumption in order for the Wald-DID estimand to capture the SLATET.
+When Assumption 25 is violated, dCDH consider the following assumption for the Wald-DID estimand to capture the SLATET.
 Assumption 29 (Homogeneous treatment effect between exposed and unexposed groups).
 $SLATET=SLATET^{\prime},$
-where we define161616dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
+where we define171717dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
 When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@
 E.2   Strong and asymmetric restrictions in Fuzzy DID designs
-In this section, we show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups.
-To do so, we first prove some lemmas implied by dCDH’s assumptions.
-Based on these lemmas, we then create the tables as in Table 1 and Table 2 in section 2.4 in this article, which present all the types in exposed and unexposed groups respectively in Fuzzy DID designs, and confirm that dCDH’s assumptions ex ante exclude many types in exposed and unexposed groups and its restrictions is heavily assigned on the latter group.
-At the end of this section, we provide empirical implications of the distinction between dCDH’s restrictions and ours in this article.
+We show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups.
+First, we prove some lemmas implied by dCDH’s assumptions.
+Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs.
+These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group.
+At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours in DID-IV designs.
 First, we show that Assumption 26 (Treatment participation equation) implies the following assumption.
@@
 Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively.
-We note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption.
+Note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption.
 Indeed, in this article, we do not assume this restriction.
@@
 From this observation, we can view that in an unexposed group ($z=0$), Assumption 30 corresponds to a monotonicity assumption with respect to time, formally defined below.
-Assumption 31 (Monotonicity Assumption w.r.t. time in an unexposed group).
+Assumption 31 (Monotonicity assumption with respect to time in an unexposed group).
 […]
 Assumption 31 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$.
-In terms of the notation introduced in section 2.4 in this article, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability.
+In terms of the notation introduced in Subsection 2.4, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability.
 Hereafter, we consider the type of Assumption 31 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
@@
 $\iff$ $Pr(CM^{T}|Z=0)=0$ (40)
-The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (40) implies that there are no time comliers $CM^{T}$ in an unexposed group ($z=0$). ∎
-Finally, we show that the condition (38) in Assumption 24 (Fuzzy design) and Assumption 26 imply that in an exposed group ($z=1$), there are no units switching from treatment to non treatment between period 00 and period $1$.
+The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26. Equation (40) implies that there are no time compliers $CM^{T}$ in an unexposed group ($z=0$). ∎
+Finally, we show that the condition (38) in Assumption 24 (Fuzzy design) and Assumption 26 imply that in an exposed group ($z=1$), there are no units switching from treatment to non-treatment between period 00 and period $1$.
 In other words, in Assumption 30, we should have:
@@
 We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-Table 9 and Table 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^17] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: there exist five and three types in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
-The distinction between dCDH’s restrictions and ours is apparent when we compare Table 9 and Table 10 to Table 1 and Table 2 in this article [^18] We note that Tables 1-2 in this article is constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings.
+Tables 9 - 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^18] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
+The distinction between dCDH’s restrictions and ours is apparent when we compare Tables 9 - 10 to Table 1 - 2191919Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings.
 The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
@@
 $D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-These differences between dCDH’s restrictions and ours seem critical in many empirical settings: unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, thus the practitioners may want to avoid imposing these restrictions in empirical work.
+These differences between dCDH’s restrictions and ours seem critical in many empirical settings.
+Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, and the practitioners may want to avoid imposing these restrictions in empirical work.
 For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time.
-Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
+Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
 Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications.
 However, these assumptions give rise to two additional issues.
-First, when we impose the condition (38) in Assumption 24, we should test the plausibility of this condition from the data in practice because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$.
-Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group in this article) in our fuzzy setting”, they do not explain how we should proceed when this condition is not met in practice.
-Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time, and this complicates our estimation procedure in practice.
-Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility is not clear, especially when we have multiple periods and multiple groups.
+First, when we impose the condition (38) in Assumption 24, we should test its plausibility from the data because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$.
+Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group) in our fuzzy setting”, they do not explain how we should proceed when this condition is violated in practice.
+Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time.
+Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility are not clear, especially for multiple period and multiple group settings.
 E.3   Decomposing the SLATET
-In the previous section, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups.
-In this section, under these restrictions, we show the decomposition result for dCDH’s target parameter, the SLATET, and state the concerning implications.
+In the previous subsection, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups.
+In this subsection, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
 The following theorem presents the decomposition result for the SLATET under Fuzzy DID designs, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
@@
 The proof directly follows from Table 9 in an exposed group. ∎
-Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other parameter measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
+Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
 This theorem has two concerning implications.
 First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters.
-Although the SLATET indeed captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
-Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for the treatment.
-Suppose that the instrument represents the important policy lever for the researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group.
-In this situation, the LATET is a policy-relevant parameter by construction.
-On the other hand, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group.
-As the fraction of the time compliers $CM^{T}$ increases in that group, the SLATET puts more weight on the treatment effects among the time compliers, and this would contaminate the interpretation of the magnitude of this parameter even in such policy-relevant context.
+Although the SLATET captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
+Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for treatment.
+Suppose that the instrument represents an important policy lever for researchers, who are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group.
+In this situation, our target parameter, the LATET, is a policy-relevant parameter by construction.
+By contrast, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group.
+As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this parameter even in such policy-relevant contexts.
 Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 29.
-When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 in order for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
-We show that this assumption can not be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups.
-To do so, we first note that if we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2.
+When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
+However, we show that this assumption cannot be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups.
+If we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2.
 Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
@@
 Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group.
-This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than one might expect.
+This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than what one might expect.
 E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
-Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that while their argument is correct, it is misguided.
-During this discussion, we also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2.
+Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that their argument is misguided.
+We also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2.
 Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
-dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) in order to identify the SLATET, and propose alternative estimands that do not rely on this assumption.
-We first show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article.
-We next argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
+dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) to identify the SLATET, and propose alternative estimands that do not rely on this assumption.
+We show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article.
+Next, we argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
 First, we clarify the interpretation of Assumption 28 by introducing the additional notation.
-This assumption is actually being conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population.
+This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population.
 For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$.
 Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
-Then, from Table 9 and Table 10, we can decompose Assumption 28 in the following:
+Then, from Table 9 and Table 10, we can decompose Assumption 28 as follows:
 $E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
@@
 where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 26.
-Equation (44) shows that in an unexposed group ($z=0$), Assumption 28 requires the stable treatment effect in the time always-takers $AT^{T}$.
+Equation (44) shows that in an unexposed group ($z=0$), Assumption 28 requires the stable treatment effect for the time always-takers $AT^{T}$.
 Equation (45) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
@@
 First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand.
-We can decompose this quantity into two terms:
+We can decompose this quantity into two terms.
 $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$
@@
 The first term disappears if we assume condition (44).
-Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows:
+Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows.
 $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$
@@
 $+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1),$
-where the type SW (switcher) includes the units who switch from non treatment to treatment during the two periods.
+where the type SW (switcher) includes the units who switch from non-treatment to treatment during the two periods.
 The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term.
 The first term disappears if we assume condition (45).
-Here, we note that we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
+Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
 Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 25 and Assumption 26, the Wald-DID estimand captures the SLATET under Assumption 27 (parallel trends assumption in the untreated outcome) and conditions (44) and (45).
-From the discussions so far, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
-Here, we add few remarks to clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units in order for the Wald-DID estimand to identify the SLATET.
-It turns out that this also stems from their reliance on the parallel trends assumption in the untreated outcome.
+From the discussions, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
+Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET.
+This also stems from their reliance on the parallel trends assumption in the untreated outcome.
 Remark E.4.1.
 dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10).
-This identification result seems not intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
+This identification result does not seem intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
 Remark E.4.2.
 dCDH exclude the time compliers $CM^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for that type, we additionally capture the treatment effects among the time compliers $CM^{T}$ from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which implies that the SLATET in an exposed group is offset by the SLATET’ in an unexposed group.
-So far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome.
+We have confirmed that dCDH’s argument regarding the use of the Wald-DID estimand rests on the parallel trends assumption in the untreated outcome in DID designs.
 Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
-First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group; in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and thus we can not observe the untreated outcomes for these types in an unexposed group.
+First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group.
+This is because in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and we cannot observe the untreated outcomes for these types in an unexposed group.
 Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general.
@@
 (2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment.
-In the author’s setting, the parallel trends assumption in the untreated outcome requires that the data contain the parents with zero education attainment, which is unrealistic in practice.
+In the authors’ setting, the parallel trends assumption in the untreated outcome requires the data to contain parents with zero education attainment, which is unrealistic and not what the authors would have in mind in practice.
 Finally, this assumption is generally not testable using the pre-exposed period data in a given application.
-This seems apparent because some units can already adopt the treatment before period 00, and we can not compare the pretrends of the untreated outcome between exposed and unexposed groups.
-Although dCDH propose the placebo test in their appendix, we show that it is incomplete for checking the plausibility of the parallel trends assumption in the untreated outcome.
-dCDH’s test consists of two steps.
-First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$:
+This is apparent, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
+dCDH propose the placebo test in their appendix, but we show that it is not enough for checking the plausibility of the parallel trends assumption in the untreated outcome.
+dCDH’s test has two steps.
+First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
 $E[D|Z=1,T=0]-E[D|Z=1,T=-1]=0,$ (46)
@@
 $E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]=E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (48)
-Although dCDH ”assume” the conditions (46) and (47), dCDH also note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27.
-We therefore describe the dCDH’s test in two steps.
+Although dCDH ”assume” the conditions (46) and (47), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27.
+We, therefore, describe the dCDH’s test in two steps.
 dCDH’s test has some drawbacks.
-First, the conditions (46) and (47), which we should check in the first step, seem strong, but dCDH do not explain how we should proceed when these conditions are violated in practice.
-Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods; it does not compare the evolution of the mean untreated outcome between exposed and unexposed group before period $T=0$.
+First, the conditions (46) and (47), which should be checked in the first step, are strong, but dCDH do not explain the way ahead when these conditions are violated in practice.
+Second, in this test, the second step does not directly test Assumption 27 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$.
 Indeed, dCDH describe the equality (48) as a necessary condition for Assumption 27 and Assumption 28 to be satisfied under the conditions (46) and (47).
-Finally, it is unclear how we should conduct this test when we have multiple pre-exposed period data.
-Although dCDH assume ”for instance that data is available for the period $T=-1$”, they do not provide any guidance on how to perform this test in multiple pre-exposed period settings.
-Surprisingly, dCDH’s test corresponds to our pretrends test in section 5.3 in this article if we replace the conditions (46)- (47) in the first step with the weaker one:
+Finally, it is unclear how this test should be conducted with multiple pre-exposed period data.
+dCDH assume ”for instance that data is available for the period $T=-1$”, but they do not indicate how this test can be performed in multiple pre-exposed period settings.
+Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (46)- (47) in the first step with the weaker one.
 $E[D|Z=1,T=0]-E[D|Z=1,T=-1]=E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (49)
-Under repeated cross section settings, one can figure out that the conditions (48) - (49) coincide with the conditions (3) -(4) in section 5.3 in this article.
+Under repeated cross section settings, the conditions (48) - (49) coincide with the conditions (3) -(4).

### v5

@@
-
-Instrumented Difference-in-Differences with Heterogeneous Treatment Effectsthanks: This paper won the Kanematsu Prize from the Research Institution for Economics and Business Administration, Kobe University, in 2023.
-It also won the Best Young Scholar Award at SETA (Symposium on Econometrics and Applications) Conference, Taipei, in 2024.
-I am grateful to my advisors, Daiji Kawaguchi and Ryo Okui, for their continued guidance and support.
-I am also thankful to Alberto Abadie, Shakeeb Khan, Masamune Iwasawa, Shoya Ishimaru, Shosei Sakaguchi, Takahide Yanagi, and the three anonymous referees of the Kanematsu Prize for their insightful comments.
-I gratefully acknowledge the support provided by the JSPS KAKENHI Grant JP 24KJ0817.
-All errors are my own.
-Sho Miyaji Graduate School of Economics, The University of Tokyo, 7-3-1 Hongo, Bunkyoku, Tokyo 113-0033, Japan; Email: shomiyaji-apple@g.ecc.u-tokyo.ac.jp.
-(July 23, 2024)
-Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment.
+onlineReferences
+Abstract
+Many studies exploit variation in policy adoption timing across units as an instrument for treatment.
 This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
 In this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the local average treatment effect on the treated (LATET).
-In contrast to Fuzzy DID design considered in de Chaisemartin and D’Haultfœuille (2018), our DID-IV design does not ex ante require strong restrictions on treatment adoption behavior across units.
-Additionally, our target parameter, the LATET, is policy-relevant if the instrument is based on the policy change of interest to the researcher.
-We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design, and propose an estimation method that is robust to treatment effect heterogeneity.
-We illustrate our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom.
-In this application, the two-way fixed effects instrumental variable regression, which is the conventional approach to implement a staggered DID-IV design, yields a negative estimate, whereas our estimation method indicates a substantial gain from schooling.
+We extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
+Moreover, we propose a credible estimation method in this design that is robust to treatment effect heterogeneity.
+We illustrate the empirical relevance of our findings, estimating returns to schooling in the United Kingdom.
+In this application, the two-way fixed effects instrumental variable regression, the conventional approach to implement DID-IV designs, yields a negative estimate.
+By contrast, our estimation method indicates a substantial gain from schooling.
+Abstract
+[…]
 Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
-JEL Classification: C21, C23, C26
 1   Introduction
-Instrumental variable (IV) strategies are widely used for causal inference across many fields in economics.
-Despite its popularity, empirical researchers commonly face two challenges in employing the IV method in practice.
-The first challenge lies in the internal validity of the IV estimate; finding an instrument that is effectively randomized across units, especially in observational studies, proves to be difficult.
-The second challenge lies in the external validity of the IV estimate; under heterogeneous treatment effects, a linear IV estimand captures the average treatment effect, for those who are affected by the instrument (Imbens and Angrist (1994)).
-However, this local average treatment effect (LATE) may not be policy-relevant if the instrument does not represent the policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
-To enhance at least the external validity of the IV estimate, many studies have exploited variation in the timing of policy adoption across units as an instrument for treatment, and use instrumental variable techniques.
-For example, Duflo (2001) estimates returns to schooling in Indonesia, leveraging variation arising from the new school construction program across regions and cohorts as an instrument for education attainment.
-As another example, Black et al.
-(2005) estimates the causal link between parents’ and children’s education attainment, using variation occurring from the different timing of implementation of school reform across municipalities and cohorts as an instrument for parents’ education attainment.
-Notably, however, the underlying identification strategy behind these studies is inconsistent with IV designs: the instrument based on these policy changes is not randomized across units in reality, potentially threatening the internal validity of the IV estimate.
-Rather, the underlying identification strategy is similar to difference-in-differences (DID) designs, under where the effect of the policy shock (treatment) on the outcome is identified by exploiting the variation of that policy change across units and over time.
+To identify the effect of a treatment on an outcome, many studies have exploited variation in policy adoption timing across units as an instrument for treatment.
+For example, to estimate returns to schooling in Indonesia, Duflo (2001) leverages variation in the timing of introduction of school construction program across regions as an instrument for education attainment.
+Similarly, to estimate the causal link between parents’ and children’s education attainment, Black et al.
+(2005) leverage variation in the timing of implementation of school reforms across municipalities as an instrument for parents’ education attainment.
+Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that the timing variation of a policy shock is used for identifying the treatment effects.
+Meanwhile, it also differs from DID designs, in that the variation is used to construct an instrument rather than a treatment.
 In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
-We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity in multiple time periods settings.
-Importantly, our DID-IV design does not require the randomization assumption of the instrument.
-In this design, we can achieve the internal and external validity of the IV estimate if the parallel trends assumptions in the treatment and the outcome are plausible in a given application, and the instrument is based on the policy shift of interest to the researcher.
-First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument over two periods (unexposed group), whereas some units become exposed in the second period (exposed group).
-In this setting, our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period.
-Our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups.
-This framework comes from a simple and motivating observation: in the two-group/two-period ($2\times 2$) settings, a popular estimand, which we call the Wald-DID estimand, scales the DID estimand of the outcome by the DID estimand of the treatment between the two groups.
-We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
-In DID-IV designs, the interpretation of the parallel trends assumption in the outcome is less clear than that of the parallel trends assumption in DID designs because it does not depend on the untreated outcome.
-Rather, it depends on the outcome under no exposure to the instrument, which we call an unexposed outcome.
-Motivated by this observation, we then clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
-Specifically, we first uncover the selection mechanism behind DID-IV designs, showing that, in DID-IV settings, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of the policy shock.
+We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity.
+We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom.
+In this empirical application, we demonstrate that the estimation method matters in this design in practice.
+First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument during the two periods (unexposed group), while some units become exposed in the second period (exposed group).
+[^1] Note that this two-period/two-group ($2\times 2$) setting is already considered in de Chaisemartin (2010) and Hudson et al.
+(2017), and our $2\times 2$ DID-IV design is mainly based on these studies.
+In this paper, we revisit $2\times 2$ DID-IV designs for two reasons.
+First, we aim to complement de Chaisemartin (2010) and Hudson et al.
+(2017) along two dimensions.
+Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs, and uncover the additional identifying assumption not give in the previous literature.
+In addition, we investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in DID-IV designs.
+Second, we aim to compare $2\times 2$ DID-IV designs to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
+While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) treat the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
+In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups.
+Our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period.
+We show that in this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the LATET.
+In DID-IV designs, the parallel trends assumption in the outcome requires careful assessment because it does not depend on the potential outcomes, but depends on the outcome under no exposure to the instrument, which we call an unexposed outcome.
+Give this, we next clarify the interpretation of this assumption in DID-IV settings.
+Specifically, we first point out that, in DID-IV designs, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of a policy shock (instrument).
 Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers.
-For instance, the time compliers do not adopt the treatment in the first period, but become adopt it in the second period without instrument.
-Subsequently, we show that the average time trend in the unexposed outcome equals a weighted average of the indirect effect of time on the outcome through treatment in each time group.
-We call this an expected time gain, and show that the parallel trends assumption in the outcome in DID-IV designs requires the same expected time gain between exposed and unexposed groups.
-de Chaisemartin and D’Haultfœuille (2018) (hereafter, ”dCDH”) also consider the same setting as in this article, but formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period.
-Accordingly, the identifying assumptions in dCDH are also different from those used here.
-Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET under Fuzzy DID designs if the treatment effect is stable over time, and propose the alternative estimands that are free from the strong restriction on treatment effect heterogeneity.
-Next, motivated by the disagreement between this article and dCDH, we compare DID-IV designs to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
-First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units.
-These restrictions turn out to be asymmetric between exposed and unexposed groups, and difficult to defend from the institutional knowledge of the problem in a given application.
-Then, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers, who are affected by time but not affected by instrument.
-This decomposition result for the SLATET has two concerning implications.
-First, the interpretation of the SLATET is less clear than that of the LATET in DID-IV designs.
-Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
-We also revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and demonstrate that their argument is correct but misguided.
-First, we show that their argument stems from their reliance on the parallel trends assumption in DID designs.
-We then argue that this assumption is unsuitable for DID-IV settings for three reasons.
-First, this assumption generally fails to capture the average time trends of the outcome even in an unexposed group.
-This is because in DID-IV settings, units are allowed to receive the treatment in the absence of the policy shock (instrument).
-Second, in most applications of the DID-IV method, it is generally infeasible to impute the untreated potential outcomes (e.g., Duflo (2001), Black et al.
-(2005)).
-For instance, in Black et al.
-(2005), this assumption requires that the data contain the units with zero education attainment, which is unrealistic in practice.
-Finally, in DID-IV settings, we cannot directly test this assumption using the pre-exposed period data in general.
-Indeed, we show that the placebo test proposed by dCDH is incomplete for assessing the plausibility of this assumption.
+Subsequently, we show that the average time trend in an unexposed outcome equals a weighted average of the effect of time on an outcome through treatment in each time group.
+We call this an expected time gain, and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
+de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
+The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period.
+Because the target parameter differs from each other, the identifying assumptions in dCDH are also different from those used here.
+Motivated by the differences between this article and dCDH, we next examine the detailed connections between DID-IV and Fuzzy DID, and point out the issues inherent in Fuzzy DID.
+First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units, and these restrictions are asymmetric between unexposed and exposed groups.
+Second, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers.
+This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher.
+This is because the SLATET is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by instrument.
 Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
-In this setting, units start being exposed to the instrument at a particular point in time and remain exposed to that instrument afterward.
-Indeed, in most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation.
+In most DID-IV applications, researchers exploit variation in policy adoption timing across units in more than two periods, instrumenting for the treatment with the natural variation.
 The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g.
@@
 We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
-Specifically, in this design, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument.
+Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument.
 We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument.
-Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings and state the interpretation of each assumption.
+Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings, and state the interpretation of each assumption.
 We extend our DID-IV designs in several directions.
 We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment.
-We also consider extensions to repeated cross sections, and triple DID-IV designs.
-Lastly, we consider the case when we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
+We also consider extensions to repeated cross sections.
+Lastly, we consider the case where we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
 Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects.
-In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable regressions.
-In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs.
-Specifically, Miyaji (2024) shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
+In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable (TWFEIV) regressions (e.g.
+Johnson and Jackson (2019), Lundborg et al.
+(2014), Black et al.
+(2005), Akerman et al.
+(2015), and Bhuller et al.
+(2013)).
+In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time.
 Our proposed method avoids this issue and is robust to treatment effect heterogeneity.
@@
 We call this a stacked two stage least squares (STS) regression and ensure its validity.
-Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects under staggered DID-IV designs, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$.
-We also discuss the procedure of pre-trend tests to check the validity of the parallel trends assumptions in the treatment and the outcome in DID-IV designs.
-We illustrate our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment.
-In this application, we first assess the plausibility of DID-IV identification assumptions implicitly imposed by Oreopoulos (2006).
+Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$.
+We also discuss the procedure of pre-trends tests to check the validity of the parallel trends assumptions in DID-IV designs.
+We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment.
+In this application, we first assess the plausibility of the DID-IV identification strategy implicitly imposed by Oreopoulos (2006).
 We then estimate the TWFEIV regression in the author’s setting.
 We find that the TWFEIV estimate is strictly negative and not significantly different from zero.
-The TWFEIV regression performs the ”bad comparisons” (c.f.
-Goodman-Bacon (2021)) in the author’s setting: we treat the already exposed units as controls during the periods after both regions were affected by the policy change.
 Finally, we use our estimation method to reassess returns to schooling in the UK.
@@
 Our DID-IV design can be applied to various empirical settings.
-First, it can cope with the general adoption process of the treatment when we are interested in the effect of the treatment on the outcome in DID designs.
-For instance, consider that some units already receive the treatment in period 00 before the policy shock, which distributes the treatment in period $1$.
+First, it can cope with the general adoption process of the treatment, which the canonical DID designs cannot, when we are interested in the effect of a treatment on an outcome.
+For instance, suppose that some units already receive the treatment in period 00 before the policy shock that distributes the treatment in period $1$.
 In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data.
 By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET.
-In other words, we can view that DID-IV is DID with the noncompliance of the assigned treatment.
+In other words, we can view that DID-IV is DID with noncompliance of the assigned treatment.
 Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs.
@@
 In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality.
-However, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument).
-Once these groups are created under this design, one can identify the LATET by comparing the evolution of the treatment and the outcome between the two groups.
-Finally, our DID-IV design can overcome the endogeneity issue in empirical work.
-In practice, due to the endogenous adoption of the treatment, researchers often fail to identify the treatment effect via simple linear regressions or two-way fixed effects regressions.
-Using our DID-IV design allows addressing the common identification challenge by leveraging variation in the timing of policy adoption across units as an instrument for treatment.
-Importantly, in this design, the identifying variation comes not from the random assignment of the policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome.
-This identification strategy would be less demanding than IV designs in some applications, and enhances the internal and external validity of the IV estimate in practice.
+By contrast, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument).
+Once these groups are constructed, one can identify the local average treatment effect by comparing the evolution of the treatment and the outcome between the two groups.
+Finally, our DID-IV design allows a researcher to address the endogeneity issue in a given application.
+In many empirical settings, a researcher often fails to identify the treatment effects due to the endogenous adoption of the treatment or the potential violation of the parallel trends assumption under DID designs.
+In this situation, our DID-IV design offers an alternative identification strategy, constructing an instrument based on the timing variation of a policy shock.
+Notably, in this design, the identifying variation comes not from the random assignment of a policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome.
+This identification strategy would be less demanding than IV designs in certain applications, and enhance the internal and external validity of the IV estimate in practice (Imbens and Angrist (1994); Heckman and Vytlacil (2001); Heckman and Vytlacil (2005)).
 The rest of the paper is organized as follows.
 The following subsection discusses the related literature.
-Section 2 establishes DID-IV designs in a simple setting with two periods and two groups.
+Section 2 establishes DID-IV designs in two periods and two groups settings.
 Section 3 formalizes the target parameter and identifying assumptions in staggered DID-IV designs.
@@
 1.1   Related literature
-Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al. (2017); de Chaisemartin and D’Haultfœuille (2018)) and contributes to this literature in three important ways.
-First, this paper investigates the detailed connections between DID-IV and Fuzzy DID designs proposed by de Chaisemartin and D’Haultfœuille (2018).
-In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that the Wald-DID estimand identifies the local average treatment effect on the treated (LATET) under parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption [^1] In biometrics, Ye et al.
-(2023) also consider the ”instrumented difference-in-differences” design.
-However, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to identify this parameter..
-Hudson et al.
+Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al.
+(2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al.
+(2024)) and contributes to this literature in three important ways.
+[^2] While Ye et al.
+(2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to capture this parameter.
+We therefore see that Ye et al.
+(2023) are distinct from the recent DID-IV literature.
+First, this paper investigates the detailed connections between DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
+In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that if parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption hold, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET).
+Following de Chaisemartin (2010), Hudson et al.
 (2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings.
 Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs.
-The settings are the same between the two designs, but the target parameter and identifying assumptions are different from each other.
 In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al.
+(2017) in two ways.
+Specifically, the framework we consider is mainly based on de Chaisemartin (2010) and Hudson et al.
 (2017).
-Specifically, we introduce the path of the instrument in $2\times 2$ DID-IV designs and uncover the ”hidden assumption” in the previous literature.
-This assumption requires the no anticipatory behavior of the treatment adoption before the exposure to the instrument.
-We also investigate the selection mechanism behind $2\times 2$ DID-IV designs, and clarify the interpretation of the parallel trends assumption in the outcome.
-Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID designs, and point out the issues inherent in Fuzzy DID designs.
-Specifically, we show that the identifying assumptions in Fuzzy DID designs ex ante requires strong restrictions on treatment adoption behavior across units, and the target parameter in Fuzzy DID designs can be decomposed into a weighted average of two different causal parameters.
-The discussion organizes the relationship between the two designs in the previous literature, which essentially treat the same setting, and indicates the relative advantages of DID-IV over Fuzzy DID.
-Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units.
-In practice, to estimate the treatment effects, empirical researchers often leverage variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g., Black et al.
+Nevertheless, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover the additional identifying assumption not given in the previous literature.
+We also investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
+Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID, and point out the issues inherent in Fuzzy DID.
+The discussion organizes the relationship between the two designs, which essentially treat the same setting, in the previous literature, and indicates the relative advantages of DID-IV over Fuzzy DID.
+Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design.
+In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al.
 (2005), Lundborg et al.
@@
 No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings.
-Our staggered DID-IV designs ensure the theoretical validity of the underlying identification strategy behind these studies, and allow the practitioners to estimate the local average treatment effect on the treated even when the treatment adoption is endogenous over time.
+In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions.
+Our staggered DID-IV designs allow the practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
 Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects.
 When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g.
-Black et al.
-(2005), Lundborg et al.
+Johnson and Jackson (2019), Lundborg et al.
+(2014), Black et al.
+(2005), Akerman et al.
+(2015), and Bhuller et al.
 (2013)).
-In companion paper (Miyaji (2024)), however, we show that the TWFEIV estimand generally fails to summarize the treatment effects under staggered DID-IV designs if the effect of the instrument on the treatment or the outcome evolves over time.
-The proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
-Our paper is also related to a growing DID literature.
-Several recent studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications in the presence of heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al.
+In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or the treatment evolves over time.
+[^3] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix.
+Our proposed estimation method would serve as an alternative to the TWFEIV estimator and make DID-IV designs more credible in a given application.
+In work related to this paper, de Chaisemartin et al.
+(2024) also consider DID-IV designs following de Chaisemartin (2010) and Hudson et al.
+(2017).
+This paper is distinct from theirs in three ways.
+First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument.
+Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument.
+Third, they consider the Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
+Our paper is also related to a growing DID literature in two ways.
+First, this paper provides an alternative identification strategy to DID designs, addressing cases where the parallel trends assumption is not satisfied or no control group exists in a given application.
+When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often concern that the parallel trends assumption is not plausible in practice.
+To address this, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g.
+Miller and Segal (2019)).
+In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences.
+Our DID-IV design is a natural extension of DID design, in that the timing variation is used to construct an instrument rather than a treatment.
+Second, in line with the DID literature, this paper develops a reliable estimation method in DID-IV designs in the presence of heterogeneous treatment effects.
+Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects.
+(Athey and Imbens (2022); Borusyak et al.
 (2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)).
-All of them have documented that regression coefficients by conventional approaches may fail to properly summarize treatment effects under DID designs.
-Some of them have proposed an alternative estimation method that delivers a sensible estimand and is robust to treatment effect heterogeneity.
-None of them, however, have explicitly considered the causal interpretation of the TWFEIV estimand when researchers exploit variation in policy adoption timing across units as an instrument for treatment.
-This paper formalizes the rationale behind using the TWFEIV regression as an instrumented difference-in-differences design, and provide a reliable estimation method in multiple periods settings.
-Miyaji (2024) builds on this work to study the properties of the TWFEIV estimand under staggered DID-IV design and points out the issue of implementing this design via TWFEIV regressions.
+Some of them have proposed a credible estimation method that delivers a sensible estimand and is robust to treatment heterogeneity.
+Along with the recent DID literature, this paper proposes an alternative estimation method to the TWFEIV regression, and illustrate its usefulness through an empirical application.
 2   DID-IV in two time periods
@@
 We first establish the target parameter and the identifying assumptions in this design.
-We then uncover the selection mechanism behind this design, and clarify the interpretation of the parallel assumption in the outcome.
-We next describe the relationship between DID and DID-IV, indicating that DID-IV is DID with the noncompliance of the assigned treatment.
-At the end of this section, we also describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
+We then investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel assumption in the outcome.
+Finally, we describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
 2.1   Set up
@@
 We denote $E_{i}\in\{0,1\}$ as the group variable: $E_{i}=1$ if unit $i$ is exposed to the instrument in period $t=1$ (exposed group) and $E_{i}=0$ if unit $i$ is not exposed to the instrument in period $t=1$ (unexposed group).
-In contrast to the sharp assignment of the instrument, we allow the general adoption process for the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
-In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock, which occurs in period $t=1$ and encourages people to adopt the treatment.
-For instance, Duflo (2001) estimates returns to schooling in Indonesia, using a new school construction program across regions as an instrument for education attainment.
+In contrast to the sharp assignment of the instrument, we allow the general adoption process of the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
+In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock that encourages people to adopt the treatment in period $t=1$.
+For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
 Next, we introduce the potential outcomes framework.
@@
 We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
+Since the treatment and the instrument takes only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
+$D_{i,t}=\sum_{z\in\mathcal{S}(Z)}\mathbf{1}\{Z_{i}=z\}D_{i,t}(z),\hskip 8.53581ptY_{i,t}=\sum_{z\in\mathcal{S}(Z)}\sum_{d\in\mathcal{S}(D)}\mathbf{1}\{Z_{i}=z,D_{i}=d\}Y_{i,t}(d,z).$
 We make a no carryover assumption on the potential outcomes $Y_{i,t}(d,z)$.
@@
 Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$.
-In the DID literature, several studies have used this assumption in non-staggered treatment settings (e.g.
+In the DID literature, several studies adopt this assumption in non-staggered treatment settings (e.g.
 de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)).
-[^2] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes because it takes only four values.
-We therefore consider this case in Appendix B as an extension.
-One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatments, or multiple time periods, but it would require cumbersome notations and complicate the definition of the target parameter, going beyond scope of this paper.
-Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$, which describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$.
-The first element $D_{i,1}((0,0))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an unexposed group, and the second element $D_{i,1}((0,1))$ represents the potential treatment choice in period $t=1$ if unit $i$ belongs to an exposed group.
+[^4] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes, as it takes only four values.
+We therefore consider this case in online Appendix Section A.3 as an extension.
+One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatment, or multiple time periods, but it requires cumbersome notations and complicates the definition of the target parameter, being not practical for empirical applications.
+Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$ that describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$.
 Following the terminology in Imbens and Angrist (1994), we define $G_{i}^{Z}=(0,0)\equiv NT^{Z}$ as the never-takers, $G_{i}^{Z}=(1,1)\equiv AT^{Z}$ as the always-takers, $G_{i}^{Z}=(0,1)\equiv CM^{Z}$ as the compliers, and $G_{i}^{Z}=(1,0)\equiv DF^{Z}$ as the defiers.
-Henceforth, we keep Assumption 1. In the next subsection, we use the notations developed so far to define our target parameter in $2\times 2$ DID-IV designs.
-2.2   Target parameter in two time periods
+Henceforth, we keep Assumption 1. In the next subsection, we define the target parameter in $2\times 2$ DID-IV designs.
+2.2   Target parameter in $2\times 2$ DID-IV designs
 In $2\times 2$ DID-IV designs, our target parameter is the local average treatment effect on the treated (LATET) in period $t=1$ defined below.
@@
 This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$.
-In the DID-IV literature, de Chaisemartin (2010) consider the same target parameter, and Hudson et al.
+In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, and Hudson et al.
 (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$.
@@
 (2022)).
-We define LATET as the target parameter for two reasons.
-First, this parameter would be particularly of interest if the instrument reflects a policy change of interest (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
-Duflo (2001), for instance, is interested in ”whether investments in infrastructure can cause an increase in education attainment, and whether an increase in education attainment causes an increase in earnings”.
-To answer this question, the author exploits a new school construction program across regions as an instrument for education attainment.
-In this context, the LATET would be a policy-relevant parameter in the presence of heterogeneous treatment effects.
-Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT).
-Subsection 2.5 shows that both causal parameters measure the treatment effects among the units affected by instrument (policy shock) and belonging to an exposed group $(E_{i}=1)$.
-Remark 1.
-de Chaisemartin and D’Haultfœuille (2018) define the switcher local average treatment effect on the treated (SLATET) as their target parameter in the same setting as here333de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings.
-Therefore, we rewrite their target parameter by the notation used in this article..
-Definition.
-The switcher local average treatment effect on the treated (SLATET) is
-$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
-$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
-where the switcher (SW) is the units who become treated in period $1$.
-This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW).
-As shown in Appendix E, this parameter turns out to be different from the LATET and can be decomposed into a weighted average of two different causal parameters.
-In section 2.6, we provide a detailed discussion of the connections between this article and de Chaisemartin and D’Haultfœuille (2018).
-2.3   Identification assumptions in two time periods
-This subsection establishes DID-IV designs in a simple setting with two groups and two periods.
-In this setting, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
+We define the LATET as the target parameter for two reasons.
+First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
+Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
+2.3   Identification assumptions in $2\times 2$ DID-IV designs
+This subsection formalizes the identification assumptions in $2\times 2$ DID-IV designs.
+In two periods and two groups settings, a popular estimand is the ratio between the DID estimand of the outcome and the DID estimand of the treatment (Duflo (2001), Field (2007)):
 $w_{DID}$ $=\frac{E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]}{E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]}.$
 Following the terminology in de Chaisemartin and D’Haultfœuille (2018), we call this the Wald-DID estimand.
-This estimand is obtained from the IV regression of the outcome on the treatment where the group and post-time dummies are included instruments and the interaction of the two is excluded instrument.
-In practice, the Wald-DID estimand $w_{DID}$ is commonly interpreted as measuring the local average treatment effect (LATE) in the presence of heterogeneous treatment effects, but the underlying identifying assumptions are not stated formally.
-Duflo (2001), for instance, notes that ”if returns to education are not constant, the 2SLS estimates are a weighted average of the returns to education for people who are affected by the instruments (Angrist and Imbens (1995))”.
 We consider the following identifying assumptions for the Wald-DID estimand to capture the LATET.
-Henceforth, we assume that the denominator of the Wald-DID estimand is different from zero without loss of generality.
 Assumption 2 (Exclusion restriction for potential outcomes).
@@
 Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes.
-For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the potential outcome if the instrument path is $z$:
+For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
 $Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
 Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes.
-Note that we can not observe different potential outcomes for the same unit at the same time.
-For instance, when unit $i$ is assigned to $z=(0,0)$, that is, unit $i$ is not exposed to the instrument during the two periods, we can observe only $Y_{i,t}(D_{i,t}((0,0)))$, and can not observe $Y_{i,t}(D_{i,t}((0,1)))$ for that unit.
+[^5] Exposed and unexposed outcomes are not new concepts in econometrics.
+For example, under IV designs, we compare the expectation of the outcome with instrument and the one without instrument in the numerator of the wald estimand: $E[Y|Z=1]-E[Y|Z=0]$.
+This comparison corresponds to the one between the expectation of the exposed outcome and the expectation of the unexposed outcome; that is, $E[Y|Z=1]-E[Y|Z=0]=E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0]$.
 Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
-Assumption 3 (Monotonicity assumption in period $t=1$).
+Assumption 3 (Monotonicity assumption at period $t=1$).
 $Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 8.53581pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
 This assumption requires that the instrument path affects the treatment choice at period $t=1$ in a monotone (uniform) way.
-It implies that the group variable $G_{i}^{Z}=(D_{i,1}((0,0)),D_{i,1}((0,1)))$ can take three values with non-zero probability.
+It implies that the group variable $G_{i}^{Z}$ can take three values with non-zero probability.
 In the DID-IV literature, de Chaisemartin (2010) and Hudson et al.
@@
 Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group.
-This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex-ante not known for all the units in an exposed group.
-This is ”hidden assumption” in the DID-IV literature: de Chaisemartin (2010) and Hudson et al.
-(2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
-In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways.
+[^6] In the DID literature, recent studies impose the no anticipation assumption on untreated potential outcomes in several ways.
 Callaway and Sant’Anna (2021) and Sun and Abraham (2021) assume the average version of the no anticipation assumption, whereas Athey and Imbens (2022) assume it for all units $i$.
@@
 (2023) take the intermediate approach: they adopt the no anticipation assumption for the treated units.
-In this paper, the no anticipation assumption on potential treatment choices is in line with that of Roth et al.
+The no anticipation assumption on potential treatment choices in Assumption 4 is in line with that of Roth et al.
 (2023).
+[^7] This assumption is not provided in the previous DID-IV literature: de Chaisemartin (2010) and Hudson et al.
+(2017) implicitly impose this assumption by writing observed treatment choice in period $t=0$ as $D_{i,0}(0)$.
+This assumption restricts the anticipatory behavior and would be plausible if the instrument path is ex ante not known for all the units in an exposed group.
 Next, we impose the parallel trends assumptions in the treatment and the outcome. de Chaisemartin (2010) and Hudson et al. (2017) make the similar assumptions.
@@
 This assumption requires that the expectation of the treatment between exposed and unexposed groups would have followed the same path if the assignment of the instrument had not occurred.
-For instance, in Duflo (2001), this assumption requires that the evolution of mean education attainment would have been the same between exposed and unexposed groups if the policy shock had not occurred during two periods.
+For instance, in Duflo (2001), this assumption requires that the evolution of mean education attainment would have been the same between exposed and unexposed groups if the policy shock had not occurred during the two periods.
 Assumption 6 (Parallel Trends Assumption in the outcome).
-$E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]=E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
+$E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
+$=$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
 Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.
 For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
-When empirical researchers exploit variation arising from policy shock as an instrument for the treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6.
-For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on ”the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
-The theorem below shows that if Assumptions 1 - 6 hold, the Wald-DID estimand captures the LATET in period $1$.
+Note that Assumption 6 depends on the unexposed outcome, and does not depend on the potential outcomes directly.
+In Subsection 2.4, however, we show that the average time trends of the unexposed outcome in Assumption 6 can be written as the weighted average of the expectation of the potential outcomes, and this assumption can be interpreted as the restriction on the time effect between exposed and unexposed groups.
+When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6.
+For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
+Finally, we assume a relevance condition.
+This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined.
+[^8] Note that Assumption 3, 4, 5, and 7 imply that the share of compliers in period $1$ is non-zero.
+$E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]=Pr(CM^{Z}|E_{i}=1)>0.$ This ensures that the LATET in period $t=1$ is also well defined.
+Assumption 7 (Relevance condition).
+$E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]>0.$
+The theorem below shows that if Assumptions 1-7 hold, the Wald-DID estimand captures the LATET in period $1$.
 Theorem 1.
-If Assumptions 1-6 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
+If Assumptions 1-7 hold, the Wald-DID estimand $w_{DID}$ is equal to the LATET in period $t=1$; that is,
 $w_{DID}=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
@@
 Proof.
-See Appendix A. ∎
-We now provide the intuition behind $2\times 2$ DID-IV designs.
-As we already noted, the Wald-DID estimand scales the DID estimand of the outcome by the DID estimand of the treatment.
-Motivated by this observation, under $2\times 2$ DID-IV designs, we rely on the common DID identification strategy both in the numerator and the denominator of the Wald-DID estimand (see Figure 1).
-First, in the denominator of the Wald-DID estimand, we use the DID identification strategy to identify the direct effect of the instrument on the treatment.
-We recover the expectation of the unexposed treatment in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 4 (no anticipation in the first stage) and Assumption 5 (parallel trends assumption in the treatment).
-Next, in the numerator of the Wald-DID estimand, we use the DID identification strategy to identify the indirect effect of the instrument on the outcome through treatment.
-We recover the expectation of the unexposed outcome in period $t=1$ in an exposed group $(z=(0,1))$, using Assumption 6 (parallel trends assumption in the outcome).
-Finally, by scaling the indirect effect obtained from the numerator by the direct effect obtained from the denominator, we can capture the LATET in period $t=1$.
-Figure 1: DID-IV identification strategy in two periods and two groups.
-Notes: This figure plots the evolution of the expectation of the treatment (First-stage DID) and the outcome (Reduced-form DID) for exposed and unexposed groups, respectively.
-The dotted line represents the counterfactual trends of the expectation of the treatment and the outcome in an exposed group.
-The effects of the instrument on the treatment and the outcome are $0.2$ and $2$ respectively, and the local average treatment effect on the treated is $10$.
+See Appendix. ∎
 2.4   Interpreting the parallel trends assumption in the outcome
-Thus far, we have formalized the target parameter and the identifying assumptions in $2\times 2$ DID-IV designs.
-Overall, in this design, we rely on the common DID identification strategy in both the denominator and the numerator of the Wald-DID estimand to identify the effects of the instrument on the treatment and the outcome.
-The interpretation of the parallel trends assumption in the outcome, however, is less clear than that of the parallel trends assumption in the canonical DID designs because it does not depend on the untreated outcome, but depends on the unexposed outcome.
-In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in $2\times 2$ DID-IV designs.
-To do so, we first uncover the selection mechanism behind the canonical DID-IV designs.
-Specifically, we investigate how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups.
-Then, we show that the average time trend in the outcome is equal to a weighted average of the indirect effect of time on the outcome through treatment.
-We call this an expected time gain and show that the parallel trends assumption in the outcome requires the homogeneous expected time gain between exposed and unexposed groups.
+In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
+It is less clear than that in DID designs because this assumption does not depend on the untreated outcome, but depends on the unexposed outcome.
 First, we introduce the additional notation.
-Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$, which represents the treatment adoption process from period 00 to period 1 if the instrument path is $z=(0,0)$: the first element $D_{i,0}((0,0))$ represents the potential treatment choice without instrument in period 00 and the second element $D_{i,1}((0,0))$ represents the potential treatment choice without instrument in period $1$.
-Recall that in $2\times 2$ DID-IV designs, the units are allowed to take the treatment without instrument during the two periods.
+Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$ that represents the treatment adoption process during the two periods if the instrument path is $z=(0,0)$.
+Recall that in DID-IV designs, the units are allowed to take the treatment both period 00 and $1$ without instrument.
 In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods.
 Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
-The time never-takers $NT^{T}$ and the time always-takers $AT^{T}$ do not change their treatment status without instrument during the two periods.
-In other words, their treatment choices are time-invariant in the absence of the policy shock (instrument).
-However, the time compliers $CM^{T}$ are induced to treatment in period $1$ without instrument, and the time defiers $DF^{T}$ leave the treatment in period $1$ without instrument; that is, their treatment choices are affected by time in the absence of a policy shock.
-By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs.
-Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
-In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1).
-We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$.
-In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2).
-If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time.
-If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
 Table 1: Exposed group (assigned to $z=(0,1)$)
@@
 $D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
+Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 3 (monotonicity assumption).
+By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
+Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs.
+[^9] In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1).
+We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$.
+In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2).
+If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time.
+If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
+Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
 We now consider the interpretation of the parallel trends assumption in the outcome.
-Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in the exposed and unexposed groups, respectively.
+Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively.
 $\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
 $\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
-Note that the parallel trends assumption in the outcome requires $\Delta_{0}=\Delta_{1}$.
-At the same time, $\Delta_{0}$ is estimable by the observed change in outcomes in an unexposed group during the two periods, whereas $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
+Note that $\Delta_{0}$ is estimable by the observed change in the outcome in an unexposed group during the two periods.
+By contrast, $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
 To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation.
@@
 Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
-$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$
-$\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
-$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$
-$\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
-For instance, $w_{CM^{T},0}=Pr[CM^{T}|E_{i}=0]$ and $\Delta_{CM^{T},0}=E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=0]$ represent the population share and the expectation of the outcome path without instrument, respectively, for the time compliers in an unexposed group.
-Each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are estimable in an unexposed group ($e=0$).
-We can identify the time group of unit $i$ in an unexposed group and can observe the evolution of the outcome without instrument.
-For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$ for unit $i$ in an unexposed group, we can figure out that unit $i$ is the time compliers $CM^{T}$, and can observe the potential outcome path $(Y_{i,0}(0),Y_{i,1}(1))$ without instrument for that unit.
-However, each $w_{g^{t},e}$ and $\Delta_{g^{t},e}$ are unobservable in an exposed group ($e=1$) because we cannot specify the time group of unit $i$ in that group.
-For instance, when the observed treatment choices are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$ for unit $i$ in an exposed group, we cannot distinguish whether unit $i$ belongs to the time compliers $CM^{T}$ or the time never takers $NT^{T}$ (see Table 1).
-We can interpret each $\Delta_{g^{t},e}$ as the indirect effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$.
+$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$ $\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
+$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$ $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
+We can interpret each $\Delta_{g^{t},e}$ as the effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$.
 First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
-$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},$
-$\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
+$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},\hskip 5.69054pt\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
 This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods.
-If the exclusion restriction of time on the outcome holds, that is, $Y_{i,1}(1)=Y_{i,0}(1)$ and $Y_{i,1}(0)=Y_{i,0}(0)$ hold for all units $i$ (the stationary conditions hold), these quantities are equal to zero.
 Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
@@
 $\Delta_{DF^{T},e}$ $=E[Y_{i,1}(1)-Y_{i,0}(1)|DF^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(0)-Y_{i,1}(1)|DF^{T},E_{i}=e]}_{\text{Selection gain among $DF^{T}$}}.$
-The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the expectation of the time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively.
+The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the average time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively.
 The second term represents the selection gain in each time group in group $E_{i}=e$, respectively.
@@
 Proof.
-See Appendix A. ∎
-Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ is equal to a weighted average of the indirect effect of time on the outcome through treatment for each time group $g^{t}$ in group $e$.
+This directly follows from the Law of Iterated Expectation:
+$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
+$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
+$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
+$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
+$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
+Clearly, the weight $w_{g^{t},e}$ sums to one by construction. ∎
+Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ can be written as the weighted average of the expectation of the potential outcomes, and can be interpreted as measuring the convex average of the time effect for each time group $g^{t}$ in group $e$.
 The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$.
-In this paper, we call $\Delta_{e}$ an expected time gain in group $e$ hereafter.
-Based on the above discussions, we conclude that the parallel trends assumption in the outcome (Assumption 6), that is, $\Delta_{0}=\Delta_{1}$ requires the same expected time gain between exposed and unexposed groups; the whole indirect effects of time on the outcome through treatment are the same on average between the two groups.
-Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the indirect effect of the instrument on the outcome through treatment from that of time in an exposed group.
+Hereafter, we call $\Delta_{e}$ an expected time gain in group $e$.
+From Theorem 2, we can interpret that the parallel trends assumption in the outcome (Assumption 6), $\Delta_{0}=\Delta_{1}$, requires the same expected time gain between exposed and unexposed groups; that is, the effects of time on the outcome through treatment are the same between the two groups.
+Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the effect of the instrument on the outcome through treatment from that of time in an exposed group.
 The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
-In practice, if we have pre-exposed period data, we can indirectly check the validity of Assumption 6 by comparing the evolution of the outcome between exposed and unexposed groups.
-The formal discussion for the pre-trend test is provided in Subsection 5.3.
-2.5   Relationship between DID and DID-IV
-In this subsection, we describe the relationship between DID and DID-IV designs, and provide the view that DID-IV is DID with noncompliance of the assigned treatment.
-Following Subsection 2.1, we introduce the notation in DID designs.
-Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$.
-In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group).
-Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group.
-We refer to treatment and control groups as exposed and unexposed groups, respectively.
-Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units.
-First, we assume that all units are not allowed to receive the treatment in period 00.
-$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (1)
-Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
-$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (2)
-These restrictions imply that we have Tables 3-4 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
-Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions.
-First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs.
-Assumption 3 follows from (2) and Assumption 4 follows from (1).
-Assumption 6 follows from combining (1) with (2).
-Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 3 - 4:
-Table 3: Exposed group (assigned to $z=(0,1)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
-$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
-Table 4: Treatment group (assigned to $z=(0,0)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
-$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
-$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
-$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
-$=LATET.$
-Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$).
-Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
-Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups.
-From Tables 3- 4, we can rewrite the parallel trends assumption in the untreated outcome as:
-$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
-$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
-$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
-$\iff$ $\Delta_{0}=\Delta_{1},$
-where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
-$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
-Overall, this subsection clarifies that DID-IV is a natural generalization of DID, and the differences between the two designs arise because DID-IV allows the noncompliance of the assigned treatment: $D_{i,t}\neq Z_{i,t}$.
-In the next section, we compare DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
+Finally, we provide an empirical example to illustrate the interpretation of each notation and Assumption 6 in practice.
+Example.
+Suppose that the treatment $D_{i,t}\in\{0,1\}$ represents participation in a job training program, the outcome $Y_{i,t}$ represents real earnings, and the instrument $Z_{i,1}\in\{0,1\}$ represents a policy shock that encourages people to receive the treatment in the second period.
+[^10] Assume that the treatment is non-staggered and the treatment and the instrument are not randomized across units.
+In this setting, the instrumental variable strategy nor the canonical DID designs are directly applicable.
+In this setting, the time compliers $CM^{T}$ are the units who start adopting the program in the second period in the absence of the policy shock, and $w_{CM^{T},0}$ measures its share in an unexposed group.
+$\Delta_{CM^{T},0}$ represents the average time trends of real earnings among the time compliers in an unexposed group, and can be decomposed into the time effects on real earnings without policy shock, and the causal effects of program participation on real earnings in the second period, as their treatment adoption behavior is affected by time.
+Assumption 6 requires that the average time trends of real earnings are the same without policy shock between exposed and unexposed groups.
+This assumption can be interpreted as the restriction on the effect of time on real earnings through program participation between the two groups.
 2.5   Comparing DID-IV to Fuzzy DID
-de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture some causal effects under the same setting as in this article.
+de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting as in this article.
+[^11] The Wald-DID estimand considered in dCDH corresponds to ours in this article.
+In their empirical application, dCDH revisit Duflo (2001), in which the group variable $E_{i}$ ($G_{i}$ in dCDH) takes one if unit $i$ belongs to the region that are highly exposed to the policy shock.
 However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID.
-The detailed discussion is given in Appendix E, and this section provides the overview of that discussion.
-The main difference between dCDH and this article is in the definition of target parameter: we focus on the LATET, whereas dCDH focus on the SLATET (see Subsection 2.2).
-Accordingly, the identifying assumptions in Fuzzy DID designs are different from those in $2\times 2$ DID-IV designs.
+In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID in an overview of the detailed discussion provided in online Appendix Section E.
+The main difference between dCDH and this article lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.121212de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings.
+Therefore, we rewrite their target parameter by the notation used in this article.
+Definition.
+The switcher local average treatment effect on the treated (SLATET) is
+$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
+$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
+where the switcher (SW) is the units who become treated in period $1$.
+This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW).
+Because the target parameter is different from each other, the identifying assumptions in Fuzzy DID designs are also different from those in $2\times 2$ DID-IV designs.
 Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
-In Appendix E, we first show that the identifying assumptions in Fuzzy DID designs ex-ante impose the strong restrictions on the treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
-Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 5-6444Because dCDH consider repeated cross section settings, we can observe either $D_{0}(0)$ or $D_{1}(1)$ for each unit in Tables 5-6., which are in line with Tables 1-2.
-From these tables, one can see that dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
-Note that in our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption.
-The difference between dCDH’s restrictions and ours is critical for empirical application; in contrast to the monotonicity assumption, which excludes the defiers in both groups, dCDH’s additional restrictions can be difficult to defend from the institutional knowledge of the problem in a given application.
-This is because these restrictions ex ante exclude the units who change the treatment adoption in the absence of a policy shock (instrument).
+In online Appendix Section E, we first show that the identifying assumptions in Fuzzy DID designs ex ante impose strong restrictions on treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
+Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 3-4, which are in line with Tables 1-2.131313Because dCDH consider repeated cross section settings, we can observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.
+Comparing Tables 1-2 with Tables 3-4, one can figure out the difference between dCDH’s restrictions and ours.
+In our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption.
+By contrast, dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
 Table 3: Exposed group ($z=1$)
@@
 $D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-Next, under Tables 5-6, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
+Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by dCDH’s identifying assumptions.
+Next, under Tables 3-4, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
 The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers.
 The other causal parameter measures the treatment effects among the time compliers $CM^{T}$.
-The time compliers switch from non-treatment to treatment in period $1$ in the absence of the policy shock (instrument); that is, they are affected by time but not affected by the instrument.
-This decomposition result for the SLATET has two concerning implications.
-First, the interpretation of the SLATET is less clear than that of the LATET.
-Second, the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher because this parameter is always contaminated by the treatment effects among the time compliers.
-Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
-During this discussion, we also clarify why dCDH ex ante exclude so many types in exposed and unexposed groups for the Wald-DID estimand to identify the SLATET.
-First, we show that dCDH’s argument regarding the use of the Wald-DID estimand is based on their reliance on the parallel trends assumption in DID designs.
-Recall that in our DID-IV design, we did not make the parallel trends assumption in the untreated outcome, but made the parallel trends assumption in the unexposed outcome.
+[^14] Intuitively, we cannot identify this parameter in DID-IV settings because the untreated potential outcomes in period $t=1$ among the time compliers are not observed in both groups (see Tables 3-4).
+This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher, as this parameter is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by the instrument.
+Finally, we revisit dCDH’s argument regarding the use of the Wald-DID estimand, and show that this problem is less serious than one might expect, and does not arise once a researcher imposes Assumption 6 instead of the parallel trends assumption in the untreated outcome.
+First, we show that their argument is based on the parallel trends assumption in DID designs.
+Recall that in our DID-IV design, we do not make the parallel trends assumption in the untreated outcome, but make the parallel trends assumption in the unexposed outcome.
 Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group.
 We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
-In the proof, we also demonstrate that imposing the parallel trends assumption in DID designs yields the strong restrictions on the treatment adoption behavior across units under Fuzzy DID designs.
-Next, after confirming that dCDH’s argument is based on the parallel trends assumption in DID designs, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
-First, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group.
+Next, we argue that the different type of the parallel trends assumption, Assumption 6, is suitable for DID-IV settings, and the parallel trends assumption in the untreated outcome would be difficult to justify in practice for three reasons.
+First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group.
 Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al.
 (2005)).
-For instance, in Black et al.
-(2005), this assumption requires the data to contain parents with zero education attainment, which is unrealistic in practice.
+For instance, in Duflo (2001), this assumption requires the data to contain units with zero education attainment, which is unrealistic in practice.
 Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.
-We show that dCDH’s placebo test is incomplete for assessing the validity of this assumption.
-Surprisingly, we also show that their placebo test coincides with our pretrends test with a slight modification, introduced in section 5.3, to assess the plausibility of Assumption 5 and Assumption 6.
 3   DID-IV in multiple time periods
-We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
-In reality, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment in more than two periods (e.g.
-Black et al.
-(2005), Bhuller et al.
-(2013), Lundborg et al.
-(2014), and Meghir et al.
+We now extend the $2\times 2$ DID-IV design to multiple period settings with the staggered adoption of the instrument across units (Black et al.
+(2005); Bhuller et al.
+(2013) Lundborg et al.
+(2014); and Meghir et al.
 (2018)).
-We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
+We call it a staggered DID-IV design, and establish the target parameter and identifying assumptions.
 3.1   Set up
@@
 Let $D_{i}=(D_{i,1},\dots,D_{i,T})$ and $Z_{i}=(Z_{i,1},\dots,Z_{i,T})$ denote the path of the treatment and the instrument for unit $i$, respectively.
-Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are independent and identically distributed (i.i.d).
+Throughout Section 3 to Section 5, we assume that $\{Y_{i,t},D_{i,t},Z_{i,t}\}_{t=1}^{T}$ are i.i.d.
 We make the following assumption about the assignment process of the instrument.
@@
 Assumption 8 requires that once units start getting exposed to the instrument, units remain exposed to that instrument.
-In the DID literature, several studies make a similar assumption for the treatment and call it the ”staggered treatment adoption” (e.g.
+In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g.
 Athey and Imbens (2022); Callaway and Sant’Anna (2021); and Sun and Abraham (2021)).
@@
 If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$.
-Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$: all the units in cohort $e$ start being exposed to the instrument in period $E_{i}=e$.
+Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$.
 Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time.
@@
 Let $D_{i,t}^{\infty}$ denote the potential treatment status in period $t$ if unit $i$ is never exposed to the instrument.
-We call $D_{i,t}^{\infty}$ the ”never exposed treatment”.
+We call $D_{i,t}^{\infty}$ the “never exposed treatment”.
 Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
 $D_{i,t}=D_{i,t}^{\infty}+\sum_{1\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
-We define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$, which is the difference between the observed treatment status $D_{i,t}$ to the never exposed treatment status $D_{i,t}^{\infty}$.
+We define the effect of an instrument on treatment for unit $i$ in period $t$ as the difference between the observed treatment status to the never exposed treatment status: $D_{i,t}-D_{i,t}^{\infty}$.
 We refer to $D_{i,t}-D_{i,t}^{\infty}$ as the individual exposed effect in the first stage.
-In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
+[^15] In the DID literature, Callaway and Sant’Anna (2021) and Sun and Abraham (2021) define the effect of a treatment on an outcome in the same fashion.
 Next, we introduce the group variable that describes the type of unit $i$ in period $t$, based on the reaction of potential treatment choices in period $t$ to the instrument path $z$.
 Let $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ ($t\geq e$) be the group variable in period $t$ for unit $i$ and the initial exposure date $e$.
-The first element $D_{i,t}^{\infty}$ represents the treatment status in period $t$ if unit $i$ is never exposed to the instrument $E_{i}=\infty$ and the second element $D_{i,t}^{e}$ represents the treatment status in period $t$ if unit $i$ is first exposed to the instrument at $E_{i}=e$.
 Following the terminology in section 2.1, we define $G_{i,e,t}=(0,0)\equiv NT_{e,t}$ as the never-takers, $G_{i,e,t}=(1,1)\equiv AT_{e,t}$ as the always-takers, $G_{i,e,t}=(0,1)\equiv CM_{e,t}$ as the compliers and $G_{i,e,t}=(1,0)\equiv DF_{e,t}$ as the defiers in period $t$ and the initial exposure date $e$.
 Finally, we make a no carryover assumption on potential outcomes $Y_{i,t}(d,z)$.
-Assumption 8 (No carryover assumption).
+Assumption 9 (No carryover assumption in multiple time periods).
 $\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d_{t},z).$
-Henceforth, we keep Assumption 7 and 8. In the next section, we define the target parameter in staggered DID-IV designs.
+Henceforth, we keep Assumption 8 and Assumption 9. In the next section, we define the target parameter in staggered DID-IV designs.
 3.2   Target parameter in staggered DID-IV designs
@@
 Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$.
-Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$.
+[^16] Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$.
 For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
 This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher.
-For example, suppose that the treatment represents participation in a job training program, the outcome represents real earnings, and the instrument represents a policy shock that encourages people to receive the treatment.
-In this setting, if the instrument is an important policy lever for the researchers (e.g., nudge), each CLATT is of particular interest; the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
+For example, in a job training program setting considered in Subsection 2.5, each CLATT would be of particular interest if the instrument $Z_{i,t}$ represents an important policy lever for the researchers (e.g., nudge); the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
 3.3   Identification assumptions in staggered DID-IV designs
@@
 The potential outcome $Y_{i,t}(D_{i,t}^{e})$ represents the outcome status in period $t$ if unit $i$ is first exposed to the instrument in period $e$ and the potential outcome $Y_{i,t}(D_{i,t}^{\infty})$ represents the outcome status in period $t$ if unit $i$ is never exposed.
-We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the ”never exposed outcome”.
+We refer to $Y_{i,t}(D_{i,t}^{\infty})$ as the “never exposed outcome”.
 Assumption 11 (Monotonicity assumption in multiple time periods).
@@
 It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument.
-Assumption 12 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated outcomes in settings of multiple periods and cohorts.
+Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
 Assumption 14 (Parallel trends assumption in the outcome in multiple time periods).
@@
 It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned.
-From the discussion in Subsection 2.4, this assumption requires the same expected time gain across cohorts and over time: the effects of time on the outcome through treatment are the same on average across cohorts and over time.
+From the discussion in Subsection 2.4, this assumption can be interpreted as requiring the same expected time gain across cohorts and over time; that is, the effects of time on the outcome through treatment are the same on average across cohorts and over time.
+Assumption 15 (Relevance condition in multiple time periods).
+$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]>0.$
+Assumption 15 is a relevance condition in multiple period settings.
+With a binary treatment, Assumption 11 and 15 implies that the fraction of the compliers should be non-negative for all $e$ and $l\geq 0$:
+$E[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]=Pr(CM_{e,l}|E_{i}=e)>0.$
+This condition ensures that each $CLATT_{e,l}$ is well defined.
 4   Extensions
-This section contains extensions to non-binary, ordered treatments, triple DID-IV designs, and repeated cross sections.
-It includes the case in which the treatment path was introduced in potential outcomes in $2\times 2$ DID-IV designs.
-For more details and proofs, see Appendix B.
+This section contains extensions to non-binary, ordered treatments and repeated cross sections.
+It also includes the case that the treatment path is introduced in potential outcomes in $2\times 2$ DID-IV designs.
+For more details and proofs, see online Appendix Section A.
 Non-binary, ordered treatment
 Up to now, we have considered only the case of a binary treatment.
-However the same idea can be applied when treatment takes a finite number of ordered values; that is, we have $D_{i,t}\in\{0,1,\dots,J\}$.
-When we have only two periods and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
+However the same idea can be applied when treatment takes a finite number of ordered values: $D_{i,t}\in\{0,1,\dots,J\}$.
+When only two periods exist and treatment is non-binary, our target parameter is the average causal response on the treated (ACRT) defined below.
 Definition.
 The average causal response on the treated (ACRT) is
-$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
+$ACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E[Y_{i,1}(j)-Y_{i,1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1].$
 where the weight $w_{j}$ is:
-$w_{j}=\frac{Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E_{i}=1)}.$
-The ACRT is a weighted average of the effects of an increase of one unit in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument.
-The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in this parameter are conditional on $E_{i}=1$.
-Theorem 3 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under the same assumptions in Theorem 1.
-Theorem 3.
-If Assumptions 1-6 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
-$W_{DID}$ $=\sum_{j=1}^{J}w_{j}\cdot E[Y_{1}(j)-Y_{1}(j-1)|D_{i,1}((0,1))\geq j>D_{i,1}((0,0)),E_{i}=1]$
-$=ACRT.$
-Proof.
-See Appendix B.1. ∎
+$w_{j}=\frac{Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}{\sum_{j=1}^{J}Pr(D_{i,1}((0,1))\geq j>D_{i,1}((0,0))|E_{i}=1)}.$
+The ACRT is a weighted average of the effect of one unit increase in the treatment on the outcome, for those who belong to an exposed group and are induced to increase the treatment in period $t=1$ by instrument.
+The ACRT is similar to the average causal response (ACR) considered in Angrist and Imbens (1995), with the difference that each weight $w_{j}$ and the associated causal parameter in the ACRT are conditional on $E_{i}=1$.
+Online Appendix Section A.1 shows that if we have a non-binary, ordered treatment, the Wald-DID estimand captures the ACRT under the same assumptions in Theorem 1.
 With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT.
@@
 $w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|E_{i}=e)}.$
-Triple DID-IV designs
-[…]
 Repeated cross sections
-In some applications, researchers have no access to panel data and use repeated cross section data [^6] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for the treatment as in Duflo (2001)..
-In Appendix B.2, we consider repeated cross section data settings and present the identification assumptions in DID-IV designs.
+In some applications, researchers have only access to repeated cross section data.
+[^17] It also includes the case that researchers use the cross section data, and exploit a policy shock across cohorts as an instrument for treatment as in Duflo (2001).
+Online Appendix Section A.2 presents the identification assumptions in DID-IV designs under repeated cross section settings.
 Introducing the treatment path in $2\times 2$ DID-IV designs
-In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$ as we consider the non-staggered adoption of the treatment.
-We can extend our results to the case where we introduce the treatment path in potential outcomes with binary treatment.
-In this case, our target parameter changes slightly with some complex notation.
-Definition.
-The local average treatment effect on the treated with the treatment path (LATET’) is
-$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
-where $\Delta_{0}$ and $\Delta_{1}$ are:
-$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
-$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
-The weights $w_{0}$ and $w_{1}$ are:
-$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
-Appendix B.3 shows that the similar assumptions to Assumptions 1-6 in Subsection 2.3 are sufficient for the Wald-DID to capture the LATET’.
+In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$, as we consider the non-staggered adoption of the treatment.
+We can extend our results to the case that we introduce the treatment path in potential outcomes with binary treatment.
+Online Appendix Section A.3 shows that the similar assumptions to Assumptions 1-6 are sufficient for the Wald-DID to capture the LATET-like parameter.
 5   Estimation and inference
-When researchers leverage variation in policy adoption timing across units as an instrument for treatment and implicitly rely on a staggered DID-IV design, they usually implement this design via TWFEIV regressions (e.g., Johnson and Jackson (2019), Lundborg et al.
+In this section, we propose a credible estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity.
+First, we propose a simple regression-based method for estimating each CLATT.
+Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT.
+We also discuss the pre-trends tests for checking the plausibility of parallel trends assumptions in DID-IV designs.
+Remark 1.
+In practice, when researchers implicitly rely on a staggered DID-IV design, they commonly implement this design via two-way fixed instrumental variable (TWFEIV) regressions (e.g., Johnson and Jackson (2019), Lundborg et al.
 (2014), Black et al.
@@
 (2015), and Bhuller et al.
-(2013)).
-The companion paper (Miyaji (2024)), however, shows that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects under staggered DID-IV designs in the presence of heterogeneous treatment effects.
-Specifically, it shows that under staggered DID-IV designs, the TWFEIV estimand is a weighted average of all possible CLATTs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
-[^7] See also the discussions in de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator in the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally.
-In this section, we propose an alternative estimation method in staggered DID-IV designs that is robust to treatment effect heterogeneity.
-First, we propose a simple regression-based method for estimating each CLATT.
-Following Callaway and Sant’Anna (2021), we then propose a weighting scheme to construct the summary causal parameters from each CLATT.
-At the end of this section, we also discuss the pre-trend tests for checking the plausibility of parallel trends assumption in the treatment and the outcome in DID-IV designs.
+(2013)):
+$Y_{i,t}=\phi_{i.}+\lambda_{t.}+\beta_{IV}D_{i,t}+v_{i,t},$
+$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\pi Z_{i,t}+\eta_{i,t}.$
+In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand potentially fails to summarize the treatment effects in the presence of heterogeneous treatment effects.
+Specifically, we show that the TWFEIV estimand is a weighted average of all possible CLATTs under staggered DID-IV designs, but some weights can be negative if the effect of the instrument on the treatment or the outcome is not stable over time.
 5.1   Stacked two stage least squares regression
-We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs.
-If we have a binary treatment, the target parameter is the $CLATT_{e,e+l}$ for each cohort $e$ and a relative period $l\geq 0$:
-$CLATT_{e,e+l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}].$
-Our estimation method consists of two steps.
+We propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs. Our estimation method consists of two steps.
 Step 1.
@@
 If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
-Note that the units in an already exposed cohort $e=1$ are not included in any data sets because the pre-exposed period $e-1$ should be included for estimating each $CLATT_{e,e+l}$.
 Step 2.
 In each data set, we run the following IV regression and obtain the IV estimator $\hat{\beta}^{e,l}_{IV}$:
-$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
-The first stage regression is:
-$D_{i,t}=\pi^{e,l}_{0}+\pi^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\pi^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\pi^{e,l}(\mathbf{1}\{E_{i}=e\}\cdot\mathbf{1}\{T_{i}=e+l\})+\eta^{e,l}_{i,t},$
+$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t},$
 where the group indicator $\mathbf{1}\{E_{i}=e\}$ and the post-period indicator $\mathbf{1}\{T_{i}=e+l\}$ are the included instruments and the interaction of the two is the excluded instrument.
@@
 From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator.
-Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
+[^18] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021).
 The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group.
-Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, and our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
+Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
 Theorem 3 below guarantees the validity of our STS estimator.
 Theorem 3.
-Suppose Assumptions 7-13 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
+Suppose Assumptions 8-14 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
 […]
@@
 Proof.
-See Appendix C. ∎
+See Appendix. ∎
 From Theorem 3, we can also construct the standard error of the STS estimator, using the sample analogue of the asymptotic variance $V(\psi_{i,e,l})$.
 Remark 2.
-Our estimation procedure is identical if we have a non-binary, ordered treatment or use repeated cross section data.
-Appendix C presents the influence function of the SLS estimator in repeated cross section settings.
-In triple DID-IV designs, one can replace the IV regression in step $2$ with the following one ($g_{i}=A,B$):
-$Y_{i,t}$ $=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\beta^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\beta^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\beta^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\beta^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\beta^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
-$+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
-The first stage regression is:
-$D_{i,t}$ $=\pi^{e,l}_{0}+\pi^{e,l}_{1}\mathbf{1}_{E_{i}=e}+\pi^{e,l}_{2}\mathbf{1}_{T_{i}=e+l}+\pi^{e,l}_{3}\mathbf{1}_{g_{i}=A}+\pi^{e,l}_{4}\mathbf{1}_{E_{i}=e,T_{i}=e+l}+\pi^{e,l}_{5}\mathbf{1}_{E_{i}=e,g_{i}=A}+\pi^{e,l}_{6}\mathbf{1}_{T_{i}=e+l,g_{i}=A}$
-$+\pi^{e,l}_{7}\mathbf{1}_{E_{i}=e,T_{i}=e+l,g_{i}=A}+\eta^{e,l}_{i,t},$
-where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true. Appendix C presents the influence function of the triple DID-IV estimator $\beta_{IV}^{e,l}$.
+Our estimation procedure is the same if we have a non-binary, ordered treatment or use repeated cross section data.
+Online Appendix Section B presents the influence function of the STS estimator in repeated cross section settings.
 5.2   Weighting scheme
-The previous subsection provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs.
-In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$.
-However, researchers may also want to estimate summary causal parameters that capture the overall effect of a treatment on an outcome.
-In staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$.
-In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
-In this section, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
-Specifically, we consider the following weighting scheme as in Callaway and Sant’Anna (2021):
+In this subsection, we explain how one can construct the summary causal parameters from each CLATT in staggered DID-IV designs, based on the weighting scheme proposed by Callaway and Sant’Anna (2021).
+We consider the following weighting scheme as in Callaway and Sant’Anna (2021):
 $\theta^{IV}=\sum_{e}\sum_{t=1}^{T}w(e,t)\cdot CLATT_{e,t},$
@@
 $CAET_{e,e+l}^{1}=E[D_{i,e+l}-D_{i,e+l}^{\infty}|E_{i}=e].$
-If treatment is binary, each $CETT_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
+If treatment is binary, each $CAET_{e,l}^{1}$ is equal to the share of the compliers in cohort $e$ in period $e+l$:
 $CAET^{1}_{e,e+l}$ $=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]$
@@
 where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)).
-Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start being exposed to it.
 Table 5: Weights for a variety of summary causal parameters
@@
 Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters.
-Each target parameter corresponds with that of Callaway and Sant’Anna (2021).
-IV is added to associate each parameter with staggered DID-IV designs.
+Each target parameter without the superscript IV is defined in Callaway and Sant’Anna (2021).
+The superscript IV is added to associate each parameter with staggered DID-IV designs.
 In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
-$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t},$
-where the weights are non-negative and sum to one by construction.
+$\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t}.$
 This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$.
 This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
-By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021).
-In Table 7, we summarize the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter.
-Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs.
+By similar arguments, we can specify the weighting functions for various summary measures, which correspond with those in Callaway and Sant’Anna (2021).
+Table 5 summarizes the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter.
 Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
@@
 $\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
-where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the stacked IV regression in section 5.1.
+where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1.
 Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
 $l_{i}^{\theta_{IV}}=\sum_{e}\sum_{t=1}^{T}\left(w(e,t)\cdot\psi_{i,e,t}+\zeta_{i,e,t}^{w}\cdot{CLATT}_{e,t}\right).$
-5.3   Pre-trend test
-We now describe how one can conduct pre-trend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
-Suppose that data are available for period $t=-1$; one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
+5.3   Pre-trends test
+In most DID applications, researchers assess the plausibility of the parallel trends assumption by testing for pre-treatment differences in the outcome between treatment and control groups (pre-trends test).
+In this section, we describe the procedure of the pre-trends test in DID-IV designs to check the validity of the parallel trends assumption in the treatment and the outcome.
+Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
 $E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
@@
 $E[Y_{i,0}-Y_{i,-1}|E_{i}=0]=E[Y_{i,0}-Y_{i,-1}|E_{i}=1]$
-$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]=E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (4)
-These tests can be generalized to multiple pre-exposed periods settings by using the pre-trend tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al.
-(2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
-Remark 3.
-When researchers implement the staggered DID-IV design via TWFEIV regressions, they often use the dynamic specifications in the first stage and the reduced form:
-$Y_{i,t}=\mu_{i.}+\delta_{t.}+\sum_{l}\alpha_{l}\mathbf{1}\{t-E_{i}=l\}+\epsilon_{i,t},$ (5)
-$D_{i,t}=\gamma_{i.}+\zeta_{t.}+\sum_{l}\beta_{l}\mathbf{1}\{t-E_{i}=l\}+\eta_{i,t},$ (6)
-where $E_{i}$ represents the initial exposure date.
-They plot the coefficients $\alpha_{l}$ in equation (5) and $\beta_{l}$ in equation (6) in each relative period $l$, and test whether the coefficients before the initial exposure to the instrument are significantly different from zero or not (e.g., Akerman et al.
-(2015) and Bhuller et al.
-(2013)).
-However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous.
-Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and parallel trends assumption in the untreated outcome are plausible.
-Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
+$\iff$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=0]$
+$=$ $E[Y_{i,0}(D_{i,0}((0,0)))-Y_{i,-1}(D_{i,-1}((0,0)))|E_{i}=1].$ (2)
+These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al.
+(2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumption 13 and 14 in pre-exposed periods.
 6   Application
+We illustrate the empirical relevance of our findings in the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom.
+We first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006).
+We then estimate the TWFEIV regression in the author’s setting.
+Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
+6.1   Setting
 Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15.
-Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland.
+Specifically, Oreopoulos (2006) exploits variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland.
+[^19] In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately.
+Due to the imprecision of their standard errors, Oreopoulos (2006) then moves to “a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K.
+data”.
 The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K.
 General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
-In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately.
-Due to the imprecision of their standard errors [^8] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K.
-data”.
-Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
+Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
 $Y_{i,t}=\mu_{n.}+\delta_{.t}+\beta_{IV}D_{i,t}+\epsilon_{i,t},$ (3)
@@
 The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit.
-This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the IV design.
+This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of exploiting the random variation of the policy change.
 Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
-In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, potentially varying across cohorts.
-That the author was conscious of the heterogeneous treatment effects is clear from the title of ”Estimating Average and Local Average Treatment Effects of Education when Compulsory Schooling Laws Really Matter”.
-The TWFEIV estimator, therefore, can potentially fail to aggregate the treatment effects, yielding a misleading conclusion under staggered DID-IV designs.
-In this section, we first assess the identifying assumptions in staggered DID-IV designs implicitly imposed by Oreopoulos (2006).
-We then estimate the TWFEIV regression in the author’s setting.
-Finally, we estimate the target parameter and its summary measure by employing our proposed method and weighting scheme.
 6.2   Assessing the identifying assumptions in staggered DID-IV designs
-We discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006).
+We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006).
 In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment.
-We can identify each CACRT and its summary measure if the underlying DID-IV identifying assumptions seem plausible.
 Exclusion restriction.
@@
 Panel (b) in Figure 1 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval.
-The pre-exposed estimates seems consistent with Assumption 13: though an upward pre-trend exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
-Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post-exposed periods.
+The pre-exposed estimates seems consistent with Assumption 14: though an upward pre-trends exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
+Figure 1 also sheds light on the dynamic effects of the school reform on education attainment and log annual earnings in post-exposed periods.
 In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant.
 In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
-Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
-In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
+[^20] Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage.
 6.3   Illustrating our estimation method
@@
 The result is shown in Table 6.
-The TWFEIV estimate is $-0.009$, not significantly different from 00, which indicates that, on the whole, the returns to schooling in the UK is nearly zero.
+The TWFEIV estimate is $-0.009$ and not significantly different from 00.
+This indicates that, on the whole, the returns to schooling in the UK is nearly zero.
 However, this may be a misleading conclusion.
 We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
-The TWFEIV regression performs the ”bad comparisons” (c.f.
-Goodman-Bacon (2021)) in the author’s setting.
-When we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods.
-In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006).
-Specifically, we show that all the bias terms are positive, whereas all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
+In online Appendix Section C, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006).
+Specifically, we show that all the bias terms are positive, while all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
+[^21] When we run the TWFEIV regression in the author’s setting, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods.
+This procedure performed by the TWFEIV regression is called the “bad comparisons” in the recent DID literature (c.f.
+Goodman-Bacon (2021)), and yields the bias in the TWFEIV estimand from the properly weighted average of all possible CACRTs.
 Figure 2: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006).
@@
 We then run TSLS regression in each data set.
-The standard error is calculated by using the influence function shown in equation (C.2) in Appendix C.
+The standard error is calculated by using the influence function shown in equation (B) in online Appendix Section LABEL:ApeC.1.
 Table 6: Returns to schooling in Oreopoulos (2006)
@@
 7   Conclusion
-Instrumental variable strategies are pervasive in economics to estimate the effect of a treatment on an outcome.
-In this design, however, we commonly face two challenges: the internal and external validity of the IV estimate.
-To overcome at least external validity, many studies have leveraged variation occurring from the different timing of policy adoption across units as an instrument for treatment.
-The instrument based on these policy changes, however, generally fails to satisfy the randomization assumption in IV designs in practice, casting doubt on the internal validity of the IV estimate.
-Rather, the underlying identification strategy is parallel to DID designs, under which we identify the effect of the policy shock on the outcome, leveraging the variation of the exposure to that policy shock across units and over time.
-In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV) by combining the IV techniques with DID designs.
-We start with a simple two periods and two groups setting.
-In this setting, our DID-IV design mainly consists of a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups.
-We show that the Wald-DID estimand captures the LATET under $2\times 2$ DID-IV designs.
-We also clarify the interpretation of the parallel trends assumption in the outcome.
-Specifically, we show that in DID-IV settings, time also plays the role of instrument, and the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
-Finally, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
-Next, we consider the DID-IV design in more than two time periods with units being exposed to the instrument at different times.
+In this paper, we formalize an instrumented difference-in-differences (DID-IV).
+First, we consider a simple setting with two periods and two groups.
+In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups.
+We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET).
+After establishing $2\times 2$ DID-IV designs, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
+Next, we consider the DID-IV design in more than two periods with units being exposed to the instrument at different times.
 We call this a staggered DID-IV design, and formalize the target parameter and identifying assumptions.
-Our target parameter in this design is the cohort specific local average treatment effects on the treated (CLATT).
+Specifically, in this design, our target parameter is the cohort specific local average treatment effects on the treated (CLATT).
 The identifying assumptions are the natural generalization of those in $2\times 2$ DID-IV designs.
-We also provide the estimation method in staggered DID-IV designs, which does not require strong restrictions on treatment effect heterogeneity.
-In practice, most DID-IV applications implement the staggered DID-IV design by running TWFEIV regressions.
-In companion paper (Miyaji (2024)), however, we show that the conventional approach is inadequate for summarizing the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time.
+We also provide the estimation method in staggered DID-IV designs that does not require strong restrictions on treatment effect heterogeneity.
 Our estimation method carefully chooses the comparison groups and does not suffer from the bias arising from the time-varying exposed effects.
-We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trend tests to assess the validity of the parallel trends assumption in the treatment and the outcome.
-Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006) who estimate returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland.
-First, we assess the underlying DID-IV identification strategy in the author’s setting and check whether the identifying assumptions are plausible.
-Next, we estimate a TWFEIV regression and our stacked two stage least squares regressions in each dataset.
-The TWFEV estimate turns out to be strictly negative, whereas our STS estimates are all positive.
-Finally, we estimate the summary causal parameter by exploiting our STS estimates.
-The estimate for the summary measure is also positive and indicates that returns to schooling are substantial in the UK.
-Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can use variation in the timing of policy adoption across units as an instrument for the treatment.
-To avoid the issue of using the TWFEIV estimator under this design, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity.
-Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV research design in practice.
-Appendix A Proofs in Section 2
+We also propose the weighting scheme in staggered DID-IV designs, and explain how one can conduct pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
+Finally, we illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting the timing variation of the introduction of school reforms between British and North Ireland.
+In this application, the TWFEV regression, the conventional approach to implement a staggered DID-IV design, yields a negative estimate.
+By contrast, our STS regression and weighting scheme indicate the substantial gain from schooling.
+This empirical application illustrates that the estimation method matters in staggered DID-IV designs in practice.
+Overall, this paper provides a new econometric framework for estimating the causal effects when the treatment adoption is potentially endogenous over time, but researchers can exploit variation in policy adoption timing as an instrument for treatment.
+To avoid the issue of using the TWFEIV estimator, we also provide a reliable estimation method that is free from strong restrictions on treatment effect heterogeneity.
+Further developing alternative estimation methods and diagnostic tools will be a promising area for future research, facilitating the credibility of DID-IV design in practice.
+Appendix
 Proof of Theorem 1.
 Proof.
-First we consider the denominator of the Wald-DID estimand. For the denominator, we have
-$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
-$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
-$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$ (9)
-+$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
-$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
-+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
-$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1)$
-$=$ $Pr(CM^{Z}|E=1).$ (10)
-where the second equality follows from the simple manipulation.
-The third equality follows from Assumption 4 (No anticipation in the first stage) and the forth equality follows from Assumption 3 (Monotonicity assumption) and Assumption 5 (Parallel trends assumption in the treatment).
-Next, to show that the Wald-DID estimand equals the LATET, it suffices to show that
-$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$ $=E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1)$
-$=LATET\cdot Pr(CM^{Z}|E=1).$
-For the numerator, we have
+First we consider the numerator of the Wald-DID estimand. For the numerator, we have
 $E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
@@
 $=$ $E[(D_{1}((0,1))-D_{1}((0,0)))\cdot(Y_{1}(1)-Y_{1}(0))|E=1]$
-$=$ $E[Y_{1}(1)-Y_{1}(0)|E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1).$
 $=$ $LATET\cdot Pr(CM^{Z}|E=1).$ (5)
-The first equality follows from Assumption 1 (No carryover assumption) and Assumption 2 (Exclusion restriction).
+The first equality follows from Assumption 1 and Assumption 2.
 The second equality follows from the simple manipulation.
 The third equality follows from Assumption 4.
-The fourth equality follows from Assumption 6 (Parallel Trends Assumption in the outcome) and simple calculation.
+The fourth equality follows from Assumption 6 and simple calculation.
 The fifth equality follows from the Law of Iterated Expectations and Assumption 3.
-Combining the result (11) with (10), we obtain
-$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=1]}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1)}{Pr(CM^{Z}|E=1)}$
-$=LATET.$
-This completes the proof. ∎
-Proof of Theorem 2.
-Recall that we define $\Delta_{e}$ as follows:
-$\Delta_{e}\equiv E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=e].$
-Then, we obtain:
-$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
-$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
-$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
-$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
-$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
-The first equality follows from the Law of Iterated Expectation and the second equality follows from the definition of $w_{g^{t},e}$ and $\Delta_{g^{t},e}$.
-Clearly, the weight $w_{g^{t},e}$ sums to one by construction.
-This completes the proof.
-Appendix B Extensions
-B.1   Non-binary, ordered treatment
+For the denominator, we have
+$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
+$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
+$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
++$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
+$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
++$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
+$=$ $Pr(CM^{Z}|E=1).$ (6)
+where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
+Combining the result (5) with (6), we obtain the desirable result. ∎
 Proof of Theorem 2.
 Proof.
+First, we prove the consistency of our SLS estimator.
+We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined.
+More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$.
+This assumption guarantees that each Wald-DID estimator $\widehat{CLATT}_{e,e+l}$ is also well defined.
+First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and continuous mapping theorem, we have
+[…]
+Next, we show that the right hand side of (7) is equal to $CLATT_{e,e+l}$. For the numerator, we have
+$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
+$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
+$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$ (8)
+$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
+$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
+$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
+$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
+$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
+$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (9)
+The third equality follows from Assumption 9, 10 and 12. The fourth equality follows from Assumption 14 and the fifth follows from Assumption 11. For the denominator, by similar calculations,
+$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
+$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]$ (10)
+$-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
+$=Pr(CM_{e,e+l}|E_{i}=e),$ (11)
+where the final equality follows from Assumption 11 and Assumption 13. Combining (9) with (11), we obtain the desirable result.
+Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. We define $\theta_{e,l}$ to be
+$\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
+$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
+The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation of the influence function.
+Fact 1.
+If
+$\sqrt{n}(\hat{A}-A)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}+o_{p}(1),\hskip 8.53581pt\sqrt{n}(\hat{B}-B)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}b_{i}+o_{p}(1),$
+we have
+$\sqrt{n}\left(\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{a_{i}-(A/B)b_{i}}{B}+o_{p}(1).$
+Then, by some calculations, we obtain
+$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1).$
+The influence function $\psi_{i,e,l}$ is:
+$\psi_{i,e,l}$
+$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (12)
+where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
+Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
+$V(\psi_{i,e,l})$ $=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}$
+$+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
+Appendix A Extensions
+This section considers the non-binary, ordered treatment and repeated cross section settings in DID-IV designs.
+It also includes the case where we introduce the treatment path in $2\times 2$ DID-IV designs.
+A.1   Non-binary, ordered treatment
+In this subsection, we present the identification results in DID-IV designs with non-binary, ordered treatment.
+Theorem 4 below shows that if we have a non-binary, ordered treatment, the Wald-DID estimand is equal to the ACRT under Assumptions 1-7, which replace the binary treatment with a non-binary one.
+Theorem 4.
+If Assumptions 1-7 hold, the Wald-DID estimand $W_{DID}$ is equal to the ACRT.
+Proof.
 Let $\lambda_{Z,j}^{t}=\mathbf{1}(D_{t}((0,z))\geq j)$ be the indicator function for $t\in\{0,1\},z\in\{0,1\}$ and $j\in\{0,\dots,J+1\}$.
-We note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction.
+Note that $\lambda_{Z,0}^{t}=1$ and $\lambda_{Z,J+1}^{t}=0$ hold for all $t$ and $z$ by construction.
 Here, we can rewrite the observed outcome $Y_{t}$ as follows:
@@
 where we use Assumption 1 and 2.
-For the numerator in the Wald DID estimand, we have
+For the numerator of the Wald DID estimand, we have
 $E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
-$=E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
-+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
 $=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1].$
-The first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
+This follows from Assumption 4 and Assumption 6.
 In terms of the $\lambda_{Z,j}^{t}$, we write the previous expression as:
@@
 ×$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (13)
-For the denominator in the Wald-DID estimand, by similar argument, we obtain:
+For the denominator of the Wald-DID estimand, by similar argument, we obtain:
 $E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
-$=E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
-+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
 $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
-The first equality follows from simple algebra and Assumption 4. The second equality follows from Assumption 5.
+This follows from Assumption 4 and 5.
 Then, we have:
@@
 Combining the result (14) with (13), we obtain the desired result. ∎
-B.2   Repeated cross sections
+A.2   Repeated cross sections
 Here, we present our identification results under repeated cross section settings.
-B.2.1 Two time periods
+First, we consider two groups and two periods settings.
+Next, we consider multiple period settings with staggered instrument.
+A.2.1 Two time periods
 Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$.
@@
 Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$.
-We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
-In a two-period and two-group setting, our target parameter is the local average treatment effect on the treated in period $1$:
+We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
+In two periods and two groups settings, our target parameter is the local average treatment effect on the treated in period $1$:
 $LATET$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=1,E_{i}=1,D_{i,1}((0,1))>D_{i,1}((0,0))]$
@@
 $=$ $E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=1]-E[Y_{i}(D_{i}(0,0))|E_{i}=0,T_{i}=0].$
-Theorem 5 below shows that the Wald-DID estimand is equal to the LATET in period $1$ under Assumptions 14-18.
-Here, we assume the binary treatment, but the non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus, we omit it for brevity).
-In the proof, we omit the index $i$ to ease the notation.
-Theorem 5.
-Suppose Assumptions 14-18 hold. If we assume a binary treatment, the Wald-DID estimand corresponds to the local average treatment effect on the treated (LATET) in period $1$:
-$w_{DID}$ $\equiv\frac{E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}}{E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}}$
-$=E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]$
-$=LATET.$
-Proof.
-First we consider the denominator of the Wald-DID estimand. For the denominator, we have
-$E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}$
-$=$ $E[D((0,1))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
-$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}$
-$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
-+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
-$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
-$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
-+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,0))|E=1,T=0]$
-$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
-$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1,T=1)$
-$=$ $Pr(CM^{Z}|E=1,T=1).$ (14)
-The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the fourth equality follows from Assumption 15 and Assumption 17.
-Next, for the numerator, we have
-$E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}$
-$=$ $E[Y(D((0,1)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
-$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\}$
-$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
-+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
-$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
-$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
-+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,0)))|E=1,T=0]$
-$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
-$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
-$=$ $E[(D((0,1))-D((0,0)))\cdot(Y(1)-Y(0))|E=1,T=1]$
-$=$ $E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1,T=1).$
-$=$ $LATET\cdot Pr(CM^{Z}|E=1,T=1).$ (15)
-The second equality follows from the simple manipulation.
-The third equality follows from Assumption 16.
-The fourth equality follows from Assumption 18 and the simple calculation.
-The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
-Combining the result (14) with (15), we obtain
-$w_{DID}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1,T=1)}{Pr(CM^{Z}|E=1,T=1)}$
-$=LATET.$
-B.2.2 Multiple time periods
+Assumption 19 (Relevance condition).
+$E[D_{i}|E_{i}=1,T_{i}=1]-E[D_{i}|E_{i}=1,T_{i}=0]$
+-- $(E[D_{i}|E_{i}=0,T_{i}=1]-E[D_{i}|E_{i}=0,T_{i}=0])>0.$
+By similar arguments in the proof of Theorem 1 (Theorem 4), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period $1$ under Assumptions 14-19 (thus, we omit the proof for brevity).
+A.2.2 Multiple time periods
 Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$.
@@
 Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$.
-We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
-In multiple time period and multiple cohort settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
+We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
+In multiple periods and cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
 $CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
 $=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
-When repeated cross section data are available and we have multiple time periods, the staggered DID-IV designs consist of Assumptions 7, 10, 11 (in Section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
+When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 14, and the following parallel trends assumptions and relevance condition.
 Assumption 20 (Parallel trends assumption in the treatment in multiple time periods).
-[…]
+$\text{For all}\hskip 2.84526pts<t,\hskip 2.84526ptE[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=s]\hskip 2.84526pt\text{is identical for all}\hskip 2.84526pte\in\mathcal{S}(E_{i}).$
 Assumption 21 (Parallel trends assumption in the outcome in multiple time periods).
-[…]
-B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
-We discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$.
-We consider two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption).
-We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting.
-Finally, we show that a Wald-DID estimand identifies the LATET’ (see Section 4) under suitable assumptions similar to the ones in Subsection 2.3.
-We consider the same setting as in Subsection 2.1 except that we do not adopt Assumption 1.
-First, we make the following no anticipation assumptions on the outcome.
+For $\text{For all}\hskip 5.69054pts<t,\hskip 5.69054ptE[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=s]$
+is identical for $\text{is identical for all}\hskip 5.69054pte\in\mathcal{S}(E_{i}).$
+Assumption 22 (Relevance condition in multiple time periods).
+$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i}-D^{\infty}_{i}|E_{i}=e,T_{i}=e+l]>0.$
+A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
+We discuss the DID-IV identification strategy in two periods and two groups with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$.
+In this setting, our target parameter is the LATET’ defined below.
+Definition.
+The local average treatment effect on the treated with the treatment path (LATET’) is
+$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
+where $\Delta_{0}$ and $\Delta_{1}$ are:
+$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
+$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
+The weights $w_{0}$ and $w_{1}$ are:
+$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
+We consider the identifying assumptions for the Wald-DID estimand to identify the LATET’.
+First, we make the following no anticipation assumptions on the outcome instead of imposing Assumption 1 (No carryover assumption).
 Assumption 23 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
-[…]
+$\forall z\in\mathcal{S}(Z),Y_{i,0}((0,1),z)=Y_{i,0}((0,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land NT^{T}).$
 This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
-Note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al.
-(2023) (see Table 3 and Table 4 in Subsection 2.5) [^9] It is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
 Assumption 24 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
-[…]
+$\forall z\in\mathcal{S}(Z),Y_{i,0}((1,1),z)=Y_{i,0}((1,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land DF^{T}).$
 This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
-Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
-We can rationalize Assumption 21 and 22 by noting that both of them and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units belonging to the exposed group and the compliers $CM^{Z}$.
-We now make the parallel trends assumption in the outcome suitable for the set up considered here.
+We can rationalize Assumptions 23-24 by considering the similarity between these assumptions and the no anticipation assumption made in the recent DID literature; all of them restrict the anticipatory behavior for those who belong to an exposed group (treatment group) and are affected by a policy shock (the treatment in DID designs).
+Next, we make the parallel trends assumption in the outcome suitable for the set up considered here.
 We first define the similar notation as in Subsection 2.4.
@@
 Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
-$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},$
-$\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
+$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},\hskip 5.69054pt\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
 where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
-As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups.
+As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the same expected time gain between exposed and unexposed groups.
 The following assumption makes such a restriction in the current set up.
 Assumption 25 (Parallel trend assumption in the outcome).
-The expected time gain between exposed and unexposed groups is the same; that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
-The theorem below shows that a Wald-DID estimand is equal to the target parameter $LATET^{\prime}$ under Assumption 2-5 and Assumption 21-23.
+The expected time gain between exposed and unexposed groups is the same: $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
+The theorem below shows that the Wald-DID estimand is equal to the LATET’ under Assumptions 2-5, 7 and Assumptions 23-25.
 Theorem 5.
-Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$; that is, we have
-$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]}=LATET^{\prime}.$
+Suppose Assumptions 2-5, 7 and Assumptions 23-25 hold. Then, the Wald-DID estimand corresponds to the LATET’.
 Proof.
@@
 $=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (15)
-The second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
-Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
+Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
+$E[Y_{1}-Y_{0}|E=0]=\bar{\Delta}_{0}.$ (16)
+This follows from Table 2 and the definition of the expected time gain $\bar{\Delta}_{0}$.
+Finally, we decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
 $E[Y_{1}-Y_{0}|E=1]=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
@@
 The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
-Next, we decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
-$E[Y_{1}-Y_{0}|E=0]$
-$=E[Y_{1}((1,1))-Y_{0}((1,1))|E=0,AT^{T}]Pr(AT^{T}|E=0)$
-$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=0,CM^{T}]Pr(CM^{T}|E=0)$
-$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=0,NT^{T}]Pr(NT^{T}|E=0)$
-$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=0,DF^{T}]Pr(DF^{T}|E=0)$
-$=\bar{\Delta}_{0}.$ (18)
-The first equality follows from Table 2 and the second follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
 From the result (16) with (17) and Assumption 25, we obtain
@@
 Combining the result (18) with (15), we obtain the desired result. ∎
-Appendix C Proofs in Section 4
-C.1   Proof of theorem 4
-First, we prove the consistency of our SLS estimator. $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
-$\hat{\alpha}^{e,l}$ $=\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
-$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
-$\hat{\pi}^{e,l}$ $=\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]}{E_{N}[\mathbf{1}\{E_{i}=e\}]}-\frac{E_{N}[(D_{i,e+l}-D_{i,e-1})\cdot\mathbf{1}\{E_{i}\in U\}]}{E_{N}[\mathbf{1}\{E_{i}\in U\}]}$
-$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
-We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined.
-More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$.
-This assumption guarantees that each Wald-DID estimator
-$\widehat{CLATT}_{e,e+l}$ $=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}}$
-$=\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}.$
-is also well defined.
-First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and Slutsky’s theorem, we have
+Appendix B Deriving the influence function in repeated cross section settings
 […]
-$=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e],$
-where the second equality follows from the Law of Iterated Expectations.
-By similar argument, we have
-[…]
-[…]
-[…]
-From Slutsky’s theorem, we have
-[…]
-Lastly, we show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
-For the numerator, we have
-$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
-$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
-$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
-$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
-$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (21)
-The third equality follows from Assumption 8, 9 and 11. The fourth equality follows from Assumption 13 and the fifth follows from Assumption 10.
-For the denominator, by similar calculations,
-$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
-$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
-$=Pr(CM_{e,e+l}|E_{i}=e),$ (22)
-where the final equality follows from Assumption 10 and Assumption 12.
-Combining (21) with (22), we obtain the desirable result:
-$\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}$
-$=\frac{E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e)}{Pr(CM_{e,e+l}|E_{i}=e)}$
-$=CLATT_{e,e+l}.$
-Next, we prove that our STS estimator is asymptotically normal, deriving its influence function.
-We define $\theta_{e,l}$ to be
-$\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
-$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
-The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation.
-Fact 1.
-If
-$\sqrt{n}(\hat{A}-A)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}+o_{p}(1),\hskip 8.53581pt\sqrt{n}(\hat{B}-B)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}b_{i}+o_{p}(1),$
-we have
-$\sqrt{n}\left(\frac{\hat{A}}{\hat{B}}-\frac{A}{B}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{a_{i}-(A/B)b_{i}}{B}+o_{p}(1).$
-First we derive the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{1}+\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}+\alpha_{e,l}^{2}))$. We note that
-$\sqrt{n}(E_{N}[\mathbf{1}\{E_{i}=e\}]-E[\mathbf{1}\{E_{i}=e\}])=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\mathbf{1}\{E_{i}=e\}-E[\mathbf{1}\{E_{i}=e\}]).$
-and
-$\sqrt{n}(E_{N}[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}]-E[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}])$
-$=\frac{1}{\sqrt{n}}\sum_{i}^{n}((Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}-E[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]).$
-From the observations and using Fact 1, we obtain
-$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\alpha_{e,l}^{1})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}+o_{p}(1).$
-$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\zeta_{i,e,l}+o_{p}(1).$
-By symmetry, we also obtain the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$:
-$\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}+o_{p}(1).$
-$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\phi_{i,e,l}+o_{p}(1).$
-This implies that
-$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}-\alpha_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\zeta_{i,e,l}-\phi_{i,e,l})+o_{p}(1)$
-$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}a_{i,e,l}+o_{p}(1).$ (23)
-Next we derive the influence function of $\sqrt{n}(\hat{\pi}_{e,l}^{1}+\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}+\pi_{e,l}^{2}))$. From the symmetry, by replacing $Y$ with $D$, we obtain the following:
-$\sqrt{n}(\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}-\pi_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\eta_{i,e,l}-\xi_{i,e,l})+o_{p}(1)$
-$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}b_{i,e,l}+o_{p}(1),$ (24)
-where we define
-$\eta_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]},$
-$\xi_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}.$
-Using (23), (24) and Fact 1, by some calculations, we obtain
-$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\sqrt{n}(\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}})$
-$=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1),$
-where the influence function $\psi_{i,e,l}$ is:
-$\psi_{i,e,l}$
-$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\left(\zeta_{i,e,l}-\phi_{i,e,l}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}(\eta_{i,e,l}-\xi_{i,e,l})\right)$
-$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (25)
-where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
-Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
-$V(\psi_{i,e,l})$
-$=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
-C.2   Repeated cross sections
-We also present the influence function of our STS estimator when repeated cross section data are available.
-[^10] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who presented the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups..
-Let $\theta_{e,l}^{r}$ define
 $\theta_{e,l}^{r}\equiv\frac{\alpha_{e,e+l}-\alpha_{e,e-1}-[\beta_{U,e+l}-\beta_{U,e-1}]}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]},$
@@
 In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$.
-Using similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
-$\psi_{i,e,l}^{r}$ $=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}-[\gamma_{e,l}^{1}-\gamma_{e,l}^{2}]}$
-$\times$ $\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
+Using similar arguments in the proof of Theorem 3, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 14, 20-22, and have the following influence function:
+$\psi_{i,e,l}^{r}=\frac{1}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]}$
+×$\times\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
 $-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg{)},$ (19)
-where $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
-C.3   Triple DID-IV designs
-We also present the influence function of our triple DID-IV estimator when panel or repeated cross section data are available. Let $\theta_{e,l,g}^{p}$ and $\theta_{e,l,g}^{r}$ define
-$\theta_{e,l,g}^{p}\equiv\frac{\alpha_{e,A}-\alpha_{e,B}-[\beta_{U,A}-\beta_{U,B}]}{\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]},$
-$\theta_{e,l,g}^{r}\equiv\frac{[(\alpha_{e,e+l,A}-\alpha_{e,e-1,A})-(\alpha_{e,e+l,B}-\alpha_{e,e-l,B})]-[(\beta_{U,e+l,A}-\beta_{U,e-1,A})-(\beta_{U,e+l,B}-\beta_{U,e-1,B})]}{[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]},$
-where
-$\alpha_{e,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}],$
-$\pi_{e,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B),$
-and
-$\alpha_{e,t,g^{\prime}}=E[Y_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,t,g^{\prime}}=E[Y_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}],$
-$\pi_{e,t,g^{\prime}}=E[D_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,t,g^{\prime}}=E[D_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B,\hskip 5.69054ptt=e+l,e-1).$
-In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function.
-$\psi_{i,e,l,g}^{p}=\frac{1}{DID_{e,l,g}^{p}}(\zeta_{i,e,A}-\zeta_{i,e,B}-\zeta_{i,U,A}+\zeta_{i,U,B}),$
-where $DID_{e,l,g}^{p}=\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]$, $\delta_{i,e,l,g}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l,g}^{p}\cdot(D_{i,e+l}-D_{i,e-1})$, and
-$\zeta_{i,e,g^{\prime}}=\frac{\mathbf{1}_{e,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}=e,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,g^{\prime}}]},\hskip 5.69054pt\zeta_{i,U,g^{\prime}}=\frac{\mathbf{1}_{U,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}\in U,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,g^{\prime}}]}.$
-In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function.
-$\psi_{i,e,l,g}^{r}$
-$=\frac{1}{DID_{e,l,g}^{r}}([\zeta_{i,e,e+l,A}-\zeta_{i,e,e-1,A}-\zeta_{i,e,e+l,B}+\zeta_{i,e,e-1,B}]-[\zeta_{i,U,e+l,A}-\zeta_{i,U,e-1,A}-\zeta_{i,U,e+l,B}+\zeta_{i,U,e-1,B}]),$
-where $DID_{e,l,g}^{r}=[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]$, $\delta_{i,e,l,g}^{r}=Y_{i}-\theta_{e,l,g}^{r}\cdot D_{i}$, and
-$\zeta_{i,e,t,g^{\prime}}=\frac{\mathbf{1}_{e,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,t,g^{\prime}}]},$
-$\zeta_{i,U,t,g^{\prime}}=\frac{\mathbf{1}_{U,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,t,g^{\prime}}]}.$
-Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
-Here, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006).
-Let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$.
-Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
-$R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}.$
-First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^11] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024).
-In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
+where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true, and $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
+Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
+[…]
+[…]
 $\hat{\beta}_{IV}$ $=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]}$
@@
 and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
-Using similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
 […]
-where $CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e]$. Similarly, we have
+$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (20)
+where we define:
+$CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e],$
+$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
+and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
+$w_{e,t}^{1}$ (21)
+$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (22)
+$w_{e,t}^{2}$ (23)
+$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (24)
+where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
 […]
-Combining the result (30) with (34) and by Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
-$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (35)
-where $\Delta_{e,t}$ is:
-$\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
-and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
-$w_{e,t}^{1}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (36)
-$w_{e,t}^{2}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (37)
-where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
-In Oreopoulos (2006), we can only identify $CLATT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in North Ireland in $1957$.
-This implies that each $\Delta_{e,t}$ in equation (35) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
 Figure 3 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term).
-The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, using equations (30) - (34).
-Figure 4 shows that all the CLATTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted.
-This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
+The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
+[…]
 Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006).
@@
 The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
+Appendix D Relationship between DID and DID-IV
+In this appendix, we describe the relationship between DID and DID-IV designs, and establish that DID-IV is DID with noncompliance of the assigned treatment.
+Following Subsection 2.1, we introduce the notation in DID designs.
+Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$.
+In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group).
+Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group.
+We refer to treatment and control groups as exposed and unexposed groups, respectively.
+Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units.
+First, we assume that all units are not allowed to receive the treatment in period 00.
+$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (25)
+Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
+$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (26)
+These restrictions imply that we have Tables 5-6 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
+Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions.
+First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs.
+Assumption 3 follows from (26) and Assumption 4 follows from (25).
+Assumption 5 follows from combining (25) with (26).
+Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 5-6:
+Table 5: Exposed group (assigned to $z=(0,1)$)
+observed counterfactual
+$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
+$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
+Table 6: Treatment group (assigned to $z=(0,0)$)
+observed counterfactual
+$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
+$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
+$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
+$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
+$=LATET.$
+Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$).
+Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
+Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups.
+From Tables 5-6, we can rewrite the parallel trends assumption in the untreated outcome as:
+$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
+$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
+$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
+$\iff$ $\Delta_{0}=\Delta_{1},$
+where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
+$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
+[…]
 Appendix E Comparing DID-IV to Fuzzy DID
-In this appendix we compare DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH”).
-We assume that the reader has briefly read dCDH.
-We first explain how dCDH’s identifying assumptions are related to ours in this article.
-We then point out the issues inherent in Fuzzy DID designs.
-First, we show that dCDH’s assumptions ex ante impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups.
-Next, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
-At the end of this appendix, we also revisit the ”issue” pointed out by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
+[…]
 E.1   Comparing the identifying assumptions in dCDH and this article
-In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours.
-Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-18.
-For now, we ignore the difference in the definition of target parameter between dCDH and this article.
-dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^13] In repeated cross section settings, the formal definition is provided in Assumption 29..
-In dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^14] This follows from two observations in dCDH.
+In this subsection, we investigate the detailed connections between dCDH’s identifying assumptions and ours.
+Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-19.
+For now, we ignore the difference in the definition of target parameter.
+dCDH consider the following identification assumptions for the Wald-DID estimand to capture the SLATET.
+[^25] In repeated cross section settings, the formal definition is provided in Assumption 31.
+Note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation.
+[^26] This follows from two observations in dCDH.
 First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively.
-Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation)..
+Second, the group variable $G$ is included in Assumption 28 (Treatment participation equation).
 We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively.
-[^15] Note that in Appendix B.2, we construct the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage), the hidden assumption in the previous literature.
+[^27] Note that in Appendix A.2, we distinguish the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage).
 For the detailed explanations for each assumption, see dCDH.
@@
 Assumption 29 (Parallel trends assumption in the untreated outcome).
-$E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]=E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0].$
+$E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]$
+$=$ $E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0].$
 Assumption 30 (Stable treatment effect over time).
@@
 In Assumption 30, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
-We now state the relationships between dCDH’s assumptions and ours.
-We first explain the similarities.
-First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 imply this condition [^16] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
-Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article.
-The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002).
-We point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26; that is, the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
-Finally, dCDH implicitly impose Assumption 14 (Exclusion restriction) and Assumption 16 (No anticipation in the first stage) when they formalize potential outcomes and treatment participation equations.
+[…]
+Finally, dCDH implicitly impose Assumption 14 and Assumption 16 when they formalize potential outcomes and treatment participation equations.
 Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14.
@@
 Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 30) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET.
-Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
+Finally, dCDH assume the condition (27) in Assumption 26, which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
 Remark E.1.1.
@@
 $SLATET=SLATET^{\prime},$
-where we define171717dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
+where we define282828dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
 When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@
 $SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)].$
-This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the ”homogeneous” treatment effect assumption.
-In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the ”homogeneous” treatment effect between the two groups.
-In the following discussion, we therefore treat Assumptions 24-28 as dCDH’s main identifying assumptions.
-E.2   Strong and asymmetric restrictions in Fuzzy DID designs
-We show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups.
+This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption.
+In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups.
+In the following discussion, we therefore treat Assumptions 26-30 as dCDH’s main identifying assumptions.
+E.2   Strong restrictions under Fuzzy DID designs
+In this subsection, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units.
 First, we prove some lemmas implied by dCDH’s assumptions.
@@
 These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group.
-At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours in DID-IV designs.
-First, we show that Assumption 26 (Treatment participation equation) implies the following assumption.
+At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours.
+First, we show that Assumption 28 implies the following assumption.
 Assumption 32 (Monotone change in treatment in exposed and unexposed groups).
@@
 […]
-because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$.
-Completing the proof. ∎
-Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively.
+[…]
+because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$. ∎
+Lemma 1 shows that under Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively.
 Note that the treatment participation equation itself does not imply Assumption 32: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption.
 Indeed, in this article, we do not assume this restriction.
-As we have seen in section 2.4 in this article, time also plays the role of instrument in DID-IV settings.
+As we have seen in Subsection 2.4, time also plays the role of instrument in DID-IV settings.
 From this observation, we can view that in an unexposed group ($z=0$), Assumption 32 corresponds to a monotonicity assumption with respect to time, formally defined below.
@@
 We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-Tables 9 - 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^18] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
-The distinction between dCDH’s restrictions and ours is apparent when we compare Tables 9 - 10 to Table 1 - 2191919Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings.
-The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
-Table 9: Exposed group ($z=1$)
-observed counterfactual
-$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
-$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
-$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
-$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
-$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
-Table 10: Unexposed group ($z=0$)
-observed counterfactual
-$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
-$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
-$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
-$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
-$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-These differences between dCDH’s restrictions and ours seem critical in many empirical settings.
-Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, and the practitioners may want to avoid imposing these restrictions in empirical work.
+Tables 3-4 in Subsection 2.5 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
+Comparing Tables 3-4 to Tables 1-2, one can see the difference between dCDH’s restrictions and ours [^29] Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings.
+The only modification is that as in Tables 3-4, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
+The distinction between dCDH’s restrictions and ours appears to be important in many empirical settings.
+Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions might be difficult to defend in some applications, and the practitioners may want to avoid imposing these restrictions in practice.
 For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time.
 Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
-Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications.
-However, these assumptions give rise to two additional issues.
-First, when we impose the condition (38) in Assumption 24, we should test its plausibility from the data because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$.
-Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group) in our fuzzy setting”, they do not explain how we should proceed when this condition is violated in practice.
-Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time.
-Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility are not clear, especially for multiple period and multiple group settings.
 E.3   Decomposing the SLATET
-In the previous subsection, we investigate how dCDH’s identifying assumptions ex ante restrict the treatment adoption behavior across units between exposed and unexposed groups.
-In this subsection, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
-The following theorem presents the decomposition result for the SLATET under Fuzzy DID designs, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
+In this subsection, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs.
+The following theorem presents the decomposition result for the SLATET, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
 Theorem 6.
-Suppose Assumptions 24 - 28 hold. Then, we can decompose the SLATET into two different causal parameters:
+Suppose Assumptions 26-30 hold. Then, we can decompose the SLATET into two different causal parameters:
 $SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)]$
@@
 The proof directly follows from Table 3 in an exposed group. ∎
-Theorem 7 shows that we can decompose the SLATET into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
-This theorem has two concerning implications.
-First, the interpretation of the SLATET is less clear than that of the LATET because it consists of two different causal parameters.
-Although the SLATET captures the ”local” average treatment effect, it may be difficult to interpret economically in a given application because one of the two causal parameters captures the treatment effects, for those who are induced to treatment by time.
-Second, the SLATET may not be policy-relevant even when the researchers exploit the policy change of their interest as an instrument for treatment.
-Suppose that the instrument represents an important policy lever for researchers, who are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group.
-In this situation, our target parameter, the LATET, is a policy-relevant parameter by construction.
-By contrast, the SLATET only captures the treatment effects among the type $(CM^{Z}\land NT^{T})$, which is the sub-population of the compliers $CM^{Z}$, even when the proportion of the time compliers $CM^{T}$ is zero within that group.
-As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this parameter even in such policy-relevant contexts.
+Theorem 6 shows that the SLATET can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
+This theorem has a concerning implications; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher.
+Suppose that the instrument represents an important policy lever for researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group.
+In this setting, our target parameter, the LATET, is a policy-relevant by construction.
+However, the SLATET may not be policy-relevant because this parameter captures the treatment effects among the sub-population of the compliers and the time compliers, who are affected by time but not affected by the instrument.
+As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this paramter even in such policy-relevant contexts.
 Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 31.
-When Assumption 25 (Stable percentage of treated units in an unexposed group) is violated, dCDH additionally consider Assumption 29 for the Wald-DID estimand to identify the SLATET, and call this the ”homogeneous” treatment effect assumption between exposed and unexposed groups.
-However, we show that this assumption cannot be interpreted as requiring the ”homogeneous” treatment effect among the same sub-population between exposed and unexposed groups.
-If we do not assume Assumption 25, we have Table 11 in an unexposed group instead of Table 10 in Appendix E.2.
-Here, we assume that Assumption 26 and 31 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
+We show that this assumption cannot be interpreted as requiring the homogeneous treatment effect among the same sub-population between exposed and unexposed groups.
+If we do not assume Assumption 27, we have Table 5 in an unexposed group.
+[^30] Here, we assume that Assumption 28 and 33 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
 Table 5: Unexposed group ($z=0$)
@@
 $=E[Y(1)-Y(0)|Z=0,T=1,CM^{T}].$ (32)
-By combining equations (42) and (43), we can conclude that Assumption 29 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups.
+By combining equation (31) with (32), one can see that Assumption 31 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups.
 Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group.
-This interpretation implies that the requirement of the ”homogeneous” treatment effect in Assumption 29 is less clear than what one might expect.
-E.4   Revisiting the ”issue” regarding the use of the Wald-DID estimand
-Finally, we revisit the ”issue” raised by dCDH regarding the use of the Wald-DID estimand in Fuzzy DID designs, and show that their argument is misguided.
-We also clarify why dCDH exclude so many types in exposed and unexposed groups as we see in Appendix E.2.
-Hereafter, we assume Assumptions 24 - 28 in Fuzzy DID designs.
-dCDH argue that the Wald-DID estimand requires Assumption 28 (stable treatment effect over time) to identify the SLATET, and propose alternative estimands that do not rely on this assumption.
-We show that their argument stems from their reliance on the parallel trends assumption in the untreated outcome, which is common in DID designs, instead of on the parallel trends assumption in the unexposed outcome as in this article.
-Next, we argue that the former type of parallel trends assumption is unsuitable for DID-IV settings.
+This interpretation implies that the requirement of the homogeneous treatment effect in Assumption 31 is less clear than what one might expect.
+E.4   Revisiting the issue regarding the use of the Wald-DID estimand
+Finally, we revisit the issue pointed by dCDH regarding the use of the Wald-DID estimand, and show that this problem is not serious than one might expect, and does not arise under our DID-IV designs.
+We also clarify why dCDH exclude so many types in exposed and unexposed groups as seen in Tables 3-4.
+Hereafter, we assume Assumptions 26-30 in Fuzzy DID designs.
+dCDH argue that the Wald-DID estimand requires Assumption 30 to identify the SLATET, and propose alternative estimands that do not rely on this assumption.
+We first show that their argument depends on the parallel trends assumption in DID designs.
+Next, we argue that this assumption would be difficult to justify in DID-IV settings.
 First, we clarify the interpretation of Assumption 30 by introducing the additional notation.
-This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but dCDH do not provide the reason why this strong restriction is required only for the sub-population.
+This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why this strong restriction is required only for the sub-population.
 For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$.
 Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
-Then, from Table 9 and Table 10, we can decompose Assumption 28 as follows:
-$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
+Then, from Tables 3-4, we can decompose Assumption 30 as follows:
+$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$
+$=$ $E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
 $\iff$ $\delta_{0,1,AT^{T}}=\delta_{0,0,AT^{T}},$ (33)
-$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]=E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
+$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$
+$=$ $E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
 $\iff$ $\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
@@
 Equation (34) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
-Conditions (44) and (45) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcomes.
-In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods (see Table 9 and Table 10).
+Conditions (33) and (34) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcome.
+In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods.
 Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (33) and (34).
-To see this formally, we use our notation to briefly sketch the proof of dCDH on the Wald-DID estimand identifying SLATET under fuzzy DID designs.
+To see this formally, we use our notation to briefly sketch dCDH’s proof about how the Wald-DID estimand identifies the SLATET under fuzzy DID designs.
 First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand.
@@
 Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 27 and Assumption 28, the Wald-DID estimand captures the SLATET under Assumption 29 (parallel trends assumption in the untreated outcome) and conditions (33) and (34).
-From the discussions, one can conclude that the parallel trends assumption in the untreated outcome leads dCDH to impose conditions (44) and (45).
+From the above discussion, one can see that imposing the parallel trends assumption in the untreated outcome leads dCDH to assume conditions (33) and (34).
 Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET.
@@
 Remark E.4.1.
-dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 9-10).
-This identification result does not seem intuitive because the identification variation comes from the policy shock (instrument) in period $1$.
+dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 3-4).
+This identification result, however, is not intuitive because the identification variation comes from a policy shock (instrument) in period $1$.
 Remark E.4.2.
-dCDH exclude the time compliers $CM^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for that type, we additionally capture the treatment effects among the time compliers $CM^{T}$ from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which implies that the SLATET in an exposed group is offset by the SLATET’ in an unexposed group.
-We have confirmed that dCDH’s argument regarding the use of the Wald-DID estimand rests on the parallel trends assumption in the untreated outcome in DID designs.
-Finally, we argue that this assumption is unsuitable for DID-IV settings for three reasons.
-First, this assumption is not sufficient to capture the average time trends of the outcome even in an unexposed group.
-This is because in DID-IV settings, the units are allowed to take the treatment in the absence of the instrument, and we cannot observe the untreated outcomes for these types in an unexposed group.
-Second, in most of the DID-IV applications, we can not impute the untreated potential outcomes in general.
-For instance, Black et al.
-(2005) estimates the causal link between parents’ and children’s education attainment, exploiting variation in the timing of the implementation of the school reform across municipalities and cohorts as an instrument for parents’ education attainment.
-In the authors’ setting, the parallel trends assumption in the untreated outcome requires the data to contain parents with zero education attainment, which is unrealistic and not what the authors would have in mind in practice.
-Finally, this assumption is generally not testable using the pre-exposed period data in a given application.
-This is apparent, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
-dCDH propose the placebo test in their appendix, but we show that it is not enough for checking the plausibility of the parallel trends assumption in the untreated outcome.
+dCDH exclude the time compliers $CM^{T}$ in an unexposed group in order to avoid the SLATET in an exposed group being offset by the SLATET’ in an unexposed group.
+If we add and subtract the average time trends of the untreated outcome for the time compliers, we additionally capture the SLATET’ in an unexposed group from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$.
+Thus far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome.
+Finally, we argue that this assumption would be difficult to rely on in DID-IV settings for three reasons.
+[…]
+Finally, this assumption is generally not testable using the pre-exposed period data, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
+Although dCDH propose the placebo test in their appendix, it has shortcomings for checking the plausibility of the parallel trends assumption in the untreated outcome.
 dCDH’s test has two steps.
+[^31] Although dCDH ”assume” the conditions (35) and (36), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 29.
+We, therefore, describe the dCDH’s test in two steps.
 First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
@@
 Second, we test the following null hypothesis:
-$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]=E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (48)
-Although dCDH ”assume” the conditions (46) and (47), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 27.
-We, therefore, describe the dCDH’s test in two steps.
-dCDH’s test has some drawbacks.
-First, the conditions (46) and (47), which should be checked in the first step, are strong, but dCDH do not explain the way ahead when these conditions are violated in practice.
+$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$ (37)
+$=$ $E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (38)
+dCDH’s test has some weaknesses.
+First, the conditions (35) and (36), which should be checked in the first step, are strong, but no guidance is provided when these conditions are violated in practice.
 Second, in this test, the second step does not directly test Assumption 29 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$.
 Indeed, dCDH describe the equality (37) as a necessary condition for Assumption 29 and Assumption 30 to be satisfied under the conditions (35) and (36).
-Finally, it is unclear how this test should be conducted with multiple pre-exposed period data.
-dCDH assume ”for instance that data is available for the period $T=-1$”, but they do not indicate how this test can be performed in multiple pre-exposed period settings.
+Finally, it is not clear how this test should be conducted with multiple pre-exposed period data.
+Although dCDH assume “for instance that data is available for the period $T=-1$”, it is not indicated how this test can be performed in multiple pre-exposed period settings.
 Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (35)-(36) in the first step with the weaker one.
-$E[D|Z=1,T=0]-E[D|Z=1,T=-1]=E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (49)
+$E[D|Z=1,T=0]-E[D|Z=1,T=-1]$ (39)
+$=$ $E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (40)
 Under repeated cross section settings, the conditions (37)-(39) coincide with the conditions (1)-(2).
+This indicates that the rationale behind dCDH’s test is more suitable for the assessment of parallel trends assumptions in DID-IV designs.
+E.5   Other connections
+In this subsection, we document the relationship between DID-IV and Fuzzy DID under non-binary, ordered treatment, and multiple groups settings.
+We also compare the Wald-DID estimand to the time-corrected Wald estimand considered in dCDH.
+E.5.1 Non-binary, ordered treatment
+In this subsection, we investigate the relationship between DID-IV to Fuzzy DID under non-binary, ordered treatment settings.
+First, we compare the target parameter between the two designs.
+With a non-binary, ordered treatment $D\in\{0,1,\dots,J\}$, dCDH consider the following target parameter, which we call the switcher average causal response (SACR).
+$SACR\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(0)<j<D_{1}(1)),$
+where $w_{j}$ is
+$w_{j}=\frac{P(D\geq j|Z=1,T=1)-P(D\geq j|Z=1,T=0)}{E[D|Z=1,T=1]-E[D|Z=1,T=0]}.$
+[…]
+Next, we compare the identifying assumptions between the two designs. In order for the Wald-DID estimand to identify the SACR, dCDH replace Assumption 28 with the following one.
+Assumption 34 (Ordered treatment equation).
+$D=\sum_{j=1}^{J}\mathbf{1}\{V\geq v^{j}_{ZT}\},\hskip 5.69054pt\text{with}\hskip 5.69054pt-\infty=v^{0}_{ZT}<v^{1}_{ZT}\dots<v^{J+1}_{ZT}=\infty\hskip 5.69054pt\text{and}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
+In addition, dCDH replace the condition (27) in Assumption 26 with the condition (41), and Assumption 27 with the condition (42), respectively.
+$D_{11}\succeq D_{10},$ (41)
+$D_{01}\sim D_{00},$ (42)
+where $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
+Note that the condition (42) is stronger than Assumption 27 because this assumption requires that the distribution of the treatment should be stable during the two periods in an unexposed group.
+In our DID-IV designs, the identifying assumptions required for the Wald-DID estimand to capture the ACRT are unchanged under non-binary, ordered treatment settings.
+In addition, Assumptions 17-18 are indirectly testable using the pre-exposed period data in practice.
+E.5.2 Multiple groups
+In multiple groups settings, our DID-IV design is not directly compatible to Fuzzy DID design because dCDH exploit the group structure in the data ($G\in\{0,1,\dots,\bar{g}\}$) for the Wald-DID estimand to identify the following target parameter.
+$\Delta^{\ast}=E[Y(1)-Y(0)|\cup_{g=0}^{\bar{g}}S_{g},T=1],$
+where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ represents the switchers in group $G=g$ [^32] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
+[…]
+[…]
+In some applications, a researcher might prefer our approach to dCDH’s approach for three reasons.
+First, our approach yields a policy-relevant parameter by construction if the instrument represents a policy change of interest to the researcher.
+Second, in multiple period settings, our approach avoids the need to estimate the super groups in each consecutive period $(t-1,t)$, which simplifies both estimation and inference, making our method easier to implement in practice.
+Finally, in our approach, the monotonicity assumption can be easy to justify if the instrument is based on a policy shock.
+In dCDH’s approach, this assumption might be harder to defend in practice because the group variable $G$ typically represents the geographic feature in the data such as municipality, state, or country.
+[^34] Note that, in multiple groups settings, dCDH also maintain the monotonicity assumption on each group pair $(g,g^{\prime})$ through Assumption 28.
+Remark E.4.3 (Multiple groups settings with an ordered treatment).
+[…]
+E.5.3 Alternative estimand
+In this subsection, we compare the Wald-DID estimand to the alternative estimand proposed by dCDH that does not rely on Assumption 29 and 30.353535In this subsection, we do not consider Wald-CIC estimand proposed by dCDH, as we focus on the mean treatment effect in this article.
+As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand defined below.
+$W_{TC}\equiv\frac{E[Y|Z=1,T=1]-E[Y+(1-D)\delta_{0}+D\delta_{1}|Z=1,T=0]}{E[D|Z=1,T=1]-E[D|Z=1,T=0]},$
+where $\delta_{d}=E[Y|Z=0,T=1,D_{1}(0)=d]-E[Y|Z=0,T=0,D_{0}(0)=d]$.
+In order for the TC-Wald estimand to identify the SLATET, dCDH consider the following identifying assumption instead of imposing Assumption 29 and 30.
+Assumption 35 (Conditional common trends).
+For all $d\in\mathcal{S}(D),$
+$E[Y(d)|Z=1,T=1,D_{0}(0)=d]-E[Y(d)|Z=1,T=1,D_{0}(0)=d]$
+$=$ $E[Y(d)|Z=0,T=1,D_{0}(0)=d]-E[Y(d)|Z=0,T=1,D_{0}(0)=d].$
+Note that with a binary treatment, this assumption exploits the observed outcome path in an unexposed groups similar to Assumption 18 in our DID-IV designs (see Table 4).
+While the target parameter differs between the two designs, we see that using the Wald-DID estimand under DID-IV designs would be more appealing than using the TC-Wald estimand under Fuzzy DID designs for two reasons.
+First, when using the Wald-DID estimand in DID-IV designs, researchers can easily assess the plausibility of parallel trends assumptions in practice, employing the pre-trends tests described in Section 5.3.
+Although the TC-Wald estimand avoids the stable treatment effect assumption under Fuzzy DID designs, it presents challenges in testing the plausibility of Assumption 35 in pre-exposed period data: it is conditional on $D_{0}(0)=d$ and imposes restrictions on the mean evolution of $Y(d)$ for all $d\in\mathcal{S}(D)$.
+[…]

### v6

@@
-onlineReferences
+
 Abstract
-Many studies exploit variation in policy adoption timing across units as an instrument for treatment.
+Many studies exploit variation in the timing of policy adoption across units as an instrument for treatment.
 This paper formalizes the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
@@
 Abstract
-[…]
+In this online appendix, we first present additional identification results.
+Second, we derive the influence function of our STS estimator in repeated cross section settings.
+Third, we decompose the TWFEIV estimand in the setting of Oreopoulos (2006).
+Finally, we investigate the connections between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
 Keywords: difference-in-differences, instrumental variable, local average treatment effect, returns to education
 1  Introduction
-To identify the effect of a treatment on an outcome, many studies have exploited variation in policy adoption timing across units as an instrument for treatment.
-For example, to estimate returns to schooling in Indonesia, Duflo (2001) leverages variation in the timing of introduction of school construction program across regions as an instrument for education attainment.
-Similarly, to estimate the causal link between parents’ and children’s education attainment, Black et al.
-(2005) leverage variation in the timing of implementation of school reforms across municipalities as an instrument for parents’ education attainment.
-Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that the timing variation of a policy shock is used for identifying the treatment effects.
-Meanwhile, it also differs from DID designs, in that the variation is used to construct an instrument rather than a treatment.
-In this paper, we formalize the underlying identification strategy as an instrumented difference-in-differences (DID-IV).
-We define the target parameter and identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity.
+To identify the effect of a treatment on an outcome, many studies exploit variation in the timing of policy adoption across units as an instrument for treatment.
+For example, to estimate returns to schooling in Indonesia, Duflo (2001) exploits variation in the timing of the introduction of a school construction program across regions as an instrument for educational attainment.
+Similarly, to estimate the causal link between parents’ and children’s educational attainment, Black et al.
+(2005) exploit variation in the timing of the implementation of school reforms across municipalities as an instrument for parents’ educational attainment.
+Notably, the identification strategy underlying these studies is similar to difference-in-differences (DID) designs, in that variation in the timing of a policy shock is used to identify treatment effects.
+At the same time, it differs from DID designs in that this variation is used to construct an instrument rather than a treatment.
+In this paper, we formalize this identification strategy as an instrumented difference-in-differences (DID-IV).
+We define the target parameter and the identifying assumptions in this design, and develop a credible estimation method that is robust to treatment effect heterogeneity.
 We illustrate the empirical relevance of our findings with the setting of Oreopoulos (2006), estimating returns to schooling in the United Kingdom.
-In this empirical application, we demonstrate that the estimation method matters in this design in practice.
-First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument during the two periods (unexposed group), while some units become exposed in the second period (exposed group).
-[^1] Note that this two-period/two-group ($2\times 2$) setting is already considered in de Chaisemartin (2010) and Hudson et al.
-(2017), and our $2\times 2$ DID-IV design is mainly based on these studies.
+In this application, we show that the choice of estimation method matters in practice.
+First, we consider a simple setting with two periods and two groups: some units are not exposed to the instrument in either period (the unexposed group), while others become exposed in the second period (the exposed group).
+[^1] Note that this two-period/two-group ($2\times 2$) setting has already been considered in de Chaisemartin (2010) and Hudson et al.
+(2017), and our $2\times 2$ DID-IV design builds on these studies.
 In this paper, we revisit $2\times 2$ DID-IV designs for two reasons.
 First, we aim to complement de Chaisemartin (2010) and Hudson et al.
-(2017) along two dimensions.
-Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs, and uncover the additional identifying assumption not give in the previous literature.
-In addition, we investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in DID-IV designs.
-Second, we aim to compare $2\times 2$ DID-IV designs to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
-While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) treat the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between exposed and unexposed groups.
-Our target parameter is the local average treatment effect on the treated (LATET); this parameter measures the treatment effects, for those who belong to an exposed group and are induced to treatment by instrument in the second period.
-We show that in this design, a Wald-DID estimand, which scales the DID estimand of the outcome by the DID estimand of the treatment, captures the LATET.
-In DID-IV designs, the parallel trends assumption in the outcome requires careful assessment because it does not depend on the potential outcomes, but depends on the outcome under no exposure to the instrument, which we call an unexposed outcome.
-Give this, we next clarify the interpretation of this assumption in DID-IV settings.
-Specifically, we first point out that, in DID-IV designs, time also plays the role of instrument; that is, it affects one’s treatment choices over time in the absence of a policy shock (instrument).
-Following the terminology in Imbens and Angrist (1994), we divide the population into time groups defined by the potential treatment choices without instrument during the two periods: time always-takers, time never-takers, time compliers, and time defiers.
-Subsequently, we show that the average time trend in an unexposed outcome equals a weighted average of the effect of time on an outcome through treatment in each time group.
-We call this an expected time gain, and show that the parallel trends assumption in the outcome requires the same expected time gain between exposed and unexposed groups.
+(2017).
+Specifically, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption that is not made explicit in the previous literature.
+Second, we compare $2\times 2$ DID-IV designs to the Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018).
+While both de Chaisemartin (2010) and de Chaisemartin and D’Haultfœuille (2018) study the same setting, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
+In this setting, our DID-IV design relies on a monotonicity assumption and parallel trends assumptions for both the treatment and the outcome between the exposed and unexposed groups.
+Our target parameter is the local average treatment effect on the treated (LATET), which measures the treatment effect for units that belong to the exposed group and are induced to receive the treatment by the instrument in the second period.
+We show that, in this design, a Wald-DID estimand—defined as the ratio of the DID estimand of the outcome to that of the treatment—identifies the LATET.
 de Chaisemartin and D’Haultfœuille (2018) (hereafter, “dCDH”) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-The main difference between this article and dCDH lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period.
-Because the target parameter differs from each other, the identifying assumptions in dCDH are also different from those used here.
-Motivated by the differences between this article and dCDH, we next examine the detailed connections between DID-IV and Fuzzy DID, and point out the issues inherent in Fuzzy DID.
-First, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units, and these restrictions are asymmetric between unexposed and exposed groups.
-Second, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters: one measures the treatment effects among the sub-population of the compliers and the other measures the treatment effects among the time compliers.
-This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher.
-This is because the SLATET is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by instrument.
+The main difference between this paper and dCDH lies in the definition of the target parameter.
+While we focus on the LATET, dCDH focus on the switcher local average treatment effect on the treated (SLATET); this parameter measures the treatment effects, for those who belong to an exposed group and start receiving the treatment in the second period.
+Because the target parameters differ, the identifying assumptions adopted in dCDH also differ from those used in this paper.
+Motivated by these differences, we next examine the detailed connections between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter.
+We first show that the identifying assumptions under Fuzzy DID impose stronger restrictions on treatment adoption behavior than those under DID-IV.
+Under these restrictions, we then demonstrate that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two distinct causal parameters.
+One parameter captures the treatment effects among the subpopulation of the compliers in the sense of Imbens and Angrist (1994), while the other captures the treatment effects for time compliers—units whose treatment status is affected by time but not by the instrument.
+This decomposition result has an important implication: even when the instrument is directly linked to the policy change of interest, the SLATET may fail to be a policy-relevant parameter (Heckman and Vytlacil (2001)).
 Next, we extend the canonical DID-IV design to multiple period settings with the staggered adoption of the instrument across units.
-In most DID-IV applications, researchers exploit variation in policy adoption timing across units in more than two periods, instrumenting for the treatment with the natural variation.
-The instrument is constructed, for instance from the staggered adoption of school reforms across municipalities or across countries (e.g.
+In most DID-IV applications, researchers exploit variation in the timing of policy adoption across units in more than two periods, instrumenting for the treatment with the natural variation.
+The instrument is constructed, for example, from the staggered adoption of school reforms across municipalities or countries (e.g.
 Oreopoulos (2006), Lundborg et al.
 (2014), and Meghir et al.
-(2018)), phased-in introduction of head starts across states (e.g.
-Johnson and Jackson (2019)), or gradual adoption of broadband internet programs across municipalities (e.g.
+(2018)), the phased-in introduction of Head Start across states (e.g.
+Johnson and Jackson (2019)), or the gradual adoption of broadband internet programs across municipalities (e.g.
 Akerman et al.
@@
 (2013)).
-We refer to the underlying identification strategy as a staggered DID-IV design, and establish the target parameter and identifying assumptions.
-Specifically, we first partition units into mutually exclusive and exhaustive cohorts by the initial exposure date of the instrument.
-We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT); this parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs, and measures the treatment effects, for those who belong to cohort $e$ and are the compliers at a given relative period $l$ starting from the initial exposure to the instrument.
-Finally, we extend the identification assumptions in $2\times 2$ DID-IV designs to multiple period settings, and state the interpretation of each assumption.
-We extend our DID-IV designs in several directions.
-We first show that our DID-IV framework can be applied when we have a non-binary, ordered treatment.
-We also consider extensions to repeated cross sections.
-Lastly, we consider the case where we introduce the treatment path in potential outcomes in $2\times 2$ DID-IV designs.
+We refer to this identification strategy as a staggered DID-IV design, and establish the corresponding target parameter and identifying assumptions.
+Specifically, we first partition units into mutually exclusive and exhaustive cohorts based on the initial exposure date of the instrument.
+We then define our target parameter as the cohort specific local average treatment effect on the treated (CLATT).
+This parameter is a natural generalization of the LATET in $2\times 2$ DID-IV designs and measures the treatment effects for units that belong to cohort $e$ and are the compliers at a given relative period $l$ following the initial exposure to the instrument.
+Finally, we introduce two Wald–DID estimands that use either never-exposed cohorts or not-yet-exposed cohorts as control groups.
+For each estimand, we extend the identification assumptions from the $2\times 2$ DID–IV setting to multiple time periods, and show that, under these assumptions, the corresponding Wald–DID estimand identifies the CLATT at each relative period following the initial exposure to the instrument.
+We extend our DID-IV framework along two dimensions. First, we show that it can be applied to settings with a non-binary, ordered treatment. Second, we consider extensions to repeated cross sections.
 Finally, we propose a regression-based method to consistently estimate our target parameter in staggered DID-IV designs under heterogeneous treatment effects.
-In practice, when researchers implicitly rely on a staggered DID-IV design, they usually implement this design via two-way fixed effects instrumental variable (TWFEIV) regressions (e.g.
+In practice, when researchers implicitly rely on a staggered DID-IV design, they typically implement it using two-way fixed effects instrumental variable (TWFEIV) regressions (e.g.
 Johnson and Jackson (2019), Lundborg et al.
@@
 (2013)).
-In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the treatment or the outcome is not stable over time.
+In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize treatment effects if the effect of the instrument on the treatment or on the outcome is not stable over time.
 Our proposed method avoids this issue and is robust to treatment effect heterogeneity.
@@
 We call this a stacked two stage least squares (STS) regression and ensure its validity.
-Following Callaway and Sant’Anna (2021), we propose the weighting scheme to summarize the treatment effects, in which the weight reflects the share of compliers in a given relative period $l$ in cohort $e$.
-We also discuss the procedure of pre-trends tests to check the validity of the parallel trends assumptions in DID-IV designs.
-We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of implementation of school reforms between Britain and North Ireland as an instrument for education attainment.
-In this application, we first assess the plausibility of the DID-IV identification strategy implicitly imposed by Oreopoulos (2006).
+Following Callaway and Sant’Anna (2021), we propose a weighting scheme to summarize the treatment effects, where the weight reflects the share of compliers in a given relative period $l$ in cohort $e$.
+We also discuss the procedure of pre-trends tests to assess the validity of the parallel trends assumptions in DID-IV designs.
+We illustrate our findings with the setting of Oreopoulos (2006), who estimates returns to schooling in the UK, exploiting variation in the timing of the implementation of school reforms between Britain and Northern Ireland as an instrument for education attainment.
+In this application, we first assess the plausibility of the DID-IV identification strategy implicitly adopted by Oreopoulos (2006).
 We then estimate the TWFEIV regression in the author’s setting.
@@
 We find that our STS estimates are all positive in each relative period after school reform, and our weighting scheme yields a more plausible estimate than the TWFEIV estimate.
-Specifically, our weighting estimate indicates roughly a $20\%$ gain from schooling in the UK.
-Overall, this paper provides a new econometric framework by combining the IV techniques with DID designs.
-Our DID-IV design can be applied to various empirical settings.
-First, it can cope with the general adoption process of the treatment, which the canonical DID designs cannot, when we are interested in the effect of a treatment on an outcome.
-For instance, suppose that some units already receive the treatment in period 00 before the policy shock that distributes the treatment in period $1$.
-In this setting, the canonical DID design is infeasible in practice unless these units are discarded ex ante from the data.
-By contrast, our DID-IV design can be directly applied to this setting, allowing us to identify the LATET.
-In other words, we can view that DID-IV is DID with noncompliance of the assigned treatment.
-Second, our DID-IV design can yield valid counterfactuals to estimate the treatment effect when no control group exists under DID designs.
-For instance, suppose that the treatment is education attainment and we are interested in returns to schooling as in Duflo (2001).
-In this setting, the canonical DID design generally fails to construct the control group, as the data would not contain the units with zero education attainment in reality.
-By contrast, our DID-IV design can partition units into exposed and unexposed groups depending on whether the unit is exposed to the policy shock (instrument).
-Once these groups are constructed, one can identify the local average treatment effect by comparing the evolution of the treatment and the outcome between the two groups.
-Finally, our DID-IV design allows a researcher to address the endogeneity issue in a given application.
-In many empirical settings, a researcher often fails to identify the treatment effects due to the endogenous adoption of the treatment or the potential violation of the parallel trends assumption under DID designs.
-In this situation, our DID-IV design offers an alternative identification strategy, constructing an instrument based on the timing variation of a policy shock.
-Notably, in this design, the identifying variation comes not from the random assignment of a policy shock across units, but mainly comes from the parallel trends assumption in the treatment and the outcome.
-This identification strategy would be less demanding than IV designs in certain applications, and enhance the internal and external validity of the IV estimate in practice (Imbens and Angrist (1994); Heckman and Vytlacil (2001); Heckman and Vytlacil (2005)).
+Specifically, our weighted estimate indicates roughly a $20\%$ gain from schooling in the UK.
 The rest of the paper is organized as follows.
@@
 All proofs are given in the Appendix.
-1.1   Related literature
-Our paper is related to recent DID-IV literature (de Chaisemartin (2010); Hudson et al.
+Related literature
+Our paper is related to the recent DID-IV literature (de Chaisemartin (2010); Hudson et al.
 (2017); de Chaisemartin and D’Haultfœuille (2018); de Chaisemartin et al.
-(2024)) and contributes to this literature in three important ways.
+(2024); Chen et al.
+(2025); Helmers et al.
+(2025)) and contributes to this literature in three ways.
 [^2] While Ye et al.
-(2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they make strong assumptions for the Wald-DID estimand to capture this parameter.
-We therefore see that Ye et al.
-(2023) are distinct from the recent DID-IV literature.
-First, this paper investigates the detailed connections between DID-IV and Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
-In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that if parallel trends assumptions in the treatment and the outcome, and a monotonicity assumption hold, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET).
+(2023) also consider the “instrumented difference-in-differences”, their target parameter is the average treatment effect (ATE), and they impose strong assumptions for the Wald-DID estimand to identify this parameter.
+We therefore view Ye et al.
+(2023) as distinct from the recent DID-IV literature.
+First, this paper investigates the detailed connections between DID-IV and the Fuzzy DID considered in de Chaisemartin and D’Haultfœuille (2018).
+In econometrics, a pioneering work formalizing $2\times 2$ DID-IV designs is de Chaisemartin (2010), who shows that, under parallel trends assumptions for both the treatment and the outcome, together with a monotonicity assumption, the Wald-DID estimand identifies the local average treatment effect on the treated (LATET).
+[^3] Blundell and Dias (2009) also consider a DID–IV setting with a binary treatment and a binary instrument in Subsection F.2 of Section III.
+There are two differences between de Chaisemartin (2010) and Blundell and Dias (2009).
+First, while de Chaisemartin (2010) focus on the LATET as the target parameter, Blundell and Dias (2009) focus on the switcher local average treatment effect (SLATET), as in dCDH.
+Second, while de Chaisemartin (2010) impose a parallel trends assumption on unexposed outcomes, Blundell and Dias (2009) impose a parallel trends assumption on untreated outcomes.
 Following de Chaisemartin (2010), Hudson et al.
-(2017) also consider $2\times 2$ DID-IV designs with non-binary, ordered treatment settings.
-Building on the work in de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently and call them Fuzzy DID designs.
+(2017) also study $2\times 2$ DID-IV designs with a non-binary, ordered treatment.
+[^4] Recently, Chen et al.
+(2025) extend our DID-IV framework by allowing for covariates.
+Helmers et al.
+(2025) consider an extension to the continuous treatment case.
+Building on de Chaisemartin (2010), however, de Chaisemartin and D’Haultfœuille (2018) formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
 In this paper, we first formalize $2\times 2$ DID-IV designs, complementing de Chaisemartin (2010) and Hudson et al.
-(2017) in two ways.
-Specifically, the framework we consider is mainly based on de Chaisemartin (2010) and Hudson et al.
 (2017).
-Nevertheless, we introduce the instrument path into $2\times 2$ DID-IV designs and uncover the additional identifying assumption not given in the previous literature.
-We also investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
-Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID, and point out the issues inherent in Fuzzy DID.
-The discussion organizes the relationship between the two designs, which essentially treat the same setting, in the previous literature, and indicates the relative advantages of DID-IV over Fuzzy DID.
-Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, calling it a staggered DID-IV design.
+Specifically, while the framework we consider here is mainly based on de Chaisemartin (2010) and Hudson et al.
+(2017), we explicitly introduce the instrument path into $2\times 2$ DID-IV designs and uncover an additional identifying assumption not given in the previous literature.
+Given the identification results and the terminology developed in this paper, we then compare DID-IV to Fuzzy DID.
+Specifically, we clarify the differences between DID-IV and Fuzzy DID and discuss their implications for treatment adoption behavior and the interpretation of the target parameter.
+Second, this paper extends $2\times 2$ DID-IV designs to multiple period settings with the staggered adoption of the instrument across units, which we refer to as staggered DID-IV designs.
 In practice, empirical researchers often leverage variation in policy adoption timing across units as an instrument for treatment in more than two periods (e.g., Black et al.
@@
 (2014), and Johnson and Jackson (2019)).
-No previous study, however, has extended $2\times 2$ DID-IV designs to such important settings.
+However, no previous study has extended $2\times 2$ DID-IV designs to such important settings.
 In this paper, we formalize the underlying identification strategy as staggered DID-IV designs, and establish the target parameter and the identifying assumptions.
-Our staggered DID-IV designs allow the practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
+Our staggered DID-IV designs allow practitioners to estimate the local average treatment effect even when the treatment adoption is endogenous over time.
 Finally, this paper provides a credible estimation method in staggered DID-IV designs under heterogeneous treatment effects.
-When empirical researchers implicitly rely on the staggered DID-IV design in practice, they commonly implement this design via TWFEIV regressions (e.g.
+When empirical researchers implicitly rely on the staggered DID-IV designs in practice, they commonly implement this design via TWFEIV regressions (e.g.
 Johnson and Jackson (2019), Lundborg et al.
@@
 (2013)).
-In companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or the treatment evolves over time.
+In a companion paper (Miyaji (2024)), however, we show that in more than two periods, the TWFEIV estimand generally fails to summarize the treatment effects if the effect of the instrument on the outcome or on the treatment evolves over time.
 [^5] See also de Chaisemartin and D’Haultfœuille (2020), who decompose the numerator and denominator of the TWFEIV estimand separately, and point out the issue of interpreting this estimand causally in Section 3.4 of their Web Appendix.
@@
 In work related to this paper, de Chaisemartin et al.
-(2024) also consider DID-IV designs following de Chaisemartin (2010) and Hudson et al.
+(2024) also study DID-IV designs building on de Chaisemartin (2010) and Hudson et al.
 (2017).
-This paper is distinct from theirs in three ways.
+Our paper is distinct from theirs in three ways.
 First, they consider continuous treatment with continuous but static instrument, while we consider binary or ordered treatment with binary but dynamic instrument.
 Second, they consider multiple period settings with potentially non-staggered instrument, while we consider multiple periods settings with staggered instrument.
-Third, they consider the Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
-Our paper is also related to a growing DID literature in two ways.
-First, this paper provides an alternative identification strategy to DID designs, addressing cases where the parallel trends assumption is not satisfied or no control group exists in a given application.
-When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often concern that the parallel trends assumption is not plausible in practice.
-To address this, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g.
-Miller and Segal (2019)).
+Third, they view Fuzzy DID as a special case of DID-IV, while we formally discuss the relationship between the two designs.
+Our paper is also related to the recent DID literature in two ways.
+First, this paper provides an alternative identification strategy to DID designs, addressing settings in which the parallel trends assumption is unlikely to hold or no credible control group exists.
+When empirical researchers rely on DID designs and run two-way fixed effects regressions, they often worry that the parallel trends assumption is implausible in practice.
+To address this concern, they often turn to TWFEIV regressions, exploiting the timing variation of a policy shock as an instrument for treatment (e.g., Miller and Segal (2019)).
 In this paper, we formalize the underlying identification strategy as instrumented difference-in-differences.
-Our DID-IV design is a natural extension of DID design, in that the timing variation is used to construct an instrument rather than a treatment.
-Second, in line with the DID literature, this paper develops a reliable estimation method in DID-IV designs in the presence of heterogeneous treatment effects.
-Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects.
-(Athey and Imbens (2022); Borusyak et al.
-(2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021)).
-Some of them have proposed a credible estimation method that delivers a sensible estimand and is robust to treatment heterogeneity.
-Along with the recent DID literature, this paper proposes an alternative estimation method to the TWFEIV regression, and illustrate its usefulness through an empirical application.
+Our DID-IV design can be viewed as a natural extension of DID designs, in that the timing variation is used to construct an instrument rather than a treatment.
+Second, in line with the DID literature, this paper develops a reliable estimation method for DID-IV designs in the presence of heterogeneous treatment effects.
+Recently, several studies have pointed out the issue of implementing DID designs via two-way fixed effects regressions, or its dynamic specifications under heterogeneous treatment effects (Athey and Imbens (2022); Borusyak et al.
+(2024); Callaway and Sant’Anna (2021); de Chaisemartin and D’Haultfœuille (2020); Goodman-Bacon (2021); Imai and Kim (2021); Sun and Abraham (2021); Callaway et al.
+(2024)).
+Some of these studies propose credible estimation methods that deliver a sensible estimand and are robust to treatment heterogeneity.
+In the same spirit as the recent DID literature, this paper proposes an alternative to the TWFEIV regression and illustrates its usefulness through an empirical application.
 2  DID-IV in two time periods
 In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings.
-We first establish the target parameter and the identifying assumptions in this design.
-We then investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel assumption in the outcome.
-Finally, we describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
+Specifically, we establish the target parameter and the identifying assumptions in this design.
+At the end of this section, we also discuss the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
 2.1  Set up
@@
 In contrast to the sharp assignment of the instrument, we allow the general adoption process of the treatment; that is, we assume the treatment path can take four values with non-zero probability: $\{(0,0),(0,1),(1,0),(1,1)\}\in\mathcal{S}(D)$.
-In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock that encourages people to adopt the treatment in period $t=1$.
+In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a policy shock that encourages people to adopt the treatment in period $t=1$.
 For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
@@
 We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment.
-Since the treatment and the instrument takes only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
+Since the treatment and the instrument take only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
 $D_{i,t}=\sum_{z\in\mathcal{S}(Z)}\mathbf{1}\{Z_{i}=z\}D_{i,t}(z),\hskip 8.53581ptY_{i,t}=\sum_{z\in\mathcal{S}(Z)}\sum_{d\in\mathcal{S}(D)}\mathbf{1}\{Z_{i}=z,D_{i}=d\}Y_{i,t}(d,z).$
@@
 de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)).
-[^4] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes, as it takes only four values.
-We therefore consider this case in online Appendix Section A.3 as an extension.
-One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatment, or multiple time periods, but it requires cumbersome notations and complicates the definition of the target parameter, being not practical for empirical applications.
 Next, we introduce the group variable $G_{i}^{Z}\equiv(D_{i,1}((0,0)),D_{i,1}((0,1)))$ that describes the type of unit $i$ according to the response of $D_{i,1}$ on the instrument path $z$.
@@
 $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}].$
-This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$.
-In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, and Hudson et al.
-(2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$.
+This parameter measures the treatment effects in period $1$ for units that belong to the exposed group ($E_{i}=1$) and are induced to receive the treatment by the instrument in that period.
+In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, while Hudson et al.
+(2017) define the local average treatment effect (LATE) in period t=1t=1—unconditional on $E_{i}$—as their target parameter.
 The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g.
@@
 (2022)).
-We define the LATET as the target parameter for two reasons.
-First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)).
+We focus on the LATET for two reasons.
+First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), J.
+and Vytlacil,Edward (2005)).
 Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
@@
 (2017) impose a similar assumption without introducing the instrument path.
-Given Assumption 1 and Assumption 2, we can write the observed outcomes $Y_{i,t}$ as
+Given Assumptions 1-2, the observed outcome $Y_{i,t}$ can be written as
 $Y_{i,t}=D_{i,t}Y_{i,t}(1)+(1-D_{i,t})Y_{i,t}(0).$
-Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes.
-For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
+Assumptions 1-2 also allow us to introduce the notions of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
 $Y_{i,t}(D_{i,t}(z))\equiv D_{i,t}(z)Y_{i,t}(1)+(1-D_{i,t}(z))Y_{i,t}(0).$
-Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes.
+Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as the unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as the exposed outcomes.
 [^6] Exposed and unexposed outcomes are not new concepts in econometrics.
-For example, under IV designs, we compare the expectation of the outcome with instrument and the one without instrument in the numerator of the wald estimand: $E[Y|Z=1]-E[Y|Z=0]$.
-This comparison corresponds to the one between the expectation of the exposed outcome and the expectation of the unexposed outcome; that is, $E[Y|Z=1]-E[Y|Z=0]=E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0]$.
+For example, in IV designs, the numerator of the Wald estimand compares expected outcomes across instrument values: E[Y|Z=1]E[Y|Z=0]E[Y|Z=1]-E[Y|Z=0].
+This difference can be written as E[Y(D(1))|Z=1]E[Y(D(0))|Z=0]E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0], which corresponds to a comparison between exposed and unexposed outcomes.
 Next, we make the following monotonicity assumption as in Imbens and Angrist (1994).
@@
 Assumption 4 (No anticipation in the first stage).
-[…]
+$D_{i,0}((0,1))=D_{i,0}((0,0))\hskip 5.69054pta.s.\hskip 8.53581pt\text{for all units $i$ with $E_{i}=1$}.$
 Assumption 4 requires that the potential treatment choice before the exposure to instrument is equal to the baseline treatment choice $D_{i,0}((0,0))$ in an exposed group.
@@
 $=$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
-Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.
-For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
-Note that Assumption 6 depends on the unexposed outcome, and does not depend on the potential outcomes directly.
-In Subsection 2.4, however, we show that the average time trends of the unexposed outcome in Assumption 6 can be written as the weighted average of the expectation of the potential outcomes, and this assumption can be interpreted as the restriction on the time effect between exposed and unexposed groups.
-When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6.
+Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.999de Chaisemartin et al.
+(2024) note that this assumption imposes restrictions on treatment effect heterogeneity when the standard parallel trends assumption holds between exposed and unexposed groups.
+For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 0 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
+When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumptions 5 and 6.
 For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
-Finally, we assume a relevance condition.
-This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined.
-[^8] Note that Assumption 3, 4, 5, and 7 imply that the share of compliers in period $1$ is non-zero.
-$E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]=Pr(CM^{Z}|E_{i}=1)>0.$ This ensures that the LATET in period $t=1$ is also well defined.
+Finally, we assume a relevance condition. This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined.
 Assumption 7 (Relevance condition).
@@
 See Appendix. ∎
-2.4   Interpreting the parallel trends assumption in the outcome
-In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs.
-It is less clear than that in DID designs because this assumption does not depend on the untreated outcome, but depends on the unexposed outcome.
-First, we introduce the additional notation.
-Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$ that represents the treatment adoption process during the two periods if the instrument path is $z=(0,0)$.
-Recall that in DID-IV designs, the units are allowed to take the treatment both period 00 and $1$ without instrument.
-In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods.
-Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
-Table 1: Exposed group (assigned to $z=(0,1)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=1$ $D_{1}((0,0))=0$
-$D_{0}((0,0))=1,D_{1}((0,1))=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
-$D_{0}((0,0))=1,D_{1}((0,1))=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
-$D_{0}((0,0))=0,D_{1}((0,1))=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
-$D_{0}((0,0))=0,D_{1}((0,1))=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
-Table 2: Unexposed group (assigned to $z=(0,0)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$ $D_{1}((0,1))=0$
-$D_{0}((0,0))=1,D_{1}((0,0))=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
-$D_{0}((0,0))=1,D_{1}((0,0))=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
-$D_{0}((0,0))=0,D_{1}((0,0))=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
-$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 3 (monotonicity assumption).
-By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs.
-[^9] In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1).
-We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$.
-In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2).
-If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time.
-If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument.
-Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
-We now consider the interpretation of the parallel trends assumption in the outcome.
-Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively.
-$\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
-$\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
-Note that $\Delta_{0}$ is estimable by the observed change in the outcome in an unexposed group during the two periods.
-By contrast, $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
-To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation.
-For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$.
-Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
-$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$ $\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
-$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$ $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
-We can interpret each $\Delta_{g^{t},e}$ as the effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$.
-First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
-$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},\hskip 5.69054pt\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
-This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods.
-Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
-$\Delta_{CM^{T},e}$ $=E[Y_{i,1}(0)-Y_{i,0}(0)|CM^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(1)-Y_{i,1}(0)|CM^{T},E_{i}=e]}_{\text{Selection gain among $CM^{T}$}},$
-$\Delta_{DF^{T},e}$ $=E[Y_{i,1}(1)-Y_{i,0}(1)|DF^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(0)-Y_{i,1}(1)|DF^{T},E_{i}=e]}_{\text{Selection gain among $DF^{T}$}}.$
-The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the average time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively.
-The second term represents the selection gain in each time group in group $E_{i}=e$, respectively.
-The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
-The following theorem establishes the relationship between $\Delta_{e}$ and $\Delta_{g^{t},e}$.
-Theorem 2.
-Suppose Assumptions 1-4 hold, then
-$\Delta_{e}=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e},$
-where the weight $w_{g^{t},e}$ is non-negative and sums to one; that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
-Proof.
-This directly follows from the Law of Iterated Expectation:
-$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
-$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
-$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
-$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
-$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
-Clearly, the weight $w_{g^{t},e}$ sums to one by construction. ∎
-Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ can be written as the weighted average of the expectation of the potential outcomes, and can be interpreted as measuring the convex average of the time effect for each time group $g^{t}$ in group $e$.
-The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$.
-Hereafter, we call $\Delta_{e}$ an expected time gain in group $e$.
-From Theorem 2, we can interpret that the parallel trends assumption in the outcome (Assumption 6), $\Delta_{0}=\Delta_{1}$, requires the same expected time gain between exposed and unexposed groups; that is, the effects of time on the outcome through treatment are the same between the two groups.
-Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the effect of the instrument on the outcome through treatment from that of time in an exposed group.
-The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
-Finally, we provide an empirical example to illustrate the interpretation of each notation and Assumption 6 in practice.
-Example.
-Suppose that the treatment $D_{i,t}\in\{0,1\}$ represents participation in a job training program, the outcome $Y_{i,t}$ represents real earnings, and the instrument $Z_{i,1}\in\{0,1\}$ represents a policy shock that encourages people to receive the treatment in the second period.
-[^10] Assume that the treatment is non-staggered and the treatment and the instrument are not randomized across units.
-In this setting, the instrumental variable strategy nor the canonical DID designs are directly applicable.
-In this setting, the time compliers $CM^{T}$ are the units who start adopting the program in the second period in the absence of the policy shock, and $w_{CM^{T},0}$ measures its share in an unexposed group.
-$\Delta_{CM^{T},0}$ represents the average time trends of real earnings among the time compliers in an unexposed group, and can be decomposed into the time effects on real earnings without policy shock, and the causal effects of program participation on real earnings in the second period, as their treatment adoption behavior is affected by time.
-Assumption 6 requires that the average time trends of real earnings are the same without policy shock between exposed and unexposed groups.
-This assumption can be interpreted as the restriction on the effect of time on real earnings through program participation between the two groups.
-2.5   Comparing DID-IV to Fuzzy DID
-de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting as in this article.
-[^11] The Wald-DID estimand considered in dCDH corresponds to ours in this article.
-In their empirical application, dCDH revisit Duflo (2001), in which the group variable $E_{i}$ ($G_{i}$ in dCDH) takes one if unit $i$ belongs to the region that are highly exposed to the policy shock.
-However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
-In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID in an overview of the detailed discussion provided in online Appendix Section E.
-The main difference between dCDH and this article lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.121212de Chaisemartin and D’Haultfœuille (2018) consider the same situation in repeated cross section settings.
-Therefore, we rewrite their target parameter by the notation used in this article.
+2.4  Comparing DID-IV with Fuzzy DID
+de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”) also investigate the identifying assumptions for the Wald-DID estimand to capture the causal effects under the same setting considered here.101010dCDH also consider the DID-IV setting as in this paper.
+This follows from two observations.
+First, the group variable $G$ is included in their treatment participation equation $D=\mathbf{1}\{V\geq v_{GT}\}$ (see Assumption 3 in dCDH), implying that $G$ plays the role of the instrument $Z$ in our framework.
+Second, dCDH also assume the sharp assignment of the instrument: in their treatment participation equation, they impose $v_{10}=v_{00}$ (see page $1019$ of dCDH).
+However, dCDH formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs.
+Moreover, dCDH point out that the Wald-DID estimand requires the stable treatment effect assumption to identify their target parameter.
+In this section, we clarify the differences between this paper and dCDH, and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
+The detailed discussions and proofs are provided in Online Appendix D.
+Since dCDH formalize Fuzzy DID designs under repeated cross section data, we first introduce notation for the repeated cross section setting.
+Let $Y$ and $D$ denote the outcome and the treatment, respectively.
+Let $Z\in\{0,1\}$ denote the group indicator, where $Z=1$ indicates the exposed group.
+Let $T\in\{0,1\}$ denote the time indicator.
+Let $Y(0),Y(1)$ and $D(0),D(1)$ denote the potential outcomes and the potential treatment choices, respectively.
+[^11] Here, we impose the no carryover assumption and the exclusion restriction on potential outcomes, and the no anticipation assumption on potential treatment choices, as in dCDH.
+Let $D_{t}(0)$ and $D_{t}(1)$ denote the potential treatment choices in period $T=t$ when the group indicator takes value $Z=0$ and $Z=1$, respectively.121212dCDH introduce the treatment participation equation $D=\mathbf{1}\{V\geq v_{ZT}\}$ and define the potential treatment choice in period $t$ as $D(t)=\mathbf{1}\{V\geq v_{Zt}\}$.
+In this paper, we instead adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to facilitate interpretation.
+In dCDH, the group variable $G$ plays the role of the instrument $Z$ in this paper.
+Accordingly, we relabel $G$ as $Z$ in the following discussion.
+The main difference between dCDH and this paper lies in the definition of the target parameter: we focus on the LATET, while dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
 Definition.
 The switcher local average treatment effect on the treated (SLATET) is
-$SLATET$ $\equiv E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,D_{i,1}((0,1))>D_{i,0}((0,1))]$
-$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,SW],$
-where the switcher (SW) is the units who become treated in period $1$.
-This parameter measures the treatment effects among the units who belong to an exposed group ($E_{i}=1$) and switch to receive the treatment in period $1$ (SW).
-Because the target parameter is different from each other, the identifying assumptions in Fuzzy DID designs are also different from those in $2\times 2$ DID-IV designs.
-Furthermore, dCDH argue that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose the alternative estimands that do not require the strong restriction on treatment effect heterogeneity.
-In online Appendix Section E, we first show that the identifying assumptions in Fuzzy DID designs ex ante impose strong restrictions on treatment adoption behavior across units and these restrictions are asymmetric between exposed and unexposed groups.
-Specifically, we examine how dCDH’s identifying assumptions exclude the types in the two groups and construct Tables 3-4, which are in line with Tables 1-2.131313Because dCDH consider repeated cross section settings, we can observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.
-Comparing Tables 1-2 with Tables 3-4, one can figure out the difference between dCDH’s restrictions and ours.
-In our DID-IV design, we exclude only the defiers $DF^{Z}$ in exposed and unexposed groups by a monotonicity assumption.
-By contrast, dCDH exclude the defiers $DF^{Z}$ and the type $NT^{Z}\land DF^{T}$ in an exposed group, and the defiers $DF^{Z}$, the time defiers $CM^{T}$ and the time compliers $DF^{T}$ in an unexposed group.
+$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
+$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
+where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
+This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers, SW).
+Because the target parameters differ, the identifying assumptions underlying Fuzzy DID designs also differ from those in DID-IV designs, though there are similarities as well.
+[^13] Specifically, dCDH also impose Assumptions 1-4 and 7 in this paper.
+The main differences are as follows.
+First, dCDH assume the stable treatment rate assumption in an unexposed group (see Assumption 22 in dCDH), whereas we assume the parallel trends assumption in the treatment (Assumption 5).
+Second, dCDH assume the treatment participation equation (see Assumption 33 in dCDH), which imposes the monotonicity with respect to time $T$ in addition to the monotonicity with respect to instrument $Z$ (which corresponds to Assumption 3 in this paper).
+Finally, while we impose a parallel trends (PT) assumption on unexposed outcomes, dCDH impose a PT assumption on untreated outcomes (see Assumption $4$ in dCDH).
+In Online Appendix D, we first examine how these differences imply different restrictions on treatment adoption behavior across the two designs.
+We begin by describing the heterogeneity in treatment adoption behavior under the setting considered here.
+Specifically, we introduce the group variable $G^{T}=(D_{0}(0),D_{1}(0))$, which characterizes treatment adoption over time when the instrument is $Z=0$ in the second period.
+We define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers.
+Using the group variables $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case) and $G^{T}$, we then partition units into eight types within each group, as summarized in Tables 1-2.
+Next, we investigate which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively.
+Specifically, Tables 12 and Tables 34 summarize all latent treatment adoption types under DID-IV and Fuzzy DID respectively, where the types painted in gray color are excluded by the identifying assumptions of each design.
+Here, in Table 3, we exclude the $DF^{Z}$ and the $DF^{T}$ by the monotonicity assumptions with respect to instrument $Z$ and time $T$, which are implied by the treatment participation equation (see Lemma 1 in Online Appendix D.2.3).
 Table 1: Exposed group ($z=1$)
 observed counterfactual
-$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
+[…]
 $D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
@@
 observed counterfactual
-$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
+[…]
 $D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
@@
 $D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by dCDH’s identifying assumptions.
-Next, under Tables 3-4, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters.
-The one parameter measures the treatment effects among the type $CM^{Z}\land AT^{T}$, who are the sub-population of the compliers.
-The other causal parameter measures the treatment effects among the time compliers $CM^{T}$.
-[^14] Intuitively, we cannot identify this parameter in DID-IV settings because the untreated potential outcomes in period $t=1$ among the time compliers are not observed in both groups (see Tables 3-4).
-This decomposition result has a concerning implication; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher, as this parameter is always contaminated by the treatment effects among the time compliers, who are affected by time but not affected by the instrument.
-Finally, we revisit dCDH’s argument regarding the use of the Wald-DID estimand, and show that this problem is less serious than one might expect, and does not arise once a researcher imposes Assumption 6 instead of the parallel trends assumption in the untreated outcome.
-First, we show that their argument is based on the parallel trends assumption in DID designs.
-Recall that in our DID-IV design, we do not make the parallel trends assumption in the untreated outcome, but make the parallel trends assumption in the unexposed outcome.
-Because some units are allowed to adopt the treatment without instrument in DID-IV settings, the former type of parallel trends assumption is generally not sufficient to capture the average time trend of the outcome in an unexposed group.
+Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by the monotonicity assumption with respect to instrument $Z$.
+Table 3: Exposed group ($z=1$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
+$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
+$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
+Table 4: Unexposed group ($z=0$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
+$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
+$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
+Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs. The types painted in gray color are excluded by the identifying assumptions in dCDH.
+By comparing Tables 12 with Tables 34, we obtain two implications.
+First, the restrictions imposed by Fuzzy DID designs are stronger than those imposed by DID-IV designs.
+Second, while the restrictions under DID-IV designs are symmetric across exposed and unexposed groups, those under Fuzzy DID designs are asymmetric.
+Building on Tables 34, we next clarify the difference in target parameter between the two designs.
+Specifically, we show that dCDH’s target parameter, the SLATET, can be expressed as a weighted average of two distinct causal parameters.
+One parameter captures the treatment effects for units of type $CM^{Z}\land NT^{T}$, a subpopulation of the compliers $CM^{Z}$, while the other captures the treatment effects among the type $AT^{Z}\land CM^{T}$ (case (i)).
+Importantly, this decomposition depends on the direction of the two monotonicity assumptions.
+If $DF^{Z}$ and $CM^{T}$ are excluded by the two monotonicity conditions in dCDH, the SLATET identifies the treatment effects for units of type $CM^{Z}\land NT^{T}$ (case (ii)).
+If $CM^{Z}$ and $DF^{T}$ are excluded, the SLATET identifies the treatment effects for units of type $AT^{Z}\land CM^{T}$ (case (iii)), as formally established in Theorem 6 in Appendix D.3.
+This decomposition result has several important implications.
+First, empirical researchers should specify the direction of the two monotonicity conditions ex ante if they wish to understand which latent type’s causal effect is identified by the SLATET.
+Second, the SLATET may fail to be the policy-relevant parameter (Heckman and Vytlacil (2001)), as it is generally contaminated by treatment effect for units whose treatment status is affected by time but not affected by the instrument (see cases (i) and (iii)).
+Finally, even when the two monotonicity conditions ensure that the SLATET is policy relevant, it identifies the treatment effects for a narrower population than the LATET (see case (ii)).
+Finally, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs.
+In DID-IV, we view the Wald–DID estimand as a natural estimand for identifying the LATET.
+By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption to identify the SLATET.141414dCDH also show that the Wald–DID estimand additionally requires homogeneity of the SLATET between the exposed and unexposed groups in order to identify the SLATET in the exposed group when the stable treatment rate condition in the unexposed group is not satisfied.
+In earlier work, Blundell and Dias (2009) make a similar argument in Subsection F.2 of Section III.
+In Remark 4 of Section D.3 in Online Appendix D, however, we show that this homogeneity condition cannot be straightforwardly interpreted as requiring homogeneous treatment effects between the two groups.
+We demonstrate that this difference arises because Fuzzy DID and DID-IV rely on different types of the PT assumption.
+Given this, we also discuss which PT assumption is more suitable for DID-IV settings.
+Specifically, we first show that the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH impose the PT assumption on untreated outcomes.
+Recall that in DID-IV settings, units are allowed to adopt the treatment without instrument during the two periods.
+As a result, the PT assumption on untreated outcomes is insufficient to capture the average time trends of the outcome even in the unexposed group.
 We show that this fact leads dCDH to impose the stable treatment effect assumption for the Wald-DID estimand to identify the SLATET.
-Next, we argue that the different type of the parallel trends assumption, Assumption 6, is suitable for DID-IV settings, and the parallel trends assumption in the untreated outcome would be difficult to justify in practice for three reasons.
-First, as already noted, this assumption is not sufficient to identify the average time trends of the outcome even in an unexposed group.
-Second, in most applications of the DID-IV method, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al.
+By contrast, this issue does not arise in DID-IV designs because we impose the PT assumption on unexposed outcomes, which is sufficient to capture the average time trends of the outcome for the unexposed group.
+Next, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings for three reasons.
+First, the PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes.
+This is because the latter relies on variation in treatment, while the former exploits variation in instrument.
+Second, in most applications of DID-IV methods, we cannot impute the untreated potential outcomes in general (e.g., Duflo (2001), Black et al.
 (2005)).
-For instance, in Duflo (2001), this assumption requires the data to contain units with zero education attainment, which is unrealistic in practice.
-Finally, in most of the DID-IV settings, we cannot test this assumption using pre-exposed period data, as some units can potentially adopt the treatment before period 00.
+For instance, in Duflo (2001), the PT assumption on untreated outcomes would require the data to include units with zero educational attainment, which is unrealistic in practice.
+Finally, in DID-IV settings, while the PT assumption on unexposed outcomes is indirectly testable (see Section 5.3 in this paper), the PT assumption on untreated outcomes is difficult to assess using pre-exposed period data, as some units may already adopt the treatment before period 0.
+We conclude this section by providing guidance for empirical researchers choosing between DID-IV and Fuzzy DID designs in practice.
+First, when the goal is to identify a policy-relevant parameter, DID-IV designs are more attractive.
+Under Fuzzy DID designs, the target parameter may not be policy relevant in general.
+Second, when researchers are interested in identifying the SLATET, they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed by Fuzzy DID designs.
+If these restrictions appear questionable, researchers may instead target the LATET and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument.
+Finally, if researchers wish to assess the plausibility of the PT assumption, DID-IV designs are more appealing because the PT assumption imposed under Fuzzy DID is generally difficult to test using the pre-exposed data.
 3  DID-IV in multiple time periods
@@
 Assumption 8 (Staggered adoption for $Z_{i,t}$).
-For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
-Assumption 8 requires that once units start getting exposed to the instrument, units remain exposed to that instrument.
+$Z_{i,1}=0$ for all $i$. For $s<t$, $Z_{i,s}\leq Z_{i,t}$ where $s,t\in\{1,\dots T\}$.
+Assumption 8 requires that no one is exposed to the instrument in time $t=1$ and once units start getting exposed to the instrument, units remain exposed to that instrument.
 In the DID literature, several studies make a similar assumption for the treatment and call it the “staggered treatment adoption” (e.g.
@@
 If unit $i$ is not exposed to the instrument for all time periods, we define $E_{i}=\infty$.
-Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{1,2,\dots,T,\infty\}$.
+Based on the initial exposure period $E_{i}$, one can uniquely partition units into mutually exclusive and exhaustive cohorts $e$ for $e\in\{2,3,\dots,T,\infty\}$.
+Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$.
+Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset.
+Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
 Similar to the two periods setting in Subsection 2.1, we allow the general adoption process for the treatment: the treatment can potentially turn on/off repeatedly over time.
@@
 Since the adoption date of the instrument uniquely pins down one’s instrument path, we can write the observed treatment status $D_{i,t}$ for unit $i$ in period $t$ as
-$D_{i,t}=D_{i,t}^{\infty}+\sum_{1\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
+$D_{i,t}=D_{i,t}^{\infty}+\sum_{2\leq e\leq T}(D_{i,t}^{e}-D_{i,t}^{\infty})\cdot\mathbf{1}\{E_{i}=e\}.$
 We define the effect of an instrument on treatment for unit $i$ in period $t$ as the difference between the observed treatment status to the never exposed treatment status: $D_{i,t}-D_{i,t}^{\infty}$.
@@
 Each CLATT can potentially vary across cohorts and over time because it depends on cohort $e$, relative period $l$, and the compliers $CM_{e,e+l}$.
-[^16] Note that the composition of the compliers can overlap over the post-exposure periods in cohort $e$.
-For instance, some units in cohort $e$ may belong to $CM_{e,e+l}$ for all the relative periods ($l\geq e$): they are affected by the instrument for all post-exposure periods.
-This parameter would be policy-relevant if the instrument is based on the policy change of interest to the researcher.
-For example, in a job training program setting considered in Subsection 2.5, each CLATT would be of particular interest if the instrument $Z_{i,t}$ represents an important policy lever for the researchers (e.g., nudge); the share of compliers in each relative period after the policy shock represents the dynamic effect of that policy change on the take-up rate of the program, and the corresponding CLATT can be interpreted as estimating the effect of participation in job training on real earnings among those affected.
 3.3  Identification assumptions in staggered DID-IV designs
-In this subsection, we establish the identification assumptions in staggered DID-IV designs. These assumptions are the natural generalization of Assumptions 1-6 in $2\times 2$ DID-IV designs.
+In this subsection, we establish the identification assumptions in staggered DID-IV designs.
+In staggered DID-IV designs, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
+$w^{DID}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=\infty])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])},$
+for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$.
+Note that this estimand is the Wald-DID estimand, where the pre-exposed period is e1e-1 and the control group is the never exposed cohort.
+Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
+We consider the following identification assumptions for each $w^{DID}_{e,l}$ to capture the $CLATT_{e,e+l}$.
 Assumption 10 (Exclusion restriction in multiple time periods).
-$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d)\hskip 8.53581pta.s.$
+$\forall z\in\mathcal{S}(Z),\forall d\in\mathcal{S}(D),\forall t\in\{1,\dots,T\},Y_{i,t}(d,z)=Y_{i,t}(d).$
 Assumption 10 extends the exclusion restriction in two time periods (Assumption 2) to multiple period settings.
@@
 Assumption 11 (Monotonicity assumption in multiple time periods).
-$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{S}(E_{i})\hskip 5.69054pt\text{and for all}\hskip 5.69054ptl\geq 0.$
+$Pr(D_{i,e+l}^{e}\geq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,e+l}^{e}\leq D_{i,e+l}^{\infty})=1\hskip 5.69054pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl\geq 0.$
 This assumption requires that the instrument path affects the treatment adoption behavior in a monotone way for all relative periods after the initial exposure date $E_{i}=e$.
 Recall that we define $D_{i,t}-D_{i,t}^{\infty}$ to be the effect of an instrument on treatment for unit $i$ in period $t$.
-Assumption 11 requires that the individual exposed effect in the first stage should be non-negative (or non-positive) during the periods after the initial exposure to the instrument for all $i$.
-This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e$ and all $t\geq e$.
-Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e$.
+Assumption 11 requires that the individual exposed effect in the first stage should be non-negative or non-positive during the periods after the initial exposure to the instrument for all $i$.
+This assumption implies that the group variable $G_{i,e,t}\equiv(D_{i,t}^{\infty},D_{i,t}^{e})$ can take three values with non-zero probability for all $e\in\mathcal{E}$ and all $t\geq e$.
+Hereafter, we consider the type of the monotonicity assumption that rules out the existence of the defiers $DF_{e,t}$ for all $t\geq e$ in any cohort $e\in\mathcal{E}$.
 Assumption 12 (No anticipation in the first stage).
-[…]
+$D_{i,e+l}^{e}=D_{i,e+l}^{\infty}\hskip 8.53581pt\text{for all}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and all}\hskip 5.69054ptl<0.$
 Assumption 12 requires that potential treatment choices in any $l$ period before the initial exposure to the instrument is equal to the never exposed treatment.
 This assumption is a natural generalization of Assumption 4 to multiple period settings and restricts the anticipatory behavior before the initial exposure to the instrument.
-Assumption 13 (Parallel trends assumption in the treatment in multiple time periods).
-[…]
-Assumption 13 is a parallel trends assumption in the treatment in multiple periods and cohorts.
-It requires that the trends of the treatment across cohorts would have followed the same path, on average, if there was no exposure to the instrument.
+Assumption 13 (Parallel trends assumption in the treatment based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
+$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=\infty].$
+Assumption 13 is a parallel trends assumption in the treatment based on the never exposed cohort.
+It requires that, in the absence of exposure to the instrument, the average evolution of the treatment would have followed the same path between cohort $e$ and the never exposed cohort $\infty$.
 Assumption 13 is analogous to that of Callaway and Sant’Anna (2021) and Sun and Abraham (2021) in DID designs; these studies impose the same type of parallel trends assumption on untreated potential outcomes.
-Assumption 14 (Parallel trends assumption in the outcome in multiple time periods).
-[…]
-Assumption 14 is a parallel trends assumption in the outcome with settings of multiple periods and cohorts.
-It requires that the expectation of the never exposed outcome across cohorts would have followed the same evolution if the instrument had not been assigned.
-From the discussion in Subsection 2.4, this assumption can be interpreted as requiring the same expected time gain across cohorts and over time; that is, the effects of time on the outcome through treatment are the same on average across cohorts and over time.
-Assumption 15 (Relevance condition in multiple time periods).
-$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]>0.$
-Assumption 15 is a relevance condition in multiple period settings.
-With a binary treatment, Assumption 11 and 15 implies that the fraction of the compliers should be non-negative for all $e$ and $l\geq 0$:
-$E[D_{i,e+l}-D^{\infty}_{i,e+l}|E_{i}=e]=Pr(CM_{e,l}|E_{i}=e)>0.$
-This condition ensures that each $CLATT_{e,l}$ is well defined.
+Assumption 14 (Parallel trends assumption in the outcome based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
+$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=\infty].$
+Assumption 14 is a parallel trends assumption in the outcome based on the never exposed cohort.
+It requires that, in the absence of exposure to the instrument, the expected evolution of the outcome under no exposure to the instrument would have been the same, on average, between cohort $e$ and the never exposed cohort $\infty$.
+Assumption 15 (Relevance condition based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
+$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|E_{i}=\infty])>0.$
+Assumption 15 is a relevance condition in multiple period settings, ensuring that the Wald-DID estimand $w^{DID}_{e,l}$ is well defined.
+The following theorem shows that under Assumptions 815, each Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
+Theorem 2.
+If Assumptions 815 hold, the Wald-DID estimand $w^{DID}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
+$w^{DID}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
+for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$.
+Proof.
+See Appendix. ∎
+In some applications, however, there may be no never exposed cohort, or its sample size may be too small to serve as a reliable control group.
+Moreover, the parallel trends assumptions based on the never exposed cohort may be viewed as less credible in certain empirical settings.
+In such cases, it is natural to instead use the not-yet-exposed cohorts as the control group.
+Accordingly, we consider the following alternative Wald-DID estimand:
+$w^{DID,ny}_{e,l}=\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-(E[Y_{i,e+l}-Y_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])},$
+for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
+In this Wald-DID estimand, the control group is the not-yet-exposed cohorts, that is, the cohorts not exposed to the instrument by time $t=e+l$.
+If we adopt the above estimand $w^{DID,ny}_{e,l}$, we can replace Assumptions 1315 with Assumptions 1618 below.
+Assumption 16 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
+$E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|E_{i}=e]=E[D_{i,t}^{\infty}-D_{i,t-1}^{\infty}|Z_{i,t}=0,E_{i,e}=0].$
+Assumption 17 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
+$E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|E_{i}=e]=E[Y_{i,t}(D_{i,t}^{\infty})-Y_{i,t-1}(D_{i,t-1}^{\infty})|Z_{i,t}=0,E_{i,e}=0].$
+Assumption 18 (Relevance condition based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
+$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-(E[D_{i,e+l}-D_{i,e-1}|Z_{i,e+l}=0,E_{i,e}=0])>0.$
+The following theorem shows that under Assumptions 812 and 1618, each Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$.
+Theorem 3.
+If Assumptions 812 and 1618 hold, the Wald-DID estimand $w^{DID,ny}_{e,l}$ identifies the corresponding $CLATT_{e,e+l}$:
+$w^{DID,ny}_{e,l}=E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}],$
+for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
+Proof.
+See Appendix. ∎
 4  Extensions
-This section contains extensions to non-binary, ordered treatments and repeated cross sections.
-It also includes the case that the treatment path is introduced in potential outcomes in $2\times 2$ DID-IV designs.
-For more details and proofs, see online Appendix Section A.
+This section contains extensions to non-binary, ordered treatments and repeated cross sections. For more details and proofs, see online Appendix Section A.
 Non-binary, ordered treatment
@@
 With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT.
-The identifying assumptions are the same as in Subsection 3.3.
+The estimand and the associated identifying assumptions are the same as in Subsection 3.3.
 Definition.
@@
 Online Appendix Section A.2 presents the identification assumptions in DID-IV designs under repeated cross section settings.
-Introducing the treatment path in $2\times 2$ DID-IV designs
-In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$, as we consider the non-staggered adoption of the treatment.
-We can extend our results to the case that we introduce the treatment path in potential outcomes with binary treatment.
-Online Appendix Section A.3 shows that the similar assumptions to Assumptions 1-6 are sufficient for the Wald-DID to capture the LATET-like parameter.
 5  Estimation and inference
@@
 We create the data sets for each $CLATT_{e,e+l}$.
-Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$).
-Researchers should carefully specify the choice set $U$ in practice.
-If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$).
-If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
+Each data set includes the units of time t=e1t=e-1 and $t=e+l$, who are either in cohort $e\in\mathcal{E}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$).
+If we adopt the parallel trends assumption based on a never-exposed cohort (Assumptions 1314), we can set $U=\{\infty\}$.
+In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$.
+If we adopt the parallel trends assumption based on not-yet-exposed cohorts (Assumptions 1617), we can set $U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$.
+In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
 Step 2.
 In each data set, we run the following IV regression and obtain the IV estimator $\hat{\beta}^{e,l}_{IV}$:
-$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t},$
+$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
+The first stage regression is:
+$D_{i,t}=\delta^{e,l}_{0}+\delta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\delta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\delta^{e,l}_{3}\mathbf{1}\{E_{i}=e\}\mathbf{1}\{T_{i}=e+l\}+\eta^{e,l}_{i,t}.$
 where the group indicator $\mathbf{1}\{E_{i}=e\}$ and the post-period indicator $\mathbf{1}\{T_{i}=e+l\}$ are the included instruments and the interaction of the two is the excluded instrument.
@@
 Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation.
-Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period.
 From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator.
@@
 Theorem 4.
-Suppose Assumptions 8-14 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
-[…]
-[…]
-Proof.
-See Appendix. ∎
+  • (i)
+    Suppose Assumptions 815 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
+    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
+  • (ii)
+    Suppose that Assumptions 812 and 1618 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
+    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
 From Theorem 4, we can also construct the standard error of the STS estimator, using the sample analogue of the asymptotic variance $V(\psi_{i,e,l})$.
@@
 The superscript IV is added to associate each parameter with staggered DID-IV designs.
-In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
+[…]
 $\theta_{sel}(\tilde{e})^{IV}=\sum_{t=\tilde{e}}^{T}\frac{CAET^{1}_{\tilde{e},t}}{\sum_{t=\tilde{e}}^{T}CAET^{1}_{\tilde{e},t}}CLATT_{\tilde{e},t}.$
@@
 $\hat{\theta}^{IV}=\sum_{e}\sum_{t=1}^{T}\hat{w}(e,t)\cdot\widehat{CLATT}_{e,t},$
-where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1.
+where $\hat{w}(e,t)$ is the sample analog of each w(e,t){w}(e,t) and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1.
 Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
 If the assumptions of Theorem 4 hold,
-[…]
+$\sqrt{n}(\hat{\theta}_{IV}-\theta_{IV})\xrightarrow{d}\mathcal{N}(0,V(l_{i}^{\theta_{IV}})),$
 where the influence function $l_{i}^{\theta_{IV}}$ takes the following form:
@@
 In this section, we describe the procedure of the pre-trends test in DID-IV designs to check the validity of the parallel trends assumption in the treatment and the outcome.
-Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
+Suppose that data are available for period t=1t=-1. Then, one can assess the plausibility of Assumption 5 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
 $E[D_{i,0}-D_{i,-1}|E_{i}=0]=E[D_{i,0}-D_{i,-1}|E_{i}=1]$
 $\iff$ $E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=0]=E[D_{i,0}((0,0))-D_{i,-1}((0,0))|E_{i}=1].$ (1)
-Similarly, one can assess the plausibility of Assumption 6 between period $t=-1$ and $t=0$ by testing the following null hypothesis
+Similarly, one can assess the plausibility of Assumption 6 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
 $E[Y_{i,0}-Y_{i,-1}|E_{i}=0]=E[Y_{i,0}-Y_{i,-1}|E_{i}=1]$
@@
 These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al.
-(2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumption 13 and 14 in pre-exposed periods.
+(2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumptions 1314 or Assumptions 1617 in pre-exposed periods.
 6  Application
@@
 The result is shown in Table 6.
-The TWFEIV estimate is $-0.009$ and not significantly different from 00.
+The TWFEIV estimate is 0.009-0.009 and not significantly different from 0.
 This indicates that, on the whole, the returns to schooling in the UK is nearly zero.
+[^20] Young (2024) revisits Oreopoulos (2006) and shows that the 2SLS estimates in this setting can be numerically unstable due to near collinearity among cohort indicators.
+Following Young (2024), we conduct stability checks based on variable and observation ordering, and find that our TWFEIV estimate does not suffer from such numerical instability.
 However, this may be a misleading conclusion.
@@
 Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland.
-We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons.
-We then run TSLS regression in each data set.
-The standard error is calculated by using the influence function shown in equation (B) in online Appendix Section LABEL:ApeC.1.
+We define North Ireland as an unexposed group $U$.
+We then run the STS regression in each data set.
+The standard error is calculated by using the influence function shown in equation (B) in Online Appendix Section B.
 Table 6: Returns to schooling in Oreopoulos (2006)
@@
 We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET).
-After establishing $2\times 2$ DID-IV designs, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
+After establishing $2\times 2$ DID-IV designs, we clarify the differences between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
 Next, we consider the DID-IV design in more than two periods with units being exposed to the instrument at different times.
@@
 Appendix
-Proof of Theorem 1.
+Proof of Theorem 1
 Proof.
 First we consider the numerator of the Wald-DID estimand. For the numerator, we have
-$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
-$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]$
-$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
-+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
-$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
-+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
-$=$ $E[(D_{1}((0,1))-D_{1}((0,0)))\cdot(Y_{1}(1)-Y_{1}(0))|E=1]$
-$=$ $LATET\cdot Pr(CM^{Z}|E=1).$ (5)
+$E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]$
+$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
+- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
+$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
+$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
+- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
+$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
+$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1]$
+- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
+$=$ $E[(D_{i,1}((0,1))-D_{i,1}((0,0)))\cdot(Y_{i,1}(1)-Y_{i,1}(0))|E_{i}=1]$
+$=$ $LATET\cdot Pr(CM^{Z}|E_{i}=1).$ (5)
 The first equality follows from Assumption 1 and Assumption 2.
@@
 For the denominator, we have
-$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
-$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
-$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
-+$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
-$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
-+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
-$=$ $Pr(CM^{Z}|E=1).$ (6)
-where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
+$E[D_{i,1}-D_{i,0}\mid E_{i}=1]-E[D_{i,1}-D_{i,0}\mid E_{i}=0]$
+$=$ $E[D_{i,1}((0,1))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]$
+$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
+$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
+$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
+$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
+$=$ $\Pr(CM^{Z}\mid E_{i}=1).$ (6)
+Here, the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
 Combining the result (5) with (6), we obtain the desirable result. ∎
-Proof of Theorem 2.
+Proof of Theorem 2
 Proof.
-First, we prove the consistency of our SLS estimator.
-We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined.
-More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$.
-This assumption guarantees that each Wald-DID estimator $\widehat{CLATT}_{e,e+l}$ is also well defined.
-First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and continuous mapping theorem, we have
-[…]
-Next, we show that the right hand side of (7) is equal to $CLATT_{e,e+l}$. For the numerator, we have
-$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
-$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$ (8)
-$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
-$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
-$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
-$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
-$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (9)
-The third equality follows from Assumption 9, 10 and 12. The fourth equality follows from Assumption 14 and the fifth follows from Assumption 11. For the denominator, by similar calculations,
-$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
-$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]$ (10)
-$-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
-$=Pr(CM_{e,e+l}|E_{i}=e),$ (11)
-where the final equality follows from Assumption 11 and Assumption 13. Combining (9) with (11), we obtain the desirable result.
-Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. We define $\theta_{e,l}$ to be
+Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. Note that Assumption 15 guarantees that the Wald-DID estimand $w_{e,l}^{DID}$ is well defined.
+First, we consider the numerator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
+$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=\infty]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{e}_{i,e-1})\mid E_{i}=e]$
+- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
+- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
+$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
+- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
+$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
+$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (7)
+The first equality follows from Assumptions 810. The second equality follows from Assumption 12. The third equality follows from Assumption 14. The final equality follows from Assumption 11.
+Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
+$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=\infty]$
+$=$ $E[D^{e}_{i,e+l}-D^{e}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
+- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
+$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (8)
+The first equality follows from Assumption 8. The second equality follows from Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
+Combining (7) with (8), we obtain the desirable result. ∎
+Proof of Theorem 3
+Proof.
+Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. Note that Assumption 18 guarantees that the Wald-DID estimand $w_{e,l}^{DID,ny}$ is well defined.
+First, we consider the numerator of $w_{e,l}^{DID,ny}$. From Assumptions 812 and Assumption 17, we have:
+$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
+- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
+$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
+- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
+$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
+$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (9)
+The first equality follows from Assumptions 810 and Assumption 12. The third equality follows from Assumption 17. The final equality follows from Assumption 11.
+Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 812 and Assumption 16, we have:
+$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
+- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
+$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
+$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (10)
+The first equality follows from Assumption 8 and Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
+Combining (9) with (10), we obtain the desirable result. ∎
+Proof of Theorem 4
+Proof.
+First, we show that $\widehat{CLATT}_{e,e+l}$ is consistent for the corresponding $CLATT_{e,e+l}$.
+Consider case (i), where a never-exposed cohort is used as the control group.
+By the Law of Large Numbers (LLN) and the Continuous Mapping Theorem (CMT), $\widehat{CLATT}_{e,e+l}\xrightarrow[p]{}w^{DID}_{e,l}$ as defined in Section 3.3.
+Then, by Theorem 2, we have $w^{DID}_{e,l}=CLATT_{e,e+l}$.
+The same argument, together with Theorem 3, establishes consistency under case (ii), where not-yet-exposed cohorts are used as the control group.
+Next, we prove that our STS estimator is asymptotically normal, deriving its influence function.
+This derivation does not depend on whether we are in case (i) or case (ii).
+We define $\theta_{e,l}$ to be
 $\theta_{e,l}$ $\equiv\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}(=CLATT_{e,e+l})$
@@
 $\psi_{i,e,l}$
-$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (12)
+$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg(\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U])\right]}{E[\mathbf{1}\{E_{i}\in U\}]}\Bigg),$ (11)
 where we define $\delta_{i,e,l}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l}\cdot(D_{i,e+l}-D_{i,e-1})$.
@@
 $V(\psi_{i,e,l})$ $=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg(\frac{E\big[\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big]}{E[\mathbf{1}\{E_{i}=e\}]^{2}}$
-$+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
+$+\frac{E\big[\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U]\right]^{2}\big]}{E[\mathbf{1}\{E_{i}\in U\}]^{2}}\Bigg).$
 Appendix A Extensions
@@
 $Y_{t}$ $=Z_{1}\cdot Y_{t}(D_{t}((0,1)))+(1-Z_{1})\cdot Y_{t}(D_{t}((0,0)))$
-=$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
+$=\left\{Z_{1}\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{1,j}^{t}-\lambda_{1,j+1}^{t})\right\}+\left\{(1-Z_{1})\cdot\sum_{j=0}^{J}Y_{t}(j)\cdot(\lambda_{0,j}^{t}-\lambda_{0,j+1}^{t})\right\}.$
 where we use Assumption 1 and 2.
@@
 $=\sum_{j=1}^{J}E[(Y_{1}(j)-Y_{1}(j-1)|D_{1}((0,1))\geq j>D_{1}((0,0)),E=1]$
-×$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (13)
+$\times Pr(D_{1}((0,1))\geq j>D_{1}((0,0))|E=1).$ (12)
 For the denominator of the Wald-DID estimand, by similar argument, we obtain:
@@
 $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1].$
-This follows from Assumption 4 and 5.
+This follows from Assumptions 45.
 Then, we have:
@@
 Assumption 20 (Monotonicity).
-$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1.$
+$Pr(D_{i,1}((0,1))\geq D_{i,1}((0,0)))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{i,1}((0,1))\leq D_{i,1}((0,0)))=1.$
+Here, we consider the monotonicity assumption that rules out the defiers, as our target parameter measures the treatment effects among the compliers.
 Assumption 21 (No anticipation in the first stage).
-[…]
+$D_{i,0}((0,1))=D_{i,0}((0,0))\hskip 5.69054pta.s.\hskip 8.53581pt\text{for all units $i$ with $E_{i}=1$}.$
 Assumption 22 (Parallel trends assumption in the treatment).
@@
 - $(E[D_{i}|E_{i}=0,T_{i}=1]-E[D_{i}|E_{i}=0,T_{i}=0])>0.$
-By similar arguments in the proof of Theorem 1 (Theorem 4), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period $1$ under Assumptions 14-19 (thus, we omit the proof for brevity).
+By similar arguments in the proof of Theorem 1 (Theorem 5), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period 11 under Assumptions 1924 (thus, we omit the proof for brevity).
 A.2.2 Multiple time periods
@@
 In addition, we adopt Assumption 19 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$.
-Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to.
-Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$.
+Let $E_{i}=\min\{t:Z_{i,t}=1\}\in\{2,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to.
+Let $E_{i,e}=\mathbf{1}\{E_{i}=e\}$ denote the binary indicator that takes one if unit $i$ belongs to cohort $e$.
+Let $\bar{e}=\max_{i=1,\dots,n}E_{i}$ denote the largest cohort value in the dataset.
+Let $\mathcal{E}=\mathcal{S}(E_{i})\setminus\{\bar{e}\}\subseteq\{2,3,\dots,T\}$ denote the support of $E_{i}$ excluding $\bar{e}$.
+Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote outcomes and potential treatment choices when unit $i$ belongs to cohort $e$.
 Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$.
 We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
-In multiple periods and cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
+In multiple periods and cohorts settings with a binary treatment, our target parameter is the $CLATT_{e,e+l}$ in period $T=e+l$:
 $CLATT_{e,e+l}$ $\equiv E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}>D_{i,e+l}^{\infty}]$
 $=E[Y_{i}(1)-Y_{i}(0)|T_{i}=e+l,E_{i}=e,CM_{e,e+l}].$
-When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 14, and the following parallel trends assumptions and relevance condition.
-Assumption 20 (Parallel trends assumption in the treatment in multiple time periods).
-$\text{For all}\hskip 2.84526pts<t,\hskip 2.84526ptE[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=s]\hskip 2.84526pt\text{is identical for all}\hskip 2.84526pte\in\mathcal{S}(E_{i}).$
-Assumption 21 (Parallel trends assumption in the outcome in multiple time periods).
-For $\text{For all}\hskip 5.69054pts<t,\hskip 5.69054ptE[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=s]$
-is identical for $\text{is identical for all}\hskip 5.69054pte\in\mathcal{S}(E_{i}).$
-Assumption 22 (Relevance condition in multiple time periods).
-$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i}-D^{\infty}_{i}|E_{i}=e,T_{i}=e+l]>0.$
-A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
-We discuss the DID-IV identification strategy in two periods and two groups with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$.
-In this setting, our target parameter is the LATET’ defined below.
-Definition.
-The local average treatment effect on the treated with the treatment path (LATET’) is
-$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
-where $\Delta_{0}$ and $\Delta_{1}$ are:
-$\Delta_{0}\equiv E[Y_{1}((1,1))-Y_{1}((1,0))|E_{i}=1,(CM^{Z}\land DF^{T})],$
-$\Delta_{1}\equiv E[Y_{1}((0,1))-Y_{1}((0,0))|E_{i}=1,(CM^{Z}\land NT^{T})].$
-The weights $w_{0}$ and $w_{1}$ are:
-$w_{0}=\frac{Pr((CM^{Z}\land DF^{T})|E_{i}=1)}{Pr((CM^{Z}\land DF^{T})|E_{i}=1)+Pr((CM^{Z}\land NT^{T})|E_{i}=1)},\hskip 5.69054ptw_{1}=1-w_{0}.$
-We consider the identifying assumptions for the Wald-DID estimand to identify the LATET’.
-First, we make the following no anticipation assumptions on the outcome instead of imposing Assumption 1 (No carryover assumption).
-Assumption 23 (No anticipation of the outcome among the $CM^{Z}\land NT^{T}$ in an exposed group).
-$\forall z\in\mathcal{S}(Z),Y_{i,0}((0,1),z)=Y_{i,0}((0,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land NT^{T}).$
-This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
-Assumption 24 (No anticipation of the outcome among the $CM^{Z}\land DF^{T}$ in an exposed group).
-$\forall z\in\mathcal{S}(Z),Y_{i,0}((1,1),z)=Y_{i,0}((1,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land DF^{T}).$
-This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
-We can rationalize Assumptions 23-24 by considering the similarity between these assumptions and the no anticipation assumption made in the recent DID literature; all of them restrict the anticipatory behavior for those who belong to an exposed group (treatment group) and are affected by a policy shock (the treatment in DID designs).
-Next, we make the parallel trends assumption in the outcome suitable for the set up considered here.
-We first define the similar notation as in Subsection 2.4.
-Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
-$\bar{\Delta}_{AT^{T},e}\equiv E[Y_{i,1}((1,1))-Y_{i,0}((1,1))|AT^{T},E_{i}=e],$
-$\bar{\Delta}_{NT^{T},e}\equiv E[Y_{i,1}((0,0))-Y_{i,0}((0,0))|NT^{T},E_{i}=e],$
-$\bar{\Delta}_{CM^{T},e}\equiv E[Y_{i,1}((0,1))-Y_{i,0}((0,1))|CM^{T},E_{i}=e],$
-$\bar{\Delta}_{DF^{T},e}\equiv E[Y_{i,1}((1,0))-Y_{i,0}((1,0))|DF^{T},E_{i}=e].$
-Let $\bar{\Delta}_{0}$ and $\bar{\Delta}_{1}$ denote the expected time gain in unexposed and exposed groups, respectively:
-$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},\hskip 5.69054pt\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
-where the definition of the weight $w_{g^{t},e}$ is same as in Subsection 2.4.
-As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the same expected time gain between exposed and unexposed groups.
-The following assumption makes such a restriction in the current set up.
-Assumption 25 (Parallel trend assumption in the outcome).
-The expected time gain between exposed and unexposed groups is the same: $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
-The theorem below shows that the Wald-DID estimand is equal to the LATET’ under Assumptions 2-5, 7 and Assumptions 23-25.
-Theorem 5.
-Suppose Assumptions 2-5, 7 and Assumptions 23-25 hold. Then, the Wald-DID estimand corresponds to the LATET’.
-Proof.
-First, we consider the denominator of the Wald-DID estimand. From the argument in the proof of Theorem 1, one can show the following result:
-$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]=Pr(CM^{Z}|E=1)$
-$=Pr(CM^{Z}\land DF^{T}|E=1)+Pr(CM^{Z}\land NT^{T}|E=1).$ (15)
-Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
-$E[Y_{1}-Y_{0}|E=0]=\bar{\Delta}_{0}.$ (16)
-This follows from Table 2 and the definition of the expected time gain $\bar{\Delta}_{0}$.
-Finally, we decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
-$E[Y_{1}-Y_{0}|E=1]=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
-$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((1,1))-Y_{0}((1,1))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
-$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
-$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
-$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
-$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
-$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
-$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=1,(NT^{Z}\land DF^{T})]Pr((NT^{Z}\land DF^{T})|E=1)$
-$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=1,(NT^{Z}\land NT^{T})]Pr((NT^{Z}\land NT^{T})|E=1)$
-$=\bar{\Delta}_{CM^{T},1}w_{CM^{T},1}+\bar{\Delta}_{AT^{T},1}w_{AT^{T},1}$
-$+\bar{\Delta}_{NT^{T},1}w_{NT^{T},1}+\bar{\Delta}_{DF^{T},1}w_{DF^{T},1}$
-$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$
-$=\bar{\Delta}_{1}$
-$+E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1)$ (17)
-The first equality follows from Table 1 and Assumption 2.
-The second equality follows from Assumption 23, Assumption 24 and simple manipulation.
-The third equality follows from Table 1.
-The final equality follows from the definition of the expected time gain $\bar{\Delta}_{1}$.
-From the result (16) with (17) and Assumption 25, we obtain
-$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
-$=E[Y_{1}((0,1))-Y_{1}((0,0))|E=1,(CM^{Z}\land NT^{T})]Pr((CM^{Z}\land NT^{T})|E=1)$
-$+E[Y_{1}((1,1))-Y_{1}((1,0))|E=1,(CM^{Z}\land DF^{T})]Pr((CM^{Z}\land DF^{T})|E=1).$ (18)
-Combining the result (18) with (15), we obtain the desired result. ∎
+If we have an ordered, non-binary treatment, our target parameter is the $CACRT_{e,e+l}$ in period $T=e+l$:
+$CACRT_{e,e+l}\equiv\sum_{j=1}^{J}w^{e}_{e+l,j}\cdot E[Y_{i,e+l}(j)-Y_{i,e+l}(j-1)|T_{i}=e+l,E_{i}=e,D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}]$
+where the weight $w^{e}_{e+l,j}$ is:
+$w^{e}_{e+l,j}=\frac{Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}{\sum_{j=1}^{J}Pr(D_{i,e+l}^{e}\geq j>D_{i,e+l}^{\infty}|T_{i}=e+l,E_{i}=e)}.$
+Similar to Section 3.3, we first consider the following estimand to identify each $CLATT_{e,e+l}$:
+$w^{DID,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,E=\infty}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}},\qquad\text{for }e\in\mathcal{E}\text{ and }l\in\{0,\dots,T-e\}.$
+where, for any $e\in\mathcal{S}(E_{i})$,
+$\Delta^{l}_{Y,E=e}$ $=E[Y_{i}\mid E_{i}=e,T_{i}=e+l]-E[Y_{i}\mid E_{i}=e,T_{i}=e-1],$
+$\Delta^{l}_{D,E=e}$ $=E[D_{i}\mid E_{i}=e,T_{i}=e+l]-E[D_{i}\mid E_{i}=e,T_{i}=e-1].$
+Note that in this Wald-DID estimand, the pre-exposed period is e1e-1 and the control group is the never exposed cohort.
+Here, we assume that the largest cohort value in the dataset is $\bar{e}=\infty$.
+When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 19, and the following parallel trends assumptions and the relevance condition.
+Assumption 25 (Parallel trends assumption in the treatment based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
+$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
+$E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=\infty,T_{i}=t-1].$
+Assumption 26 (Parallel trends assumption in the outcome based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptt\geq e,$
+$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
+$E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=\infty,T_{i}=t-1].$
+Assumption 27 (Relevance condition based on a never exposed cohort).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\},$
+$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,E=\infty}>0.$
+By similar arguments in the proof of Theorem 2, one can show that each Wald-DID estimand $w^{DID,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ under Assumptions 8, 11, 12 in Section 3, Assumption 19, and Assumptions 2527 (thus, we omit the proof for brevity).
+When there is no never-exposed cohort, or its sample size is too small, one can instead consider the following estimand:
+$w^{DID,ny,r}_{e,l}=\frac{\Delta^{l}_{Y,E=e}-\Delta^{l}_{Y,ny}}{\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}},$
+where
+$\Delta^{l}_{Y,ny}$ $=E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[Y_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1],$
+$\Delta^{l}_{D,ny}$ $=E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e+l]-E[D_{i}\mid Z_{i}=0,E_{i,e}=0,T_{i}=e-1].$
+for $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. In this Wald-DID estimand, the control group is the not-yet-exposed cohorts.
+If we adopt the above estimand $w^{DID,ny,r}_{e,l}$, we can replace Assumptions 2527 with Assumptions 2830 below.
+Assumption 28 (Parallel trends assumption in the treatment based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
+$E[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=t-1]$
+$E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[D_{i}^{\infty}|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
+Assumption 29 (Parallel trends assumption in the outcome based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054pt\text{each}\hskip 5.69054ptt\in\{2,\dots,T\}\hskip 5.69054pt\text{such that}\hskip 5.69054pte\leq t<\bar{e},$
+$E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t-1]$
+$E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|Z_{i}=0,E_{i,e}=0,T_{i}=t-1].$
+Assumption 30 (Relevance condition based on not-yet-exposed cohorts).
+$\text{For each}\hskip 5.69054pte\in\mathcal{E}\hskip 5.69054pt\text{and}\hskip 5.69054ptl\in\{0,\dots,T-e\}\hskip 5.69054pt\text{such that}\hskip 5.69054ptl<\bar{e}-e,$
+$\Delta^{l}_{D,E=e}-\Delta^{l}_{D,ny}>0.$
+Again, by similar arguments in the proof of Theorem 3, one can show that each Wald-DID estimand $w^{DID,ny,r}_{e,l}$ is equal to the $CLATT_{e,e+l}$ ($CACRT_{e,e+l}$ under an ordered, non-binary treatment) for each $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$ under Assumptions 8, 11, 12 in Section 3), 19, and Assumptions 2830 (thus, we omit the proof for brevity).
 Appendix B Deriving the influence function in repeated cross section settings
-[…]
+In this appendix, we present the influence function of our STS estimator when repeated cross section data are available.
+[^22] The derivation here is essentially the same as in de Chaisemartin and D’Haultfœuille (2018), who present the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups.
+Let $\theta_{e,l}^{r}$ define
 $\theta_{e,l}^{r}\equiv\frac{\alpha_{e,e+l}-\alpha_{e,e-1}-[\beta_{U,e+l}-\beta_{U,e-1}]}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]},$
@@
 In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$.
-Using similar arguments in the proof of Theorem 3, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 14, 20-22, and have the following influence function:
+Then, if we set $U=\{\infty\}$ ($U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$), one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 19 and Assumptions 2527 (Assumptions 2830), and have the following influence function:
 $\psi_{i,e,l}^{r}=\frac{1}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]}$
-×$\times\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
-$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg{)},$ (19)
+$\times\Bigg(\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$
+$-\frac{\mathbf{1}_{U,e+l}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e+l]\right]}{E[\mathbf{1}_{U,e+l}]}+\frac{\mathbf{1}_{U,e-1}\cdot\left[\delta^{r}_{i,e,l}-E[\delta^{r}_{i,e,l}|E_{i}\in U,T_{i}=e-1]\right]}{E[\mathbf{1}_{U,e-1}]}\Bigg),$ (14)
 where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true, and $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
 Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
-[…]
-[…]
+In this appendix, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006).
+Let $e=1947$ and $U=1957$ denote England and Northern Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$.
+Let $R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e\in\{1947,1957\}$) and cohort $t$.
+Here, $R_{e(i),t}$ indicates that the unit $i$ belongs to cohort $e$.
+First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^23] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024).
+In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
 $\hat{\beta}_{IV}$ $=\frac{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\sum_{t}N_{e,t}\hat{Z}_{e,t}\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]}$
@@
 $\widehat{WDID}_{e,t}\equiv\frac{\left[Y_{e,t}-Y_{e,1946}-(Y_{U,t}-Y_{U,1946})\right]}{\left[D_{e,t}-D_{e,1946}-(D_{U,t}-D_{U,1946})\right]},$
-and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and North Ireland fixed effects [^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions..
-[…]
-$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CLATT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (20)
+and $\hat{Z}_{e,t}$ is the residuals from regressing $Z_{i,t}$ on the cohort and Northern Ireland fixed effects.
+[^24] We can write $\hat{Z}_{i,t}=\hat{Z}_{e,t}$ because $Z_{i,t}$ only varies across cohorts and regions.
+Here, to clarify the interpretation of the TWFEIV estimand, we assume the parallel trends assumption in the treatment and the outcome over the entire sample period from $t=1936$ to $t=1965$.
+Then, by using the similar arguments in the proof of Theorem 5 for multiple time periods repeatedly, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
+$\beta_{IV}=$ $\sum\limits_{U-1\geq t\geq e}w_{e,t}^{1}\cdot CACRT_{e,t}+\sum\limits_{t\geq U}w_{e,t}^{2}\cdot\Delta_{e,t},$ (15)
 where we define:
 $CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e],$
-$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
+$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CACRT_{e,t}-CAET^{1}_{U,t}\cdot CACRT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
 and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
@@
 $w_{e,t}^{2}$ (18)
-$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (24)
-where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$ respectively.
-[…]
-Figure 3 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term).
-The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
-[…]
+$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (19)
+where $n_{e,t}$ and $E[\hat{Z}_{i,t}|E_{i}=e]$ represent the population share and residuals for England in cohort $t$, respectively.
+In Oreopoulos (2006), we can only identify $CACRT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in Northern Ireland in $1957$.
+This implies that each $\Delta_{e,t}$ in equation (15) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
+Figure 3 plots the weight and the corresponding estimate for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term).
+The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CACRT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
+Figure 3 shows that all the CACRTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted.
+This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
 Figure 3: Decomposition result for the TWFEIV estimand in Oreopoulos (2006).
-Notes: This figure plots the estimated weights and the corresponding estimates for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term).
-The closed squares are the weights and the corresponding estimates for each $CLATT_{e,t}$.
+Notes: This figure plots the estimated weights and the corresponding estimates for each $CACRT_{e,t}$ and $\Delta_{e,t}$ (bias term).
+The closed squares are the weights and the corresponding estimates for each $CACRT_{e,t}$.
 The closed triangles are the weights and the corresponding estimates for each $\Delta_{e,t}$ (bias term).
-Appendix D Relationship between DID and DID-IV
-In this appendix, we describe the relationship between DID and DID-IV designs, and establish that DID-IV is DID with noncompliance of the assigned treatment.
-Following Subsection 2.1, we introduce the notation in DID designs.
-Let $Z_{i}=(Z_{i,0},Z_{i,1})$ denote the path of a program that distributes the treatment $D_{i,t}$, and let $D_{i}=(D_{i,0},D_{i,1})$ denote the path of the treatment for unit $i$.
-In the canonical DID set up, $Z_{i}=(Z_{i,0},Z_{i,1})$ takes only two values: $Z_{i}=(0,1)$ if unit $i$ is assigned to the program in period $1$ (treatment group) and $Z_{i}=(0,0)$ if unit $i$ is not assigned to the program in period $1$ (control group).
-Let $E_{i}\in\{0,1\}$ denote the group variable: $E_{i}=1$ if unit $i$ belongs to the treatment group and $E_{i}=0$ if unit $i$ belongs to the control group.
-We refer to treatment and control groups as exposed and unexposed groups, respectively.
-Note that in the common DID set up, the treatment status $D_{i,t}$ is equal to the exposure to the program $Z_{i,t}$ because we ex ante impose two restrictions on the treatment adoption behavior across units.
-First, we assume that all units are not allowed to receive the treatment in period 00.
-$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (25)
-Second, we assume that unit $i$ should receive the treatment in period $1$ if that unit belongs to an exposed group, and that it should not receive the treatment in period $1$ if that unit belongs to an unexposed group.
-$1=D_{i,1}((0,1))>D_{i,1}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (26)
-These restrictions imply that we have Tables 5-6 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
-Then, under these restrictions, we can show that DID-IV designs nest DID designs as a special case along with three dimensions.
-First, Assumptions 3-5 in DID-IV designs are automatically satisfied in DID designs.
-Assumption 3 follows from (26) and Assumption 4 follows from (25).
-Assumption 5 follows from combining (25) with (26).
-Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 5-6:
-Table 5: Exposed group (assigned to $z=(0,1)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
-$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
-Table 6: Treatment group (assigned to $z=(0,0)$)
-observed counterfactual
-$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
-$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
-$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
-$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
-$=LATET.$
-Note that both ATT and LATET measure the treatment effects, for those who are in an exposed group (assigned to $z=(0,1)$) and are induced to treatment by instrument ($CM^{Z}$).
-Therefore, we define the LATET as our target parameter in DID-IV designs in Subsection 2.2.
-Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups.
-From Tables 5-6, we can rewrite the parallel trends assumption in the untreated outcome as:
-$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
-$\iff$ $E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0,NT^{T}]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1,NT^{T}]$
-$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
-$\iff$ $\Delta_{0}=\Delta_{1},$
-where the last equivalence holds because we have only the time never takers $NT^{T}$ in exposed and unexposed groups respectively under DID designs.
-$w_{NT^{T},0}=Pr(NT^{T}|E_{i}=e)=1,\hskip 8.53581pt\text{for all}\hskip 8.53581pte=0,1.$
-[…]
-Appendix E Comparing DID-IV to Fuzzy DID
-[…]
-E.1   Comparing the identifying assumptions in dCDH and this article
-In this subsection, we investigate the detailed connections between dCDH’s identifying assumptions and ours.
-Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-19.
-For now, we ignore the difference in the definition of target parameter.
-dCDH consider the following identification assumptions for the Wald-DID estimand to capture the SLATET.
-[^25] In repeated cross section settings, the formal definition is provided in Assumption 31.
-Note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation.
-[^26] This follows from two observations in dCDH.
-First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively.
-Second, the group variable $G$ is included in Assumption 28 (Treatment participation equation).
-We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively.
-[^27] Note that in Appendix A.2, we distinguish the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage).
-For the detailed explanations for each assumption, see dCDH.
+Appendix D Comparing DID-IV with Fuzzy DID
+In this section, we compare the DID-IV with the Fuzzy DID proposed in de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”).
+Although both papers consider the same empirical setting and the same Wald-DID estimand, they differ in their target parameters and associated identifying assumptions.
+This section clarifies the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
+D.1  Similarities and differences between this paper and dCDH
+In this subsection, we first clarify the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter, the estimand, and the identifying assumptions.
+This subsection provides the foundation for the subsequent comparison.
+Empirical setting
+In this paper, we mainly consider a two-period, two-group setting in which one group becomes exposed to the instrument in the second period (the exposed group), while the other group is never exposed to the instrument in either period (the unexposed group).
+The key features of this setting are that (i) the instrument is assigned only in the second period (sharp assignment of the instrument), and (ii) the group indicator captures exposure to the instrument.
+We refer to this setup as the “DID-IV” setting.
+Under this definition, dCDH also consider the DID-IV setting.
+In particular, the group variable $G$ in dCDH plays the role of the instrument $Z$ in this paper.
+This follows from the fact that $G$ is included in their treatment participation equation:
+$D=\mathbf{1}\{V\geq v_{GT}\}$
+(see Assumption 3 in dCDH), which corresponds to condition (ii) of the DID-IV setting.
+Condition (i) is also assumed in dCDH: they impose $v_{10}=v_{00}$ throughout their proofs (see page 1019 of dCDH), implying that the treatment participation thresholds at time $T=0$ are identical across groups $G=0$ and $G=1$.
+Here, it is useful to clarify the distinction between the empirical setting and the design in dCDH.
+In dCDH, the Fuzzy DID design is described as a situation in which the treatment rate in one group increases more than that in another group (see Assumption 1 in dCDH).
+However, this characterization corresponds to an identifying assumption, rather than a primitive feature of the empirical setting.
+Independently of Assumption $1$, dCDH also define the Fuzzy DID setting as $D\neq G\times T$.
+This definition corresponds to the DID-IV setting once combined with the restriction $v_{10}=v_{00}$.
+Based on the above discussion, in what follows, we replace $G$ with $Z$ and refer to the groups $Z=1$ and $Z=0$ as the exposed and unexposed groups, respectively.
+The target parameter
+In this paper, we consider the LATET as our target parameter. By contrast, dCDH focus on the switcher local average treatment effect on the treated (SLATET) defined below.
+Definition.
+The switcher local average treatment effect on the treated (SLATET) is
+$SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
+$=E[Y(1)-Y(0)|Z=1,T=1,SW],$
+where we use the restriction $v_{10}=v_{00}$ in dCDH, which implies $D_{0}(1)=D_{0}(0)$.
+Here, to ease interpretation, we slightly modify the notation in dCDH.
+Specifically, we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH.
+This parameter measures the treatment effects for units who belong to the exposed group ($Z=1$) and switch into treatment at time $T=1$ (switchers).
+In subsection D.3, we clarify the differences between the SLATET and the LATET considered in this paper.
+The estimand
+Both papers consider the same Wald-DID estimand.
+However, the stance regarding the use of the Wald-DID estimand differs between this paper and dCDH.
+Specifically, this paper views the Wald-DID estimand as the natural estimand for identifying the LATET.
+By contrast, dCDH point out that the Wald-DID estimand identifies the SLATET only if the treatment effect is stable over time, and propose alternative estimands that avoid this restriction.
+We return to this point in subsection D.4.
+The identifying assumptions
+dCDH impose the following identifying assumptions for the Wald-DID estimand to identify the SLATET. For detailed discussions of each assumption, see dCDH.
 Assumption 31 (Fuzzy design).
@@
 Assumption 32 (Stable percentage of treated units in an unexposed group).
-$1\displaystyle 0<E[D|Z=0,T=1]=E[D|Z=0,T=0]<$ 1 .
+$0<E[D|Z=0,T=1]=E[D|Z=0,T=0]<1.$
 Assumption 33 (Treatment participation equation).
-[…]
+$D=\mathbf{1}\{V\geq v_{ZT}\},\hskip 2.84526pt\text{with}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
 Assumption 34 (Parallel trends assumption in the untreated outcome).
@@
 $E[Y(1)-Y(0)|Z,T=1,D_{0}(Z)=1]=E[Y(1)-Y(0)|Z,T=0,D_{0}(Z)=1].$
-In Assumption 30, we modify the dCDH’s notation $D(t)$ to ease its interpretation: we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under time $T=t$ in group $Z$ instead of defining it as $D(t)$.
-[…]
-Finally, dCDH implicitly impose Assumption 14 and Assumption 16 when they formalize potential outcomes and treatment participation equations.
-Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14.
-As one can see in Assumption 28, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
+Again, in Assumption 35, we adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$, rather than $D(t)$ used in dCDH, to facilitate interpretation.
+We now state the detailed connections between the identifying assumptions in dCDH and those in this paper.
+Because dCDH consider a repeated cross-section setting, when comparing the identifying assumptions across the two papers, we refer to Assumptions 19-24.
+First, both dCDH and this paper impose condition (21) in Assumption 31, which corresponds to the standard relevance condition.
+Next, both dCDH and this paper impose Assumption 20.
+Specifically, Assumption 33 (treatment participation equation) in dCDH implies Assumption 20 in this paper (Vytlacil (2002)).
+Moreover, dCDH implicitly impose Assumptions 19 and 21 when formalizing potential outcomes and treatment participation.
+Specifically, dCDH do not introduce instrument paths in their framework and define potential outcomes as $Y(D)$, which implies Assumption 19.
+In addition, as shown in Assumption 33, dCDH assume that the treatment participation equation depends only on the current instrument status.
+Combined with sharp assignment of the instrument ($v_{10}=v_{00}$), this implies Assumption 21.
 The differences between dCDH’s assumptions and ours are as follows.
-First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 29), whereas we assume the parallel trends assumption in the unexposed outcome.
-Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 27), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups.
-Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 28, whereas we do not impose such restriction.
-Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 30) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET.
-Finally, dCDH assume the condition (27) in Assumption 26, which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
-Remark E.1.1.
+First, dCDH impose the parallel trends assumption in untreated outcomes (Assumption 34), whereas we impose the parallel trends assumption in unexposed outcomes.
+Second, dCDH assume stable treatment rates in the unexposed group (Assumption 32), whereas we assume the parallel trends in treatment between exposed and unexposed groups.
+Third, dCDH impose an independence assumption between unobserved heterogeneity $V$ and time $T$ conditional on $Z$ in Assumption 33, whereas we do not impose such a restriction.
+Fourth, the treatment participation equation in Assumption 33 is stronger than Assumption 20 in this paper.
+In particular, Assumption 33 imposes the monotonicity with respect to time $T$, in addition to the monotonicity with respect to instrument $Z$ (see Lemma 1 in Subsection D.2).
+Finally, dCDH impose condition (20) in Assumption 31, which requires that the treatment rate in the exposed group increases between the two periods, whereas we do not impose this condition.
+In the next subsection, we examine how these differences imply different restrictions on treatment adoption behavior across the two designs.
+Remark 3.
 When Assumption 32 is violated, dCDH consider the following assumption for the Wald-DID estimand to capture the SLATET.
@@
 $SLATET=SLATET^{\prime},$
-where we define282828dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
+where we define252525dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively.
 When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
-$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)],$
-$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)].$
+$SLATET$ $=E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)],$
+$SLATET^{\prime}$ $=E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)].$
 This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption.
-In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups.
-In the following discussion, we therefore treat Assumptions 26-30 as dCDH’s main identifying assumptions.
-E.2   Strong restrictions under Fuzzy DID designs
-In this subsection, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units.
-First, we prove some lemmas implied by dCDH’s assumptions.
-Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs.
-These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group.
-At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours.
-First, we show that Assumption 28 implies the following assumption.
-Assumption 32 (Monotone change in treatment in exposed and unexposed groups).
-For all $z\in\{0,1\}$,
+In Appendix D.3, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups.
+In the following discussion, we therefore treat Assumptions 31-35 as dCDH’s main identifying assumptions.
+D.2  Restrictions on treatment adoption behavior
+In this subsection, we compare the restrictions on treatment adoption behavior between the two designs.
+Since both this paper and dCDH impose Assumption 21 (see the previous section), we adopt the notation $(D_{0}(0),D_{1}(0),D_{1}(1))$ for potential treatment choices in the following.
+Note that while we do not use $D_{0}(1)$ in DID-IV designs due to the sharp assignment of the instrument, we can define $D_{0}(1)=\mathbf{1}\{V\geq v_{10}\}$ in dCDH’s notation.
+However, since dCDH assume $v_{10}=v_{00}$, we have $D_{0}(1)=D_{0}(0)$ almost surely.
+We use this fact repeatedly in the following discussion.
+D.2.1 Preparation
+We first describe the heterogeneity in treatment adoption behavior under the DID-IV setting.
+This allows us to clarify which latent adoption types are ruled out by the identifying assumptions in Fuzzy DID, and how these restrictions differ from those imposed in DID-IV.
+To do so, we introduce the following notation.
+Let $G^{T}=(D_{0}(0),D_{1}(0))$ denote the group variable that represents the treatment adoption process across the two periods when the instrument is $Z=0$ in the second period.
+That is, $G^{T}$ captures whether an individual would start, stop, or never adopt treatment over time in the absence of the instrument.
+Analogous to $G^{Z}=(D_{1}(0),D_{1}(1))$ (which was introduced in Section 2.1 for the panel data case), we define $G^{T}=(0,0)\equiv NT^{T}$ as time never-takers, $G^{T}=(1,1)\equiv AT^{T}$ as time always-takers, $G^{T}=(0,1)\equiv CM^{T}$ as time compliers, and $G^{T}=(1,0)\equiv DF^{T}$ as time defiers.
+Using the group variables $G^{Z}$ and $G^{T}$, we can partition units into eight types within each group.
+Specifically, we obtain the following tables for the exposed and unexposed groups, respectively.
+Table 7: Exposed group ($z=1$)
+observed counterfactual
 […]
-holds.
-Here, recall that we define $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to be the treatment status under $T=t$ in group $Z$.
+$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
+$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
+$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
+Table 8: Unexposed group ($z=0$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
+$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
+$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
+Notes: These tables represent mutually exclusive and exhaustive types under the DID-IV setting.
+In the following, we examine which types are excluded by the identifying assumptions under DID-IV and Fuzzy DID, respectively.
+D.2.2 DID-IV
+In DID-IV designs, the only restriction on treatment adoption behavior is the monotonicity assumption (Assumption 20), which rules out either $DF^{Z}$ or $CM^{Z}$.
+Therefore, if we rule out $DF^{Z}$, we obtain the following table.
+Table 9: Exposed group ($z=1$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
+$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
+$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
+Table 10: Unexposed group ($z=0$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
+$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
+$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
+Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 20.
+D.2.3 Fuzzy DID
+Next, we consider the treatment adoption behavior under Fuzzy DID.
+We first prove some lemmas implied by dCDH’s identifying assumptions.
+Based on these lemmas, we then construct tables that summarize all latent types in the exposed and unexposed groups under Fuzzy DID.
+We begin with Lemma 1 below.
 Lemma 1.
-Assumption 28 implies Assumption 32.
+Assumption 33 implies the following two monotonicity conditions:
+Assumption 37 (IV monotonicity at time $T=1$).
+$Pr(D_{1}(1)\geq D_{1}(0))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{1}(1)\leq D_{1}(0))=1.$
+Assumption 38 (Time monotonicity for each $z\in\{0,1\}$).
+$Pr(D_{0}(z)\geq D_{1}(z))=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(z)\leq D_{1}(z))=1$
+for each $z\in\{0,1\}$.
 Proof.
-Fix $z\in\{0,1\}$. Then we have $v_{z1}\leq v_{z0}$ or $v_{z1}\geq v_{z0}$.
-Therefore, conditioning $Z=z$, we have
-[…]
-[…]
-because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$. ∎
-Lemma 1 shows that under Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively.
-Note that the treatment participation equation itself does not imply Assumption 32: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption.
-Indeed, in this article, we do not assume this restriction.
-As we have seen in Subsection 2.4, time also plays the role of instrument in DID-IV settings.
-From this observation, we can view that in an unexposed group ($z=0$), Assumption 32 corresponds to a monotonicity assumption with respect to time, formally defined below.
-Assumption 33 (Monotonicity assumption with respect to time in an unexposed group).
-[…]
-Assumption 33 requires that time affects the treatment adoption process from period 00 to period $1$ in a monotone (uniform) way in an unexposed group $z=0$.
-In terms of the notation introduced in Subsection 2.4, this assumption implies that the group variable $G^{T}$ can take three values with non-zero probability.
-Hereafter, we consider the type of Assumption 33 which rules out the existence of the time-defiers $DF^{T}$ in an unexposed group.
-Next, we show that dCDH’s assumptions imply that there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
+We show that Assumption 33 implies Assumption 37.
+Note that we should have $v_{11}\geq v_{01}$ or $v_{11}\leq v_{01}$.
+Without loss of generality, suppose $v_{11}\geq v_{01}$.
+Then, we obtain $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}\leq D_{1}(0)=\mathbf{1}\{V\geq v_{01}\}$ almost surely.
+By comparing $v_{z0}$ and $v_{z1}$ for each $z\in\{0,1\}$, we can also show that Assumption 33 implies Assumption 38. ∎
+Next, we prove Lemma 2 below.
 Lemma 2.
-Suppose Assumption 27, Assumption 28 and Assumption 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group ($z=0$).
+Suppose Assumptions 32 and 33 hold. Then, there are no time compliers $CM^{T}$ and time defiers $DF^{T}$ in an unexposed group $Z=0$.
 Proof.
-Without loss of generality, suppose that Assumption 33 excludes the time defiers $DF^{T}$ in an unexposed group ($z=0$).
+Note that Assumption 33 also implies the time monotonicity given $Z=0$:
+$Pr(D_{0}(0)\geq D_{1}(0)|Z=0)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(0)\leq D_{1}(0)|Z=0)=1.$
+This implies that at most one of the two switching types ($CM^{T}$ or $DF^{T}$) can exist in an unexposed group $Z=0$.
+Without loss of generality, suppose that $DF^{T}$ is ruled out in the unexposed group.
 Then, Assumption 32 implies:
@@
 $\iff$ $Pr(CM^{T}|Z=0)=0$ (22)
-The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28. Equation (29) implies that there are no time compliers $CM^{T}$ in an unexposed group ($z=0$). ∎
-Finally, we show that the condition (27) in Assumption 26 (Fuzzy design) and Assumption 28 imply that in an exposed group ($z=1$), there are no units switching from treatment to non-treatment between period 00 and period $1$.
-In other words, in Assumption 32, we should have:
-$Pr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
+The first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33. Equation (22) shows that there are no time compliers $CM^{T}$ in an unexposed group $Z=0$. ∎
+Next, we prove Lemma 3 below.
+This lemma shows that in an exposed group $Z=1$, there are no units switching from treatment to non-treatment between period 0 and period $1$.
+This lemma also implies that the condition (20) in Assumption 31 and Assumption 33 ensures the existence of the switchers in an exposed group (that is, the SLATET is well defined).
 Lemma 3.
-The condition (27) in Assumption 26 and Assumption 28 imply $Pr(D_{0}(1)\leq D_{1}(1)|Z=1)$=1.
+The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1 and $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
 Proof.
+Note that Assumption 33 also implies the time monotonicity given $Z=1$:
+$Pr(D_{0}(1)\geq D_{1}(1)|Z=1)=1\hskip 5.69054pt\text{or}\hskip 5.69054ptPr(D_{0}(1)\leq D_{1}(1)|Z=1)=1.$
+In what follows, we replace $D_{0}(1)$ with $D_{0}(0)$.
 From the condition (20) in Assumption 31, we have
@@
 $\iff$ $Pr(D_{1}(1)>D_{0}(0)|Z=1)-Pr(D_{0}(0)>D_{1}(1)|Z=1)>0,$ (23)
-where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28.
-Recall that Assumption 28 implies Assumption 32 from Lemma 1.
+where the first equivalence follows from $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 33 and $D_{0}(1)=D_{0}(0)$.
 Suppose that $Pr(D_{0}(0)\geq D_{1}(1)|Z=1)=1$ holds.
@@
 This contradicts with the condition (23).
+Therefore, we should have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)=1$.
+From the condition (23), we also have $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
+∎
+Note that this lemma also implies that we cannot impose Assumptions 37 and 38 that rule out both $CM^{Z}$ and $CM^{T}$, which contradicts with $Pr(D_{0}(0)<D_{1}(1)|Z=1)>0$.
+Finally, we prove Lemma 4 below.
+Lemma 4.
+The condition (20) in Assumption 31 and Assumption 33 imply $Pr(D_{0}(0)\leq D_{1}(1)|Z=0)$=1. Equivalently, the type $(DF^{Z}\land AT^{T})$ is ruled out in an unexposed group.
+Proof.
+From Lemma 3, we have $Pr(D_{0}(0)\leq D_{1}(1)|Z=1)$=1.
+Under the threshold representation $D_{0}(0)=\mathbf{1}\{V\geq v_{00}\}$ and $D_{1}(1)=\mathbf{1}\{V\geq v_{11}\}$, this implies $v_{11}\leq v_{00}$.
+Hence, $D_{0}(0)\leq D_{1}(1)$ holds almost surely, and therefore $Pr(D_{0}(0)\leq D_{1}(1)\mid Z=0)=1$ as well.
 ∎
 We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively.
-Tables 3-4 in Subsection 2.5 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
-Comparing Tables 3-4 to Tables 1-2, one can see the difference between dCDH’s restrictions and ours [^29] Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings.
-The only modification is that as in Tables 3-4, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
-The distinction between dCDH’s restrictions and ours appears to be important in many empirical settings.
-Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions might be difficult to defend in some applications, and the practitioners may want to avoid imposing these restrictions in practice.
-For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time.
-Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
-E.3   Decomposing the SLATET
-In this subsection, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs.
-The following theorem presents the decomposition result for the SLATET, and clarifies the difference between dCDH’s target parameter (SLATET) and ours (LATET).
-Theorem 6.
-Suppose Assumptions 26-30 hold. Then, we can decompose the SLATET into two different causal parameters:
-$SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{1}(1)>D_{0}(0)]$
-$=E[Y(1)-Y(0)|Z=1,T=1,CM^{T}]Pr(CM^{T}|Z=1,SW)$
-$+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW),$ (31)
-where the switcher (SW) in an exposed group ($Z=1$) is the units who become treated in time $T=1$.
-Proof.
-The proof directly follows from Table 3 in an exposed group. ∎
-Theorem 6 shows that the SLATET can be decomposed into a weighted average of two different causal parameters under Fuzzy DID designs: one parameter measures the treatment effects among the time compliers $CM^{T}$ in an exposed group and the other measures the treatment effects among the type $(CM^{Z}\land NT^{T})$ in an exposed group.
-This theorem has a concerning implications; the SLATET may not be policy-relevant even when the instrument is based on the policy change of interest to the researcher.
-Suppose that the instrument represents an important policy lever for researchers, and they are intrinsically interested in the treatment effects among the compliers $CM^{Z}$ in an exposed group.
-In this setting, our target parameter, the LATET, is a policy-relevant by construction.
-However, the SLATET may not be policy-relevant because this parameter captures the treatment effects among the sub-population of the compliers and the time compliers, who are affected by time but not affected by the instrument.
-As the share of the time compliers $CM^{T}$ increases, the SLATET puts more weight on the treatment effects among the time compliers, which contaminates the interpretation of the magnitude of this paramter even in such policy-relevant contexts.
-Using the decomposition result for the SLATET, we finally clarify the interpretation of Assumption 31.
-We show that this assumption cannot be interpreted as requiring the homogeneous treatment effect among the same sub-population between exposed and unexposed groups.
-If we do not assume Assumption 27, we have Table 5 in an unexposed group.
-[^30] Here, we assume that Assumption 28 and 33 exclude the defiers $DF^{Z}$ and the time defiers $DF^{T}$ in an unexposed group, respectively.
+Tables 11-12 show the heterogeneity in treatment adoption behavior under Fuzzy DID.
+Here, $DF^{Z}$ and $DF^{T}$ are excluded by Assumptions 37 and 38 (under $z=0$), respectively.
+Since these restrictions depend on the direction of the two monotonicity conditions, we paint these types by blue.
+$DF^{Z}\land AT^{T}$ and $NT^{Z}\land DF^{T}$ in an exposed group are excluded by Lemma 3 and $DF^{T}$, $CM^{T}$ and $DF^{Z}\land AT^{T}$ in an unexposed group are excluded by Lemma 2 and Lemma 4.
+Since these restrictions do not depend on the direction of the two monotonicity conditions, we paint these types by gray.
+These tables show that four types are excluded in an exposed group and five types are excluded in an unexposed group under Fuzzy DID.
+Note that if we exclude $DF^{Z}$ and $CM^{T}$ by Assumption 37 and Assumption 38 (under $z=0$), the same number of types are excluded in exposed and unexposed groups, respectively.
+If we exclude $CM^{Z}$ and $DF^{T}$ by Assumptions 37 and 38 (under $z=0$), four types are excluded in an exposed group and six types are excluded in an unexposed group.
+Table 11: Exposed group ($z=1$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
+$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
+$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
 Table 12: Unexposed group ($z=0$)
 observed counterfactual
-$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
+[…]
 $D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
@@
 $D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
-Then, from Table 5, we can interpret the SLATET’ (see Assumption 31) as the treatment effects among the time compliers in an unexposed group in period $1$:
-$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{1}(0)\neq D_{0}(0)]$
-$=E[Y(1)-Y(0)|Z=0,T=1,CM^{T}].$ (32)
-By combining equation (31) with (32), one can see that Assumption 31 ($STATET=SLATET^{\prime}$) does not require the homogeneous treatment effects in the same sub-population between the two groups.
-Rather, this assumption requires that a weighted average of the treatment effects among the time compliers $CM^{T}$ and the type $(CM^{Z}\land NT^{T})$ in an exposed group should be equal to the treatment effect among the time compliers $CM^{T}$ in an unexposed group.
-This interpretation implies that the requirement of the homogeneous treatment effect in Assumption 31 is less clear than what one might expect.
-E.4   Revisiting the issue regarding the use of the Wald-DID estimand
-Finally, we revisit the issue pointed by dCDH regarding the use of the Wald-DID estimand, and show that this problem is not serious than one might expect, and does not arise under our DID-IV designs.
-We also clarify why dCDH exclude so many types in exposed and unexposed groups as seen in Tables 3-4.
-Hereafter, we assume Assumptions 26-30 in Fuzzy DID designs.
-dCDH argue that the Wald-DID estimand requires Assumption 30 to identify the SLATET, and propose alternative estimands that do not rely on this assumption.
-We first show that their argument depends on the parallel trends assumption in DID designs.
-Next, we argue that this assumption would be difficult to justify in DID-IV settings.
-First, we clarify the interpretation of Assumption 30 by introducing the additional notation.
-This assumption is actually conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why this strong restriction is required only for the sub-population.
-For each time group $g^{t}$, let $w_{g^{t},z}=Pr(G^{T}=g^{t}|Z=z)$ denote the population share conditional on group $Z=z$, and let $w_{g^{t},z,D_{0}(0)=1}=Pr(G^{T}=g^{t}|Z=z,D_{0}(0)=1)$ denote the population share conditional on $D_{0}(0)=1$ in addition to group $Z=z$.
-Additionally, let $\delta_{z,t,g^{t}}=E[Y(1)-Y(0)|Z=z,T=t,G^{T}=g^{t}]$ denote the average treatment effect conditional on group $Z=z$, time $T=t$, and time group $G^{T}=g^{t}$.
-Then, from Tables 3-4, we can decompose Assumption 30 as follows:
-$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$
-$=$ $E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
-$\iff$ $\delta_{0,1,AT^{T}}=\delta_{0,0,AT^{T}},$ (33)
-$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$
-$=$ $E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
-$\iff$ $\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
-$=$ $\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1},$ (34)
-where we use the Law of Iterated Expectation, $D_{0}(0)=D_{0}(1)$ (implied by the sharp assignment of the instrument), and $V\mathop{\perp\!\!\!\!\perp}T|Z$ in Assumption 28.
-Equation (33) shows that in an unexposed group ($z=0$), Assumption 30 requires the stable treatment effect for the time always-takers $AT^{T}$.
-Equation (34) shows that in an exposed group ($z=1$), this assumption requires that a weighted average of the treatment effects among the time defiers $DF^{T}$ and the time always takers $AT^{T}$ is stable during the two periods.
-Conditions (33) and (34) are required for Fuzzy DID designs because dCDH rely on the parallel trends assumption in the untreated outcome.
-In DID-IV settings, we can not observe the average time trends of untreated potential outcomes for the time always takers $AT^{T}$ in an unexposed group and the time always takers $AT^{T}$ and the time defiers $DF^{T}$ in an exposed group because they always adopt the treatment during the two periods.
-Therefore, if we assume the parallel trends assumption in the untreated outcomes, we should recover them by using conditions (33) and (34).
-To see this formally, we use our notation to briefly sketch dCDH’s proof about how the Wald-DID estimand identifies the SLATET under fuzzy DID designs.
-First, consider the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$, which appears in the numerator of the Wald-DID estimand.
-We can decompose this quantity into two terms.
+Notes: These tables represent mutually exclusive and exhaustive types under Fuzzy DID designs.
+The types painted in gray color are excluded by Lemmas 2 and 3.
+The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
+D.2.4 Comparison
+By comparing Tables 9-10 with Tables 11-12, we obtain two implications.
+First, the restrictions imposed under Fuzzy DID are stronger than those under DID-IV.
+Second, while the restrictions under DID-IV are symmetric between exposed and unexposed groups, they are asymmetric under Fuzzy DID.
+Note that these differences are driven by differences in the identifying assumptions between the two designs.
+Assumption 33, which is stronger than Assumption 20, imposes time monotonicity and thus rules out one of the two time-switching types ($CM^{T}$ or $DF^{T}$).
+The condition (20) in Assumption 31, which DID-IV does not impose, rules out the type $NT^{Z}\land DF^{T}$ in the exposed group and the type $DF^{Z}\land AT^{T}$ in the unexposed group.
+Finally, Assumption 32, combined with Assumption 20, eliminates both $DF^{T}$ and $CM^{T}$ in the unexposed group under Fuzzy DID.
+D.3  The difference in target parameter
+In this subsection, building on Tables 11-12, we next clarify the difference between dCDH’s target parameter (SLATET) and ours (LATET).
+Specifically, we decompose the SLATET under the identifying assumptions in Fuzzy DID and discuss its implications.
+We prove Theorem 6 below.
+Theorem 6.
+Suppose Assumptions 31-35 hold. Then, we can decompose the SLATET as follows, which depends on the direction of the two monotonicity assumptions 37 and 38 (under $z=0$).
+  • (i)
+    Under the monotonicity assumptions that exclude $DF^{Z}$ and $DF^{T}$, we have:
+    $SLATET\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
+    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}]Pr(AT^{Z}\land CM^{T}|Z=1,SW)$
+    $+E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}]Pr(CM^{Z}\land NT^{T}|Z=1,SW).$ (24)
+  • (ii)
+    Under the monotonicity assumptions that exclude $DF^{Z}$ and $CM^{T}$, we have:
+    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
+    $=E[Y(1)-Y(0)|Z=1,T=1,CM^{Z}\land NT^{T}].$
+  • (iii)
+    Under the monotonicity assumptions that exclude $CM^{Z}$ and $DF^{T}$, we have:
+    $SLATET$ $\equiv E[Y(1)-Y(0)|Z=1,T=1,D_{0}(0)<D_{1}(1)]$
+    $=E[Y(1)-Y(0)|Z=1,T=1,AT^{Z}\land CM^{T}].$
+Proof.
+Regarding the case (i), the proof directly follows from Table 11.
+We can also construct the table in an exposed group for the cases (ii) and (iii), which ensures the above interpretation of the SLATET.
+∎
+Theorem 6 has several important implications.
+First, the interpretation of the SLATET depends on the direction of the two monotonicity conditions 37 and 38 (under $z=0$).
+This implies that, if empirical researchers wish to understand which latent type’s causal effect is identified by the SLATET, they need to specify the direction of these monotonicity conditions ex ante.
+Second, the SLATET may fail to be a policy-relevant treatment parameter (Heckman and Vytlacil (2001)), even when the instrument represents the policy change of interest to the researcher.
+Specifically, under case (i), the SLATET is contaminated by the treatment effects among the type $AT^{Z}\land CM^{T}$, who are affected by time but not affected by the instrument.
+Under case (iii), the SLATET captures only the treatment effects among the type $AT^{Z}\land CM^{T}$.
+Third, the SLATET may be defined on a strictly smaller population than the LATET.
+In particular, under case (ii), the SLATET identifies the causal effect among the type $CM^{Z}\land NT^{T}$, which is a strict subpopulation of $CM^{Z}$.
+This implies that even when empirical researchers choose the direction of the two monotonicity conditions so that the SLATET is policy relevant, it will measure the treatment effects for a narrower population than the LATET.
+Remark 4.
+Using the decomposition result for the SLATET, we can also clarify the interpretation of Assumption 36, which dCDH introduce as an alternative to Assumption 32.
+First, we describe the heterogeneity in treatment adoption behavior in the unexposed group without Assumption 32.
+If we do not impose Assumption 32, we obtain Table 13 for the unexposed group.
+Here, we consider the case in which Assumptions 37 and 38 exclude $DF^{Z}$ and $DF^{T}$ in the unexposed group, respectively.
+Table 13: Unexposed group ($z=0$)
+observed counterfactual
+[…]
+$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
+$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
+$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
+$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
+Notes: This table represents mutually exclusive and exhaustive types in an unexposed group under Fuzzy DID without Assumption 32.
+The types painted in gray color are excluded by Lemma 4.
+The types painted in blue color are excluded by the two monotonicity assumptions 37 and 38 (under $z=0$).
+Then, from Table 13, we can interpret the SLATET’ (see Assumption 36) as the treatment effects among the type $AT^{Z}\land CM^{T}$ in an unexposed group:
+$SLATET^{\prime}$ $\equiv E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)\neq D_{1}(0)]$
+$=E[Y(1)-Y(0)|Z=0,T=1,AT^{Z}\land CM^{T}].$ (25)
+Combining (25) with (24), we see that Assumption 36 ($SLATET=SLATET^{\prime}$) does not require homogeneous treatment effects for the same latent type across exposed and unexposed groups.
+Rather, it equates a weighted average of treatment effects among types $AT^{Z}\land CM^{T}$ and $CM^{Z}\land NT^{T}$ in the exposed group with the treatment effect for type $AT^{Z}\land CM^{T}$ in the unexposed group.
+D.4  The use of the Wald-DID estimand
+In this section, we explain why the role of the Wald–DID estimand differs between Fuzzy DID and DID-IV designs.
+In DID-IV, we adopt the Wald–DID estimand as a natural estimand for identifying the LATET.
+By contrast, dCDH point out that, under Fuzzy DID, the Wald-DID estimand requires the stable treatment effect assumption (Assumption 35) to identify the SLATET.
+In the following discussion, we show that this difference arises because Fuzzy DID and DID-IV rely on different types of the parallel trends assumption.
+D.4.1 Why the Wald-DID estimand requires Assumption 35 under Fuzzy DID
+We begin by explaining why the Wald-DID estimand requires Assumption 35 under Fuzzy DID.
+We show that this requirement does not stem from the Wald-DID estimand itself, but rather from imposing the parallel trends assumption on untreated outcomes.
+To make this point precise, we first clarify Assumption 35 using latent treatment adoption types.
+This assumption is in fact conditional on $D_{0}(Z)=1$ in addition to $Z$ and $T$, but it is not immediately clear from dCDH why such a strong restriction is imposed only on this subpopulation.
+From Tables 11-12, we can reinterpret Assumption 35 as follows [^26] Under the identifying assumptions in Fuzzy DID, the type $DF^{Z}\land AT^{T}$ is eliminated in both groups.
+Therefore, we write the type $AT^{Z}\land AT^{T}$ as $AT^{T}$.:
+$E[Y(1)-Y(0)|Z=0,T=1,D_{0}(0)=1]$ $=E[Y(1)-Y(0)|Z=0,T=0,D_{0}(0)=1]$
+$\iff E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=0,T=0,AT^{T}],$ (26)
+$E[Y(1)-Y(0)|Z=1,T=1,D_{0}(1)=1]$ $=E[Y(1)-Y(0)|Z=1,T=0,D_{0}(1)=1]$
+$\iff E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]$ $=E[Y(1)-Y(0)|Z=1,T=0,AT^{T}].$ (27)
+Here, we use $D_{0}(0)=D_{0}(1)$ from the sharp assignment of the instrument $v_{10}=v_{00}$.
+Equations (26)-(27) show that Assumption 35 actually requires the stable treatment effect only for the time always-takers $AT^{T}$ within each group.
+We now show that Assumption 35 is not required by the Wald-DID estimand per se, but arises because dCDH impose the parallel trends assumption on untreated outcomes.
+The key point is that imposing the parallel trends assumption in untreated outcomes under the DID-IV setting forces us to impute the counterfactual untreated potential outcome for the time always-takers $AT^{T}$, which in turn requires imposing Assumption 35.
+To see this formally, we focus on the numerator of the Wald-DID estimand.
+Under Fuzzy DID, we can decompose the conditional expectations E[Y|Z=0,T=1]E[Y|Z=0,T=0]E[Y|Z=0,T=1]-E[Y|Z=0,T=0] and E[Y|Z=1,T=1]E[Y|Z=1,T=0]E[Y|Z=1,T=1]-E[Y|Z=1,T=0] as follows.
 $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$
-$=$ $(\delta_{0,1,AT^{T}}-\delta_{0,0,AT^{T}})\cdot w_{AT^{T},0}+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
-The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ in an unexposed group, and we should add and subtract this quantity to construct the second term.
-The first term disappears if we assume condition (33).
-Similarly, one can decompose the conditional expectation $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$ as follows.
+$=$ $(E[Y(1)-Y(0)|Z=0,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=0,T=0,AT^{T}])$
+$\times$ $Pr(AT^{T}|Z=0)+\{E[Y(0)|Z=0,T=1]-E[Y(0)|Z=0,T=0]\}.$
 $E[Y|Z=1,T=1]-E[Y|Z=1,T=0]$
-$=$ $\Big{\{}\delta_{1,1,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}+\delta_{1,1,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}$
-$-\delta_{1,0,AT^{T}}\cdot w_{AT^{T},1,D_{0}(0)=1}-\delta_{1,0,DF^{T}}\cdot w_{DF^{T},1,D_{0}(0)=1}\Big{\}}$
-$+$ $\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
-$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1),$
-where the type SW (switcher) includes the units who switch from non-treatment to treatment during the two periods.
-The first term arises because we can not observe the average time trends of the untreated potential outcome for the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ in an exposed group, and we should add and subtract this quantity to construct the second term.
-The first term disappears if we assume condition (34).
-Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term.
-Then, by noticing that the denominator of the Wald-DID estimand identifies $Pr(SW|Z=1)$ under Assumption 27 and Assumption 28, the Wald-DID estimand captures the SLATET under Assumption 29 (parallel trends assumption in the untreated outcome) and conditions (33) and (34).
-From the above discussion, one can see that imposing the parallel trends assumption in the untreated outcome leads dCDH to assume conditions (33) and (34).
-Here, we clarify why dCDH ex ante impose strong restrictions on treatment adoption behavior across units for the Wald-DID estimand to identify the SLATET.
-This also stems from their reliance on the parallel trends assumption in the untreated outcome.
-Remark E.4.1.
-dCDH exclude the type $NT^{Z}\land DF^{T}$ in an exposed group and the time defiers $DF^{T}$ in an unexposed group because if we add and subtract the average time trends of the untreated outcome for these types, we additionally identify the treatment effects among them in period 00 (not in period $1$), as they adopt the treatment in period 00, but leave it in period $1$ (see Tables 3-4).
-This identification result, however, is not intuitive because the identification variation comes from a policy shock (instrument) in period $1$.
-Remark E.4.2.
-dCDH exclude the time compliers $CM^{T}$ in an unexposed group in order to avoid the SLATET in an exposed group being offset by the SLATET’ in an unexposed group.
-If we add and subtract the average time trends of the untreated outcome for the time compliers, we additionally capture the SLATET’ in an unexposed group from the conditional expectation $E[Y|Z=0,T=1]-E[Y|Z=0,T=0]$.
-Thus far, we have confirmed that dCDH’s argument rests on the parallel trends assumption in the untreated outcome.
-Finally, we argue that this assumption would be difficult to rely on in DID-IV settings for three reasons.
-[…]
-Finally, this assumption is generally not testable using the pre-exposed period data, as some units can already adopt the treatment before period 00, and we cannot compare the pre-trend of the untreated outcome between exposed and unexposed groups.
-Although dCDH propose the placebo test in their appendix, it has shortcomings for checking the plausibility of the parallel trends assumption in the untreated outcome.
-dCDH’s test has two steps.
-[^31] Although dCDH ”assume” the conditions (35) and (36), they note that when one of the two conditions fails, the second test can ”no longer be used to test” Assumption 29.
-We, therefore, describe the dCDH’s test in two steps.
-First, we test whether the share of treated units is zero in both groups from $T=-1$ to $T=0$.
+$=$ $(E[Y(1)-Y(0)|Z=1,T=1,AT^{T}]-E[Y(1)-Y(0)|Z=1,T=0,AT^{T}])$
+$\times$ $Pr(AT^{T}|Z=1)+\{E[Y(0)|Z=1,T=1]-E[Y(0)|Z=1,T=0]\}$
+$+$ $E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1).$
+These decompositions clarify why Assumption 35 becomes necessary.
+Imposing parallel trends on untreated outcomes requires constructing the average time trends of the untreated potential outcome in both groups (the second term in each conditional expectation).
+Since these trends are not observed for time always-takers $AT^{T}$, this construction necessarily introduces the first term in the above decompositions, which captures changes in the average treatment effect among $AT^{T}$.
+For the Wald-DID estimand to identify the SLATET, one must therefore impose stable treatment effects for $AT^{T}$, as in Assumption 35.272727Here, we do not require the stable treatment assumption for the time compliers $CM^{T}$ and the type $CM^{Z}\land NT^{T}$ in an exposed group: by adding and subtracting the expectation of the untreated potential outcome path for these types, we have the third term E[Y(1)Y(0)|SW,Z=1,T=1]Pr(SW|Z=1)E[Y(1)-Y(0)|SW,Z=1,T=1]Pr(SW|Z=1), which is necessary for identifying the SLATET.
+Importantly, this restriction is not intrinsic to the Wald-DID estimand itself, but is a consequence of imposing parallel trends on untreated potential outcomes.
+D.4.2 Why the issue does not arise under DID-IV
+In contrast to Fuzzy DID, the issue described above does not arise under DID-IV.
+The key difference is that, in DID-IV designs, the parallel trends assumption is imposed on unexposed outcomes rather than untreated potential outcomes.
+As a result, the Wald-DID estimand does not require imputing the counterfactual untreated potential outcomes for time always-takers, and hence does not require Assumption 35.
+D.5  Which parallel trends assumption is suitable for the DID-IV setting
+In Subsection D.4, we showed that differences in the parallel trends (PT) assumption lead to different stances toward the use of the Wald-DID estimand across Fuzzy DID and DID-IV designs.
+In this subsection, we argue that the PT assumption on unexposed outcomes is more suitable for DID-IV settings in terms of (i) its alignment with the source of identifying variation, (ii) empirical relevance, and (iii) testability.
+Which PT assumption aligns with the source of identifying variation
+The PT assumption on unexposed outcomes is more consistent with the source of identifying variation in DID-IV settings than the PT assumption on untreated potential outcomes.
+This is because the latter relies on variation in treatment, while the former exploits variation in instrument.
+Note that in the canonical DID designs, we rely on the PT assumption in untreated outcomes because the units are partitioned into treatment and control groups based on the exposure to the treatment in the second period.
+In the DID-IV settings, however, the units are partitioned into exposed and unexposed groups based on the exposure to the instrument in the second period.
+Therefore, imposing the PT assumption on unexposed outcomes would be more natural than imposing the PT assumption on untreated potential outcomes in DID-IV settings.
+Empirical relevance
+In DID-IV settings, the parallel trends (PT) assumption on untreated potential outcomes is often implausible in practice, while the PT assumption on unexposed outcomes is more appealing.
+More precisely, because imposing parallel trends on untreated potential outcomes is often not attractive, empirical researchers exploit variation in exposure to an instrument and rely on parallel trends in unexposed outcomes in practice.
+A canonical example is Duflo (2001), who study the returns to schooling in Indonesia.
+The outcome is log annual earnings and the treatment is educational attainment.
+In this context, imposing PT on untreated potential outcomes would require that the average time trends of log earnings be the same across individuals in the absence of schooling.
+Given the strong selection into education, such an assumption is difficult to justify empirically.
+To address this concern, Duflo (2001) exploit the staggered rollout of a school construction program as an instrument for schooling and instead rely on the identifying assumption that, in the absence of the program, the evolution of wages and education across cohorts would not have differed systematically across regions.
+This identifying restriction corresponds precisely to the parallel trends assumptions on treatment and on unexposed outcomes in DID-IV designs.
+Testability
+In DID-IV settings, the PT assumption on unexposed outcomes is indirectly testable using pre-exposed period data as we discussed in Subsection 5.3.
+By contrast, the PT assumption on untreated potential outcomes is not testable in general.
+This is because some units may already adopt the treatment before period 0, which prevents a comparison of pre-trends in untreated outcomes between exposed and unexposed groups.
+From this perspective, the PT assumption on unexposed outcomes is more attractive to empirical researchers.
+Although dCDH propose a placebo test in their appendix, its scope for assessing the plausibility of the parallel trends assumption in untreated potential outcomes is somewhat limited.
+Their procedure consists of two steps.
+[^28] Although dCDH assume conditions (28) and (29), they note that when either condition fails, the second step can no longer be used to assess Assumption 34.
+We therefore describe their procedure in two steps.
+First, we examine whether the share of treated units remain unchanged in both groups from T=1T=-1 to $T=0$:
 $E[D|Z=1,T=0]-E[D|Z=1,T=-1]=0,$ (28)
@@
 Second, we test the following null hypothesis:
-$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$ (37)
+$E[Y|Z=1,T=0]-E[Y|Z=1,T=-1]$
 $=$ $E[Y|Z=0,T=0]-E[Y|Z=0,T=-1].$ (30)
-dCDH’s test has some weaknesses.
-First, the conditions (35) and (36), which should be checked in the first step, are strong, but no guidance is provided when these conditions are violated in practice.
-Second, in this test, the second step does not directly test Assumption 29 in the pre-exposed periods, as it does not compare the evolution of the mean untreated outcome between exposed and unexposed groups before period $T=0$.
-Indeed, dCDH describe the equality (37) as a necessary condition for Assumption 29 and Assumption 30 to be satisfied under the conditions (35) and (36).
-Finally, it is not clear how this test should be conducted with multiple pre-exposed period data.
-Although dCDH assume “for instance that data is available for the period $T=-1$”, it is not indicated how this test can be performed in multiple pre-exposed period settings.
-Surprisingly, dCDH’s test corresponds to our pre-trend test in Subsection 5.3 if we replace the conditions (35)-(36) in the first step with the weaker one.
+This testing strategy is informative under the maintained conditions (28) and (29), but its applicability may be limited in practice.
+First, these first-step conditions are strong, and the procedure does not provide guidance when they are violated.
+Second, the second step does not directly assess the evolution of the mean untreated outcome between exposed and unexposed groups in pre-exposed periods.
+Indeed, dCDH interpret (30) as a necessary condition for Assumptions 34 and 35 to hold under (28) and (29).
+Finally, while the discussion focuses on the case in which data are available for a single pre-exposed period, it is not clear how the procedure should be extended to settings with multiple pre-exposed periods.
+Notably, dCDH’s procedure corresponds to our pre-trend test in Subsection 5.3 once the first-step conditions (28)–(29) are replaced by the weaker condition:
 $E[D|Z=1,T=0]-E[D|Z=1,T=-1]$ (31)
 $=$ $E[D|Z=0,T=0]-E[D|Z=0,T=-1].$ (32)
-Under repeated cross section settings, the conditions (37)-(39) coincide with the conditions (1)-(2).
-This indicates that the rationale behind dCDH’s test is more suitable for the assessment of parallel trends assumptions in DID-IV designs.
-E.5   Other connections
-In this subsection, we document the relationship between DID-IV and Fuzzy DID under non-binary, ordered treatment, and multiple groups settings.
-We also compare the Wald-DID estimand to the time-corrected Wald estimand considered in dCDH.
-E.5.1 Non-binary, ordered treatment
-In this subsection, we investigate the relationship between DID-IV to Fuzzy DID under non-binary, ordered treatment settings.
-First, we compare the target parameter between the two designs.
-With a non-binary, ordered treatment $D\in\{0,1,\dots,J\}$, dCDH consider the following target parameter, which we call the switcher average causal response (SACR).
-$SACR\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(0)<j<D_{1}(1)),$
+Under repeated cross-section settings, conditions (30)–(31) coincide with conditions (1)–(2) in Subsection 5.3.
+This observation suggests that the underlying logic of dCDH’s placebo test is more aligned with the assessment of parallel trends assumptions in DID-IV designs.
+D.6  Other comparisons
+In this subsection, we document the relationship between DID-IV and Fuzzy DID in settings with non-binary, ordered treatments and multiple groups.
+We also discuss the alternative estimands proposed in dCDH that avoid Assumption 35.
+D.6.1 Non-binary, ordered treatment
+Again, we first document the similarities and differences between this paper and dCDH with respect to the empirical setting, the target parameter and the identifying assumptions.
+Empirical setting
+In the case of non-binary, ordered treatment, the condition (ii) of the DID-IV setting is satisfied in dCDH because the group variable $G$ is included in their ordered treatment equation (see Assumption 3’ in dCDH).
+By contrast, to the best of our knowledge, the condition (i) of the DID-IV setting is not imposed in dCDH: the threshold $v^{d}_{GT}$ in their ordered treatment equation is allowed to depend on $G$ in period $T=0$ for all $d\in\{0,\dots,\bar{d}\}$.
+This implies that the DID-IV setting differs from the setting considered in dCDH.
+Nevertheless, based on their empirical application, it is reasonable to infer that dCDH have the DID-IV setting in mind.
+In particular, dCDH revisit Duflo (2001), in which the outcome $Y$ is log annual earnings, the treatment $D$ is education attainment (an ordered treatment), and the group variable $G\in\{0,1\}$ indicates the regions that are highly exposed to the policy shock in the second period.
+Given this, the absence of condition (i) in dCDH can be interpreted as reflecting an additional restriction on treatment adoption behavior in period $T=0$.
+From the above discussion, in what follows, we replace $G$ with $Z$.
+The target parameter
+In this paper, we define the ACRT as our target parameter. By contrast, dCDH consider the following target parameter, which we call the switcher average causal response on the treated (SACRT).
+$SACRT\equiv\sum_{j=1}^{J}w_{j}\cdot E(Y(j)-Y(j-1)|Z=1,T=1,D_{0}(1)<j<D_{1}(1)),$
 where $w_{j}$ is
 $w_{j}=\frac{P(D\geq j|Z=1,T=1)-P(D\geq j|Z=1,T=0)}{E[D|Z=1,T=1]-E[D|Z=1,T=0]}.$
-[…]
-Next, we compare the identifying assumptions between the two designs. In order for the Wald-DID estimand to identify the SACR, dCDH replace Assumption 28 with the following one.
+Here, to ease the interpretation, we slightly modify the notation in dCDH.
+Specifically, we define $D_{t}(Z)=\sum_{j=1}^{J}\mathbf{1}\{V\geq v_{Zt}\}$ as the treatment status at time $T=t$ for group $Z$, rather than using the notation $D(t)$ in dCDH (see Assumption 39 below).
+The estimand
+Both papers consider the same Wald-DID estimand.
+This paper regards the Wald-DID estimand as the natural estimand for identifying the ACRT.
+By contrast, dCDH point out that this estimand identifies the SACRT only if the stable treatment effect assumption (see Assumption 40 below) is satisfied.
+The identifying assumptions
+In this paper, we consider the same identifying assumptions under non-binary, ordered treatment settings.
+By contrast, while dCDH maintain Assumption 31 and Assumption 34, they replace Assumption 33 with the following one.
 Assumption 39 (Ordered treatment equation).
 $D=\sum_{j=1}^{J}\mathbf{1}\{V\geq v^{j}_{ZT}\},\hskip 5.69054pt\text{with}\hskip 5.69054pt-\infty=v^{0}_{ZT}<v^{1}_{ZT}\dots<v^{J+1}_{ZT}=\infty\hskip 5.69054pt\text{and}\hskip 5.69054ptV\mathop{\perp\!\!\!\!\perp}T|Z.$
-In addition, dCDH replace the condition (27) in Assumption 26 with the condition (41), and Assumption 27 with the condition (42), respectively.
+In addition, they replace Assumption 32 with the following two conditions (33)-(34).
 $D_{11}\succeq D_{10},$ (33)
-$D_{01}\sim D_{00},$ (42)
-where $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
-Note that the condition (42) is stronger than Assumption 27 because this assumption requires that the distribution of the treatment should be stable during the two periods in an unexposed group.
-In our DID-IV designs, the identifying assumptions required for the Wald-DID estimand to capture the ACRT are unchanged under non-binary, ordered treatment settings.
-In addition, Assumptions 17-18 are indirectly testable using the pre-exposed period data in practice.
-E.5.2 Multiple groups
-In multiple groups settings, our DID-IV design is not directly compatible to Fuzzy DID design because dCDH exploit the group structure in the data ($G\in\{0,1,\dots,\bar{g}\}$) for the Wald-DID estimand to identify the following target parameter.
+$D_{01}\sim D_{00}.$ (34)
+Here, $D_{ZT}$ denotes the treatment conditional on $Z$ and $T$, $\succeq$ denotes stochastic dominance between two random variables, and $\sim$ denotes equality in distribution.
+Under non-binary, ordered treatment settings, Assumption 35 is formulated as follows:
+Assumption 40 (Stable treatment effect over time with non-binary, ordered treatment).
+For all $d\in\{0,1,\dots,J\}$,
+$E[Y(d)-Y(0)|Z,T=1,D_{0}(Z)=d]=E[Y(d)-Y(0)|Z,T=0,D_{0}(Z)=d].$
+Next, we compare the restrictions on treatment adoption behavior between this paper and dCDH.
+We also discuss the difference in target parameter.
+Finally, we discuss the different stances regarding the use of the Wald-DID estimand.
+Restrictions on treatment adoption behavior
+In DID-IV, we restrict the treatment adoption behavior through the monotonicity with respect to the instrument in the second period (Assumption 20).
+[^29] Note that dCDH do not impose this restriction, as Assumption 39 does not imply Assumption 20 in general.
+However, Assumption 39 implies the following restrictions: $v^{j}_{1t}\geq v^{j}_{0t}\hskip 5.69054pt\text{or}\hskip 5.69054ptv^{j}_{1t}\leq v^{j}_{0t}.$ for all $j\in\{1,\dots,J\}$ and all $t\in\{0,1\}$.
+In Fuzzy DID, dCDH impose the restrictions on treatment adoption behavior through the conditions (33)-(34).
+Specifically, as dCDH point out in their proof (see page 1025 in dCDH), these conditions combined with Assumption 39 imply:
+$v^{j}_{01}$ $=v^{j}_{00},$ (35)
+$v^{j}_{11}$ $\leq v^{j}_{10}$ (36)
+for all $j\in\{1,\dots,J\}$.
+Here, the condition (35) requires that the treatment rate should be stable over time for all units in the absence of exposure to the instrument.
+The condition (36) requires that the treatment should be weakly increasing over time for all units under exposure to the instrument.
+The difference in target parameter
+In general, the SACRT considered in dCDH differs from the ACRT defined in this paper.
+As in the case of the SLATET, the SACRT may not be policy relevant.
+This is because the type $SW^{{}^{\prime}}=\{Z=1,T=1,D_{0}(1)<j<D_{1}(1)\}$ is induced to increase the treatment by time, and not equal to the type $\{Z=1,T=1,D_{1}(0)<j<D_{1}(1)\}$ in general.
+Note that our target parameter, the ACRT, is the conditional version of the ACR defined in Angrist and Imbens (1995). As a result, the ACRT is policy-relevant parameter by construction.
+The use of the Wald-DID estimand
+dCDH point out that the stable treatment effect assumption (Assumption 40) is also required for the Wald–DID estimand to identify the SACRT.
+As in the binary treatment case, however, this requirement arises from imposing the parallel trends assumption on untreated outcomes.
+To see this formally, we first rewrite Assumption 40 as follows:
+$E[Y(d)\mid Z,T=1,D_{0}(Z)=d]-E[Y(d)\mid Z,T=0,D_{0}(Z)=d]$
+$=$ $E[Y(0)\mid Z,T=1,D_{0}(Z)=d]-E[Y(0)\mid Z,T=0,D_{0}(Z)=d].$
+From the above expression, this assumption can be interpreted as requiring that, for units with $D_{0}(Z)=d$, the average time trend of the potential outcome $Y(d)$ coincides with that of the untreated potential outcome $Y(0)$, conditional on $Z$.
+Recall that under Fuzzy DID with ordered treatments, Assumptions (33)–(34) imply that treatment adoption is stable over time in the unexposed group and weakly increasing over time in the exposed group.
+As a result, units with $D_{0}(0)=d$ in the unexposed group necessarily remain at treatment level $d$ under $T=1$, while in the exposed group such units can be partitioned into those whose treatment remains at $d$ and those whose treatment increases from $d$ to $d^{\prime}$.
+For units whose treatment remains constant over time, the left-hand side of Assumption 40 is observed.
+The assumption therefore uses this observed quantity to impute the unobserved average time trends of untreated potential outcomes for these units.
+For units in the exposed group whose treatment increases from $d$ to $d^{\prime}$, the same assumption additionally allows one to isolate the causal effect associated with the change from $d$ to $d^{\prime}$, which constitutes a component of the SACRT.
+In short, Assumption 40 is required not because of the Wald–DID estimand itself, but because imposing the parallel trends assumption on untreated outcomes forces one to impute the unobserved evolution of untreated potential outcomes for units with $D_{0}(Z)=d$.
+D.6.2 Multiple groups
+In multiple group settings, the DID-IV design considered in this paper is not directly compatible with the Fuzzy DID framework.
+There are two reasons.
+First, the group variable $G$ takes a finite number of ordered values: $G\in\{0,1,\dots,\bar{g}\}$.
+Second, dCDH construct the “super groups” (see dCDH) from the group structure in the data, $G\in\{0,1,\dots,\bar{g}\}$, and use the Wald-DID estimand to identify the following target parameter:
 $\Delta^{\ast}=E[Y(1)-Y(0)|\cup_{g=0}^{\bar{g}}S_{g},T=1],$
-where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ represents the switchers in group $G=g$ [^32] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
-[…]
-[…]
-In some applications, a researcher might prefer our approach to dCDH’s approach for three reasons.
-First, our approach yields a policy-relevant parameter by construction if the instrument represents a policy change of interest to the researcher.
-Second, in multiple period settings, our approach avoids the need to estimate the super groups in each consecutive period $(t-1,t)$, which simplifies both estimation and inference, making our method easier to implement in practice.
-Finally, in our approach, the monotonicity assumption can be easy to justify if the instrument is based on a policy shock.
-In dCDH’s approach, this assumption might be harder to defend in practice because the group variable $G$ typically represents the geographic feature in the data such as municipality, state, or country.
-[^34] Note that, in multiple groups settings, dCDH also maintain the monotonicity assumption on each group pair $(g,g^{\prime})$ through Assumption 28.
-Remark E.4.3 (Multiple groups settings with an ordered treatment).
-[…]
-E.5.3 Alternative estimand
-In this subsection, we compare the Wald-DID estimand to the alternative estimand proposed by dCDH that does not rely on Assumption 29 and 30.353535In this subsection, we do not consider Wald-CIC estimand proposed by dCDH, as we focus on the mean treatment effect in this article.
-As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand defined below.
-$W_{TC}\equiv\frac{E[Y|Z=1,T=1]-E[Y+(1-D)\delta_{0}+D\delta_{1}|Z=1,T=0]}{E[D|Z=1,T=1]-E[D|Z=1,T=0]},$
-where $\delta_{d}=E[Y|Z=0,T=1,D_{1}(0)=d]-E[Y|Z=0,T=0,D_{0}(0)=d]$.
-In order for the TC-Wald estimand to identify the SLATET, dCDH consider the following identifying assumption instead of imposing Assumption 29 and 30.
-Assumption 35 (Conditional common trends).
-For all $d\in\mathcal{S}(D),$
-$E[Y(d)|Z=1,T=1,D_{0}(0)=d]-E[Y(d)|Z=1,T=1,D_{0}(0)=d]$
-$=$ $E[Y(d)|Z=0,T=1,D_{0}(0)=d]-E[Y(d)|Z=0,T=1,D_{0}(0)=d].$
-Note that with a binary treatment, this assumption exploits the observed outcome path in an unexposed groups similar to Assumption 18 in our DID-IV designs (see Table 4).
-While the target parameter differs between the two designs, we see that using the Wald-DID estimand under DID-IV designs would be more appealing than using the TC-Wald estimand under Fuzzy DID designs for two reasons.
-First, when using the Wald-DID estimand in DID-IV designs, researchers can easily assess the plausibility of parallel trends assumptions in practice, employing the pre-trends tests described in Section 5.3.
-Although the TC-Wald estimand avoids the stable treatment effect assumption under Fuzzy DID designs, it presents challenges in testing the plausibility of Assumption 35 in pre-exposed period data: it is conditional on $D_{0}(0)=d$ and imposes restrictions on the mean evolution of $Y(d)$ for all $d\in\mathcal{S}(D)$.
-[…]
+where $S_{g}=\{D_{0}(g)\neq D_{1}(g),G=g\}$ denotes the set of switchers in group G=gG=g.
+[^30] Following $D_{t}(Z)$, we define $D_{t}(G)=\mathbf{1}\{V\geq v_{Gt}\}$ to be the treatment status under time $T=t$ in group $G$.
+However, even in multiple group settings, dCDH continue to impose Assumption 33.
+This implies that their framework still maintains two restrictions on treatment adoption behavior: monotonicity with respect to time $T$ and monotonicity with respect to the group variable $G$, in the sense of Imbens and Angrist (1994).
+Remark 5.
+While dCDH and its supplemental appendix study multiple groups and ordered treatments separately, to the best of our knowledge, we are not aware of the identification results for settings with multiple groups under ordered treatments in either the main text or the supplement.
+D.6.3 Alternative estimand
+As an alternative to the Wald-DID estimand, dCDH propose the time-corrected Wald (TC-Wald) estimand and the changes-in-changes Wald (CIC-Wald) estimand, which do not rely on Assumptions 34 and 35.
+Under both estimands, however, dCDH continue to impose Assumptions 31-33 to identify the SLATET.
+This implies that the same discussions in Subsections D.2-D.3 can be applied to both estimands regarding the restrictions on treatment adoption behavior and the target parameter, the SLATET (or SACRT).
+These estimands do not require Assumption 35 because Assumption 34 is not imposed for the time-always takers $AT^{T}$.
+D.7  Summary and guidance
+Throughout this section, we compare the DID-IV with Fuzzy DID considered in dCDH.
+To sum up, we obtain the following implications.
+First, the restrictions on treatment adoption behavior under Fuzzy DID are generally stronger than those under DID-IV.
+Second, the target parameter in Fuzzy DID, the SLATET, is not policy relevant in general, and even when it is policy relevant, it is defined on a strictly smaller population than the LATET.
+Finally, the Wald-DID estimand requires the stable treatment effect assumption under Fuzzy DID because dCDH imposes the parallel trends assumption on untreated outcomes.
+We conclude this section by summarizing practical considerations that may guide empirical researchers in choosing between DID-IV and Fuzzy DID designs under the DID-IV setting.
+Policy-relevant parameter
+If empirical researchers are interested in identifying a policy-relevant parameter, DID-IV designs are preferable.
+Under Fuzzy DID designs, the target parameter may not be policy relevant in general.
+If researchers instead focus on the SLATET (or the SACRT), it is recommended that they clarify why the causal effect for units whose treatment status is affected by time but not by the instrument is of substantive interest in their empirical context.
+Restrictions on treatment adoption behavior
+If empirical researchers are interested in identifying the SLATET (SACRT), they should carefully assess the plausibility of the restrictions on treatment adoption behavior imposed under Fuzzy DID designs.
+If these restrictions appear implausible, researchers may instead target the LATET (ACRT) and adopt DID-IV designs, under which the only restriction on treatment adoption behavior is monotonicity with respect to the instrument.
+Testability of the parallel trends assumption
+If empirical researchers wish to assess the plausibility of the parallel trends assumption, DID-IV designs are preferable.
+Under Fuzzy DID designs, Assumption 34 (and its counterparts for the alternative estimands) is typically difficult to assess using pre-exposure period data (see Subsection 1.1 of the Supplemental Appendix in dCDH).